A Complete Bibliography of Publications in IEEE Spectrum (1960–1969)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.01

Title word cross-reference

$250 [Ano65-322]. 3 [Ano66c, Raw69]. _{2} [Ano67-271].

-D [Ano66c, Raw69].

000-volt [Ano65-247].

1 [Ano64-211, Ano64-184, Ano64-267, Ano65a]. 1/4 [Ano65-391]. 1/4-inch [Ano65-391]. 10-billion-electron-volt [Ano65b]. 10.6-Micron [Ano68-297]. 11 [Ano65-420]. 12th [Ano64a]. 15 [Ano65-379]. 16 [Ano64-176]. 18 [Ano64-252, Ano64-246]. 19 [Ano65-444]. 1964 [Ano64-296, Ano64c, Ano65c, Mat65, Tow65]. 1965 [Ano65-271, Ano65-407, Ano65d, Ano65-350, Ano65-462, Cot66, Elf65]. 1966 [Ano66a, Ano66b, Ano66-312]. 1967 [Ano67-339]. 1968 [HE68]. 1969 [Tom69]. 1969-1970 [Ano69-370].

20th [Ano65e]. 21 [Ano64-293]. 22 [Ano65-279]. 24 [Ano65-336]. 25th [DR65]. 26 [Ano65e]. 27 [Ano64g]. 28 [Ano65-404]. 29 [Ano65-166].

3 [Ano65-299]. 30th [Ano65-462]. 32-year-old [Ano65-323].

4 [Ano65-235]. 4-inch [Ano65-391].

65 [Ano65-488]. ’66 [Ano65f]. 67 [Ano67-340]. 69 [Ano69-454]. 6th [Ano64g].

7 [Ano64-161, Ano65-417]. ’70 [Ano69-225]. 79th [Ano64-181].

8 [Ano64-304]. 8th [Ano64h].

9th [Ano65-406, Fri66a].

a-c [Ano65-141, Ano65-224, Ano65-225]. A. [Abr66, Ano65f, Wil65a]. ABM [Bre69, Rat69, Til69a]. abstracts [Ano64j, Ano64k, Ano64l, Ano64m, Ano64n, Ano64o, Ano64p, Ano64q, Ano64-197, Ano65g, Ano65h, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano65n, Ano65o, Ano66d, Ano66e, Ano66f, Ano66g, Ano66h, Ano66i, Ano66j, Ano66k, Ano66l, Ano66m, Ano66n, Ano66o, Ano67a, Ano67b, Ano67c, Ano67d, Ano67e, Ano67f, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l]. AC [BH64]. Accelerator [ML65, Ano65b]. access [McC67, SSC67]. accuracy [Ano64-356]. achieves [Ano64-150, Ano64-253]. Acknowledgment [Bic64a]. ACM [Ano65e]. Acoustic [Kor68]. acoustical [Fri67c]. Acoustics [Sch65]. Acoustooptical [Mal69]. Across [Ano67-273, Ano65-247]. acting [Ano67-274]. active [Hue69, Mit69, Tow69, Fow64]. active-filter [Tow69]. activities [Ano67-314]. Ad [WL64, Ano67-314]. Adaptive [Ano64i, Bic64b, Bic64a]. add [Ano65-132]. address [McC67]. adopt [Ano64-51]. Advance [Ano64j, Ano64k, Ano64l, Ano64m, Ano64n, Ano64o, Ano64p, Ano64q, Ano65g, Ano65h, Ano65i, Ano65j, Ano65k, Ano65l, Ano65m, Ano65n, Ano65o, Ano66d, Ano66e, Ano66f, Ano66g, Ano66h, Ano66i, Ano66j, Ano66k, Ano66l, Ano66m, Ano66n, Ano66o, Ano67a, Ano67b, Ano67c, Ano67d, Ano67e, Ano67f, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69l, Ano64-112, Ano64-197]. Advanced [Lin66c]. Advances [Ano69-418, Kra64]. Advertisement [Ano64r, Ano64s, Ano64t, Ano64u, Ano64v, Ano64w, Ano64x, Ano64y, Ano64z, Ano64-27, Ano64-28, Ano64-29, Ano64-30, Ano64-31, Ano64-32, Ano64-33, Ano64-34, Ano64-35, Ano64-36, Ano65-126, Ano65-127, Ano65-128, Ano65-129, Ano65-130, Ano65-131, Ano65-132, Ano65-133, Ano65-134, Ano65-135, Ano65-136, Ano65-137, Ano65-138, Ano65-140, Ano65-139, Ano65-141, Ano65p, Ano65q, Ano65-41, Ano65-42, Ano65-43, Ano65-44, Ano65-45, Ano65-46, Ano65-47, Ano65-48, Ano65-49, Ano65-50, Ano65-51, Ano65-52, Ano65-53, Ano65-54, Ano65-55, Ano65-56, Ano65-57, Ano65-58, Ano65-59, Ano65-60, Ano65-61, Ano65-62, Ano65-63, Ano65-64, Ano65r, Ano65-65, Ano65-66, Ano65-67, Ano65-68, Ano65-69, Ano65-70, Ano65-71, Ano65-72, Ano65-73, Ano65-74, Ano65-75, Ano65-76, Ano65-77, Ano65-78, Ano65-79, Ano65-80, Ano65-81, Ano65-82, Ano65-83, Ano65-84, Ano65-85, Ano65-86, Ano65-87, Ano65s, Ano65-88, Ano65-89]. Advertisement [Ano65-90, Ano65-91, Ano65-92, Ano65-93, Ano65-94, Ano65-95, Ano65-96, Ano65-97, Ano65-98, Ano65-99, Ano65-100, Ano65-101, Ano65-102, Ano65-103, Ano65-104, Ano65-105, Ano65-106, Ano65-107, Ano65-108, Ano65-109, Ano65-110, Ano65-111, Ano65t, Ano65-112, Ano65-113, Ano65-114, Ano65-115, Ano65-116, Ano65-117, Ano65-118, Ano65-119, Ano65-120, Ano65-121, Ano65-122, Ano65-123, Ano65-124, Ano65-125, Ano65u, Ano65v, Ano65w, Ano65x, Ano65y, Ano65z, Ano65-27, Ano65-28, Ano65-29, Ano65-30, Ano65-31, Ano65-32, Ano65-33, Ano65-34, Ano65-35, Ano65-36, Ano65-37, Ano65-38, Ano65-39, Ano65-40, Ano66-181, Ano66p, Ano66q, Ano66-40, Ano66-41, Ano66-42, Ano66-43, Ano66-44, Ano66-45, Ano66-46, Ano66-47, Ano66-48, Ano66-49, Ano66-50, Ano66-51, Ano66-52, Ano66-53, Ano66-54, Ano66-55, Ano66-56, Ano66-57, Ano66-58, Ano66-59, Ano66-60, Ano66-61, Ano66-62, Ano66-63, Ano66r, Ano66-64, Ano66-65, Ano66-66, Ano66-67]. Advertisement [Ano66-68, Ano66-69, Ano66-70, Ano66-71, Ano66-72, Ano66-73, Ano66-74, Ano66-75, Ano66-76, Ano66-77, Ano66-78, Ano66-79, Ano66-80, Ano66-81, Ano66-82, Ano66-83, Ano66-84, Ano66-85, Ano66-86, Ano66-87, Ano66s, Ano66-88, Ano66-89, Ano66-90, Ano66-91, Ano66-92, Ano66-93, Ano66-94, Ano66-95, Ano66-96, Ano66-97, Ano66-98, Ano66-99, Ano66-100, Ano66-101, Ano66-102, Ano66-103, Ano66-104, Ano66-105, Ano66-106, Ano66-107, Ano66-108, Ano66-109, Ano66-110, Ano66-111, Ano66-112, Ano66t, Ano66-113, Ano66-114, Ano66-115, Ano66-116, Ano66-117, Ano66-118, Ano66-119, Ano66-120, Ano66-121, Ano66-122, Ano66-123, Ano66-124, Ano66-125, Ano66-126, Ano66-127, Ano66-128, Ano66-129, Ano66-130, Ano66-131, Ano66-132, Ano66-133, Ano66-134, Ano66-135, Ano66-136, Ano66-137, Ano66-138, Ano66-139, Ano66-140, Ano66-141, Ano66-142, Ano66-143, Ano66-144, Ano66-145, Ano66-146, Ano66-147, Ano66-148, Ano66-149, Ano66-150, Ano66-151, Ano66-152, Ano66-153]. Advertisement [Ano66-154, Ano66-155, Ano66-156, Ano66-157, Ano66-158, Ano66-159, Ano66-160, Ano66-161, Ano66u, Ano66-162, Ano66-163, Ano66-164, Ano66-165, Ano66-166, Ano66-167, Ano66-168, Ano66-169, Ano66-170, Ano66-171, Ano66-172, Ano66-173, Ano66-174, Ano66-175, Ano66-176, Ano66-177, Ano66-178, Ano66-179, Ano66-180, Ano66v, Ano66w, Ano66x, Ano66y, Ano66z, Ano66-27, Ano66-28, Ano66-29, Ano66-30, Ano66-31, Ano66-32, Ano66-33, Ano66-34, Ano66-35, Ano66-36, Ano66-37, Ano66-38, Ano66-39, Ano67m, Ano67n, Ano67-37, Ano67-38, Ano67-39, Ano67-40, Ano67-41, Ano67-42, Ano67-43, Ano67-44, Ano67-45, Ano67-46, Ano67-47, Ano67-48, Ano67-49, Ano67-50, Ano67-51, Ano67-52, Ano67-53, Ano67-54, Ano67-55, Ano67-56, Ano67-57, Ano67-58, Ano67-59, Ano67-60, Ano67-61, Ano67o, Ano67-62, Ano67-63, Ano67-64, Ano67-65, Ano67-66, Ano67-67, Ano67-68, Ano67-69, Ano67-70, Ano67-71, Ano67-72, Ano67-73, Ano67-74, Ano67-75, Ano67-76]. Advertisement [Ano67-77, Ano67-78, Ano67-79, Ano67-80, Ano67-81, Ano67-82, Ano67-83, Ano67-84, Ano67-85, Ano67-86, Ano67p, Ano67-87, Ano67-88, Ano67-89, Ano67-90, Ano67-91, Ano67-92, Ano67-93, Ano67-94, Ano67-95, Ano67-96, Ano67-97, Ano67-98, Ano67-99, Ano67-100, Ano67-101, Ano67-102, Ano67-103, Ano67-104, Ano67-105, Ano67-106, Ano67-107, Ano67-108, Ano67-109, Ano67-110, Ano67q, Ano67-111, Ano67-112, Ano67-113, Ano67-114, Ano67-115, Ano67-116, Ano67-117, Ano67-118, Ano67-119, Ano67-120, Ano67-121, Ano67-122, Ano67-123, Ano67-124, Ano67-125, Ano67-126, Ano67-127, Ano67-128, Ano67-129, Ano67-130, Ano67-131, Ano67-132, Ano67-133, Ano67-134, Ano67-135, Ano67r, Ano67-136, Ano67-137, Ano67-138, Ano67-139, Ano67-140, Ano67-141, Ano67-142, Ano67-143, Ano67-144, Ano67-145, Ano67-146, Ano67-147, Ano67-148, Ano67-149, Ano67-150, Ano67-151, Ano67-152, Ano67-153, Ano67-154, Ano67-155, Ano67-156, Ano67-157, Ano67-158, Ano67-159, Ano67s, Ano67-160]. Advertisement [Ano67-161, Ano67-162, Ano67-163, Ano67-164, Ano67-165, Ano67-166, Ano67-167, Ano67-168, Ano67-169, Ano67-170, Ano67-171, Ano67-172, Ano67-173, Ano67-174, Ano67-175, Ano67-176, Ano67-177, Ano67-178, Ano67-179, Ano67-180, Ano67-181, Ano67-182, Ano67-183, Ano67t, Ano67-184, Ano67-185, Ano67-186, Ano67-187, Ano67-188, Ano67-189, Ano67-190, Ano67-191, Ano67-192, Ano67-193, Ano67-194, Ano67-195, Ano67-196, Ano67-197, Ano67-198, Ano67-199, Ano67-200, Ano67-201, Ano67-202, Ano67-203, Ano67-204, Ano67u, Ano67v, Ano67w, Ano67x, Ano67y, Ano67z, Ano67-27, Ano67-28, Ano67-29, Ano67-30, Ano67-31, Ano67-32, Ano67-33, Ano67-34, Ano67-35, Ano67-36, Ano68m, Ano68n, Ano68-38, Ano68-39, Ano68-40, Ano68-41, Ano68-42, Ano68-43, Ano68-44, Ano68-45, Ano68-46, Ano68-47, Ano68-48, Ano68-49, Ano68-50, Ano68-51, Ano68-52, Ano68-53, Ano68-54, Ano68-55, Ano68-56, Ano68-57, Ano68-58, Ano68-59, Ano68-60, Ano68-61, Ano68o]. Advertisement [Ano68-62, Ano68-63, Ano68-64, Ano68-65, Ano68-66, Ano68-67, Ano68-68, Ano68-69, Ano68-70, Ano68-71, Ano68-72, Ano68-73, Ano68-74, Ano68-75, Ano68-76, Ano68-77, Ano68-78, Ano68-79, Ano68-80, Ano68-81, Ano68-82, Ano68-83, Ano68-84, Ano68p, Ano68-85, Ano68-86, Ano68-87, Ano68-88, Ano68-89, Ano68-90, Ano68-91, Ano68-92, Ano68-93, Ano68-94, Ano68-95, Ano68-96, Ano68-97, Ano68-98, Ano68-99, Ano68-100, Ano68-101, Ano68-102, Ano68-103, Ano68-104, Ano68-105, Ano68-106, Ano68-107, Ano68-108, Ano68q, Ano68-109, Ano68-110, Ano68-111, Ano68-112, Ano68-113, Ano68-114, Ano68-115, Ano68-116, Ano68-117, Ano68-118, Ano68-119, Ano68-120, Ano68-121, Ano68-122, Ano68-123, Ano68-124, Ano68-125, Ano68-126, Ano68-127, Ano68-128, Ano68-129, Ano68-130, Ano68-131, Ano68-132, Ano68r, Ano68-133, Ano68-134, Ano68-135, Ano68-136, Ano68-137, Ano68-138, Ano68-139, Ano68-140, Ano68-141, Ano68-142, Ano68-143, Ano68-144, Ano68-145, Ano68-146]. Advertisement [Ano68-147, Ano68-148, Ano68-149, Ano68-150, Ano68-151, Ano68-152, Ano68-153, Ano68-154, Ano68-155, Ano68-156, Ano68s, Ano68-157, Ano68-158, Ano68-159, Ano68-160, Ano68-161, Ano68-162, Ano68-163, Ano68-164, Ano68-165, Ano68-166, Ano68-167, Ano68-168, Ano68-169, Ano68-170, Ano68-171, Ano68-172, Ano68-173, Ano68-174, Ano68-175, Ano68t, Ano68u, Ano68v, Ano68w, Ano68x, Ano68y, Ano68z, Ano68-27, Ano68-28, Ano68-29, Ano68-30, Ano68-31, Ano68-32, Ano68-33, Ano68-34, Ano68-35, Ano68-36, Ano68-37, Ano69m, Ano69n, Ano69-39, Ano69-40, Ano69-41, Ano69-42, Ano69-43, Ano69-44, Ano69-45, Ano69-46, Ano69-47, Ano69-48, Ano69-49, Ano69-50, Ano69-51, Ano69-52, Ano69-53, Ano69-54, Ano69-55, Ano69-56, Ano69-57, Ano69-58, Ano69-59, Ano69-60, Ano69-61, Ano69-62, Ano69-63, Ano69-64, Ano69o, Ano69-65, Ano69-66, Ano69-67, Ano69-68, Ano69-69, Ano69-70, Ano69-71, Ano69-72, Ano69-73, Ano69-74, Ano69-75, Ano69-76]. Advertisement [Ano69-77, Ano69-78, Ano69-79, Ano69-80, Ano69-81, Ano69-82, Ano69-83, Ano69-84, Ano69-85, Ano69-86, Ano69-87, Ano69-88, Ano69-89, Ano69-90, Ano69p, Ano69-91, Ano69-92, Ano69-93, Ano69-94, Ano69-95, Ano69-96, Ano69-97, Ano69-98, Ano69-99, Ano69-100, Ano69-101, Ano69-102, Ano69-103, Ano69-104, Ano69-105, Ano69-106, Ano69-107, Ano69-108, Ano69-109, Ano69-110, Ano69-111, Ano69-112, Ano69-113, Ano69-114, Ano69-115, Ano69-116, Ano69q, Ano69-117, Ano69-118, Ano69-119, Ano69-120, Ano69-121, Ano69-122, Ano69-123, Ano69-124, Ano69-125, Ano69-126, Ano69-127, Ano69-128, Ano69-129, Ano69-130, Ano69-131, Ano69-132, Ano69-133, Ano69-134, Ano69-135, Ano69-136, Ano69-137, Ano69-138, Ano69-139, Ano69-140, Ano69-141, Ano69-142, Ano69r, Ano69-143, Ano69-144, Ano69-145, Ano69-146, Ano69-147, Ano69-148, Ano69-149, Ano69-150, Ano69-151, Ano69-152, Ano69-153, Ano69-154, Ano69-155, Ano69-156, Ano69-157, Ano69-158, Ano69s, Ano69t, Ano69u]. Advertisement [Ano69v, Ano69w, Ano69x, Ano69y, Ano69z, Ano69-27, Ano69-28, Ano69-29, Ano69-30, Ano69-31, Ano69-32, Ano69-33, Ano69-34, Ano69-35, Ano69-36, Ano69-37, Ano69-38, Ano64i, Ano64-135, Ano64-148, Ano64-146, Ano64-147, Ano64-268, Ano65-233, Ano65-270, Ano65-338, Ano65-387, Ano65a, Ano65-388, Ano65-442, Ano65-487, Ano66-308, Ano66-310, Ano66-284, Ano66-285, Ano66-286, Ano66-287, Ano66-288, Ano66-289, Ano67-207, Ano67-315, Ano67-316, Ano67-317, Ano67-318, Ano67-319, Ano67-320, Ano67-321, Ano68-256, Ano68-293, Ano68-298, Ano68-299, Ano68-300, Ano69-418]. Advertisements [Ano64-40, Ano64-41, Ano64-42, Ano64-43, Ano64-44, Ano64-45, Ano64-46, Ano64-47, Ano64-48, Ano64-49, Ano64-50, Ano64-37, Ano64-38, Ano64-39, Ano65-142, Ano65-143, Ano65-144, Ano65-145, Ano64-298, Ano65-385, Ano66-354, Ano66-355, Ano66-356, Ano66-357, Ano66-358, Ano66-359, Ano66-360, Ano66-361, Ano66-362, Ano66-363, Ano66-364, Ano67-385, Ano67-386, Ano67-387, Ano67-388, Ano67-389, Ano67-390, Ano67-391, Ano67-392, Ano67-393, Ano67-394, Ano67-395, Ano67-396, Ano68-358, Ano68-359, Ano68-360, Ano68-361, Ano68-362, Ano68-363, Ano68-364, Ano68-365, Ano68-366, Ano68-367, Ano68-368, Ano68-369, Ano69-346, Ano69-347, Ano69-348, Ano69-349, Ano69-350, Ano69-351, Ano69-352, Ano69-353, Ano69-354, Ano69-355, Ano69-356, Ano69-357]. advertisers [Ano64-199, Ano64-200, Ano64-201, Ano64-202, Ano64-203, Ano64-204, Ano64-205, Ano64-206, Ano64-207, Ano64-208, Ano64-209, Ano64-210, Ano65-287, Ano65-288, Ano65-289, Ano65-290, Ano65-291, Ano65-292, Ano65-293, Ano65-294, Ano65-295, Ano65-296, Ano65-297, Ano65-298, Ano66-294, Ano66-295, Ano66-296, Ano66-297, Ano66-298, Ano66-299, Ano66-300, Ano66-301, Ano66-302, Ano66-303, Ano66-304, Ano66-305, Ano67-324, Ano67-325, Ano67-326, Ano67-327, Ano67-328, Ano67-329, Ano67-330, Ano67-331, Ano67-332, Ano67-333, Ano67-334, Ano67-335, Ano68-302, Ano68-303, Ano68-304, Ano68-305, Ano68-306, Ano68-307, Ano68-308, Ano68-309, Ano68-310, Ano68-311, Ano68-312, Ano68-313, Ano69-278, Ano69-279, Ano69-280, Ano69-281, Ano69-282, Ano69-283, Ano69-284, Ano69-285, Ano69-286, Ano69-287, Ano69-288, Ano69-289]. advertising [Ano65-199, Ano65-200, Ano65-201, Ano65-202, Ano65-203, Ano65-204, Ano65-205, Ano66-233, Ano66-234, Ano67-269, Ano67-270, Ano68-242, Ano68-243, Ano68-244, Ano68-245, Ano68-246, Ano68-247, Ano68-248, Ano68-249, Ano68-250, Ano68-251, Ano69-214, Ano69-215, Ano69-216, Ano69-217, Ano69-218, Ano69-219, Ano69-220, Ano69-221]. Aerospace [Ano65-183, Sor65]. AES [Ano65f]. affect [Ano64-346]. AFIPS [Ano64-177]. Against [Rat69]. age [Ano64-315, BL68, Lic68]. Agency [Ano64-146, Ano65-127, Ano65-128, Ano65-129]. ago [Ano64-316, Ano64-317]. ahead [Oli65a]. aided [Sid67]. AIL [Ano67-207, Ano67-205]. AIP [Ano64-136]. Air [Ano69-224, Net69, Ano64-140, Col66b, Nas68, Pea66, Sav69, Spe66]. Airborne [Fri65a]. Aircraft [FK66]. Airlines [Ano64-51]. Albany [Ano65-350]. Alessandro [Hen65]. Alfred [Ano65-407]. algebra [Nod65]. Allen [Gut65]. alloys [Esa66]. aluminum [Ano64-159]. Am [Ano64-147]. America [Mac67a]. American [Ano64-180, Ano65-363, Ano65-246, Ano65-280, Wor64]. amplification [Ano64-253]. amplifier [Ano65-338]. Amplifiers [Ano69-332, Ano66-281]. Analog [Hea64]. Analysis [CR69, FK64, Hac65, MOV66]. analysts [Ano65-140]. Analytical [Vog69]. Anatomy [Ano68-176]. ancestors [Pra68]. Angeles [Ano64g]. Angle [Ano67-255]. anniversary [DR65, Zve66]. announces [Ano64-181]. Annual [Ano64a, Ano64-53, Ano65-146, Ano66-182, Ano67-206, Ano68-177, Ano69-159, Spr65, Ano64-176, Ano65-406, Ano65-417, SSF65, Ano64-52, Ano64-138]. Annular [Ano64-268]. answer [Fri64b]. Antenna [Ano65-147, Ano66-314, Ano69-226, Bic64b, Bic64a, Chi66]. Antenna-Coupling [Ano65-147]. Antennas [Ano68-371, Ano68-476, All64, Ano64i, Ano65-443, JDDM64]. antiballistic [Til69a]. aperture [BP69b]. Apollo [Eva69]. Appliance [Ano64g]. Application [Ano67-344, Ano68-357, Bla65, Gro65, Hue69]. Applications [Har69, All64, Ano64-211, Ano64-249, Ano64-356, Ano69-300, Ano69-301, Ano69-302, Ano69-303, CB66, Dix68, Fra64a, FK64, Fra64b, Fra64c, Hau68, Hor68a, Mar66, Pri66, FK64]. Applied [Vol67, Ano65-364]. appointed [Ano64-177, Ano65-223]. approach [Fan65]. Approaches [Ano64-55, Mal69]. Approaching [CN67]. approves [Ano64-183, Ano65-232]. April [Ano64-267, Ano64-304, Ano65-404]. Area [Fri65c]. argon [Paa66]. Argonne [Ano64-269]. Array [All64]. arrays [Bic64b, Bic64a, Che68]. Art [Ano67-207, Fri69a, Lin69b, Lin69a, Ano64-56, Ano64-57, Ano64-58, Ano64-59, Ano64-60, Ano64-61, Ano64-62, Ano64-63, Ano64-64, Ano64-65, Ano64-66, Ano64-67, Fin66, Hoc69, Jam65]. Articulated [Sho66]. artificial [Ano64-134, Fei64a, Lin65a, Pag69]. ASCE [Ano65-148]. ASEE [Ano65-438, Ano68-301]. ASME [Ano65-350]. Asmus [Ano64-177]. aspects [Dom69, Hir67, Hor68b, Kee65, Pie68b]. asphyxia [Fri65a]. ASQC [Ano64-179]. assembly [Ano64h]. astronaut [Ano66-366]. Astronomical [AKAG66, Ano64-248]. astronomy [Kra64, MGC65a, MGC65b, Sha68a]. Astrophysics [Gut65]. Asynchronized [Alg64]. Atmosphere [AKAG66, Ano66-353, Coo66]. Atmospheric [AFW68, BL65a, BL65b]. Atomic [Ano68-241]. atoms [Tow65]. attend [Ano64-132]. attended [Ano64-265, Ano64h, Ano65-351]. attitude [Her68a]. audio [Hin69]. August [Ano65e]. Author [Ano64-68, Ano67-220, Ano69-160]. Authors [Ano64-72, Ano64-73, Ano64-74, Ano64-75, Ano64-76, Ano64-77, Ano64-78, Ano64-79, Ano64-80, Ano64-69, Ano64-70, Ano64-71, Ano65-150, Ano65-151, Ano65-152, Ano65-153, Ano65-154, Ano65-155, Ano65-156, Ano65-157, Ano65-158, Ano65-159, Ano65-160, Ano65-149, Ano66-185, Ano66-186, Ano66-187, Ano66-188, Ano66-189, Ano66-190, Ano66-191, Ano66-183, Ano66-192, Ano66-193, Ano66-194, Ano66-184, Ano67-210, Ano67-211, Ano67-212, Ano67-213, Ano67-214, Ano67-215, Ano67-216, Ano67-217, Ano67-218, Ano67-219, Ano67-208, Ano67-209, Ano68-179, Ano68-180, Ano68-181, Ano68-182, Ano68-183, Ano68-184, Ano68-185, Ano68-186, Ano68-187, Ano68-188, Ano68-189, Ano68-178, Ano65-308, Ano66-406]. Autobiographical [Ano69-375]. autobiography [Pra68]. Automatic [Lin65b, Rud67, AOB64, Cla65, Coh65, Lam66a, Sal65c]. Automation [Ano68-190, Fri68a, Lef64, Fri68g]. automobile [Ano65-130]. Automobiles [Ano67-268]. Automotive [Ano64-299]. available [AWSB67, Ano64-162, Ano64e, Ano64-297, Ano64-296, Ano64-322, Ano65-332]. avoidance [FK66]. award [Ano64-250, Ano65-271]. awards [Ano64-267].

B.S. [Ano64-251]. baccalaureate [Ano64-295]. Back [Ano64-81, Ano64-82, Ano64-83, Ano64-84, Ano64-85, Ano64-86, Ano64-87, Ano64-88, Ano64-89, Ano64-90, Ano64-91, Ano65-161, Ano65-162, Ano65-163, Ano65-164, Ano65-165, Ano64-212, Ano64-213, Ano64-214, Ano64-215, Ano64-216, Ano64-217, Ano64-218, Ano64-219, Ano64-220, Ano64-221, Ano64-222, Ano65-309, Ano65-310, Ano65-311, Ano65-312, Ano65-313]. balanced [Mic67, Sho66]. balances [Vog69]. balancing [Dia69]. Band [Ano68-315, Ano69-298]. bandwagon [Moo68]. Bandwidth [Ano67-370, Ano68-253, Ano68-296, Ano68-297, Ano69-299]. Based [Ano65-227, Ano65-228]. basic [Ano64-346, Ber64, SRLF64a]. Basin [Fri68f]. basis [RB64]. batch [Ano64-274]. batch-fabrication [Ano64-274]. batteries [Lie66]. battery [Hen65]. be [Ano64-158, Ano64-246, Ano64-267, Ano65-197, Ano65-222, Ano65-299, Ano65-333, Ano65-332, Ano65-350, Ano65-382, Ano65-418, Ano65-484, She66a, Sus67]. beam [Ano64-356, Hee64, Lot68, Ano66-195]. beams [LS64]. become [Ano66-366]. been [Ano65-249]. before [Sus68a]. begin [Ano64-184]. behavior [Ano64-141, Ano65-333]. being [Ano65b]. Bell [Ano65-126]. below [Ano64d, Gre67b]. Benjamin [Ano65-299]. Bern [Hen65]. Better [EC64, Ano69-374]. between [Adl67]. Beyond [BL68]. Biennial [Ano65-166]. billion [Ano65b]. binary [Len66]. binary-data [Len66]. binaural [Hir67]. biological [FKS64, MOV66]. biology [Ano64-111, Ano65-365, Lin67b]. Biomedical [Gol65, Ano64-297]. biophysics [Ano65-418]. biosphere [Mil68b]. bipolar [War67]. bird [Flo68]. Birth [Fri67a]. blanket [Fri66b]. board [Ano65-439]. boat [Ano64-131]. Boiler [Fri65b]. Boiler-turbine [Fri65b]. Bonneville [Fri68f, Fri68g]. Book [Ano64-92, Ano64-93, Ano64-94, Ano65-167, Ano65-168, Ano65-169, Ano65-170, Ano65-171, Ano65-172, Ano65-173, Ano65-174, Ano65-175, Ano66-196, Ano66-197, Ano66-198, Ano66-199, Ano66-200, Ano66-201, Ano66-202, Ano66-203, Ano66-204, Ano66-205, Ano66-206, Ano66-207, Ano67-221, Ano67-222, Ano67-223, Ano67-224, Ano67-225, Ano67-226, Ano67-227, Ano67-228, Ano67-229, Ano67-230, Ano67-231, Ano67-232, Ano68-191, Ano68-192, Ano68-193, Ano68-194, Ano68-195, Ano68-196, Ano68-197, Ano68-198, Ano68-199, Ano68-200, Ano68-201, Ano68-202, Ano69-161, Ano69-162, Ano69-163, Ano69-164, Ano69-165, Ano69-166, Ano69-167, Ano69-168, Ano69-169, Ano69-170, Ano69-171, Ano69-172, BDS^{+}64, Bon64, Boo64, Cho64, Cut64a, Dur65, FFC64, Fra64e, Gre64, Gut65, Hen65, LBCM65, Nod65, Not64, Pra68, Web64a]. Books [Ano69-418, Fri68a]. Boolean [Nod65]. born [Ano64-54]. both [Ano65-391]. bottom [Bra64]. Boulder [Ano64-342]. Boundary [Ano67-273]. brains [Lin68b]. branch [Ano64-182]. Breadboarding [Ano68-215]. breakdown [Nas68]. Breaker [Ano64-96, Ano65-177, Ano65-178, Ano65-179, Ano65-180, Ano65-181, Ano66-208, Ano66-209, Ano66-210, Ano66-211, Ano66-212, Ano66-213, Ano66-214, Ano66-215, Ano66-216, Ano66-217, Ano66-218, Ano66-219, Ano67-243, Ano67-244, Ano67-245, Ano67-246, Ano67-247, Ano67-248, Ano67-249, Ano67-250, Ano67-251, Ano67-252, Ano67-253, Ano67-254, Ano68-216, Ano68-217, Ano68-218, Ano68-219, Ano68-220, Ano68-221, Ano68-222, Ano68-223, Ano68-224, Ano68-225, Ano68-226, Ano68-227, Ano69-189, Ano69-190, Ano69-191, Ano69-192, Ano69-193, Ano69-194, Ano69-195, Ano69-196, Ano69-197, Ano69-198, Ano69-199, Ano69-200]. breeder [DBW69]. Brewster [Ano67-255]. bridge [Dia69, Oli64]. brief [Ano64d]. brightness [HS68]. Britain [Gab67]. broadband [JDDM64]. broadcasting [Gab67]. Brookyln [Ano65-182]. budget [Jac69]. building [Hoc69]. built [Ano65b]. Bulk [Cop67]. Bunk [Ano66-311]. Bureau [Har66]. bureaucratic [Ano64-300]. burst [Ano65-361]. Business [Ano65-183]. button [Ano64-303].

[Boo64, Ano65-141, Ano65-224, Ano65-225, Ano64-333, Ano69-201, LMT64, Sha68b]. C-Scan [Ano69-201]. C. [Ano64-292]. cable [Tay69]. cables [Mot65, OT66, RJ68, Wes66]. calculating [AOB64]. Calendar [Ano64-97, Ano64-109, Ano64-98, Ano64-99, Ano64-100, Ano64-101, Ano64-102, Ano64-103, Ano64-104, Ano64-105, Ano64-106, Ano64-107, Ano64-108, Ano64-263, Ano65-184, Ano65-185, Ano65-186, Ano65-187, Ano65-188, Ano65-189, Ano65-190, Ano65-191, Ano65-192, Ano65-193, Ano65-194, Ano65-195, Ano66-220, Ano66-221, Ano66-222, Ano66-223, Ano66-224, Ano66-225, Ano66-226, Ano66-227, Ano66-228, Ano66-229, Ano66-230, Ano66-231, Ano66-232, Ano67-256, Ano67-257, Ano67-258, Ano67-259, Ano67-260, Ano67-261, Ano67-262, Ano67-263, Ano67-264, Ano67-265, Ano67-266, Ano67-267, Ano68-228, Ano68-229, Ano68-230, Ano68-231, Ano68-232, Ano68-233, Ano68-234, Ano68-235, Ano68-236, Ano68-237, Ano68-238, Ano68-239, Ano69-202, Ano69-203, Ano69-204, Ano69-205, Ano69-206, Ano69-207, Ano69-208, Ano69-209, Ano69-210, Ano69-211, Ano69-212, Ano69-213]. Calibration [Ano69-275, Ano65-249]. California [Ano65-485]. call [Ano64-52, Ano64-345, Lin69b]. Camera [Ano65-196]. campus [And69]. can [Ano64-158, Hor64, Rob68b]. Canadian [Ano65-197, Kee65]. Capability [Ano68-296, DGH67]. capacitors [Ham69]. capture [Ano65-131, Ano65-360]. Carbon [Ano65-198]. Carbon-coated [Ano65-198]. Card [Ano68-412, Ano68-413, Ano69-400, Ano69-401, Ano69-402, Ano69-403, Ano64-178, Ano64-95, Ano65-272, Ano65-273, Ano65-274, Ano65-275, Ano65-276, Ano65-277, Ano69-173, Ano69-174, Ano69-175, Ano69-176, Ano69-177, Ano69-178, Ano69-179, Ano69-180, Ano69-181, Ano69-182, Ano69-183, Ano69-184, Ano69-185, Ano69-186, Ano69-187, Ano69-188]. Careers [Ano66-307, Ano66-308]. Carlo [Ano68-294]. Carrier [Ano69-201]. cars [Lin67a]. Case [Ano67-268, Atk65, Hal67]. cash [Hac65]. causes [Ano65-248]. CCITT [Lam66a]. celebrated [Ano64-342]. cell [Fry69, Pea66]. cells [Lie66]. centennial [Dib67]. central [Tsu67]. Ceramic [Ham69, SB68]. Ceramics [Ano68-240]. Challenge [Ano68-241, Smi66b, Gen64, Hof68, Wes66]. challenges [Web64b]. Chameleon [Smi66a]. change [Kop69, Whi69]. changing [And69, Bla64]. Channel [Ano69-297]. channels [Jam66, KL64]. chapters [Ano64-183, Ano65-232]. Charles [Gos68, WAS65]. Charter [Ano64-211]. Chas. [Ano64-273]. check [Ano64-140]. check-in [Ano64-140]. chemical [Ano65-248]. Circuit [Geh66, Ano64-341, Atk65, Cas64, Fow64, Jac69, LMT64]. circuitry [Hag64a]. circuits [Ano65-442, Ano65-420, Ano66-309, Ano67-313, Bro64c, CR69, Gre69, HHN^{+}64, Ham69, Hir68b, Hog64, Kno64, Mai64, Noy64, Phi64, SN69, War67]. Circulators [Ano65-335]. citizen [OFM^{+}69]. city [Fri67a, Ano64-247]. civilization [Hor66]. Classified [Ano65-199, Ano65-200, Ano65-201, Ano65-202, Ano65-203, Ano65-204, Ano65-205, Ano66-233, Ano66-234, Ano67-269, Ano67-270, Ano68-242, Ano68-243, Ano68-244, Ano68-245, Ano68-246, Ano68-247, Ano68-248, Ano68-249, Ano68-250, Ano68-251, Ano69-214, Ano69-215, Ano69-216, Ano69-217, Ano69-218, Ano69-219, Ano69-220, Ano69-221]. classroom [Ano66-366]. Clear [Col66b]. Clerk [Dib64, Boo64]. Cleveland [Ano65-166]. climate [Lau67]. close [Ano65-269]. closely [Gol66]. closer [Ano65-126]. CO [Ano67-271]. coated [Ano65-198]. coaxial [Ano66-281]. code [DR65]. coding [Len66]. Coefficients [Ano66-256]. coherent [Hun64, Las66, Tow65]. collaboration [Lin69b]. collect [Ano65-416]. collision [FK66]. Color [Hir68a, Ele68, HP67, McL66, Roi67, The66]. Columbia [Fri68f, Kee65]. come [Fri67b]. comes [Ano68-190, Fri68a]. coming [EC64]. command [Mat65]. Comments [Ano66-235]. Commercial [MM67, Ano65-417]. Commission [Kel67, Ano65-489, Oli67]. Committee [Ano64-295, Ano65-232, Ano67-314, WL64, Ano64-183]. committees [Ano64b]. Common [For64, Ano66-405]. communicate [Rob68b]. Communication [Ano66-311, WGC^{+}66, Ano65-361, Ano67-432, Cut64a, FL66, Hay64, Hir68b, McC67, McP67, Moo67, Pie66, Wes66]. Communications [Ano64-52, EK69, Hor66, Kel67, Lew66, Omi66, Spr65, Ano64-54, Ano65-249, Ano66-365, Ano68-190, BKW67, Coo66, DGH67, End66, Fra66, Fri68c, Gun66, Har69, MM67, May65, Wei66b, Wes66, Wes68]. Community [Chi66]. Compact [Ano67-271, Ano64-276]. company [NP69]. Comparing [War67]. comparison [Sel66]. Compatibility [Fra66, Wes66, Ano64-113, Koh69]. Compatible [SW67, The66]. compensates [Ano64-161]. Competition [Fri68c]. Complex [Ano67-397, Mel68]. Components [Ano66-310, Ano64-301]. comprehensive [Ano64-264]. compressed [Ano66-283]. compression [Ano64-198, Kor67]. computation [Tru64]. Computer [Com68, Fri65b, Fri65c, Fri69b, Hin69, Sid67, Ano64-300, Ano64-344, Ano67-411, Ano67-470, Ano68-319, Ano68-385, Bla65, Fan65, Hor64, Key69, Lic68, McP67, Nol67, RK65, ST67]. Computer-aided [Sid67]. Computer-controlled [Fri65b, Fri65c, Fri69b]. Computers [Nie68, Ano64-51, Ano64-149, Ano64-274, Cha67, Raj65, Sch69b, SSF65]. Computing [Ric67, Cut64b, Nie68]. Comsat [End66]. concept [Lew66, SRLF64a]. concepts [Lin66c]. concerning [Ano65-323, Ano65-380]. concert [Sch65]. conductance [Kro68]. conduction [Wei66a]. conductor [Ano64-150]. conductors [HHA^{+}66, Mon66]. Conference [Ano64-111, Ano64-113, Ano64-110, Ano64-112, Ano64-179, Ano65-206, Ano65e, Cot65, Fri66a, Hig66, Sus67, YW66, Ano64a, Ano64-132, Ano64-246, Ano64-293, Ano64-277, Ano64-278, Ano64-304, Ano64-340, Ano65-166, Ano65-229, Ano65-235, Ano65-299, Ano65-321, Ano65-351, Ano65-379, Ano65-406, Ano65-404, Ano65-417, Ano65-444, Ano65-420, Bla65, Cla65, Mir65, SSF65, Sor65, Tom64, Ano64-264, Ano64-265, Ano64-299, Ano64f, Ano64g, Ano65-350, Elf65, ML65]. Conferences [Ano67-339]. Confidence [Ano67-458]. confirms [Ano65-323]. congress [Ano64-345, Ano65-418, Lef64, Ano65-363]. Considerations [Con68]. constituent [Fra64b]. Consulting [Ano65-133, Ano65-134, Ano65-135, Ano65-136, Ano65-137, Ano65-138, Ano64-298, Ano65-385, Ano66-354, Ano66-355, Ano66-356, Ano66-357, Ano66-358, Ano66-359, Ano66-360, Ano66-361, Ano66-362, Ano66-363, Ano66-364, Ano67-385, Ano67-386, Ano67-387, Ano67-388, Ano67-389, Ano67-390, Ano67-391, Ano67-392, Ano67-393, Ano67-394, Ano67-395, Ano67-396, Ano68-358, Ano68-359, Ano68-360, Ano68-361, Ano68-362, Ano68-363, Ano68-364, Ano68-365, Ano68-366, Ano68-367, Ano68-368, Ano68-369, Ano69-346, Ano69-347, Ano69-348, Ano69-349, Ano69-350, Ano69-351, Ano69-352, Ano69-353, Ano69-354, Ano69-355, Ano69-356, Ano69-357]. consumer [Bro64c, SB68]. consumption [Fel64, VM67]. Contents [Ano64-124, Ano64-114, Ano64-115, Ano64-116, Ano64-117, Ano64-118, Ano64-119, Ano64-120, Ano64-121, Ano64-122, Ano64-123, Ano65-207, Ano65-208, Ano65-209, Ano65-210, Ano65-211, Ano64-339, Ano65-445, Ano65-446, Ano65-447, Ano65-448, Ano65-449, Ano65-450, Ano65-451, Ano66-407, Ano66-408, Ano66-409, Ano66-410, Ano66-411, Ano66-412, Ano66-413, Ano66-414, Ano66-415, Ano66-416, Ano66-417, Ano66-418, Ano67-433, Ano67-434, Ano67-435, Ano67-436, Ano67-437, Ano67-438, Ano67-439, Ano67-440, Ano67-441, Ano67-442, Ano67-443, Ano67-444, Ano68-415, Ano68-416, Ano68-417, Ano68-418, Ano68-419, Ano68-420, Ano68-421, Ano68-422, Ano68-423, Ano68-424, Ano68-425, Ano68-426, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69l, Ano69-405, Ano69-406, Ano69-407, Ano69-408, Ano69-409, Ano69-410, Ano69-411, Ano69-412, Ano69-413, Ano69-414, Ano69-415, Ano69-416]. Continuing [Bie68, EBO^{+}64, Wor64]. Continuous [Ano64-125]. contour [Ano65-461]. contourograph [WR66]. Contributed [Ano64-126]. Contributors [Ano68-316]. Control [Ano67-273, Ano69-224, Ano64-140, Ano64-338, Ano65-441, BW66, Cla65, Coh65, FS67a, Gib65, Hic67, Lau67, Mil68a, Spe66, Wag69, YW66]. controlled [Bis66, Dou66, Fri65b, Fri65c, Fri69b, Gre69]. controls [Fri65b, Fri65c, Sav69]. Controversy [Fin66]. Convention [Ano64h, Ano65c, Ano65d, Ano65f, Ano66a, Ano64-234, Ano65-148, Ano65-336, Ano64-52, HE68, Spr65]. Conventon [Ano69-277]. conversion [Ano65-299, Asi68, BFS68, Gia65, Wal67]. converters [Dob69]. Convolution [Hea69]. Cooperation [EB67]. coordination [Hic67]. copied [Sus67]. Corporate [CLM^{+}69]. Correction [Ano65-212, Ano68-252, Ano69-222]. Corrections [Ano69-223]. correlated [Ano68-294]. Correlative [Len66]. Correspondence [Ano64-127, Ano64-128, Ano64-129, Ano64-130, Ano65-213, Ano65-214, Ano65-215, Ano65-216, Ano65-217, Ano65-218, Ano65-219, Ano66-236, Ano66-237, Coo64, DMMS64, GW64, HBK64, HHE^{+}64, KDP64, ME65, MMGN65, QDH64, QBBB64, Rom65, Ano66-419, Ano66-420, Ano66-421, Ano66-422, Ano66-423, Ano66-424, Ano67-445, Ano67-446, Ano67-447, Ano67-448, Ano67-449, Ano67-450, Ano67-451, Ano67-452, Ano67-453, Ano67-454, Ano67-455, Ano67-456, Ano68-427, Ano68-428, Ano68-429, Ano68-430, Ano68-431, Ano68-432, Ano68-433, Ano68-434, Ano68-435, Ano68-436, Ano68-437, Ano68-438]. corrosion [Ano65-269]. cost [Ano64-266]. costs [Ano65-198]. Coulee [BN69]. counselors [Ano64-182]. Countries [Ano68-372, Ano64h]. Coupling [Ano65-147]. Course [Ano67-272]. courses [Ano65-182, Ano68-387, Ano69-373, Ano69-371, Ano69-370, Ano69-372]. Cover [Ano65-380, Ano64-81, Ano64-82, Ano64-83, Ano64-84, Ano64-85, Ano64-86, Ano64-87, Ano64-88, Ano64-89, Ano64-90, Ano64-91, Ano64-163, Ano64-164, Ano64-165, Ano64-166, Ano64-167, Ano64-168, Ano64-169, Ano64-170, Ano64-171, Ano64-172, Ano64-173, Ano64-174, Ano64-175, Ano64-212, Ano64-213, Ano64-214, Ano64-215, Ano64-216, Ano64-217, Ano64-218, Ano64-219, Ano64-220, Ano64-221, Ano64-222, Ano64-223, Ano64-224, Ano64-225, Ano64-226, Ano64-227, Ano64-228, Ano64-229, Ano64-230, Ano64-231, Ano64-232, Ano64-233, Ano65-161, Ano65-162, Ano65-163, Ano65-164, Ano65-165, Ano65-250, Ano65-251, Ano65-252, Ano65-253, Ano65-254, Ano65-255, Ano65-256, Ano65-257, Ano65-258, Ano65-259, Ano65-260, Ano65-261, Ano65-262, Ano65-263, Ano65-264, Ano65-265, Ano65-266, Ano65-267, Ano65-268, Ano65-309, Ano65-310, Ano65-311, Ano65-312, Ano65-313, Ano65-314, Ano65-315, Ano65-316, Ano65-317, Ano65-318, Ano66-257, Ano66-258, Ano66-259, Ano66-260, Ano66-261]. cover [Ano66-262, Ano66-263, Ano66-264, Ano66-265, Ano66-266, Ano66-267, Ano66-268, Ano66-269, Ano66-270, Ano66-271, Ano66-272, Ano66-273, Ano66-274, Ano66-275, Ano66-276, Ano66-277, Ano66-278, Ano66-279, Ano66-280, Ano67-289, Ano67-290, Ano67-291, Ano67-292, Ano67-293, Ano67-294, Ano67-295, Ano67-296, Ano67-297, Ano67-298, Ano67-299, Ano67-300, Ano67-301, Ano67-302, Ano67-303, Ano67-304, Ano67-305, Ano67-306, Ano67-307, Ano67-308, Ano67-309, Ano67-310, Ano67-311, Ano67-312, Ano68-269, Ano68-270, Ano68-271, Ano68-272, Ano68-273, Ano68-274, Ano68-275, Ano68-276, Ano68-277, Ano68-278, Ano68-279, Ano68-280, Ano68-281, Ano68-282, Ano68-283, Ano68-284, Ano68-285, Ano68-286, Ano68-287, Ano68-288, Ano68-289, Ano68-290, Ano68-291, Ano68-292, Ano69-251, Ano69-252, Ano69-253, Ano69-254, Ano69-255, Ano69-256, Ano69-257, Ano69-258, Ano69-259, Ano69-260, Ano69-261, Ano69-262, Ano69-263, Ano69-264, Ano69-265, Ano69-266, Ano69-267, Ano69-268, Ano69-269, Ano69-270, Ano69-271]. cover [Ano69-272, Ano69-273, Ano69-274]. Crawford [Ano64-137]. creation [Fri67a]. creative [Ama66, Nol67]. Crime [Kel67]. crisis [Gor68, Sie68]. critical [VSG^{+}66]. crosscorrelation [Ano65-338]. CRT [KS69]. Cryogenic [SSC67]. crystal [HB69]. Crystals [Ano68-176, Ano66-315]. current [Col66a, Jan64]. currents [DL69]. Cursive [Lin65d]. cybernetics [Mar67].

D [Ano64-250, Ano65-223, Cho64, Ano66c, Raw69]. D.C. [Ano65-336]. Dams [Fri68f]. darkness [Fri66b]. Data [Ano66-426, Jam65, Kor67, McP67, Ano65-131, Ano65-360, Ano65-361, Len66, MOV66]. date [Ano64e]. Davies [Dur65]. day [Ano64-131, Har65]. dc [Ath64, Gia66, Kai66, Lam66b, OT66, BH64]. deadline [Ano64-267]. deaf [Ano65-440]. dealing [Ano64-274]. death [Ano64-272]. decade [Sch69a, Sha67]. Deceptively [Ano66-403]. Deep [BKW67, Ano64-159, BP69a, DGH67]. deep-sea [Ano64-159]. Deep-space [BKW67, BP69a, DGH67]. deeper [Ano67-341]. deepest [Ano64-159]. Definition [Ano66-314, Ano64d]. deflection [Lot68]. Degree [Ano64-251]. Delay [Ano68-253, Dam67]. Delay-line [Ano68-253]. Delegates [Ano64-132, Tom69]. Delft [Cot66]. demand [Fri65e]. density [Cam65]. dependence [DGH67]. depths [Ano67-341, Fri66d]. desalination [Bar66a]. Design [Ano69-224, AI65, Ano64-341, Atk65, KL64, Nod65, Sid67, Tow69]. designated [Ano65-486]. designed [Ano65-249]. Designer [Ano64-133, Ano65-220, Ano65-221]. Detailed [SRLF64b]. Detection [Las66, Col66b, DeL68, Lin66a]. Detector [Ano68-356, Ano67-274, Dia69]. determined [Ano65-231]. developed [Ano64-276]. Development [Ano64-134, Ano68-297, Fri68d, Kno64, Kus68, Sha67, Sif65]. Developments [Ano65-222, JDDM64, Dou66]. Device [Sus67, Ano64-253, Ano65-196, Rob68a, ST67, Wal64]. devices [Ano65-484, Ano66-195, Cas64, Cop67, Cre65, Fow64, Fri68c, Kom67, LAT^{+}64, May65, Wei68, Cho64]. DEWLine [Lin66d]. Diagnostic [Den67]. dialing [Win66]. Dibner [Hen65]. dies [Ano64-273]. diesel [Fri68c]. diffraction [Hun64]. Digital [Ano65-227, Ano65-228, Mil68a, Cha67, FL66, HB69, MOV66, Nol67, Pie68b]. Dilemmas [Bro67a]. diode [LP67]. diodes [Gut67]. diodes-the [Gut67]. Directions [Ano68-254]. director [Ano64-177]. directorship [Ano64-136]. Directory [Ano66-290, Ano67-322, Ano65-332, Ano68-300]. Disabled [Ano65-366]. disappointments [Ano64-161]. Disaster [Hic67]. Discharge [Ano69-275]. disciplines [Fri69a]. Discounted [Hac65]. Discounted-cash-flow [Hac65]. discourse [Ano65-380]. discrete [McC67]. discrete-address [McC67]. discuss [Ano65-337]. discussion [Ano64-275, Ano65-319]. dispatch [Fri65c]. dispersive [KL64]. display [Bak68, Hem66]. distributed [Hue69]. distributed-lumped-active [Hue69]. distribution [Ano68-190, Fri68a, Law64]. Distributions [Ano69-453]. disturbances [Con65]. diving [Ano64-159]. Division [Ano65-148]. Do [Ano64-135, Ano66-238, Dun65]. does [Ano65-130]. Domb [Boo64]. double [Dia69]. Dow [Ano64-299]. Dr. [Ano64-136]. driver [Ano65-130]. dues [She66a]. during [Ano65-461]. dust [Ano65-416]. dynamic [CdKS66].

E. [Ano65-223]. each [Ano64d, Dun65]. ear [Ano68-241]. early [Sus68a, Sus68b, Sus69a, Sus69b]. earth [Dob68, Ano64-321, Coo66]. Earthquake [Lin66a]. eclipse [Ano65-461]. Economic [BFS68, Dom69, Hag64a, Lin66e]. economical [Fry69]. Economics [Ano65-224, Ano65-225, Lau67, Bar66a]. ECPD [Ano64-138]. Ed. [Dur65, Boo64]. edge [Lin68a]. Edgerton [Ano65-271]. editorial [Ryd64a]. EDP [Ano64-139]. Educating [Joh67, Web64b]. Education [Ano68-301, De 68, Moo64, She68, Ano64-55, Ano64-160, Ano64-295, Ano65-223, Ber64, Bie68, Bro67a, EB67, EBO^{+}64, Kop69, TZW^{+}67, Wee66, Wor64, Ano64-295]. Educational [Ano68-372, Ano67-314]. educator [Tel67]. effect [Ano65-248, Wal64]. effectively [Rob68b]. Effects [Ano64-158, AFW68, Hay64, Sco68]. efficiency [Ano65-237]. EHV [AJMS65, Ano65-206, Ano65-235, Ano65-247, BH64, Fri68g, LR67, Pop69]. EIA [Ano64-143, Ano64-292]. EJC [Ano64-336, Ano65-223]. elected [Ano64-137, Ano64-336]. Election [FRMH68]. Electric [Fri64a, Lin67a, UN67, Ano64-277, Ano65-166, BRP64, CLM^{+}69, Coh65, Con68, DL69, EA64, Fri65d, Hen65, Net69, Rit68, Zve66]. Electrical [Ano64-343, Ano65-454, Ano65-455, Ano65-456, Ano65-457, Ano65-458, Ano65-459, Ano65-460, Ano66-427, Ano66-428, Ano66-429, Ano66-430, Ano66-431, Ano66-432, Ano66-433, Ano66-434, Ano66-435, Ano66-436, Ano66-437, Ano66-438, Ano67-459, Ano67-338, Ano67-460, Ano67-461, Ano67-462, Ano67-463, Ano67-464, Ano67-465, Ano67-466, Ano67-467, Ano67-468, Ano67-469, Ano68-440, Ano68-441, Ano68-442, Ano68-443, Ano68-444, Ano68-445, Ano68-446, Ano68-447, Ano68-448, Ano68-449, Ano68-450, Ano68-451, Ano69-419, Ano69-420, Ano69-421, Ano69-422, Ano69-423, Ano69-424, Ano69-425, Ano69-426, Ano69-427, Ano69-428, Ano69-429, Ano69-430, BN69, Kaw66, Ano64-265, Ano68-387, Ano69-373, Ano69-371, Ano69-370, Ano69-372, Bla64, Bro69, Com68, Fin66, Ped68]. Electricity [Til69b, Til69c, Til69d, Ano65-406, Ell66]. electrification [Fri68e, Fri68c, Fri68d, Mir65]. Electrochemical [Swi68]. Electrogasdynamic [Wal67]. Electroluminescence [Str64]. Electromagnetic [Cot65, Cot66, Ano64-113, BL65a, BL65b, RB64, TA67]. Electromagnetics [Sus68a]. Electromagnetism [Gri66]. electromechanical [Ano64-305]. Electron [Kom67, Sus67, Ano65b, Ano65-323, Ano65-484, Her65, Pea67, Sal65a, Sal65b]. Electronic [Ano64-140, Ano64-141, Ano64-142, Ano65-226, Ano65-227, Ano65-228, Ano66-239, Ano66-240, Ano66-241, Ano66-242, Ano66-310, HK68, Bre65, Bro69, Cre65, FO69, KKV65, Rob65, Roi67, Sus66]. Electronically [OFM^{+}69]. Electronics [Ano64-143, Ano64-343, Ano64-264, Ano65-454, Ano65-455, Ano65-456, Ano65-457, Ano65-458, Ano65-459, Ano65-460, Ano66-427, Ano66-428, Ano66-429, Ano66-430, Ano66-431, Ano66-432, Ano66-433, Ano66-434, Ano66-435, Ano66-436, Ano66-437, Ano66-438, Ano66-312, Ano67-459, Ano67-338, Ano67-460, Ano67-461, Ano67-462, Ano67-463, Ano67-464, Ano67-465, Ano67-466, Ano67-467, Ano67-468, Ano67-469, Ano68-440, Ano68-441, Ano68-442, Ano68-443, Ano68-444, Ano68-445, Ano68-446, Ano68-447, Ano68-448, Ano68-449, Ano68-450, Ano68-451, Ano69-419, Ano69-420, Ano69-421, Ano69-422, Ano69-423, Ano69-424, Ano69-425, Ano69-426, Ano69-427, Ano69-428, Ano69-429, Ano69-430, Fri67d, Fri67c, Hag68, Omi66, Sie68, Smi67, Ano64-110, Ano65-197, Ano65-319, Ano65-336, Ano68-387, Ano69-373, Ano69-371, Ano69-370, Ano69-372, Ano69-404, Blo67, Cre65, EMK69, Eva69, Fin69, Kob69, Noy68, SSF65, Sto69, Sus68a, Sus68b]. electronics [Sus69a, Sus69b]. electrons [Ano68-317]. Electrooptic [Lot68]. Electrooptics [Sco66]. Electrotechnical [Ano65-489]. elects [Ano65-352]. eliminating [Rob65]. Elliptic [Ano69-299]. Elliptic-Function [Ano69-299]. emergence [Bro69]. EMI [SW67]. emission [Fei64b]. EMP [Hay64]. emphasis [Sho66]. End [Ano65-334]. Energy [Gia65, Hig66, LS64, Phi65, BFS68, Fel64, Fri65d, GN69, Hic66, LPK65, Mar66, Rit68, Rod65b, VM67, Vog69, Wal67]. engine [Ano65-132]. engineer [AH69, Ano65-405, Bro69, KF64, She69, Smi67]. Engineering [Ano65-229, Ano66-238, Ano68-301, Lin65a, Lin66e, Ano64a, Ano64-55, Ano64-56, Ano64-57, Ano64-58, Ano64-59, Ano64-60, Ano64-61, Ano64-62, Ano64-63, Ano64-64, Ano64-65, Ano64-66, Ano64-67, Ano64-111, Ano64-297, Ano64-355, Ano65-321, Ano65-365, Ano65-485, Ano68-387, Ano69-373, Ano69-371, Ano69-370, Ano69-372, Bie68, Bro69, Bro67a, Den67, Ebe64, For64, Fri66c, Fri66d, Fri66e, Fro69, Gib65, Gol65, Gre66, Hac65, Hor68b, Kop69, Lie66, Lin66b, Lin66c, Lin67b, Moo68, OBM^{+}64, Ped68, She68, Tel67, TZW^{+}67, Tur67, UN67, Wee66, Whi69, Ano64-146, Ano64-144, Ano64-145, Ano65-127, Ano65-128, Ano65-129, Ano65-230].

Engineering-economic [Lin66e]. Engineers [Ano64-147, Ano64-343, Ano65-139, Ano65-454, Ano65-455, Ano65-456, Ano65-457, Ano65-458, Ano65-459, Ano65-460, Ano65-387, Ano66-427, Ano66-428, Ano66-429, Ano66-430, Ano66-431, Ano66-432, Ano66-433, Ano66-434, Ano66-435, Ano66-436, Ano66-437, Ano66-438, Ano66-403, Ano67-459, Ano67-338, Ano67-460, Ano67-461, Ano67-462, Ano67-463, Ano67-464, Ano67-465, Ano67-466, Ano67-467, Ano67-468, Ano67-469, Ano68-440, Ano68-441, Ano68-442, Ano68-443, Ano68-444, Ano68-445, Ano68-446, Ano68-447, Ano68-448, Ano68-449, Ano68-450, Ano68-451, Ano69-419, Ano69-420, Ano69-421, Ano69-422, Ano69-423, Ano69-424, Ano69-425, Ano69-426, Ano69-427, Ano69-428, Ano69-429, Ano69-430, Ano64-265, Ano64h, Bla64, Dun65, EB67, Rob68b, Web64b, Wor64]. engines [Bre65]. enhancement [BL67, BL68]. enterprise [Lin65a]. entropy [Gro65]. environment [And69, Lin66b]. Environmental [AI65]. enzymatic [CPH67]. Epitaxial [Ano64-268]. equalization [Rud67]. Equations [Ano68-255]. equipment [AI65, AJMS65, And64, Mai64, Noy64]. established [Ano64-250]. Esthetics [Fri65d]. eternal [Lin67a]. European [Fri68d, McL68, The66, Wee66]. Evaluation [Ano68-295]. Eve [Ped68]. every [Ano66-365]. evolutionary [BL67]. Exact [Ano65-231]. exciting [Fri67a]. Executive [Ano64-183, Ano65-232, Ano65-405, Ano64-177]. Exemplar [Lin65a]. Exhibition [Ano66-310, Ano69-277]. exhibitor [Ano64c]. Exhibitors [Ano64-148, Ano65-233, Ano66-310, Ano68-256, Ano68-293, Ano65-488, Ano67-495, Ano68-477, Ano69-277, Ano69-454]. expanded [She66b]. expanding [OFM^{+}69]. expansion [Ano64-198]. expedite [Ano64-51]. Experience [VSG^{+}66, MM67, Pag69, TZW^{+}67]. Experienced [Ano65-387]. experiment [Ano65-323]. Experimental [Ano64-149]. Experiments [Ano67-273, Sus68b, Sus69b]. exploding [Ano64-158]. exploitation [KE68]. exploration [Ano64-337, Tra69, von65]. explosion [Lin66a]. explosions [Ano64-158, Hay64]. Expo [Ano67-340, Ano69-225]. extend [Zim69]. external [AI65]. extraterrestrial [WGC^{+}66]. eye [Fri66f].

F [Ano64-273, Ano64-292, Nod65]. fabricating [Ano66-195]. fabrication [Ano64-274]. faces [Ped68]. facilities [Fin67]. facility [Wei65]. Facsimile [Ano65-380]. factors [Lin66b, Lin66c, SSF65]. failure [Ano65-382, Ano66-235, Fri66b]. failures [Fri68b]. Fall [Ano68-387, Ano69-371, Ano69-370]. Fall-Winter [Ano69-370]. Family [Ano64-150]. fantasia [Roi67]. Faraday [Ano65-331, Dib67]. fast [Ano64-344, Ano67-274, Ber69, BM67]. fast-acting [Ano67-274]. faster [Ano67-470]. fatigue [Ano65-234]. Fault [SRLF64a, SRLF64b]. feature [Ano65-206]. featured [Ano65-222]. features [BN69, Kaw66]. February [Ano65-420]. Feedback [Ano67-275, Sav69]. fellowships [Ano64-162]. ferrimagnetics [Ano65-333]. Ferrite [Ano65-335]. field [Ano67-275, Lav68, Wal64]. field-effect [Wal64]. Fields [Ano64d, Sun69]. Fighting [Ano65-234]. Figures [Ano69-226]. File [Ano64-133, Ano65-220, Ano65-221]. film [Ano64-344, Cas64, Fow64, GJPM69, LMT64, Pri66]. films [Ano66-282, Ano67-470, Thu69]. Filter [Ano66-465, Ano68-318, Ano69-299, Tow69]. filtering [Hue69]. Filters [Ano68-370, Mit69, Rud67, SB68, SW67, Zve66]. First [Ano64-159, Ano65-235, Ano65-148, Ano65-248, Sus69b]. fiscal [BL65b]. fish [Fri66c]. fist [SRLF64a, SRLF64b]. Flexibility [Ano65-441]. flight [Ano64-51]. Floor [Ano65-236, And64, Ano65-361, FAB64]. flow [Hac65]. Flowing [Ano65-237]. Fluidics [Wag69]. flying [Bro67b]. flying-spot [Bro67b]. Focal [Ano64-158, Ano64-159, Ano64-160, Ano64-161, Ano64-151, Ano64-152, Ano64-153, Ano64-154, Ano64-155, Ano64-156, Ano64-157, Ano65-246, Ano65-247, Ano65-248, Ano65-249, Ano65-238, Ano65-239, Ano65-240, Ano65-241, Ano65-242, Ano65-243, Ano65-244, Ano65-245, Ano66-243, Ano66-244, Ano66-245, Ano66-246, Ano66-247, Ano66-248, Ano66-249, Ano66-250, Ano66-251, Ano66-252, Ano66-253, Ano66-254, Ano67-276, Ano67-277, Ano67-278, Ano67-279, Ano67-280, Ano67-281, Ano67-282, Ano67-283, Ano67-284, Ano67-285, Ano67-286, Ano67-287, Ano68-257, Ano68-258, Ano68-259, Ano68-260, Ano68-261, Ano68-262, Ano68-263, Ano68-264, Ano68-265, Ano68-266, Ano68-267, Ano68-268, Ano69-227, Ano69-228, Ano69-229, Ano69-230, Ano69-231, Ano69-232, Ano69-233, Ano69-234, Ano69-235, Ano69-236, Ano69-237, Ano69-238, Mul64]. following [Con65]. folly [Tra69]. food [Fri66c]. Force [EK69]. forecasting [NP69]. forgotten [Gos68]. Form [Ano68-414]. former [Ano64-181]. Fortescue [Ano64-162]. Forum [Ano66-284, Ano66-285, Ano66-286, Ano66-287, Ano66-288, Ano66-289, Ano67-315, Ano67-316, Ano67-317, Ano67-318, Ano67-319, Ano67-320, Ano67-321, Ano68-298, Ano68-299, Ano69-239, Ano69-240, Ano69-241, Ano69-242, Ano69-243, Ano69-244, Ano69-245, Ano69-246, Ano69-247, Ano69-248, Ano69-249, Ano69-250]. Found [Ano66-255]. foundation [Ano65-229]. Foundations [Hal67, For64]. four [Ano64-183]. Fourier [Ber69, BM67, Ano66-256]. FPC [Fri68b]. Franklin [Ano65-299, Hen65]. Free [Bro64d]. Free-space [Bro64d]. Frequencies [Yam66, Ano65-362, Ano66-309, Sel66]. Frequency [Ano67-273, Ano68-315, Ano69-333, Sei66, Ano65-419, Ano65-442]. Fritz [Ano65-486]. Front [Ano64-163, Ano64-164, Ano64-165, Ano64-166, Ano64-167, Ano64-168, Ano64-169, Ano64-170, Ano64-171, Ano64-172, Ano64-173, Ano64-174, Ano64-175, Ano65-250, Ano65-251, Ano65-252, Ano65-253, Ano65-254, Ano65-255, Ano65-256, Ano65-257, Ano65-258, Ano65-259, Ano65-260, Ano65-261, Ano65-262, Ano65-263, Ano65-264, Ano65-265, Ano65-266, Ano65-267, Ano65-268, Ano65-334, Ano66-257, Ano66-258, Ano66-259, Ano66-260, Ano66-261, Ano66-262, Ano66-263, Ano66-264, Ano66-265, Ano66-266, Ano66-267, Ano66-268, Ano66-269, Ano66-270, Ano66-271, Ano66-272, Ano66-273, Ano66-274, Ano66-275, Ano66-276, Ano66-277, Ano66-278, Ano66-279, Ano66-280, Ano67-289, Ano67-290, Ano67-291, Ano67-292, Ano67-293, Ano67-294, Ano67-295, Ano67-296, Ano67-297, Ano67-298, Ano67-299, Ano67-300, Ano67-301, Ano67-302, Ano67-303, Ano67-304, Ano67-305, Ano67-306, Ano67-307, Ano67-308, Ano67-309, Ano67-310, Ano67-311, Ano67-312, Ano68-269, Ano68-270, Ano68-271, Ano68-272, Ano68-273, Ano68-274, Ano68-275]. Front [Ano68-276, Ano68-277, Ano68-278, Ano68-279, Ano68-280, Ano68-281, Ano68-282, Ano68-283, Ano68-284, Ano68-285, Ano68-286, Ano68-287, Ano68-288, Ano68-289, Ano68-290, Ano68-291, Ano68-292, Ano69-251, Ano69-252, Ano69-253, Ano69-254, Ano69-255, Ano69-256, Ano69-257, Ano69-258, Ano69-259, Ano69-260, Ano69-261, Ano69-262, Ano69-263, Ano69-264, Ano69-265, Ano69-266, Ano69-267, Ano69-268, Ano69-269, Ano69-270, Ano69-271, Ano69-272, Ano69-273, Ano69-274, Ano64-223, Ano64-224, Ano64-225, Ano64-226, Ano64-227, Ano64-228, Ano64-229, Ano64-230, Ano64-231, Ano64-232, Ano64-233, Ano65-314, Ano65-315, Ano65-316, Ano65-317, Ano65-318]. frontier [Fri66d]. Fuel [Lie66, Fry69, Pea66]. Function [Ano67-273, Ano69-299, Ano64-159, Mor65]. Functional [Ano67-313]. Functions [Ano68-295, Ano64-249, Fin67, Fra64a, FK64, Fra64b, Fra64c, Har69, Fra64a]. Fundamentals [MC66]. fuses [Ano64-276]. fusion [Bis66, Dou66]. Future [Ano68-293, Ano69-224, Cre65, Lin67b, Nie68, Spe66, Ano64-274, Bur67, DBW69, Fri68e, Fri68c, Fri68d, Gro66, Gun66, Her65, Hic66, HHA^{+}66, Par69, Raj65, Tay69, Wal69]. fuzes [Fri67c].

G [Ano64-177, Cho64]. gain [Ano67-383]. Galvanomagnetic [Wei68]. Game [Dix68]. Game-theoretic [Dix68]. Gap [Ano68-356]. Gas [Ano69-275, Ano64-125, Ano65-237, Ano67-343]. gate [VGRM69]. General [Ano65-130, Ano65-489, Ano68-241, Gou64, Ano64-211, Ano65-462, Ano65-132]. generalized [Fra64d]. generating [Ano65-198]. generation [Ano68-294, Bro64a, BM64, Cha67, CPH67, Fri68g, Rit68, Wal69, Wil64b]. generations [Raj65]. Generators [Ano69-275, Ath64, Fri65e, Tsu67]. genesis [Ryd64a]. genius [Gos68]. Geophysical [Bra64]. George [Cut64a, Ano64-250]. Gerhard [Nod65]. Giant [Fri65e]. Gigahertz [Ano68-296]. GIT [Ano69-331]. give [Ano64-344]. glass [Smi66a]. Glideslope [Ano67-288]. Glimpse [Ano66c]. Global [Wes66]. Globecom [Ano64-52]. Goals [Ano68-301, Ano65-438, Lin67b]. golden [Zve66]. good [Ano65-269]. government [Ano64-160, For68, Hol67]. Graduate [Ber64]. Grand [BN69]. grants [Ano65-322]. Graphical [Ano65-270, Ano68-295]. Great [Gab67, Fri68d, Gor68]. greatest [Ano65-331]. green [Fri66f]. grid [LR67]. grounding [Rob65]. Group [Ano64-176, Ano64-183, Ano64-211, Ano64-184, Ano64-186, Ano65-352]. groups [Ano64d, Ano64b, Ano64e]. growing [Fri65e, Phi65]. Growth [Fel64, Fin66, NP69]. guide [Ano65-226, Ano67-220, Ano69-160]. guided [Ber69]. Gulf [Ano68-241]. gyro [Kil67]. gyrotropic [Ano64-302].

H [Pra68]. H. [Ano64-177, Ano65-271]. Half [HG65]. Half-wavelength [HG65]. halftones [HK68]. hall [Sch65]. handle [Ano65-130]. handling [Ano65-222]. handmaiden [Kna69]. Handprinting [ST67]. Handwriting [ME65]. Haraden [Ano69-375]. Hardware [Stu68]. Harris [Ano65-407]. Hartford [Ano64-246]. Haystack [Wei65]. head [Ano64-273, Ano65f]. Headquarters [Fin67, Ano64-181]. Hearing [Ano66c, Ano65-440]. heart [Ano64-134, Lin65a]. Heat [BRP64, Ano65-132]. heating [Ano64-277, Ano65-166, BRP64, Fri64a]. HEC [Ano64-336]. Hector [Ski64].

Heilweil [Nod65]. held

[Ano64a, Ano64-246, Ano65-299, Ano65-350, Ano65-351, Ano65-418, Ano65-484]. Helicon [Ano67-370]. Helicon-Wave [Ano67-370]. help [Ano65-226]. Hertz [Sus68a, Sus68b]. Heterodyne [Ano68-296, Ano68-297, DeL68]. HF [Ano67-370, Ano68-370]. High [Ano66-281, Ano66-282, Ano66-403, Ano68-296, Ano68-297, Bur67, Hig66, Jam66, Lav68, LAT^{+}64, Mon66, Ano64-253, Ano64-356, Ano65-337, HS68, Koe65, Lam66b, Sel66]. high-brightness [HS68]. High-Energy [Hig66]. High-field [Lav68]. High-performance [Ano66-281]. High-power [Bur67, LAT^{+}64, Ano65-337]. High-purity [Ano66-282]. High-Sensitivity [Ano68-296, Ano68-297]. High-speed [Jam66]. High-strength [Mon66]. high-voltage [Lam66b]. highest [Ano65-237]. highlights [HE68]. highly [Ano65-408]. Highway [Ano67-272, FO69, Fri67b]. Hill [Ano64-138]. history [Fri68e, Sus68a, Sus68b, Sus69a, Sus69b]. Hoc [WL64, Ano67-314]. Hoernes [Nod65]. hold [Ano64-176, Ano65-229]. holography [Col66a, Kor68]. homing [Fri67c]. honor [Ano64-250, Ano64-292]. hope [Lin67a]. horizons [Blo67, Esa66, LPK65]. Horne [Ano64-292]. hot [Fri64a]. Hotel [Ano65-299]. house [BRP64, Ano64-355]. Hughes [Ano65-140]. Human [BL67, BL68, Lin66b, Lin66c, Pag69, Lin65b, Lin65c, Lin65d, SSF65, Tur67]. hundred [Ano67-369]. hungry [Fri66c]. Hutchisson [Ano64-136]. HV [Kai66]. Hybrid [Tru64, Ham69, Hin69, Jac69]. hybrid-circuit [Jac69]. Hydrocarbon [Pea66].

IC [WP69]. ICMCI [Tom64]. IEC [Ano65-462]. IEDM [Ano65-279]. IEEE [Ano65c, Ano65d, Ano66a, Ano66b, Ano66-406, Fin67, HE68, Nag67, Ano64-139, Ano64-179, Ano64-180, Ano64-181, Ano64-182, Ano64b, Ano64-234, Ano64-183, Ano64e, Ano64-184, Ano64-186, Ano64-185, Ano64-197, Ano64-192, Ano64-193, Ano64-194, Ano64-195, Ano64-196, Ano64-187, Ano64-188, Ano64-189, Ano64-190, Ano64-191, Ano64-267, Ano64-263, Ano64-264, Ano64-265, Ano64-255, Ano64-256, Ano64-257, Ano64-258, Ano64-259, Ano64-260, Ano64-261, Ano64-262, Ano64-295, Ano64-275, Ano64f, Ano64-279, Ano64-318, Ano64-319, Ano65-232, Ano65-280, Ano65-308, Ano65-284, Ano65-285, Ano65-281, Ano65-282, Ano65-283, Ano65-350, Ano65-351, Ano65-340, Ano65-341, Ano65-342, Ano65-343, Ano65-344, Ano65-345, Ano65-346, Ano65-347, Ano65-348, Ano65-349, Ano65-406, Ano66-284, Ano66-285, Ano66-286, Ano66-287, Ano66-288, Ano66-289, Ano66-290, Ano66-291, Ano66-292, Ano66-293, Ano66-317, Ano66-318, Ano66-319, Ano66-320, Ano66-321, Ano66-322, Ano66-323]. IEEE [Ano66-324, Ano66-325, Ano66-326, Ano66-327, Ano67-220, Ano67-314, Ano67-315, Ano67-316, Ano67-317, Ano67-318, Ano67-319, Ano67-320, Ano67-321, Ano67-322, Ano67-323, Ano67-345, Ano67-346, Ano67-347, Ano67-348, Ano67-349, Ano67-350, Ano67-351, Ano67-352, Ano67-353, Ano67-354, Ano67-355, Ano67-356, Ano67-344, Ano68-298, Ano68-299, Ano68-300, Ano68-301, Ano68-321, Ano68-322, Ano68-323, Ano68-324, Ano68-325, Ano68-326, Ano68-327, Ano68-328, Ano68-329, Ano68-330, Ano68-331, Ano68-332, Ano68-357, Ano69-160, Ano69-276, Ano69-307, Ano69-308, Ano69-309, Ano69-310, Ano69-311, Ano69-312, Ano69-313, Ano69-314, Ano69-315, Ano69-316, Ano69-317, Ano69-318, Bar66b, Her68a, Her68b, Hof68, Lin64, Mac67a, Mac67b, ML65, Oli65a, Oli65b, She66a, She66b, She66c, Spr65, Ano64-53, Ano64-95, Ano65-146, Ano65-176, Ano67-233, Ano67-234, Ano67-235, Ano67-236, Ano67-237, Ano67-238, Ano67-239, Ano67-240, Ano67-241, Ano67-242, Ano68-203, Ano68-204]. IEEE [Ano68-205, Ano68-206, Ano68-207, Ano68-208, Ano68-209, Ano68-210, Ano68-211, Ano68-212, Ano68-213, Ano68-214, Ano69-173, Ano69-174, Ano69-175, Ano69-176, Ano69-177, Ano69-178, Ano69-179, Ano69-180, Ano69-181, Ano69-182, Ano69-183, Ano69-184, Ano69-185, Ano69-186, Ano69-187, Ano69-188, Gol64]. IEEE’69 [Ano69-277]. IEEE’s [Ano64d]. If [Ano64-198, Ano68-296, Ano68-297, SB68]. IFAC [Ano64-322, Ano64-345, Ano65-286]. IFIP [Ano65-286]. II [Ano65-228, Ano68-297, BL65b, Cas64, Fra64a, Fri65c, Fri66f, Fri67d, Fri67c, Lin65c, Lin66c, Lin69b, MGC65b, Sal65b, SRLF64b, Sus68b, Til69c]. III [Fow64, FK64, Lin65d, Sus69a, Til69d]. illness [VSG^{+}66]. Image [Ano65-334]. Image-Phasing [Ano65-334]. Images [Sch69b]. Imaginary [Ano68-295]. imaging [Ano65-196, Hor68a, Hor68b, Kor68, Raw69]. impact [Ano64-274, Hag64a, Til69d]. implantable [Ano64-134]. impressions [Tom69]. Improvements [EMK69]. Inc. [Ano64-146, Ano64-144, Ano64-145, Ano64-343, Ano64-273, Ano65-127, Ano65-128, Ano65-129, Ano65-230, Ano65-454, Ano65-455, Ano65-456, Ano65-457, Ano65-458, Ano65-459, Ano65-460, Ano66-427, Ano66-428, Ano66-429, Ano66-430, Ano66-431, Ano66-432, Ano66-433, Ano66-434, Ano66-435, Ano66-436, Ano66-437, Ano66-438, Ano67-459, Ano67-338, Ano67-460, Ano67-461, Ano67-462, Ano67-463, Ano67-464, Ano67-465, Ano67-466, Ano67-467, Ano67-468, Ano67-469, Ano68-241, Ano68-440, Ano68-441, Ano68-442, Ano68-443, Ano68-444, Ano68-445, Ano68-446, Ano68-447, Ano68-448, Ano68-449, Ano68-450, Ano68-451, Ano69-419, Ano69-420, Ano69-421, Ano69-422, Ano69-423, Ano69-424, Ano69-425, Ano69-426, Ano69-427, Ano69-428, Ano69-429, Ano69-430]. inch [Ano65-391]. including [Ano64-52]. income [Fel64]. increased [Ano64-125, She66a]. Index [Ano64-199, Ano64-200, Ano64-201, Ano64-202, Ano64-203, Ano64-204, Ano64-205, Ano64-206, Ano64-207, Ano64-208, Ano64-209, Ano64-210, Ano65-287, Ano65-288, Ano65-289, Ano65-290, Ano65-291, Ano65-292, Ano65-293, Ano65-294, Ano65-295, Ano65-296, Ano65-297, Ano65-298, Ano66-294, Ano66-295, Ano66-296, Ano66-297, Ano66-298, Ano66-299, Ano66-300, Ano66-301, Ano66-302, Ano66-303, Ano66-304, Ano66-305, Ano67-324, Ano67-325, Ano67-326, Ano67-327, Ano67-328, Ano67-329, Ano67-330, Ano67-331, Ano67-332, Ano67-333, Ano67-334, Ano67-335, Ano68-302, Ano68-303, Ano68-304, Ano68-305, Ano68-306, Ano68-307, Ano68-308, Ano68-309, Ano68-310, Ano68-311, Ano68-312, Ano68-313, Ano69-278, Ano69-279, Ano69-280, Ano69-281, Ano69-282, Ano69-283, Ano69-284, Ano69-285, Ano69-286, Ano69-287, Ano69-288, Ano69-289, Ano64-53, Ano64-68, Ano64-254, Ano64-334, Ano65-146, Ano66-182, Ano67-206, Ano68-177, Ano69-159]. indoor [Lau67]. Industrial [Ano65-299, Hoe66, Ano64-356, Ano65-417, Mai64]. industrialist [Ano65-386]. Industry [Ano64-160, Ano64-211, Ano65-183, Omi66, Ano64-246, Ano64-295, Ano66-464, Ano68-190, Bla65, Fin69, Fri68a, Gen64, Hol67]. industry-government [Hol67]. Information [Ano65-300, Ano65-301, Ano65-302, Ano65-303, Ano65-304, Ano65-305, Ano65-306, Ano65-307, Ano65-308, Ano66-306, Ano66-406, Ano68-314, Ano69-290, Ano69-291, Ano69-292, Ano69-293, Ano69-294, Ano69-295, Cam65, Mar67, Ano65-131, Ano65-320, Ano65-360, Gro65, Jam65, Jam66, Sal65c]. Infrared [Ano67-271, Ano68-296, Ano68-297, Ano68-356, Ano69-296, Ano67-274, Fri67d, Hor68a, Hor68b]. innovation [De 68]. Input [ST67]. Insertion [Ano66-465]. installation [And64]. Instant [Ano66-307, Ano66-308, Asi68]. Institute [Ano64-343, Ano65-454, Ano65-455, Ano65-456, Ano65-457, Ano65-458, Ano65-459, Ano65-460, Ano66-427, Ano66-428, Ano66-429, Ano66-430, Ano66-431, Ano66-432, Ano66-433, Ano66-434, Ano66-435, Ano66-436, Ano66-437, Ano66-438, Ano67-459, Ano67-460, Ano67-461, Ano67-462, Ano67-463, Ano67-464, Ano67-465, Ano67-466, Ano67-467, Ano67-468, Ano67-469, Ano68-440, Ano68-441, Ano68-442, Ano68-443, Ano68-444, Ano68-445, Ano68-446, Ano68-447, Ano68-448, Ano68-449, Ano68-450, Ano68-451, Ano69-419, Ano69-420, Ano69-421, Ano69-422, Ano69-423, Ano69-424, Ano69-425, Ano69-426, Ano69-427, Ano69-428, Ano69-429, Ano69-430, Ano67-338, Wor64]. instruct [Ano64-149]. Instrumentation [Hig66, Ano64-110, Fri66e, RJ68, SRLF64b]. Instrumenting [FAB64]. Insulated [HHA^{+}66]. Intact [Ano67-273]. Integrated [Ano66-309, Ano69-298, Bro64c, HHN^{+}64, Mai64, Noy64, SN69, Ano67-313, Ano69-404, Gre69, Hag64a, Ham69, Hir68b, Hog64, Kno64, Noy68, Phi64, War67]. integration [EA64, Gut65]. intelligence [Pag69, WGC^{+}66]. intelligentsia [Fei64a]. Interaction [Adl67]. interactions [Sei66]. interconnected [Con65, Hic67]. interest [Ano64d]. interface [Gro66]. interference [Rob65]. interferometer [Ano64-356]. interior [Ano64-321]. INTERMAG [Elf65, Ano64-340]. internal [Ano66-315]. International [Ada66, Ano64-234, Ano65c, Ano65d, Ano65-319, Ano65-320, Ano65-489, Ano66-310, Ano66a, Dou66, Gro66, HP67, Sel66, Wes68, End66, Fri65a, Ano67-339]. Internatonal [Ano69-277]. interpersonal [Ano64-141]. Interplanetary [Rei66]. Interscience [Ano69-418]. Interstellar [Ano66-311]. Intertie [Dom69]. Introduction [Eve64, Hag64b, Okr64a, BP69b, Nod65, Cut64a, Nod65]. inverses [Fra64d]. investment [Hac65]. invited [Ano64-274, Ano64-340, Ano65-364]. ion [Bre65]. ionized [Paa66]. ionized-argon [Paa66]. Ionosphere [Ano66-353]. Ionospheric [Dur65]. IRE [Ano64-181]. irradiation [FKS64]. ISA [Ano64-137]. isolation [SRLF64a, SRLF64b]. Isolator [Ano67-370]. issued [Ano64-341]. issues [Ano64-52, Ano64-112, Ano64-197, Ano64-192, Ano64-193, Ano64-194, Ano64-195, Ano64-196, Ano64-320, Ano64-345, Ano65-284, Ano65-285, Ano65-409, Ano65-410, Ano65-411, Ano65-412, Ano65-413, Ano65-414, Ano65-415, Ano66-367, Ano66-368, Ano66-369, Ano66-370, Ano66-371, Ano66-372, Ano66-373, Ano66-374, Ano66-375, Ano66-376, Ano66-377, Ano67-399, Ano67-400, Ano67-401, Ano67-402, Ano67-403, Ano67-404, Ano67-405, Ano67-406, Ano67-407, Ano67-408, Ano67-409, Ano67-410, Ano68-373, Ano68-374, Ano68-375, Ano68-376, Ano68-377, Ano68-378, Ano68-379, Ano68-380, Ano68-381, Ano68-382, Ano68-383, Ano68-384, Ano69-358, Ano69-359, Ano69-360, Ano69-361, Ano69-362, Ano69-363, Ano69-364, Ano69-365, Ano69-366, Ano69-367, Ano69-368, Ano69-369]. items [KDP64]. IV [Fra64b, Sus69b].

J [Gut65]. JACC [Ano64-296]. Jackson [Cut64a]. James [Dib64]. January [Ano64-211, Ano64-184]. Japan [Ano67-339, Ano67-369, Ano68-300, IN69, Kob69, Omi66, UN67, Yam66, Ano66-290, Ano66-312, Ano67-322, Ano67-340]. Jerrems [Ano65f]. jets [Ano64-294]. job [Ano64-135, Ano68-241]. John [Ano65-486]. joining [Ano66-339]. Joins [Ano64-54, Ano64-303]. Joint [Ano65-321, Cla65, Fri66a, Ano64-276]. journals [Ano64-197, Ano64-347, Ano64-348, Ano64-349, Ano64-350, Ano64-351, Ano64-352, Ano64-353, Ano64-354, Ano65-475, Ano65-476, Ano65-477, Ano65-478, Ano65-479, Ano65-480, Ano65-481, Ano65-482, Ano65-483, Ano66-451, Ano66-452, Ano66-453, Ano66-454, Ano66-455, Ano66-456, Ano66-457, Ano66-458, Ano66-459, Ano66-460, Ano66-461, Ano66-462, Ano67-483, Ano67-484, Ano67-485, Ano67-486, Ano67-487, Ano67-488, Ano67-489, Ano67-490, Ano67-491, Ano67-492, Ano67-493, Ano67-494, Ano68-464, Ano68-465, Ano68-466, Ano68-467, Ano68-468, Ano68-469, Ano68-470, Ano68-471, Ano68-472, Ano68-473, Ano68-474, Ano68-475, Ano69-441, Ano69-442, Ano69-443, Ano69-444, Ano69-445, Ano69-446, Ano69-447, Ano69-448, Ano69-449, Ano69-450, Ano69-451, Ano69-452]. Just [Ano65-132].

Kansas [Ano64-247]. Karnaugh [Ano68-414]. Kenneth [Dur65]. keyboard [Ros67]. keynotes [Ano64-299]. keys [Ros67]. kHz [Gre67b]. know [Ano64-135]. Knowledge [Lin68a]. Kodak [Ano65-322]. Krassner [Cut64a].

L [Ano68-315]. L-Band [Ano68-315]. Laboratories [Ano65-130, Ano65-126, Ano65-132]. laboratory [Ano64-159, Ano64-342, Ano66-366, Har66]. LaGuardia [Ano67-288]. land [Cox65]. Landing [Ano69-201, MW66]. language [Hal67, Lin65b, Lin65c, Lin65d]. languages [Pie68a]. Large [Ano68-297, Ano64i, Hic67, Rin67]. Laser [Ano65-323, Ano66-463, Bak68, FKS64, Hau68, Ano64-125, Ano64-305, Ano64-335, Ano64-356, Ano65-237, Ano65-248, Ano65-419, Ano68-386, Kil67]. Lasers [Ano67-271, Ano67-343, Bur67, Lax65, Paa66, Vol67]. late [Ano64-250, Ano65-332]. Latin [Mac67a]. launch [Ano64-333]. launched [Ano64-303, Ano65-249]. launches [She66b]. law [Ano64-346]. Lawrence [Ano65-247]. lead [Ano66-195]. learn [Hor64]. leaves [Ano66-366]. lecture [Tow65]. LED [HS68]. Leeds [Ano65-271]. Lethal [DL69]. level [Len66]. Lewis [Ano65-486]. Library [Ano66-238, Lef64]. life [She69]. Light [Ano66-463, Adl67, Ano65-248, Ano67-274, Hun64, LP67, Lot68]. light-beam [Lot68]. lightning [Uma66]. Limitations [FS67b, Bur67, Lig67]. limits [Bro64b, Key69]. Line [Kaw66, Ano65-247, Ano68-253, Fri64a, Gre67a, VSG^{+}66]. Linear [Hir68b]. Lines [Ano65-147, Ano64-331, Ano64-332, Ano65-438, Ano65-439, Ano65-437, Ano65-430, Ano65-431, Ano65-432, Ano65-433, Ano65-434, Ano65-435, Ano65-436, Ano66-391, Ano66-392, Ano66-393, Ano66-394, Ano66-395, Ano66-396, Ano66-397, Ano66-398, Ano66-399, Ano66-400, Ano66-401, Ano66-402, Ano67-424, Ano67-425, Ano67-426, Ano67-427, Ano67-428, Ano67-429, Ano67-430, Ano67-431, Ano68-400, Ano68-401, Ano68-402, Ano68-403, Ano68-404, Ano68-405, Ano68-406, Ano68-407, Ano68-408, Ano68-409, Ano68-410, Ano68-411, Ano69-388, Ano69-389, Ano69-390, Ano69-391, Ano69-392, Ano69-393, Ano69-394, Ano69-395, Ano69-396, Ano69-397, Ano69-398, Ano69-399, Cut67a, Cut67b, Cut67c, Cut67d, Dam67, HG65, Mac64, Par69, Ryd64b, Ryd64c, Ryd64d, Ryd64e, Ryd64f, Ryd64g, Ryd64h, Ryd64i, Wil65b, Wil65c, Wil64a]. links [Ano65-485, BP69a]. Liquid [Pet65]. Liquid-metal [Pet65]. Liquids [Ano68-240]. List [Ano64-235, Ano64-236, Ano64-237, Ano64-238, Ano64-239, Ano64-240, Ano64-241, Ano64-242, Ano64-243, Ano64-244, Ano64-245, Ano65-324, Ano65-325, Ano65-326, Ano65-327, Ano65-328, Ano66-310, Ano68-316, Ano64c, Ano65-488, Ano67-495, Ano68-477, Ano69-454]. listed [Ano64d]. living [She69]. load [Fri65c]. lock [Ano65-338]. lock-in [Ano65-338]. Lodge [Ano66-313]. Logarithmic [Dob69]. Logic [Ano67-205, Asi68, Key69, Nod65]. logical [Ano64-341]. Long [Kin65]. look [Ano65-126, Ano65-269, Ben68, Eva69, Fro69]. looking [Ano65-140]. looks [Oli65a]. loran [Fri67d, Sha68b]. Loran-C [Sha68b]. Loss [Ano66-465, Ano65-419]. Low [Ano66-314, Ano64-266]. Low-Temperature [Ano66-314]. LSIs [Hoc69]. lumped [Hue69]. lunar [MW66].

M [Cho64]. M.I.T. [Ano65-249, YW66].

MAC [Fan65]. Machine

[Ano64-246, Lin65b, Lin65c, Lin65d, AOB64, Ano64-141, BL68, Lin66c]. machines [Alg64, Pie68a]. Made [Ano64-268, Lin66b].

Maecon [Ano64-247]. magazines [Fri68a]. Magnetic [Sun69, Fri67c]. magnetics [Ano64-184, Ano65-364]. Magnetoelastic [Rod65b]. magnetohydrodynamics [Pet65]. magnets [Mon64]. Main [Ano64-273]. Maintaining [Ebe64]. maintenance [And64]. Major [Oli67]. Making [Ano67-341, Noy68, Ano65-339]. Man [Lin66b, Car69, Lin66c, Ano67-432]. man-machine [Lin66c]. man-made [Lin66b]. management [Ano64a, Ano65-321, For64]. manager [Bie68]. managers [Rob68b]. manipulation [FS67a]. Manual [YW66, Ano64-303]. manufacture [Ano68-190]. Map [Ano68-414, Ano65-461]. maps [Ano65-226]. March [Ano69-277]. marine [Jan64, Ano64-303]. Mariner [Ano64-333, Mat65]. Mariner-C [Ano64-333]. Market [Fin69, Bla64]. Marriott [Pra68]. Mars [Mat65, Ano64-333, Ano66-353]. Maser [Ano64-248]. Masers [Ano69-332]. mass [AWSB67]. Masthead [Ano65-329, Ano65-330]. materials [Ano65-222, Koh69]. Mathematics [Pur67, Gor68]. matrices [FK64, Fra64b, Fra64c]. Matrix [Ano64-249, Fra64a, FK64, Fra64b, Fra64c, Fra64d, Fra64b]. Maximum [Ano69-297]. Maxwell [Boo64, Ano65-331, Dib64]. May [Ano65-198, Ano64-346, Ano64-293]. Mcmahon [Gut65]. Meaning [Ano66-404]. means [Her68a, Pie66]. Measurements [Ano64-248, Ano65-147, Cot65, Bra64, RB64, Sel66]. measures [Ano65-338]. measuring [Ano65a]. mechanical [Ano64-265, Lig67]. mechanisms [Ano65-269, Rod65b]. mechanization [Ama66]. Medal [Ano64-292]. medalist [Ano65-486]. medals [Ano64-267]. media [Ano64-302]. medical [Lin65a]. medicine [Ano64-111, Ano65-365, Ano65-485, Lin67b, Lin69c]. medium [Nol67, WP69]. meet [Ano64-304, Whi69].

Meeting

[Ano65-489, Abr66, Ano64-186, Ano64-293, Ano64-322, Ano65-383, Ano65-462, Wil65a, Ano64-138, Ano66b, Tom69]. meetings [Ano65-462]. megalith [Ano65-246]. Melvin [Nod65]. members [Ano64-336]. Membership [Ano65-332, Ano67-344, Ano68-357, Ano64-211, Her68a]. Memorial [Ano64-250]. memoriam [Ski64]. Memories [Raj65, SSC67]. memory [Ano64-344]. Men [Pie68a, SSF65]. Merchant [Ano64-303]. Mercury [Ano66-352]. merger [Fri69a]. Merit [Ano69-226]. Message [Her68c, Mac67c]. metal [Pet65]. metallurgical [Sun69]. Metals [Ano67-336, Ano67-337, Ehr65]. meteor [Ano65-361]. meteor-burst [Ano65-361]. Method [Ano66-404]. methods [Ano65-270, CB66, The66]. metro [BW66]. MHD [Bro64a, Tsu67]. Michael [Dib67]. Michaels [Cut64a]. Michigan [Ano64-251]. Microbridges [Ano68-317]. Microelectronics [Ano64-252]. Microminiature [Ano69-298]. microminiaturized [CR69]. Micron [Ano68-296, Ano68-297]. microphone [Ano69-374]. microscope [Pea67]. microscopy [Ano66-315]. Microwave [Ada66, Ano65-333, Ano65-335, Ano65-334, Ano69-299, BM64, OBM^{+}64, Ano64-132, Ano65-337, Ano66-309, BG64, Dam67, IN69, Wei65, Cho64]. might [Ano66-238]. Milestone [Ano69-201]. Military [Ano65-336, Noy64, Wei66b]. Millimeter [Ano66-314, Ano67-342, Ano68-356, AFW68, Kin65]. Millimeter-To-Infrared [Ano68-356]. mines [Fri67c]. miniature [RJ68]. Miniaturization [Ano65-335, Ano65-334]. Minimum [Ano66-465]. minutes [Ano66-283]. mirrors [Ano65-339, Raw69]. missile [Til69a]. mission [She68]. mobile [Cox65]. model [CLM^{+}69]. modeling [Mil68b]. models [Lin65c]. modern [Ano67-369, Ano68-215, Boo64, Fri68c, KE68, Moo68, Sif65]. modification [Til69b, Til69c, Til69d]. modulation [DR65, The66]. Molecular [Ano67-343]. Molecular-gas [Ano67-343]. molecule [Ano65-231]. molecules [Tow65]. Monitoring [Flo68, VSG^{+}66]. Monolithic [Ano68-318, Phi64]. Monte [Ano68-294]. month [Ano64-318, Ano64-319]. Montreal [Ano67-340]. Moon [Ano65-461]. Morris [Ano65-271]. MOS [War67]. mosaic [WPS^{+}69]. Motion [Hem66]. Motorola [Ano64-268]. Motors [Ano65-130, Ano65-132, Ath64]. movements [Flo68]. mover [CdKS66]. Movies [Ano68-319]. moving [Ano65-141, Jam65]. much [Ano66-238]. Multielement [WPS^{+}69]. Multiple [McC67]. Multiple-access [McC67]. multiprocessing [Cri64]. Multiprogramming [Cri64]. multivariate [Ano68-294]. Munich [Ano65-286]. mystery [Ano65-246].

N [Cut64a]. Name [Ano68-371]. Narrow [Ano69-299]. NASA [YW66]. National [Ano65e, Ano65-350, Fri66a, Ber64, Fel64, Hol64, Ano64-264, Ano64-265, Ano64-299, Har66]. nations [Ano64-132]. natural [Ano67-432, Hal67]. natural-language [Hal67]. nature [EMK69, Gro65, Sal65a, Sal65b]. navigation [Jan64, MC66]. Navy [Fri67d, Fri67c]. NBS [ML65, Ano64-342, Ano65-337]. near [Ano65-416]. nearly [Ano65-339]. neat [Ano65-131, Ano65-360]. need [Ano66-365, Dun65, Mic67]. needs [Tru64]. Negative [Kro68, Cop67]. negative-resistance [Cop67]. nerve [Ano65-440, Wei66a]. Netherlands [Cot66]. Network [Ano65-387, Hue69]. networks [LMT64]. News [Ano64-142, Ano64-263, Ano64-264, Ano64-265, Ano64-255, Ano64-256, Ano64-257, Ano64-258, Ano64-259, Ano64-260, Ano64-261, Ano64-262, Ano64-254, Ano65-350, Ano65-351, Ano65-340, Ano65-341, Ano65-342, Ano65-343, Ano65-344, Ano65-345, Ano65-346, Ano65-347, Ano65-348, Ano65-349, Ano66-239, Ano66-240, Ano66-241, Ano66-242, Ano66-317, Ano66-318, Ano66-319, Ano66-320, Ano66-321, Ano66-322, Ano66-323, Ano66-324, Ano66-325, Ano66-326, Ano66-327, Ano67-345, Ano67-346, Ano67-347, Ano67-348, Ano67-349, Ano67-350, Ano67-351, Ano67-352, Ano67-353, Ano67-354, Ano67-355, Ano67-356, Ano68-321, Ano68-322, Ano68-323, Ano68-324, Ano68-325, Ano68-326, Ano68-327, Ano68-328, Ano68-329, Ano68-330, Ano68-331, Ano68-332, Ano69-307, Ano69-308, Ano69-309, Ano69-310, Ano69-311, Ano69-312, Ano69-313, Ano69-314, Ano69-315, Ano69-316, Ano69-317, Ano69-318, Ano69-305, Ano69-306]. Newspaper [Ano68-190]. next [Cre65, Sha67]. Nimbus [Ano64-266]. no [Ano64-335]. Nobel [Tow65]. Noble [Ano65-407]. Noise [Ano66-314, Ano69-275, Ano65-338]. Nominations [Ano64-267]. non [Ano66-339]. non-stick [Ano66-339]. nondomestic [For68]. nonmagnetic [Har66]. nonuniform [Ano64-302]. Northeast [Ano66-235, Fri66b]. Northwest [Dom69, LL65, LPK65]. notice [Ano65-212]. notices [Ano64-272]. November [Ano64-247, Ano64-252, Ano64-246, Ano65-299]. Nuclear [Ano64-269, Ano65-352, RJ68, Rod65a, SW64, Sha67, Ano64-158, CN67, Hay64, Kna69, Lin66a, Mar66]. Nuclei [Ano65-452]. number [Cha67]. Numerical [TA67]. NY [Hen65].

Obituaries [Ano64-272, Ano64-273, Ano64-270, Ano64-271, Ano65-353, Ano65-354, Ano65-355, Ano65-356, Ano65-357, Ano65-358, Ano65-359, Ano66-328, Ano66-329, Ano66-330, Ano66-331, Ano66-332, Ano66-333, Ano66-334, Ano66-335, Ano66-336, Ano66-337, Ano66-338, Ano67-357, Ano67-358, Ano67-359, Ano67-360, Ano67-361, Ano67-362, Ano67-363, Ano67-364, Ano67-365, Ano67-366, Ano67-367, Ano67-368, Ano68-333, Ano68-334, Ano68-335, Ano68-336, Ano68-337, Ano68-338, Ano68-339, Ano68-340, Ano68-341, Ano68-342, Ano68-343, Ano68-344, Ano69-319, Ano69-320, Ano69-321, Ano69-322, Ano69-323, Ano69-324, Ano69-325, Ano69-326, Ano69-327, Ano69-328, Ano69-329, Ano69-330, Ano64-263]. observations [Oli64]. Obsolescence [Ano66-425]. obtainable [Ano64-266]. Obviously [Ano65-131, Ano65-360]. Occupancy [Ano69-297]. Ocean [Ano65-361, Fri66c, Fri66d, Ano65-226, Bra64, Fri66e, KE68]. Oceanographic [Ano65-361]. Octave [Ano67-370]. Octave-Bandwidth [Ano67-370]. October [Ano64-176, Ano64g, Ano65-235, Ano65-286, Ano65-279, Ano65-417, Ano65-444, Ano65-462]. off [Bro64b]. offerings [Ano68-241]. offers [Ano65-182, Ano65-246]. officers [Ano65-352]. old [All64, Ano65-323, Wal64]. on-line [VSG^{+}66]. One [Ano67-369, Ano64-125]. open [Ano64-211]. Openings [Ano65-139]. operation [Ano64-269, Ano64-303, Rin67, Sco66, Ano69-331]. operations [Ano64-184, Fra64d]. Optical [Coo66, Cut64b, DeL68, Dob68, Ano65-362, BKW67, Ehr65, Las66, RK65]. Optics [Ano66-313, Kam65]. optimization [Ano64-126, BFS68, Mil68b]. orbit [Ano64-54]. Organization [Whi69]. organized [Ano65-408]. oriented [BL65b]. origin [Ano66-405, Sus66]. Original [Ano64-274]. orthogonal [Fra64c]. Oscillator [Ano65-362, Ano64-305, Gre69]. Oscillators [Ano67-398, CPH67, Hin69]. oscilloscope [Hem66, Zim69]. other [Ano64-335, Dun65, Oli64]. Our [Ano65-132]. outlook [GN69]. output [Ano64-125].

Pacific [Dom69, LL65, LPK65]. Page [Ano65-380, Ano64-96, Ano65-177, Ano65-178, Ano65-179, Ano65-180, Ano65-181, Ano65-278, Ano65-389, Ano65-390, Ano66-208, Ano66-209, Ano66-210, Ano66-211, Ano66-212, Ano66-213, Ano66-214, Ano66-215, Ano66-216, Ano66-217, Ano66-218, Ano66-219, Ano67-243, Ano67-244, Ano67-245, Ano67-246, Ano67-247, Ano67-248, Ano67-249, Ano67-250, Ano67-251, Ano67-252, Ano67-253, Ano67-254, Ano68-216, Ano68-217, Ano68-218, Ano68-219, Ano68-220, Ano68-221, Ano68-222, Ano68-223, Ano68-224, Ano68-225, Ano68-226, Ano68-227, Ano69-189, Ano69-190, Ano69-191, Ano69-192, Ano69-193, Ano69-194, Ano69-195, Ano69-196, Ano69-197, Ano69-198, Ano69-199, Ano69-200]. painter [Ano65-386]. Pan [Ano64-147, Ano65-363]. Panel [Ano64-275, Ano65-319]. Panelboard [Ano64-276]. Papers [Ano64-279, Ano64-277, Ano64-278, Ano65-364, Ano65-365, Ano65-363, Ano64-52, Ano64-126, Ano64-274, Ano64-293, Ano64-296, Ano64-337, Ano64-322, Ano64-345, Ano64-340, Ano65-206]. Parametric [Ano69-296, Ano69-332, Ano69-333, Kam65, Ano64-305, Sei66]. Part [Lin66b, Fow64, Ano65-227, Ano65-228, Ano68-296, Ano68-297, BL65a, BL65b, Cas64, Fra64a, FK64, Fra64b, Fra64c, Fra64d, Fri65b, Fri65c, LMT64, Lin65b, Lin65c, Lin65d, Lin66c, MGC65a, MGC65b, Sal65a, Sal65b, SRLF64a, SRLF64b]. participate [Ano64f]. Particle [ML65]. particles [Ano65-198, Ano65-416]. partnership [Hol67]. parts [Ano65-141]. passenger [Ano64-140]. Passivated [Ano64-268]. Passive [Ano67-370]. Past [Fri68e, Gun66, Nie68, Fri68c, Fri68d]. Patent [Oli67, CB66]. path [Ano66-255, Bis66]. Paths [Ano67-398]. Patrolling [Ano65-366]. pattern [Sus67, Nag67]. Patterns [Ano69-226, VM67, Hol67, Jac69]. Pavilion [Ano67-340]. PCM [Hua65]. peaking [Fri64b]. peculiarities [Lam66b]. People [Ano64-292, Ano64-280, Ano64-281, Ano64-282, Ano64-283, Ano64-284, Ano64-285, Ano64-286, Ano64-287, Ano64-288, Ano64-289, Ano64-290, Ano64-291, Ano65-367, Ano65-368, Ano65-369, Ano65-370, Ano65-371, Ano65-372, Ano65-373, Ano65-374, Ano65-375, Ano65-376, Ano65-377, Ano65-378, Ano66-340, Ano66-341, Ano66-342, Ano66-343, Ano66-344, Ano66-345, Ano66-346, Ano66-347, Ano66-348, Ano66-349, Ano66-350, Ano66-351, Ano67-371, Ano67-372, Ano67-373, Ano67-374, Ano67-375, Ano67-376, Ano67-377, Ano67-378, Ano67-379, Ano67-380, Ano67-381, Ano67-382, Ano68-345, Ano68-346, Ano68-347, Ano68-348, Ano68-349, Ano68-350, Ano68-351, Ano68-352, Ano68-353, Ano68-354, Ano68-355, Ano69-334, Ano69-335, Ano69-336, Ano69-337, Ano69-338, Ano69-339, Ano69-340, Ano69-341, Ano69-342, Ano69-343, Ano69-344, Ano69-345, Ano64-263, Fin67]. perception [Hem66, Lin65c]. perfect [Ano65-339]. Performance [Con65, Ano64-161, Ano66-281, Bur67, CdKS66]. Pergamon [Ano66-181]. Period [Ano66-352]. periodicals [Fri68a]. person [Ano65-440].

Personnel [Ano64-146, Ano64-144, Ano64-145, Ano65-127, Ano65-128, Ano65-129, Ano65-230]. Persons [Ano64f]. perspectives [Kin65]. PERT [KF64]. Petroleum [Ano64f, Ano65-379]. phase [Dia69]. phase-sensitive [Dia69]. Phased [Che68]. Phasing [Ano65-334]. phenomena [Ano65-364, Bre65]. Philadelphia [Ano65-299]. philosophical [Ano65-380]. philosophy [Bro64b, Gre66, Tel67]. photochromic [Smi66a]. Photodiodes [Ano67-383]. photography [Ano65-131, Ano65-360, EMK69]. Photon [Ano65-381, Ano65-248]. Photos [Ano67-384]. phototypesetting [KS69]. Physical [Bre65, Key69]. physically [Lan67]. Physics [Ano65-382, Mor65, Ano64-346, Lan67, Hig66]. Physiological [Ano67-273]. Pi [Smi66b]. Piano [Ano65-227, Ano65-228]. Piano-Tuner [Ano65-227, Ano65-228]. PICA [Ano64-293]. picture [Ano65-391, Hua65]. pictures [Ano64-266]. pinball [Ano64-141]. Pittsburgh [Ano64-304]. place [Ano64g, Ano65-279, Ano65-379]. plan [Ano65-236, Ano65-321]. Planet [Ano66-353, Ano66-352]. Planetary [Sha68a]. planned [Ano64-134, Ano64-275, Ano65-320, Ano65-404, Ano65-443]. Planning [Ben68, Rin67, Con68, DBW69, NP69, Whi69]. Plans [Ano65-383, Ano64-186, Ano64-264, Ano65-319, Til69c]. plant [BN69]. plants [Swi68]. Plasma [Ano64-294, Mil65]. Plasmas [Lev66, Sco68]. plastics [Ano64-150, Ano66-339]. plowing [Ano68-255]. PNP [Ano64-268]. point [Nas68]. Points [Ano67-282, Ano64-158, Ano64-159, Ano64-160, Ano64-161, Ano64-151, Ano64-152, Ano64-153, Ano64-154, Ano64-155, Ano64-156, Ano64-157, Ano65-246, Ano65-247, Ano65-248, Ano65-249, Ano65-238, Ano65-239, Ano65-240, Ano65-241, Ano65-242, Ano65-243, Ano65-244, Ano65-245, Ano66-243, Ano66-244, Ano66-245, Ano66-246, Ano66-247, Ano66-248, Ano66-249, Ano66-250, Ano66-251, Ano66-252, Ano66-253, Ano66-254, Ano67-276, Ano67-277, Ano67-278, Ano67-279, Ano67-280, Ano67-281, Ano67-283, Ano67-284, Ano67-285, Ano67-286, Ano67-287, Ano68-257, Ano68-258, Ano68-259, Ano68-260, Ano68-261, Ano68-262, Ano68-263, Ano68-264, Ano68-265, Ano68-266, Ano68-267, Ano68-268, Ano69-227, Ano69-228, Ano69-229, Ano69-230, Ano69-231, Ano69-232, Ano69-233, Ano69-234, Ano69-235, Ano69-236, Ano69-237, Ano69-238, Mul64]. police [Kel67]. policy [Ano64-185, Hol64, Ryd64a, WL64, EK69]. pollution [Net69, Sav69, Til69d]. Polynomial [Ano68-295]. polytech [Ano65-182]. Pont [Ano64-137]. pool [Rin67]. Popov [Abr66, Tom69, Wil65a]. portrait [Ano68-386]. position [Ano68-301, Mil68a]. positive [Nas68]. possible [Lan67]. Post [Ano64-295]. Post-baccalaureate [Ano64-295]. Potentiometric [Ano65a]. Power [Ano64-265, Ano65-148, Ano65-350, Ano66b, Bla65, BP69a, LL65, Yam66, Ano64-55, Ano64-186, Ano64-335, Ano65-141, Ano65-224, Ano65-225, Ano65-247, Ano65-299, Ano65-351, Ano65-337, Ano65-383, Ano65-417, Ano65-419, Ano66-235, BW66, BN69, Ben68, Bro64a, BM64, BG64, Bur67, CN67, Coh65, Dag68, EB67, EA64, Fri64b, Fri65b, Fri65c, Fri65e, Fri66b, Fri68b, Gre67a, Har65, HG65, Kna69, Kus68, Lam66b, LR67, LAT^{+}64, Net69, OBM^{+}64, Ric67, Rin67, Rod65a, SW64, Sha67, Sto69, Swi68, Tsu67, UN67, WP69, Wil64b, Fri68f, Fri68g]. power-system [EB67]. PPM [Ano65a]. Practical [Ano68-356, TZW^{+}67, Pie68b]. Practice [Ano66-292]. Prague [Geh66]. Pratt [Ano69-375]. Pre [Ano65-384]. Pre-Specification [Ano65-384]. Precision [Cot65]. prediction [Ano65-323, Kus68]. predicts [Ano64-143]. Prehistory [Sus69a]. Preprint [Ano64-296]. presence [Ano65-338]. Present [Har65, Fri68e, Fri68c, Fri68d, Gun66, Nie68, Raj65]. Present-day [Har65]. presented [Ano64-322]. President [Her68c, Mac67c, Oli67, Ano64-137, Ano64-138, EK69]. Press [Ano66-181]. Prevention [Fri68b]. previous [Ano64-161]. prices [Jac69]. Prime [CdKS66]. Prime-mover [CdKS66]. Principles [Ano65-183, Hor68a, Kam65]. printing [Ano68-190, Fri68a]. Privacy [Lic68]. Prize [Ano65-407]. prizes [Ano64-267]. Probability [Ano69-453]. probing [Fri66d]. Problem [Ano69-275, Fri65a]. Problems [Ano65-387, Ano67-397, Yam66, AI65, Hue69, Key69, TA67, Til69d, Wes68]. Procedure [Ano67-344, Ano68-357]. Proceedings [Ano64-297, Ano64-318, Ano64-319, Ano65-437, Ano64-279]. Process [Ano64-268, Ano66c]. processes [Ama66, Sun69]. processing [Dob68, Mal69, Mea65, Ols64]. produced [Ano65-419]. produces [Ano65-440]. Product [Ano66-239, Ano66-240, Ano66-241, Ano66-242, Ano69-300, Ano69-301, Ano69-302, Ano69-303]. Production [Fri68a, Tow65]. Products [Ano64-142, Bro64c, SB68]. profession [Lin66e]. Professional [Ano64-298, Ano65-133, Ano65-134, Ano65-135, Ano65-136, Ano65-137, Ano65-138, Ano65-385, Ano66-354, Ano66-355, Ano66-356, Ano66-357, Ano66-358, Ano66-359, Ano66-360, Ano66-361, Ano66-362, Ano66-363, Ano66-364, Ano67-385, Ano67-386, Ano67-387, Ano67-388, Ano67-389, Ano67-390, Ano67-391, Ano67-392, Ano67-393, Ano67-394, Ano67-395, Ano67-396, Ano68-358, Ano68-359, Ano68-360, Ano68-361, Ano68-362, Ano68-363, Ano68-364, Ano68-365, Ano68-366, Ano68-367, Ano68-368, Ano68-369, Ano69-346, Ano69-347, Ano69-348, Ano69-349, Ano69-350, Ano69-351, Ano69-352, Ano69-353, Ano69-354, Ano69-355, Ano69-356, Ano69-357, Ano64-185, How69, She66d, WL64]. professor [Ano65-386, Ano66-366, Ano64-299]. proficiency [Ebe64]. Program [Ano65-453, Ano67-457, Ano68-439, Ano69-417, Ano64-234, Ano64-264, Ano64-341, Ano65-223, Ano65-319, Ano65-361, Ano65-441, Nod65, She66b, Ano66-293]. Programming [Ano64-300, Ano67-397, Hal67]. programs [BL65b, Til69c]. Progress [Ano64-301, Ano68-370, Lax65, Omi66, Paa66, RK65, Sal65c, Win66, Ano64-342, Ano65-383, LL65, Wes68]. Project [SRLF64a, SRLF64b, Ano65-487]. projection [FRMH68]. promise [Wal64]. proof [Ano65-246]. Propagation [Ano64-302, AFW68, Ano65-443, Dur65]. Proper [Ano68-371]. Proper-Name [Ano68-371]. properties [Ehr65]. proportional [MC66]. Proposed [AOB64, Ano64-294]. Propulsion [Ano65-381, Bre65, Mil65]. prospect [Gol64]. prospects [Fri68c, Til69c]. Protecting [Hay64]. Proteus [WAS65]. prototype [Jac69]. provide [Ano65-247]. provides [Ano64-356]. Provisional [Ano65-462]. proximity [Fri67c]. Public [EK69]. Publications [Ano69-377, Ano64e, Ano64-197, Ano64-192, Ano64-193, Ano64-194, Ano64-195, Ano64-196, Ano64-187, Ano64-188, Ano64-189, Ano64-190, Ano64-191, Ano64-323, Ano64-324, Ano64-325, Ano64-326, Ano64-327, Ano64-328, Ano64-329, Ano64-330, Ano65-284, Ano65-285, Ano65-281, Ano65-282, Ano65-283, Ano65-439, Ano65-421, Ano65-422, Ano65-423, Ano65-424, Ano65-425, Ano65-426, Ano65-427, Ano65-428, Ano65-429, Ano66-291, Ano66-379, Ano66-380, Ano66-381, Ano66-382, Ano66-383, Ano66-384, Ano66-385, Ano66-386, Ano66-387, Ano66-388, Ano66-389, Ano66-390, Ano67-220, Ano67-412, Ano67-413, Ano67-414, Ano67-415, Ano67-416, Ano67-417, Ano67-418, Ano67-419, Ano67-420, Ano67-421, Ano67-422, Ano67-423, Ano68-388, Ano68-389, Ano68-390, Ano68-391, Ano68-392, Ano68-393, Ano68-394, Ano68-395, Ano68-396, Ano68-397, Ano68-398, Ano68-399, Ano69-160, Ano69-376, Ano69-378, Ano69-379, Ano69-380, Ano69-381, Ano69-382, Ano69-383, Ano69-384, Ano69-385, Ano69-386, Ano69-387]. Published [Ano66-292]. publishing [Ano68-190, Fri68a]. Pulse [Ano65-387, GJPM69, DR65, Koe65]. pulsed [Dag68]. Pumped [Fri64b, Kna69, Ano64-335]. purchases [For68]. Purdue [Hig66]. purity [Ano66-282]. purpose [Til69c]. Purposive [Lin68a]. Push [Ano64-303]. Push-button [Ano64-303].

quality [Ano64-338]. Quantitative

[Uma66]. quantum [Blo67, Gri66]. Quasi [Ano66-313]. Quasi-Optics [Ano66-313]. quasioptical [Ano64-301]. Quebec [Ano65-247]. question [Fri64c]. ’Quoddy [Fri64c].

R [Ski64, LMT64]. R-C [LMT64]. Radar [Gol66, Ano64-279, Ano65-140, Ano65-249, BP69b, Flo68, Fri66f, Fri67d, Mal69, MGC65a, MGC65b, Sha68a]. radars [Che68]. radiation [Ele68, Las66, Tow65]. Radio [AKAG66, Ano65-388, Ano67-273, Dur65, Moo67, Ano64-342, Ano65-442, Ano69-375, Cox65, EK69, Gre67b, Hoe66, Kra64, Pra68]. Radio-Frequency [Ano67-273, Ano65-442]. radioelectronics [Sif65]. Radiometer [Ano64-248]. radiotelegraphy [Sus69a, Sus69b]. Railroad [Ano64-304, Fri66a, Fri68e, Fri68c, Fri68d]. Railway [Fri67b]. rainstorm [Ano67-384]. Raman [Ano64-305]. Random [Cha67, SSC67]. random-access [SSC67]. Random-number [Cha67]. Ranger [Ano64-161]. rapid [AWSB67, Ano66-255]. rates [Koe65]. rational [Fra64c]. Raw [Rit68]. Rayleigh [Ano67-288]. reaction [Ano65-248]. Reactive [BW66]. reactor [Ano64-269]. reactors [DBW69]. read [Ano64-135]. Reader [Ano65-389, Ano65-390, Ano68-412, Ano68-413, Ano69-400, Ano69-401, Ano69-402, Ano69-403, Ano64-178, Ano64-95, Ano65-272, Ano65-278, Ano65-273, Ano65-274, Ano65-275, Ano65-276, Ano65-277, Ano69-173, Ano69-174, Ano69-175, Ano69-176, Ano69-177, Ano69-178, Ano69-179, Ano69-180, Ano69-181, Ano69-182, Ano69-183, Ano69-184, Ano69-185, Ano69-186, Ano69-187, Ano69-188]. Real [MOV66]. Real-time [MOV66]. reality [Lew66]. receive [Ano64-292]. received [Ano64-267]. Receiver [Ano68-297, Ano69-298, Ano64-198]. Receivers [Ano68-296, Ano68-297, Ele68, SN69]. receives [Ano65-271]. recently [Ano64-303]. receptive [Ano68-241]. recipient [Ano65-407]. Recognition [Nag67, Lin65b, Lin65c, Lin65d, ME65]. recommendations [Oli67]. Recommended [Ano66-292]. record [Ano64-143, Ano65-131, Ano65-360, Ano65c, Ano65d, Ano66a]. Recorder [Ano65-391]. Rectification [BG64]. rectifier [Gut67]. reduce [Ano65-198]. Reducer [Ano68-253]. reduces [Ano64-251]. Reduction [Ano66-426, Kor67]. reductions [Fra64c]. redundancy [Kor67]. reelects [Ano64-138]. Reflected [Ano69-418]. Reflection [Ano64-306]. Reflections [Ano64-316, Ano64-315, Ano64-317, Ano64-307, Ano64-308, Ano64-309, Ano64-310, Ano64-311, Ano64-312, Ano64-313, Ano64-314, Ano65-392, Ano65-393, Ano65-394, Ano65-395, Ano65-396, Ano65-397, Ano65-398, Ano65-399, Ano65-400, Ano65-401, Ano65-402, Ano65-403]. reflectivity [Ano64-279]. reflector [Ano64i]. regarding [Ano64-185]. Region [Ano65-404, Hir68a, Kin65, Ano67-323]. Regional [EA64]. registrants [Ano65-351]. regulator [WP69]. related [Ano65-364]. relations [Ano64-185, Sei66, WL64]. relationships [Til69b]. Relativity [Ell66]. relay [Ano66-365, IN69]. release [Ano64-355]. Reliability [Ano64-179, Ano64-275, Ano64-338, Fri68g]. reliable [Con68, Ric67]. reminiscence [Ano69-375]. remote [FS67a]. Remotely [Ano67-398]. Report [Abr66, Ano65-126, Cla64, Cot65, Elf65, Geh66, ML65, Pra64, Til69a, WL64, Wil65a, Ano64-295, Ano64-301, Ano65-438, Ano67-314, Bla65, Fri68b, Mir65, SSF65, Tom64, Ano68-301]. reported [Ano65-237]. Reports [Ano69-276, Her68b, Mac67b, She66c, Lin64, Oli65b]. requirements [Ano64-251, BP69a, McP67]. requires [Ano64-335]. Research [And69, Ano65-130, Ano65-132, Ano67-342, Fei64b, Kno64, Sus67, Ano65-229, Ano65-322, Ano65-361, Ano66-366, Ano68-254, BL65a, BL65b, RK65, Sco66, Wei65]. reservations [Ano64-51]. Residential [Law64]. resistance [Cop67]. resistors [GJPM69]. Resolution [Bro67b]. resonator [HB69]. resonator- [HB69]. resource [Ber64]. response [CdKS66]. responsibility [AH69].

Restless [Ano65-452]. retired [Ano64-273]. retires [Ano64-136]. retrieval [Sal65c]. Retrospect [Gol64]. return [Hac65]. reveals [Ano66-315]. reversal [Ano64-346]. Review [Boo64, Cho64, Cut64a, Dur65, Gut65, Hen65, Nod65, Pra68, MGC65a, MGC65b]. Reviewed [Ano64-318, Ano64-319]. reviews [Ano64-92, Ano64-93, Ano64-94, Ano65-167, Ano65-168, Ano65-169, Ano65-170, Ano65-171, Ano65-172, Ano65-173, Ano65-174, Ano65-175, Ano66-196, Ano66-197, Ano66-198, Ano66-199, Ano66-200, Ano66-201, Ano66-202, Ano66-203, Ano66-204, Ano66-205, Ano66-206, Ano66-207, Ano67-221, Ano67-222, Ano67-223, Ano67-224, Ano67-225, Ano67-226, Ano67-227, Ano67-228, Ano67-229, Ano67-230, Ano67-231, Ano67-232, Ano68-191, Ano68-192, Ano68-193, Ano68-194, Ano68-195, Ano68-196, Ano68-197, Ano68-198, Ano68-199, Ano68-200, Ano68-201, Ano68-202, Ano69-161, Ano69-162, Ano69-163, Ano69-164, Ano69-165, Ano69-166, Ano69-167, Ano69-168, Ano69-169, Ano69-170, Ano69-171, Ano69-172, BDS^{+}64, Bon64, FFC64, Fra64e, Gre64, LBCM65, Not64, Web64a]. revisited [Hea69]. Richmond [Ano65-235]. Rise [Ano65-405]. River [Fri68f, Ano65-247]. Roadblocks [Bis66]. Robert [Pra68]. rocked [Ano64-131]. rods [Ano64-344]. Role [Ano65-437, BO64, Bro67c, DBW69, End66, Rob65, Spe66]. Rotation [Ano66-352]. route [Ano66-255]. Rural [Ano65-406, Mir65].

S [Abr66, Ano64-138, Ano65f, Wil65a, Ano69-298]. S-Band [Ano69-298]. S. [Ano65-407]. saccharine [Ano65-231]. safety [EK69, Fri68c]. Salaries [Ano66-403, How69]. sale [Ano64-181]. sales [Ano64-143]. salty [Fri65f]. samples [Ano68-294]. sampling [Zim69]. sampling-oscilloscope [Zim69]. Satellite [AKAG66, Ano68-372, Ano64-294, Ano65-249, MM67, Wes66]. satellites [Ano65-249, Spe66]. Satisfaction [Ano65-408]. Saturn [Ano65-139]. Scan [Ano69-201]. scanned [WPS^{+}69]. scanner [Bro67b]. Scanning [Ano64-197, Ano64-192, Ano64-193, Ano64-194, Ano64-195, Ano64-196, Ano64-320, Ano65-284, Ano65-285, Ano65-409, Ano65-410, Ano65-411, Ano65-412, Ano65-413, Ano65-414, Ano65-415, Ano66-367, Ano66-368, Ano66-369, Ano66-370, Ano66-371, Ano66-372, Ano66-373, Ano66-374, Ano66-375, Ano66-376, Ano66-377, Ano67-399, Ano67-400, Ano67-401, Ano67-402, Ano67-403, Ano67-404, Ano67-405, Ano67-406, Ano67-407, Ano67-408, Ano67-409, Ano67-410, Ano68-373, Ano68-374, Ano68-375, Ano68-376, Ano68-377, Ano68-378, Ano68-379, Ano68-380, Ano68-381, Ano68-382, Ano68-383, Ano68-384, Ano69-358, Ano69-359, Ano69-360, Ano69-361, Ano69-362, Ano69-363, Ano69-364, Ano69-365, Ano69-366, Ano69-367, Ano69-368, Ano69-369, Pea67]. scatter [Gun66]. scattering [Ano65-323]. scheduled [Ano64-110, Ano64-247, Ano64-252, Ano64-338, Ano65-235, Ano65-444, Ano65-420]. School [Geh66]. Science [Fri65f, Por65, Ano65-352, Ano65-444, Ano66-464, Com68, Gib65, Gut65, Hol64, Jan64, Kom67, Lin66a, Sho66, Boo64]. sciences [Dob68, Gut65]. Scientific [Ano66-292, Til69b]. scientist [Ano65-246]. Scientists [Ano65-139, Ano65-416, Ano64-147, Dun65]. screens [Jac69]. script [Lin65d]. sea [And64, Ano64-159, Fri65f, FAB64, Moo67, Ano65-461]. sea-floor [And64]. Search [CB66, Wal69]. Second [AKAG66, Ano65-417, Ano65-418]. Secretary [Pra64]. secrets [Car69]. sections [Ano64b]. See [Ano67-411, Ano68-385]. seeks [Ano64-293, Ano64-337]. Seismic [Ano64-321, Fri66e]. selected [Ano64-292]. self [WPS^{+}69, Ano68-386, Nod65]. self-introduction [Nod65]. Self-portrait [Ano68-386]. self-scanned [WPS^{+}69]. Selling [Ano65-384]. semiautomatic [SRLF64a, SRLF64b]. Semiconductor [Koe65, LP67, Ano64-253, Cho64, Cop67, Gut67, Lax65]. Semiconductor-diode [LP67]. semiconductors [Kro68]. semimetal [Esa66]. seminar [FRMH68]. seminars [Ano65-337]. sensitive [Dia69]. Sensitivity [Ano68-296, Ano68-297]. sensors [WPS^{+}69]. separation [Gol66]. September [Ano65-166, Ano65-336, Ano65-379]. Service [Ano64-146, Ano64-144, Ano64-145, Ano65-127, Ano65-128, Ano65-129, Ano65-230, Ano68-412, Ano68-413, Ano69-400, Ano69-401, Ano69-402, Ano69-403, Ano64-178, Ano64-95, Ano65-272, Ano65-278, Ano65-273, Ano65-274, Ano65-275, Ano65-276, Ano65-277, Ano65-389, Ano65-390, Ano69-173, Ano69-174, Ano69-175, Ano69-176, Ano69-177, Ano69-178, Ano69-179, Ano69-180, Ano69-181, Ano69-182, Ano69-183, Ano69-184, Ano69-185, Ano69-186, Ano69-187, Ano69-188, Con68, Fri68g]. Services [Ano65-133, Ano65-134, Ano65-135, Ano65-136, Ano65-137, Ano65-138, Ano64-298, Ano65-385, Ano66-354, Ano66-355, Ano66-356, Ano66-357, Ano66-358, Ano66-359, Ano66-360, Ano66-361, Ano66-362, Ano66-363, Ano66-364, Ano67-385, Ano67-386, Ano67-387, Ano67-388, Ano67-389, Ano67-390, Ano67-391, Ano67-392, Ano67-393, Ano67-394, Ano67-395, Ano67-396, Ano68-358, Ano68-359, Ano68-360, Ano68-361, Ano68-362, Ano68-363, Ano68-364, Ano68-365, Ano68-366, Ano68-367, Ano68-368, Ano68-369, Ano69-346, Ano69-347, Ano69-348, Ano69-349, Ano69-350, Ano69-351, Ano69-352, Ano69-353, Ano69-354, Ano69-355, Ano69-356, Ano69-357, Cox65, She66b]. sessions [Ano65-319, Ano65-351]. set [Ano64-143]. SFPE [Ano64-336]. shaft [Mil68a]. Shanken [Ano65-223]. shape [Ano65-231]. Shepherd [Ano66-378]. ship [Ano64-303]. Short [Ano68-387, Ano69-373, Ano69-371, Ano69-370, Ano69-372, Kin65]. shortest [Ano66-255]. should [Ano64-267]. Show [Ano66-312]. shows [Ano65-461]. Side [Ano65-140, Tur67]. Side-looking [Ano65-140]. Signal [KL64, Ano68-386, Mal69, Mea65]. signaling [Fri68c]. signals [Ano65-338]. Silicon [Ano64-268, Rob68a, VGRM69]. Silicon-gate [VGRM69]. Similarity [Fra64c]. simple [Ano69-374]. Simply [Ano67-397]. Sims [Cho64]. simulated [Ano64-158]. simulation [Ano68-294, Hea64]. Single [Ano65-419]. Single-frequency [Ano65-419]. sinusoidal [Gre69]. Sir [Gos68]. six [Ano65-404]. Sixth [SSF65]. Sixty [Ano69-375]. sizes [Ben68]. Skifter [Ski64]. Skin [Ano67-273]. slate [Ano65-352]. slated [Ano64-111, Ano64-179, Ano65-166, Ano65-286, Ano65-336, Ano65e, Ano65-417]. social [Til69d]. Societies [Ano64-146, Ano64-144, Ano64-145, Ano65-127, Ano65-128, Ano65-129, Ano65-230, Ano66-238, Ano65-321]. Society [Ano65-300, Ano65-301, Ano65-302, Ano65-303, Ano65-304, Ano65-305, Ano65-306, Ano65-307, Ano66-306, Ano68-314, Ano69-290, Ano69-291, Ano69-292, Ano69-293, Ano69-294, Ano69-295, Tom69, Ano65-408, She66d, Abr66, Wil65a]. sociosphere [Mil68b]. sociotechnical [Hof68]. sodium [HHA^{+}66]. software [Stu68]. solar [Ano64-337]. solicited [Ano64-277, Ano64-278, Ano65-363]. Solid [Ano65-420, Dam67, IN69, Kai66, Sto69, Ano65-196, Ano66-195, FS67b, Har65, LAT^{+}64]. Solid-state [Ano65-420, Dam67, IN69, Kai66, Sto69, Ano65-196, Ano66-195, FS67b, Har65, LAT^{+}64]. Solidifying [Ano67-336, Ano67-337]. solids [Ano65-126, Ein66]. solution [Ros65, TA67]. solve [Ano65-246]. Solved [Ano65-387]. Some [Col66a, Hir67, Pie68b, Tom69]. Sonar [KE68, Fri67d]. sonic [Fri66e]. sonics [Ano64-176]. sought [Ano64-126, Ano64-279, Ano65-365]. sound [Adl67, Ano65-391, Ano65-440, Ano67-341, Hir67]. sounding [Ano64-321]. Source [Ano66-463, Ano64-335]. sources [LP67, Rit68]. Southwest [Dom69]. Soviet [Ano64-322, Pop69, She69]. Space [Ano66-404, Tra69, AI65, Ano65-319, Ano65-416, BP69a, BKW67, Bro64d, Cut64a, DGH67, Gut65, Lev66, Mil65, Por65, Sco66, Sco68, von65]. Spacecraft [Hor68a, Hor68b, Ano65-226, Rei66]. spaced [Gol66]. Spark [Nas68]. spec [Ano64-135]. Special [Ano64-323, Ano64-324, Ano64-325, Ano64-326, Ano64-327, Ano64-328, Ano64-329, Ano64-330, Ano65-421, Ano65-422, Ano65-423, Ano65-424, Ano65-425, Ano65-426, Ano65-427, Ano65-428, Ano65-429, Ano66-379, Ano66-380, Ano66-381, Ano66-382, Ano66-383, Ano66-384, Ano66-385, Ano66-386, Ano66-387, Ano66-388, Ano66-389, Ano66-390, Ano67-412, Ano67-413, Ano67-414, Ano67-415, Ano67-416, Ano67-417, Ano67-418, Ano67-419, Ano67-420, Ano67-421, Ano67-422, Ano67-423, Ano68-388, Ano68-389, Ano68-390, Ano68-391, Ano68-392, Ano68-393, Ano68-394, Ano68-395, Ano68-396, Ano68-397, Ano68-398, Ano68-399, Ano69-376, Ano69-377, Ano69-378, Ano69-379, Ano69-380, Ano69-381, Ano69-382, Ano69-383, Ano69-384, Ano69-385, Ano69-386, Ano69-387, Bla65, Mir65, SSF65, Tom64, Ano64d, Ano64-197]. specific [Ano65-248]. Specification [Ano65-384]. spectra [Ric67]. Spectral [Ano64-331, Ano64-332, Ano65-438, Ano65-439, Ano65-437, Ano65-430, Ano65-431, Ano65-432, Ano65-433, Ano65-434, Ano65-435, Ano65-436, Ano66-391, Ano66-392, Ano66-393, Ano66-394, Ano66-395, Ano66-396, Ano66-397, Ano66-398, Ano66-399, Ano66-400, Ano66-401, Ano66-402, Ano67-424, Ano67-425, Ano67-426, Ano67-427, Ano67-428, Ano67-429, Ano67-430, Ano67-431, Ano68-400, Ano68-401, Ano68-402, Ano68-403, Ano68-404, Ano68-405, Ano68-406, Ano68-407, Ano68-408, Ano68-409, Ano68-410, Ano68-411, Ano69-388, Ano69-389, Ano69-390, Ano69-391, Ano69-392, Ano69-393, Ano69-394, Ano69-395, Ano69-396, Ano69-397, Ano69-398, Ano69-399, Cut67a, Cut67b, Cut67c, Cut67d, DGH67, Mac64, Ryd64b, Ryd64c, Ryd64d, Ryd64e, Ryd64f, Ryd64g, Ryd64h, Ryd64i, Wil65b, Wil65c, Wil64a]. spectroscopy [Uma66]. Spectrum [Ano68-412, Ano68-413, Ano69-400, Ano69-401, Ano69-402, Ano69-403, Ano65-388, EK69, Gre67b, Ano64-53, Ano64-95, Ano65-146, Ano65-176, Ano67-233, Ano67-234, Ano67-235, Ano67-236, Ano67-237, Ano67-238, Ano67-239, Ano67-240, Ano67-241, Ano67-242, Ano68-203, Ano68-204, Ano68-205, Ano68-206, Ano68-207, Ano68-208, Ano68-209, Ano68-210, Ano68-211, Ano68-212, Ano68-213, Ano68-214, Ano69-173, Ano69-174, Ano69-175, Ano69-176, Ano69-177, Ano69-178, Ano69-179, Ano69-180, Ano69-181, Ano69-182, Ano69-183, Ano69-184, Ano69-185, Ano69-186, Ano69-187, Ano69-188, Gol64]. Speech [Ano67-432, Ols64, Ano65-380, Ano68-254, Lin65b, Lin65c]. speed [Jam66, Lig67, Mil68a]. speeds [Ano64-140]. sphere [Ano65-249]. Spiral [Ano68-414]. sponsor [Ano65-148]. sponsored [Ano65-350]. spot [Bro67b]. Spring [Ano69-373, Ano69-372]. springs [Lin67a]. St. [Ano65-247]. Staff [Ano64-235, Ano65-324, Ano65-325, Ano64-236, Ano64-237, Ano64-238, Ano64-239, Ano64-240, Ano64-241, Ano64-242, Ano64-243, Ano64-244, Ano64-245, Ano65-326, Ano65-327, Ano65-328]. stand [Bar66b]. Standard [Ano67-205]. standardization [Gro66]. standards [Ano64-342, HP67, Hir68a, McL66, Ano64-180, Ano65-280, Har66]. standby [Ano65-224, Ano65-225]. star [Ano64-54]. starship [Oli64]. start [Ano64-333]. Starting [Ano66-403]. starts [Ano64-211]. state [Ano65-196, Ano65-420, Ano66-195, Dam67, FS67b, Har65, IN69, Jan64, Kai66, LAT^{+}64, Sto69, Ano64-251]. statement [Ano68-301]. States [Coh65, Fri68e, GN69]. static [Ano65-224, Ano65-225, Ano65-299]. station [Ben68]. stations [And64, Tsu67]. status [Ano64-150]. steering [Ano64-294]. Steinmetz [Ros65, WAS65]. step [Tow69]. step-by-step [Tow69]. Stephenson [Cho64]. Steps [Lin69a]. stick [Ano66-339]. Stimulating [Kop69]. Stimulation [Ano65-440]. Stirling [Ano65-132]. Stonehenge [Ano65-246]. storage [BRP64, BFS68, Cam65, Fri64b, Gia65, Kna69]. Stored [Ano65-441]. Stored-program [Ano65-441]. stormproof [Gre67a]. story [Fri68f, Fri68g]. Strategy [Ano69-404]. Street [Ano64-181]. strength [Mon66]. stress [Ano65-234]. structures [Ano66-315]. studied [Ano64-55, Ano65-197, Ano65-333, Ano65-406]. Studies [AKAG66, Ein66]. Study [Ano64-333, Ano64-113]. Subject [Ano64-334, Ano64-355, Ano65-382]. Subjective [Ano66-404]. Submarine [Mot65, OT66, Ano64-159]. Submillimeter [Ano66-314]. success [Ano65-408, Ano69-331]. successful [Ano65-487]. Such [Ano66-256]. Sum [Ano69-333]. Summary [Okr64b]. Summer [Ano65-351, Ano69-373, Ano69-371, Geh66]. Summer-Fall [Ano69-371]. Sun [Ano64-335, Ano68-320, Ano69-304, Smi66a]. Sun-pumped [Ano64-335]. Superconducting [Ano65-442, Ano68-370, Ath64, Mon64, Pri66, Cas64]. superconductive [RJ68]. superconductivity [Ano66-405]. superconductor [Lav68]. Superfluidity [Ano66-405]. supermagnets [Mon66]. Supervisory [FS67a]. supplement [Ano64-112, Ano66-406]. supply [Ano65-247, Dag68, Hic66]. surfaces [Ano65-126]. survey [Her68a, Til69b]. Surveyor [MW66]. survival [Moo64]. SWE [Ano64-336]. switches [Har65]. switching [Ano64-341, Ano65-441, KKV65, Koe65]. Symbols [Ano67-205]. Symposium [AKAG66, Ada66, Ano64-337, Ano64-338, Ano65-443, Cot66, Fri66a, Mea65, Ano64-176, Ano64-252, Ano64-341, Ano65-286, Ano65-319, Ano65-320, Ano65-382, Ano65-484, EBO^{+}64, SSF65, Ano64-275]. synchronization [Sha68b]. Synchronizing [Ano67-398]. synchronous [Alg64]. synergism [Lin69a]. Synthesizer [Ano68-315]. Synthesizing [Mit69]. synthetic [BP69b]. synthetic-aperture [BP69b]. system [Ano64-140, Ano64-149, Ano64-158, Ano64-337, Ano65-196, Ano65-361, Ano65a, BW66, CdKS66, EB67, Fan65, Hor64, KKV65, Kus68, Mat65, MW66, MOV66, Sco68]. Systems [Ano65-387, Ano65-444, Hor68b, Ano64-294, Ano65-140, Ano65-224, Ano65-225, Ano65-417, Ano65-441, Bro67b, Chi66, Con65, EA64, FKS64, FK64, Fri64a, Fri65b, Fri65c, Fri68e, Fri68d, Fro69, FK66, Hay64, Hic67, Hir68b, IN69, KS69, Lin66c, Lin68a, Lin66e, McC67, McP67, Mic67, Mil68b, Ols64, Pea66, Rei66, ST67, Cut64a].

T [Ano64-273]. Table [Ano64-339, Ano65-445, Ano65-446, Ano65-447, Ano65-448, Ano65-449, Ano65-450, Ano65-451, Ano66-407, Ano66-408, Ano66-409, Ano66-410, Ano66-411, Ano66-412, Ano66-413, Ano66-414, Ano66-415, Ano66-416, Ano66-417, Ano66-418, Ano67-433, Ano67-434, Ano67-435, Ano67-436, Ano67-437, Ano67-438, Ano67-439, Ano67-440, Ano67-441, Ano67-442, Ano67-443, Ano67-444, Ano68-415, Ano68-416, Ano68-417, Ano68-418, Ano68-419, Ano68-420, Ano68-421, Ano68-422, Ano68-423, Ano68-424, Ano68-425, Ano68-426, Ano69-405, Ano69-406, Ano69-407, Ano69-408, Ano69-409, Ano69-410, Ano69-411, Ano69-412, Ano69-413, Ano69-414, Ano69-415, Ano69-416]. tables [Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69l]. tactics [Ano69-404]. take [Ano64g, Ano65-279, Ano65-379]. takes [Ano64-303, Bar66b]. Tale [Ano66-378]. talent [Ano64-355]. Tales [Ano65-452]. talk [Ano67-411, Ano68-385]. Talking [Ano67-268]. tantalum [Ano67-313]. tape [Ano65-391]. targets [Gol66]. Task [EK69]. tasks [Ano64-303]. teaching [Sho66]. Technical [Ano64-341, Ano64-340, Ano65-453, Ano66-292, Ano66-293, Ano66-419, Ano66-420, Ano66-421, Ano66-422, Ano66-423, Ano66-424, Ano66-425, Ano67-339, Ano67-445, Ano67-446, Ano67-447, Ano67-448, Ano67-449, Ano67-450, Ano67-451, Ano67-452, Ano67-453, Ano67-454, Ano67-455, Ano67-456, Ano67-457, Ano68-427, Ano68-428, Ano68-429, Ano68-430, Ano68-431, Ano68-432, Ano68-433, Ano68-434, Ano68-435, Ano68-436, Ano68-437, Ano68-438, Ano68-439, Ano69-417, HE68, Ano64-234, Ano64-345, Ano65-319, Ano65-351, Car69, Sor65]. technician [Bro69]. Technique [Ano66-426, All64, Ano66-195, KE68]. Techniques [Ano65-227, Ano65-228, Ano64i, Ano64-126, Ano65-182, Ano65-339, Cut64b, Fri67d, Hea64, Hun64, SRLF64a, SRLF64b]. Technological [Ano69-418, NP69, Web64b]. technology [Bak68, BL67, Cas64, Fow64, Fri69a, FS67b, Hol64, Joh67, Kai66, Kom67, Lav68, LMT64, Lin69b, Lin69a, Noy68, Pri66, Rob68a, VGRM69]. technosphere [Mil68b]. telecommunication [BL65a, BL65b, BP69a, Rei66]. Telecommunication-oriented [BL65b]. Telecommunications [Kob69, Fri68g]. telemetering [Ano64-278]. telemetry [Hoe66, Mat65]. telephone [Ano65-441, Mot65, Win66]. telephony [Lam66a]. Television [Ano64-161, Ano68-372, Bro67c, Chi66, Ele68, HP67, Hir68a, McL66, MW66, Roi67, SN69]. Temperature [Ano66-314, Ano65-461]. Ten [Ano64-342, Lin66d, Ano67-323]. term [Sus66]. Terminal [Ano69-224]. test [Lin66a, RJ68]. testing [Stu68]. tests [Ano64-141]. Thanks [Ano67-458]. Their [Mil68b, Bur67]. theoretic [Dix68]. Theoretical [Lin65c]. Theory [Ano67-272, Cot66, Ano64-341, Ano65-246, Ano65-320, Gri66, Lan67, Mar67, Wei66a, Geh66]. Thermal [Ano65-461, Fri68g]. Thermionic [Wil64b]. thermonuclear [Dou66]. Thick [Thu69]. Thin [Ano64-344, Ano67-470, Cas64, Fow64, LMT64, Ano66-282, GJPM69, Pri66, Thu69]. Thin-film [Ano64-344, Cas64, Fow64, LMT64, GJPM69, Pri66]. things [Fri67b]. think [Cre65]. Third [Ano64-345, BN69, Gor68]. throughout [Ano65-197, Fel64]. thrustors [Mil65]. Thyristors [Gut67]. tide [Fri64c]. Time [Ano64-346, Ano67-398, Fri64c, Sha68b, Ano65-248, MOV66, von65]. timetable [Ano65-462]. Tiros [Sch69a]. Today [Mic67, AWSB67, Rod65a, SW64]. together [Ano64d]. Tokaido [Kaw66]. Tokyo [Ano65-462, Tom64]. Tomorrow [AWSB67, Ben68, Bro67c, SW64]. tool [Wag69]. tools [Ano64-246, CR69, LS64]. torpedoes [Fri67c]. total [Ano65-322]. tour [Ber69]. Trade [Car69]. tradeoffs [Stu68]. Traffic [Ano67-272, Ano69-224, Fri69b, Smi67, Spe66]. Training [Ano64-179]. Tranquillity [Ano65-461]. transducer [HB69]. Transform [Ber69, BM67]. transformer [Gia66]. Transient [Ano68-452]. Transients [Ano65-463, Ano65-464, Ano65-465, Ano65-466, Ano65-467, Ano65-468, Ano65-469, Ano65-470, Ano65-471, Ano65-472, Ano65-473, Ano65-474, Ano66-439, Ano66-440, Ano66-441, Ano66-442, Ano66-443, Ano66-444, Ano66-445, Ano66-446, Ano66-447, Ano66-448, Ano66-449, Ano66-450, Ano67-471, Ano67-472, Ano67-473, Ano67-474, Ano67-475, Ano67-476, Ano67-477, Ano67-478, Ano67-479, Ano67-480, Ano67-481, Ano67-482, Ano68-453, Ano68-454, Ano68-455, Ano68-456, Ano68-457, Ano68-458, Ano68-459, Ano68-460, Ano68-461, Ano68-462, Ano68-463, Ano69-431, Ano69-432, Ano69-433, Ano69-434, Ano69-435, Ano69-436, Ano69-437, Ano69-438, Ano69-439, Ano69-440]. transistor [CB66, Wal64]. Transistors [Ano64-268]. transit [AWSB67, Mic67]. transit-available [AWSB67]. transition [Hag68]. Translated [Ano64-347, Ano64-348, Ano64-349, Ano64-350, Ano64-351, Ano64-352, Ano64-353, Ano64-354, Ano65-475, Ano65-476, Ano65-477, Ano65-478, Ano65-479, Ano65-480, Ano65-481, Ano65-482, Ano65-483, Ano66-451, Ano66-452, Ano66-453, Ano66-454, Ano66-455, Ano66-456, Ano66-457, Ano66-458, Ano66-459, Ano66-460, Ano66-461, Ano66-462, Ano67-483, Ano67-484, Ano67-485, Ano67-486, Ano67-487, Ano67-488, Ano67-489, Ano67-490, Ano67-491, Ano67-492, Ano67-493, Ano67-494, Ano68-464, Ano68-465, Ano68-466, Ano68-467, Ano68-468, Ano68-469, Ano68-470, Ano68-471, Ano68-472, Ano68-473, Ano68-474, Ano68-475, Ano69-441, Ano69-442, Ano69-443, Ano69-444, Ano69-445, Ano69-446, Ano69-447, Ano69-448, Ano69-449, Ano69-450, Ano69-451, Ano69-452, Ano64-197]. Transmission [Ano65-147, Ano67-273, AJMS65, Ano65-206, Ano65-235, Ano66-281, BH64, Bro64d, Fri68g, Hua65, HG65, Jam65, Kai66, Lam66b, LR67, Len66, Mea65, Par69, Pie68b, Pop69, The66]. transmit [Ano65-361]. transport [Rod65b]. Transportation [Fri66a, Sei68]. Transportation-1993 [Sei68]. transversal [Rud67]. traveling [May65, Sun69]. traveling-wave [May65]. treasurer [Cla64]. Treaty [Kee65]. trend [HHA^{+}66]. Trends [AJMS65, FL66, Sif65, Wee66, Ano65-463, Ano65-464, Ano65-465, Ano65-466, Ano65-467, Ano65-468, Ano65-469, Ano65-470, Ano65-471, Ano65-472, Ano65-473, Ano65-474, Ano66-439, Ano66-440, Ano66-441, Ano66-442, Ano66-443, Ano66-444, Ano66-445, Ano66-446, Ano66-447, Ano66-448, Ano66-449, Ano66-450, Ano67-471, Ano67-472, Ano67-473, Ano67-474, Ano67-475, Ano67-476, Ano67-477, Ano67-478, Ano67-479, Ano67-480, Ano67-481, Ano67-482, Ano68-452, Ano68-453, Ano68-454, Ano68-455, Ano68-456, Ano68-457, Ano68-458, Ano68-459, Ano68-460, Ano68-461, Ano68-462, Ano68-463, Ano69-431, Ano69-432, Ano69-433, Ano69-434, Ano69-435, Ano69-436, Ano69-437, Ano69-438, Ano69-439, Ano69-440, Fin69]. Triangulation [Ano66-426]. tribute [Ano65-331]. trimming [GJPM69]. Tropospheric [Gun66]. tube [Her65]. tubes [Ano65-337, Cho64]. Tunable [Ano66-463, Ano65-362]. Tuner [Ano65-227, Ano65-228]. tuning [Hin69]. tunnel [Fei64b]. turbine [Fri65b]. turbulence [Col66b]. Turning [Ano66-464]. Tutorial [Ano65-484]. TV [Tay69]. TVA [For68]. Two [Ano65-248, Ano65-249, Ano66-283]. Two-photon [Ano65-248]. type [Ano66-315]. Types [Hog64].

U.S. [Ano64-303, Fri67d, Fri67c, Oli67, RB64, Tom69, VM67, Wei66b]. U.S.A. [Sto69]. U.S.S.R. [LR67]. Uhl [Ano64-273]. UHV [Par69]. Ultrasonic [Ein66, May65]. Ultrasonics [Lin69c, Ano64-176]. uncertainty [Lig67]. Uncovering [Ano66-404]. Underdeveloped [Ano68-372]. underground [Law64]. underlying [For64]. understand [Lin68b]. Uninterruptible [Ano65-141]. Union [Pop69]. unique [Ano65-246]. unit [Fri65b]. United [Coh65, Fri68e, GN69]. Units [Ano66-292, Bar66b]. Universal [Roi67, Wes66]. universities [EB67]. university [BO64, Ano65-485]. Upadi [Ano64h]. Upconversion [Ano69-296]. Upconverter [Ano69-333]. urban [Mic67, Sav69, Sie68]. urd [Ano64-112]. URSI [Cot66]. USA [Hen65]. Use [Ano64-355, Ano64-356, Ano68-320, Ano69-304, Ano64-158, Ano64-178, Ano65-272, Ano65-278, Ano65-273, Ano65-274, Ano65-275, Ano65-276, Ano65-277, Ano65-361, Ano66-238]. used [CB66]. uses [Ano65-196, Ano65-391, She66d, Vog69]. using [Hem66, Rud67]. utilities [EB67]. Utility [Sco68, CLM^{+}69, DBW69, Fan65, For68]. utilization [Ano65-388]. utilizes [Ano64-149].

V [Cut64a, Ano65-139, Fra64c]. Va. [Ano65-235]. Variable [Ano68-295]. variation [Ano65-461]. Variety [Ano68-476]. varifocal [Raw69]. Varying [Ano67-398]. vehicle [Swi68]. Vehicles [And64, Ano65-366]. vehicular [Fri69b]. venture [Ano64-276]. versatile [Ano64-159]. versatility [Zim69]. via [Ano68-319]. Vibrating [Ano68-176, Raw69]. Video [Ano69-453]. Vienna [Ano65-418]. view [Lie66, Oli64]. viewpoint [McL68]. views [Col66a]. VII [Ano64-52]. VITA [Hof68]. vocoders [EC64]. voices [Ano67-341]. volt [Ano65b, Ano65-247]. Volta [Hen65]. voltage [Gre69, Lam66b, WP69]. voltage-controlled [Gre69]. volume [Ano64-53, Ano64-112, Ano64-296, Ano65-146]. Vote [Ano67-458]. vs [Fri67b].

W. [Ano64-137, Ano64-138, Ano64-250, Ano64-273, Ano65-486]. Walsh [Har69]. Wanted [Lan67]. War [Fri66f, Fri67d, Fri67c]. Was [Ano65-487]. Washington [Ano69-305, Ano69-306, Ano65-279, Ano65-336]. watt [Ano64-125]. Watts [Hen65]. Wave [Ano66-314, Ano67-342, Ano67-370, Cot66, Lig67, Ano64-125, May65, Zve66]. Wave-mechanical [Lig67]. Waveform [CPH67]. waveguide [Hee64]. Waveguides [Gou64]. wavelength [HG65]. wavelengths [AFW68]. waves [Kin65, Mel68, Rod65b]. way [Ano65-131, Ano65-360, Ano68-215]. weather [Ano64-266, Til69b, Til69c, Til69d]. welding [Dag68]. WESCON [Ano64c, Ano65-488, Ano67-495, Ano68-477, Ano69-454, Ano68-293]. WESCON/65 [Ano65-488]. WESCON/69 [Ano69-454]. West [Phi65]. Western [Ano64g]. Wheatstone [Gos68]. Which [Ano66-465]. White [Ano64-355]. Who [Tru64]. wide [Ano65-461]. Widening [LPK65]. wild [Fri66d]. Wiley [Ano69-418]. Wiley-Interscience [Ano69-418]. will [Ano64-143, Ano64-159, Ano64-333, Ano65-222, Ano65-229, Ano65-226, Ano65-247, Ano65-279, Ano65-322, Cre65]. Winter [Ano65f, Ano66b, Ano69-373, Ano69-370, Ano69-372, Ano65-383, Ano64-186]. Winter-Spring [Ano69-372]. wire [Ano64-158, FL66]. Wired [Gab67]. wireless [Ano69-375]. wisdom [Tra69]. without [Ano65-141, Ano65-419, Ros67]. Women [Ped68]. Work [Ano66-292, Noy68]. workhorses [Gut67]. Working [YW66, LS64]. Workshop [Nag67]. world [Ano65-197, Bro67c, Fel64, Fra66, Fri66c, OFM^{+}69, Rod65a, Fri66f, Fri67d, Fri67c]. Worlds [Ano66-316]. Worldwide [McL66, Lam66a, Wei66b, Win66]. Worst [Atk65]. Worst-case [Atk65]. WSE [Ano64-336].

X [Ele68]. Xerox [Ano64-135]. XXXth [Ano65-489].

YAG [Ano64-335]. year [Ano65-323, Ano65-332, Ano66-283, BL65b]. years [Ano64-316, Ano64-315, Ano64-317, Ano64-342, Ano67-369, Ano69-375, Lin66d]. yellow [Fri66f]. yellow-green [Fri66f]. York [Hen65].

zoom [Fri67b].

References

Abramson:1966:RPS

[Abr66] N. Abramson. Report on the A. S. Popov Society meeting. IEEE Spectrum, 3(8):115–117, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adams:1966:IMS

[Ada66] D. K. Adams. International microwave symposium. IEEE Spectrum, 3(9):145–148, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Adler:1967:IBL

[Adl67] R. Adler. Interaction between light and sound. IEEE Spectrum, 4(5):42–54, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Altshuler:1968:AEP

[AFW68] E. E. Altshuler, V. J. Falcone, and K. N. Wulfsberg. Atmospheric effects on propagation at millimeter wavelengths. IEEE Spectrum, 5(7):83–90, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Alger:1969:NRE

[AH69] P. L. Alger and E. H. Holt. The new responsibility of the engineer. IEEE Spectrum, 6(8): 63–65, August 1969. CODEN IEESAM. ISSN 1939-9340.

Alber:1965:EPE

[AI65] J. D. Alber and R. H. Imgram.

Environmental problems in external space equipment design. IEEE Spectrum, 2(7):106–115, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Alexander:1965:TET

[AJMS65] G. W. Alexander, I. B. Johnson, A. J. McConnell, and H. O. Simmons. Trends in EHV transmission equipment. IEEE Spectrum, 2(8):80–85, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aarons:1966:SSR

[AKAG66] J. Aarons, J. A. Klobuchar, R. S.

Allen, and D. A. Guidice. Second symposium on radio astronomical and satellite studies of the atmosphere. IEEE Spectrum, 3(3):174–180, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Alger:1964:ASM

[Alg64] P. L. Alger. Asynchronized synchronous machines. IEEE Spectrum, 1(8):154–156, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allen:1964:AAN

[All64] J. L. Allen. Array antennas:

New applications for an old technique. IEEE Spectrum, 1(11): 115–130, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Amarel:1966:MCP

[Ama66] S. Amarel. On the mechanization of creative processes. IEEE Spectrum, 3(4):112–114, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:1964:VSI

[And64]  V. C. Anderson.  Vehicles

and stations for installation and maintenance of sea-floor equipment. IEEE Spectrum, 1(11): 104–108, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:1969:RCC

[And69] C. A. Anderson. Research and the changing campus environment. IEEE Spectrum, 6(12): 61–63, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1964:AEM

[Ano64a] Anonymous. 12th annual engineering management conference held. IEEE Spectrum, 1(11): 26–27, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ICG

[Ano64b] Anonymous. 1964 IEEE committees, groups, and sections. IEEE Spectrum, 1(10):21–29, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:WEL

[Ano64c] Anonymous. 1964 WESCON exhibitors’ list. IEEE Spectrum, 1 (8):164–181, August 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FSI

[Ano64d] Anonymous. 34 fields of special interest — IEEE’s: The 34 groups are listed below, together with a brief definition of each. IEEE Spectrum, 1 (11):44, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IGAb

[Ano64e] Anonymous. 34 IEEE groups and available publications to date. IEEE Spectrum, 1(11): 45–46, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PPI

[Ano64f] Anonymous. 363 persons participate in IEEE Petroleum Conference. IEEE Spectrum, 1 (11):52, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:WAC

[Ano64g] Anonymous. 6th Western Appliance Conference to take place in Los Angeles on October 27. IEEE Spectrum, 1(10):34, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:UCA

[Ano64h] Anonymous. 8th upadi convention is attended by assembly of 400 engineers from 16 countries. IEEE Spectrum, 1(11): 34–36, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ATL

[Ano64i] Anonymous. Adaptive techniques for large reflector antennas [advertisement]. IEEE Spectrum, 1(3):5, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AAa

[Ano64j]  Anonymous.  Advance ab-

stracts. IEEE Spectrum, 1(1): 143–149, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AAb

[Ano64k] Anonymous. Advance abstracts. IEEE Spectrum, 1(6):157–165, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AAc

[Ano64l] Anonymous. Advance abstracts. IEEE Spectrum, 1(7):144–153, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AAd

[Ano64m] Anonymous. Advance abstracts. IEEE Spectrum, 1(8):116–121, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AAe

[Ano64n] Anonymous. Advance abstracts. IEEE Spectrum, 1(9): 148–153, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AAf

[Ano64o] Anonymous. Advance abstracts. IEEE Spectrum, 1(10): 125–133, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AAg

[Ano64p] Anonymous. Advance abstracts. IEEE Spectrum, 1(11): 142–151, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AAh

[Ano64q] Anonymous. Advance abstracts. IEEE Spectrum, 1(12): 103–113, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Aa

[Ano64r] Anonymous. [advertisement].

IEEE Spectrum, 1(1):1, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ab

[Ano64s] Anonymous. [advertisement].

IEEE Spectrum, 1(1):3–4, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ac

[Ano64t] Anonymous. [advertisement]. IEEE Spectrum, 1(1):9, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ad

[Ano64u] Anonymous. [advertisement]. IEEE Spectrum, 1(1):13–43, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Af

[Ano64v] Anonymous. [advertisement]. IEEE Spectrum, 1(1):44–51, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ag

[Ano64w] Anonymous. [advertisement]. IEEE Spectrum, 1(1):52–62, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ah

[Ano64x] Anonymous. [advertisement]. IEEE Spectrum, 1(1):63–69, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ai

[Ano64y] Anonymous. [advertisement]. IEEE Spectrum, 1(1):71–84, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Aj

[Ano64z] Anonymous. [advertisement].

IEEE Spectrum, 1(1):116–117, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Al

[Ano64-27] Anonymous. [advertisement].

IEEE Spectrum, 1(1):155, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Am

[Ano64-28] Anonymous. [advertisement].

IEEE Spectrum, 1(1):156–161, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:An

[Ano64-29] Anonymous. [advertisement].

IEEE Spectrum, 1(1):162–164, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ao

[Ano64-30] Anonymous. [advertisement].

IEEE Spectrum, 1(1):166–171, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ap

[Ano64-31] Anonymous. [advertisement].

IEEE Spectrum, 1(1):173, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Aq

[Ano64-32] Anonymous. [advertisement]. IEEE Spectrum, 1(1):178–192, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ar

[Ano64-33] Anonymous. [advertisement]. IEEE Spectrum, 1(1):196–217, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:As

[Ano64-34] Anonymous. [advertisement]. IEEE Spectrum, 1(1):218–226, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:At

[Ano64-35] Anonymous. [advertisement]. IEEE Spectrum, 1(1):230, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Au

[Ano64-36] Anonymous. [advertisement]. IEEE Spectrum, 1(1):4, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Aw

[Ano64-37] Anonymous. [advertisements]. IEEE Spectrum, 1(2):178, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ax

[Ano64-38] Anonymous. [advertisements].

IEEE Spectrum, 1(3):3–4, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Az

[Ano64-39] Anonymous. [advertisements].

IEEE Spectrum, 1(3):305–308, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abb

[Ano64-40] Anonymous. [advertisements].

IEEE Spectrum, 1(4):176, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abd

[Ano64-41] Anonymous. [advertisements].

IEEE Spectrum, 1(5):184, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abf

[Ano64-42] Anonymous. [advertisements].

IEEE Spectrum, 1(6):212, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abh

[Ano64-43] Anonymous. [advertisements].

IEEE Spectrum, 1(7):192, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abj

[Ano64-44] Anonymous. [advertisements]. IEEE Spectrum, 1(8):184, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abl

[Ano64-45] Anonymous. [advertisements]. IEEE Spectrum, 1(9):204, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abn

[Ano64-46] Anonymous. [advertisements]. IEEE Spectrum, 1(10):180, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abo

[Ano64-47] Anonymous. [advertisements]. IEEE Spectrum, 1(11):1–2, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abq

[Ano64-48] Anonymous. [advertisements]. IEEE Spectrum, 1(11):190, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abr

[Ano64-49] Anonymous. [advertisements]. IEEE Spectrum, 1(12):2–6, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abt

[Ano64-50] Anonymous. [advertisements].

IEEE Spectrum, 1(12):160, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AACa

[Ano64-51] Anonymous. Airlines adopt computers to expedite flight reservations. IEEE Spectrum, 1 (11):174, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ACC

[Ano64-52] Anonymous. Annual Communications Convention, including Globecom VII, issues call for papers. IEEE Spectrum, 1(10): 37–38, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AII

[Ano64-53] Anonymous. Annual index to IEEE Spectrum for volume 1, 1964. IEEE Spectrum, 1(12): 121, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ASB

[Ano64-54] Anonymous. Another star is born: Joins communications orbit. IEEE Spectrum, 1(10): 146–148, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEP

[Ano64-55] Anonymous. Approaches to education in power engineering studied. IEEE Spectrum, 1 (10):41, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEa

[Ano64-56] Anonymous. The art of engineering. IEEE Spectrum, 1(1):8, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEb

[Ano64-57] Anonymous. The art of engineering. IEEE Spectrum, 1 (2):6–9, February 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEc

[Ano64-58] Anonymous. The art of engineering. IEEE Spectrum, 1(3):6–21, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEd

[Ano64-59] Anonymous. The art of engineering. IEEE Spectrum, 1(4):6–11, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEe

[Ano64-60] Anonymous. The art of engineering. IEEE Spectrum, 1(5):7–11, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEf

[Ano64-61] Anonymous. The art of engineering. IEEE Spectrum, 1(6):7, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEg

[Ano64-62] Anonymous. The art of engineering. IEEE Spectrum, 1(7): 7, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEh

[Ano64-63] Anonymous. The art of engineering. IEEE Spectrum, 1(8):7–9, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEi

[Ano64-64] Anonymous. The art of engineering. IEEE Spectrum, 1 (9):7–11, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEj

[Ano64-65] Anonymous. The art of engineering. IEEE Spectrum, 1 (10):7–9, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEk

[Ano64-66] Anonymous. The art of engineering. IEEE Spectrum, 1 (11):9–13, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AEl

[Ano64-67] Anonymous. The art of engineering. IEEE Spectrum, 1(12):7, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:AIa

[Ano64-68] Anonymous. Author index. IEEE Spectrum, 1(12):124–126, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ak

[Ano64-69] Anonymous. Authors. IEEE Spectrum, 1(1):139–140, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Av

[Ano64-70] Anonymous. Authors. IEEE Spectrum, 1(2):114–115, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ay

[Ano64-71] Anonymous. Authors. IEEE Spectrum, 1(3):238–239, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Aba

[Ano64-72] Anonymous. Authors. IEEE Spectrum, 1(4):117–119, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abc

[Ano64-73] Anonymous. Authors. IEEE Spectrum, 1(5):123–124, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abe

[Ano64-74] Anonymous. Authors. IEEE Spectrum, 1(6):151–154, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abg

[Ano64-75] Anonymous. Authors. IEEE Spectrum, 1(7):140–141, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abi

[Ano64-76] Anonymous. Authors. IEEE Spectrum, 1(8):112–113, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abk

[Ano64-77] Anonymous. Authors. IEEE Spectrum, 1(9):145, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abm

[Ano64-78] Anonymous. Authors. IEEE Spectrum, 1(10):121–122, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abp

[Ano64-79] Anonymous. Authors. IEEE Spectrum, 1(11):138–139, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Abs

[Ano64-80] Anonymous. Authors. IEEE Spectrum, 1(12):100, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCa

[Ano64-81] Anonymous. [back cover]. IEEE Spectrum, 1(2):c4, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCb

[Ano64-82] Anonymous. [back cover]. IEEE Spectrum, 1(3):c4, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCc

[Ano64-83] Anonymous. [back cover]. IEEE Spectrum, 1(4):c4, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCd

[Ano64-84] Anonymous. [back cover]. IEEE Spectrum, 1(5):c4, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCe

[Ano64-85] Anonymous. [back cover]. IEEE Spectrum, 1(6):c4, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCf

[Ano64-86] Anonymous. [back cover]. IEEE Spectrum, 1(7):c4, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCg

[Ano64-87] Anonymous. [back cover]. IEEE Spectrum, 1(8):c4, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCh

[Ano64-88] Anonymous. [back cover]. IEEE Spectrum, 1(9):c4, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCi

[Ano64-89] Anonymous. [back cover]. IEEE Spectrum, 1(10):c4, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCj

[Ano64-90] Anonymous. [back cover]. IEEE Spectrum, 1(11):c4, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BCk

[Ano64-91] Anonymous. [back cover]. IEEE Spectrum, 1(12):c4, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BRa

[Ano64-92] Anonymous. Book reviews. IEEE Spectrum, 1(1):164–168, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BRb

[Ano64-93] Anonymous. Book reviews. IEEE Spectrum, 1(9):190–195, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BRc

[Ano64-94] Anonymous. Book reviews. IEEE Spectrum, 1(10):170– 175, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ISR

[Ano64-95] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 1(1):2–3, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:BP

[Ano64-96] Anonymous. [breaker page]. IEEE Spectrum, 1(1):2, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ca

[Ano64-97] Anonymous. Calendar. IEEE Spectrum, 1(1):44–50, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cd

[Ano64-98] Anonymous. Calendar. IEEE Spectrum, 1(2):20–31, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ce

[Ano64-99] Anonymous. Calendar. IEEE Spectrum, 1(2):32–40, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cg

[Ano64-100] Anonymous. Calendar. IEEE Spectrum, 1(3):158–167, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cj

[Ano64-101] Anonymous. Calendar. IEEE Spectrum, 1(4):42–45, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cl

[Ano64-102] Anonymous. Calendar. IEEE Spectrum, 1(5):54–61, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Co

[Ano64-103] Anonymous. Calendar. IEEE Spectrum, 1(6):36–45, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cq

[Ano64-104] Anonymous. Calendar. IEEE Spectrum, 1(7):28–31, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cs

[Ano64-105] Anonymous. Calendar. IEEE Spectrum, 1(8):44–49, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cu

[Ano64-106] Anonymous. Calendar. IEEE Spectrum, 1(9):46–52, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cw

[Ano64-107] Anonymous. Calendar. IEEE Spectrum, 1(10):42–47, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cz

[Ano64-108] Anonymous. Calendar. IEEE Spectrum, 1(11):54–59, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cbb

[Ano64-109] Anonymous. Calendar. IEEE Spectrum, 1(12):30–37, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:CSEb

[Ano64-110] Anonymous. Conference is scheduled on electronics, instrumentation. IEEE Spectrum, 1 (11):38–39, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:CEM

[Ano64-111] Anonymous. Conference on engineering in medicine and biology slated. IEEE Spectrum, 1(10):39, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:CUI

[Ano64-112] Anonymous. Conference on urd issues supplement to advance volume. IEEE Spectrum, 1 (10):40, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:CSEa

[Ano64-113] Anonymous. Conference to study electromagnetic compatibility. IEEE Spectrum, 1 (10):39, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cc

[Ano64-114] Anonymous. Contents. IEEE Spectrum, 1(2):1, February 1964.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cf

[Ano64-115] Anonymous. Contents. IEEE Spectrum, 1(3):1, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ci

[Ano64-116] Anonymous. Contents. IEEE Spectrum, 1(4):1, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ck

[Ano64-117] Anonymous. Contents. IEEE Spectrum, 1(5):1, May 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cn

[Ano64-118] Anonymous. Contents. IEEE Spectrum, 1(6):1, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cp

[Ano64-119] Anonymous. Contents. IEEE Spectrum, 1(7):1, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cr

[Ano64-120] Anonymous. Contents. IEEE Spectrum, 1(8):1, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ct

[Ano64-121] Anonymous. Contents. IEEE Spectrum, 1(9):1, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cv

[Ano64-122] Anonymous. Contents. IEEE Spectrum, 1(10):1, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cy

[Ano64-123] Anonymous. Contents. IEEE Spectrum, 1(11):3, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cba

[Ano64-124] Anonymous. Contents. IEEE Spectrum, 1(12):1, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:CWG

[Ano64-125] Anonymous. Continuous wave gas laser output increased to one watt. IEEE Spectrum, 1 (10):158, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:CPS

[Ano64-126] Anonymous. Contributed papers sought on techniques of optimization. IEEE Spectrum, 1 (11):51, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cb

[Ano64-127] Anonymous. Correspondence. IEEE Spectrum, 1(1):218–226, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ch

[Ano64-128] Anonymous. Correspondence. IEEE Spectrum, 1(3):304, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cm

[Ano64-129] Anonymous. Correspondence. IEEE Spectrum, 1(5):172–181, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Cx

[Ano64-130] Anonymous. Correspondence. IEEE Spectrum, 1(10):176– 177, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:DWR

[Ano64-131] Anonymous. The day we rocked the boat. IEEE Spectrum, 1 (10):36–37, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:DNA

[Ano64-132] Anonymous. Delegates from 20 nations attend microwave conference. IEEE Spectrum, 1 (11):168, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:DF

[Ano64-133] Anonymous. Designer’s file.

IEEE Spectrum, 1(1):6–7, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:DAI

[Ano64-134] Anonymous. Development of artificial implantable heart is planned. IEEE Spectrum, 1(10): 148–152, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:DYK

[Ano64-135] Anonymous. Do you know how to read a Xerox job spec? [advertisement]. IEEE Spectrum, 1 (10):165–166, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:DHR

[Ano64-136] Anonymous. Dr. Hutchisson retires from directorship of AIP. IEEE Spectrum, 1(11): 168, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:DPW

[Ano64-137] Anonymous. Du Pont’s W.

A. Crawford is elected president of ISA. IEEE Spectrum, 1 (11):172, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ERW

[Ano64-138] Anonymous. ECPD reelects W. S. Hill as President at Annual Meeting. IEEE Spectrum, 1 (11):43, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:EI

[Ano64-139] Anonymous. EDP at the IEEE. IEEE Spectrum, 1(12): 93–95, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ECS

[Ano64-140] Anonymous. Electronic control system speeds air passenger check-in. IEEE Spectrum, 1 (10):160–162, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:EPM

[Ano64-141] Anonymous. Electronic pinball machine tests interpersonal behavior. IEEE Spectrum, 1(11): 172–174, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:EPN

[Ano64-142] Anonymous. Electronic products news. IEEE Spectrum, 1 (1):165, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ESWb

[Ano64-143] Anonymous. Electronics sales for 1964 will set record predicts EIA. IEEE Spectrum, 1(11):

168–170, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ESPb

[Ano64-144] Anonymous. Engineering Societies Personnel Service, Inc. IEEE Spectrum, 1(10):168– 169, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ESPc

[Ano64-145] Anonymous. Engineering Societies Personnel Service, Inc. IEEE Spectrum, 1(11):178–179, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ESPa

[Ano64-146] Anonymous. Engineering Societies Personnel Service, Inc. (agency) [advertisement]. IEEE Spectrum, 1(1):194–195, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ESWa

[Ano64-147] Anonymous. Engineers scientists: What’s new at pan am? [advertisement]. IEEE Spectrum, 1(10):167, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:EA

[Ano64-148] Anonymous. Exhibitors [advertisement]. IEEE Spectrum, 1 (3):309–399, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ESU

[Ano64-149] Anonymous. Experimental system utilizes computers to instruct. IEEE Spectrum, 1(11): 174, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPAa

[Ano64-150] Anonymous. Family of plastics achieves conductor status. IEEE Spectrum, 1(10):152, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPa

[Ano64-151] Anonymous. Focal points. IEEE Spectrum, 1(1):174–216, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPb

[Ano64-152] Anonymous. Focal points. IEEE Spectrum, 1(3):260–299, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPc

[Ano64-153] Anonymous. Focal points. IEEE Spectrum, 1(4):142–161, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPd

[Ano64-154] Anonymous. Focal points. IEEE Spectrum, 1(5):140–167, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPe

[Ano64-155] Anonymous. Focal points. IEEE Spectrum, 1(6):180–203, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPf

[Ano64-156] Anonymous. Focal points. IEEE Spectrum, 1(7):166–183, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPg

[Ano64-157] Anonymous. Focal points. IEEE Spectrum, 1(12):138–153, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPEa

[Ano64-158] Anonymous. Focal points: Effects of nuclear explosions can be simulated by use of an exploding wire system. IEEE Spectrum, 1(10):144, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPFa

[Ano64-159] Anonymous. Focal points: First aluminum, deepest diving submarine will function as versatile deep-sea laboratory. IEEE Spectrum, 1(11):164, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPI

[Ano64-160] Anonymous. Focal points: Industry education government. IEEE Spectrum, 1(2):142–167, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FPT

[Ano64-161] Anonymous. Focal points: Television performance of Ranger 7 compensates for previous disappointments. IEEE Spectrum, 1(9):166–189, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FFA

[Ano64-162] Anonymous. Fortescue fellowships for 1965–1966 are available. IEEE Spectrum, 1(10):39, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCa

[Ano64-163] Anonymous. [front cover]. IEEE Spectrum, 1(1):c1, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCb

[Ano64-164] Anonymous. [front cover]. IEEE Spectrum, 1(1):c2, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCc

[Ano64-165] Anonymous. [front cover].

IEEE Spectrum, 1(2):c1, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCd

[Ano64-166] Anonymous. [front cover]. IEEE Spectrum, 1(3):c1, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCe

[Ano64-167] Anonymous. [front cover]. IEEE Spectrum, 1(4):c1, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCf

[Ano64-168] Anonymous. [front cover]. IEEE Spectrum, 1(5):c1, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCg

[Ano64-169] Anonymous. [front cover]. IEEE Spectrum, 1(6):c1, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCh

[Ano64-170] Anonymous. [front cover]. IEEE Spectrum, 1(7):c1, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCi

[Ano64-171] Anonymous. [front cover]. IEEE Spectrum, 1(8):c1, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCj

[Ano64-172] Anonymous. [front cover]. IEEE Spectrum, 1(9):c1, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCk

[Ano64-173] Anonymous. [front cover]. IEEE Spectrum, 1(10):c1, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCl

[Ano64-174] Anonymous. [front cover]. IEEE Spectrum, 1(11):c1, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:FCm

[Ano64-175] Anonymous. [front cover]. IEEE Spectrum, 1(12):c1, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:GSU

[Ano64-176] Anonymous. Group on sonics and ultrasonics to hold annual symposium October 14– 16. IEEE Spectrum, 1(10): 19–20, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:HGA

[Ano64-177] Anonymous. H. G. Asmus appointed executive director of AFIPS. IEEE Spectrum, 1(11): 52–53, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:HUR

[Ano64-178] Anonymous. How to use reader service card! IEEE Spectrum, 1(1):229, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IACa

[Ano64-179] Anonymous. IEEE–ASQC Conference on Reliability Training is slated. IEEE Spectrum, 1(11): 51–52, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IASa

[Ano64-180] Anonymous. IEEE and American Standards. IEEE Spectrum, 1(1):150–154, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IASb

[Ano64-181] Anonymous. IEEE announces sale of former IRE headquarters on 79th Street. IEEE Spectrum, 1(10):34–35, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCi

[Ano64-182] Anonymous. IEEE branch counselors. IEEE Spectrum, 1(10): 29–32, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IEC

[Ano64-183] Anonymous. IEEE Executive Committee approves four group chapters. IEEE Spectrum, 1 (11):52, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IMG

[Ano64-184] Anonymous. IEEE magnetics group to begin operations January 1. IEEE Spectrum, 1(12): 19–29, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPR

[Ano64-185] Anonymous. IEEE policy regarding professional relations. IEEE Spectrum, 1(4):113–115, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPG

[Ano64-186] Anonymous. IEEE power group plans Winter power meeting. IEEE Spectrum, 1(10):38, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPa

[Ano64-187] Anonymous. IEEE publications. IEEE Spectrum, 1(3):240–259, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPb

[Ano64-188] Anonymous. IEEE publications.

IEEE Spectrum, 1(4):120–141, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPc

[Ano64-189] Anonymous. IEEE publications.

IEEE Spectrum, 1(5):125–139, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPd

[Ano64-190] Anonymous. IEEE publications.

IEEE Spectrum, 1(6):155–156, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPe

[Ano64-191] Anonymous. IEEE publications.

IEEE Spectrum, 1(7):142–143, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPSb

[Ano64-192] Anonymous. IEEE publications: Scanning the issues. IEEE Spectrum, 1(8):114–115, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPSc

[Ano64-193] Anonymous. IEEE publications:

Scanning the issues. IEEE Spectrum, 1(9):146–147, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPSd

[Ano64-194] Anonymous. IEEE publications: Scanning the issues. IEEE Spectrum, 1(10):123–124, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPSe

[Ano64-195] Anonymous. IEEE publications: Scanning the issues. IEEE Spectrum, 1(11):140–141, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPSf

[Ano64-196] Anonymous. IEEE publications: Scanning the issues. IEEE Spectrum, 1(12):101–102, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IPSa

[Ano64-197] Anonymous. IEEE publications: Scanning the issues advance abstracts translated journals special publications. IEEE Spectrum, 1(2):116–141, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ICE

[Ano64-198] Anonymous. If compression and expansion receiver. IEEE Spectrum, 1(11):7–8, November 1964.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAa

[Ano64-199] Anonymous. Index to advertisers. IEEE Spectrum, 1(1): 228–229, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAb

[Ano64-200] Anonymous. Index to advertisers. IEEE Spectrum, 1(2): 176–177, February 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAc

[Ano64-201] Anonymous. Index to advertisers. IEEE Spectrum, 1(3): 400–404, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAd

[Ano64-202] Anonymous. Index to advertisers. IEEE Spectrum, 1(4): 174–175, April 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAe

[Ano64-203] Anonymous. Index to advertisers. IEEE Spectrum, 1 (5):182–183, May 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAf

[Ano64-204] Anonymous. Index to advertisers. IEEE Spectrum, 1 (6):210–211, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAg

[Ano64-205] Anonymous. Index to advertisers. IEEE Spectrum, 1 (7):190–191, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAh

[Ano64-206] Anonymous. Index to advertisers. IEEE Spectrum, 1(8): 182–183, August 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAi

[Ano64-207] Anonymous. Index to advertisers. IEEE Spectrum, 1(9): 202–203, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAj

[Ano64-208] Anonymous. Index to advertisers. IEEE Spectrum, 1(10): 178–179, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAk

[Ano64-209] Anonymous. Index to advertisers. IEEE Spectrum, 1(11): 188–189, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IAl

[Ano64-210] Anonymous. Index to advertisers. IEEE Spectrum, 1(12):

158–159, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IGAa

[Ano64-211] Anonymous. Industry and general applications group starts January 1; charter membership is open. IEEE Spectrum, 1(11): 28–29, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCa

[Ano64-212] Anonymous. [inside back cover].

IEEE Spectrum, 1(2):c3, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCb

[Ano64-213] Anonymous. [inside back cover].

IEEE Spectrum, 1(3):c3, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCc

[Ano64-214] Anonymous. [inside back cover].

IEEE Spectrum, 1(4):c3, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCd

[Ano64-215] Anonymous. [inside back cover].

IEEE Spectrum, 1(5):c3, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCe

[Ano64-216] Anonymous. [inside back cover]. IEEE Spectrum, 1(6):c3, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCf

[Ano64-217] Anonymous. [inside back cover]. IEEE Spectrum, 1(7):c3, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCg

[Ano64-218] Anonymous. [inside back cover]. IEEE Spectrum, 1(8):c3, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCh

[Ano64-219] Anonymous. [inside back cover]. IEEE Spectrum, 1(9):c3, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCj

[Ano64-220] Anonymous. [inside back cover]. IEEE Spectrum, 1(10):c3, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCk

[Ano64-221] Anonymous. [inside back cover]. IEEE Spectrum, 1 (11):c3, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IBCl

[Ano64-222] Anonymous. [inside back cover].

IEEE Spectrum, 1(12):c3, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCa

[Ano64-223] Anonymous. [inside front cover].

IEEE Spectrum, 1(2):c2, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCb

[Ano64-224] Anonymous. [inside front cover].

IEEE Spectrum, 1(3):c2, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCc

[Ano64-225] Anonymous. [inside front cover].

IEEE Spectrum, 1(4):c2, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCd

[Ano64-226] Anonymous. [inside front cover].

IEEE Spectrum, 1(5):c2, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCe

[Ano64-227] Anonymous. [inside front cover].

IEEE Spectrum, 1(6):c2, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCf

[Ano64-228] Anonymous. [inside front cover]. IEEE Spectrum, 1(7):c2, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCg

[Ano64-229] Anonymous. [inside front cover]. IEEE Spectrum, 1(8):c2, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCh

[Ano64-230] Anonymous. [inside front cover]. IEEE Spectrum, 1(9):c2, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCi

[Ano64-231] Anonymous. [inside front cover]. IEEE Spectrum, 1(10):c2, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCj

[Ano64-232] Anonymous. [inside front cover]. IEEE Spectrum, 1 (11):c2, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IFCk

[Ano64-233] Anonymous. [inside front cover]. IEEE Spectrum, 1(12):c2, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ICI

[Ano64-234] Anonymous. International convention: IEEE technical program. IEEE Spectrum, 1(3): 37–117, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSa

[Ano64-235] Anonymous. [list of staff]. IEEE Spectrum, 1(1):2, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSb

[Ano64-236] Anonymous. List of staff.

IEEE Spectrum, 1(2):1–5, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSc

[Ano64-237] Anonymous. List of staff. IEEE Spectrum, 1(3):2, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSd

[Ano64-238] Anonymous. List of staff. IEEE Spectrum, 1(4):2–5, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSe

[Ano64-239] Anonymous. List of staff. IEEE Spectrum, 1(5):2–6, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSf

[Ano64-240] Anonymous. List of staff. IEEE Spectrum, 1(6):2–4, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSg

[Ano64-241] Anonymous. List of staff. IEEE Spectrum, 1(7):2–4, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSh

[Ano64-242] Anonymous. List of staff. IEEE Spectrum, 1(8):2–4, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSi

[Ano64-243] Anonymous. List of staff. IEEE Spectrum, 1(9):2–4, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSj

[Ano64-244] Anonymous. List of staff. IEEE Spectrum, 1(10):2–4, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:LSk

[Ano64-245] Anonymous. List of staff. IEEE Spectrum, 1(11):4–6, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:MTI

[Ano64-246] Anonymous. Machine tools industry conference to be held in Hartford November 16–18. IEEE Spectrum, 1(10):32–33, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:MSN

[Ano64-247] Anonymous. Maecon is scheduled for November in Kansas City. IEEE Spectrum, 1(11): 40–43, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:MRA

[Ano64-248] Anonymous. A maser radiometer for astronomical measurements. IEEE Spectrum, 1(1):5, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:MFA

[Ano64-249] Anonymous. Matrix functions and applications. IEEE Spectrum, 1(3):208, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:MAE

[Ano64-250] Anonymous. Memorial award established in honor of late W. D. George. IEEE Spectrum, 1 (10):40, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:MSR

[Ano64-251] Anonymous. Michigan State reduces requirements for B.S. degree. IEEE Spectrum, 1(11): 177, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:MSS

[Ano64-252] Anonymous. Microelectronics symposium scheduled for November 17–18. IEEE Spectrum, 1 (10):162–164, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NSD

[Ano64-253] Anonymous. New semiconductor device achieves high amplification. IEEE Spectrum, 1(11): 174–177, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIi

[Ano64-254] Anonymous. News index. IEEE Spectrum, 1(12):126–131, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIa

[Ano64-255] Anonymous. News of the IEEE. IEEE Spectrum, 1(1):10–43, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIb

[Ano64-256] Anonymous. News of the IEEE. IEEE Spectrum, 1(3):118–157, March 1964. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIc

[Ano64-257] Anonymous. News of the IEEE.

IEEE Spectrum, 1(4):20–41, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NId

[Ano64-258] Anonymous. News of the IEEE.

IEEE Spectrum, 1(6):18–35, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIe

[Ano64-259] Anonymous. News of the IEEE.

IEEE Spectrum, 1(7):16–27, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIf

[Ano64-260] Anonymous. News of the IEEE.

IEEE Spectrum, 1(8):20–43, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIg

[Ano64-261] Anonymous. News of the IEEE. IEEE Spectrum, 1(10): 18–19, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIh

[Ano64-262] Anonymous. News of the IEEE. IEEE Spectrum, 1(11):

24–26, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NICa

[Ano64-263] Anonymous. News of the IEEE: Calendar people obituaries. IEEE Spectrum, 1(5):28–53, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NINa

[Ano64-264] Anonymous. News of the IEEE: National Electronics Conference for 1964 plans comprehensive program. IEEE Spectrum, 1(9): 22–45, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NINb

[Ano64-265] Anonymous. News of the IEEE: National Power Conference attended by 600 electrical and mechanical engineers. IEEE Spectrum, 1(12):19, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NWP

[Ano64-266] Anonymous. Nimbus weather pictures now obtainable at low cost. IEEE Spectrum, 1(10): 164, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NIAa

[Ano64-267] Anonymous. Nominations for IEEE awards, medals, and prizes in 1965–1966 should be received by April 1 deadline. IEEE Spectrum, 1(11):30–33, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NME

[Ano64-268] Anonymous. Now from Motorola Epitaxial, passivated PNP silicon transistors made by the annular process [advertisement]. IEEE Spectrum, 1(1):172, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:NRA

[Ano64-269] Anonymous. Nuclear reactor at Argonne now in operation. IEEE Spectrum, 1(11): 166–168, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Oa

[Ano64-270] Anonymous. Obituaries. IEEE Spectrum, 1(1):62–68, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ob

[Ano64-271] Anonymous. Obituaries. IEEE Spectrum, 1(12):43–48, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ODN

[Ano64-272] Anonymous. Obituaries and death notices. IEEE Spectrum, 1(12):134–137, December 1964.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:OWF

[Ano64-273] Anonymous. Obituaries: W. F. Uhl, retired head of Chas. T. Main, Inc., dies. IEEE Spectrum, 1(4):52, 54–, April 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:OPD

[Ano64-274] Anonymous. Original papers dealing with impact of batchfabrication on future computers invited. IEEE Spectrum, 1 (11):33, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PDP

[Ano64-275] Anonymous. Panel discussion planned for IEEE Reliability Symposium. IEEE Spectrum, 1 (11):47–48, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PCFb

[Ano64-276] Anonymous. Panelboard for compact fuses is developed in a joint venture. IEEE Spectrum, 1(11):170–172, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PSE

[Ano64-277] Anonymous. Papers are solicited for electric heating conference. IEEE Spectrum, 1 (10):41–40, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PST

[Ano64-278] Anonymous. Papers are solicited for telemetering conference. IEEE Spectrum, 1(11): 39–40, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PRR

[Ano64-279] Anonymous. Papers on radar reflectivity sought for IEEE Proceedings. IEEE Spectrum, 1 (10):33–34, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pa

[Ano64-280] Anonymous. People. IEEE Spectrum, 1(1):52–62, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pb

[Ano64-281] Anonymous. People. IEEE Spectrum, 1(3):168–180, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pc

[Ano64-282] Anonymous. People. IEEE Spectrum, 1(4):46–51, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pd

[Ano64-283] Anonymous. People. IEEE Spectrum, 1(5):62–72, May 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pe

[Ano64-284] Anonymous. People. IEEE Spectrum, 1(6):46–60, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pf

[Ano64-285] Anonymous. People. IEEE Spectrum, 1(7):32–64, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pg

[Ano64-286] Anonymous. People. IEEE Spectrum, 1(8):50–60, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ph

[Ano64-287] Anonymous. People. IEEE Spectrum, 1(9):53–64, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pi

[Ano64-288] Anonymous. People. IEEE Spectrum, 1(10):48–56, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pj

[Ano64-289] Anonymous. People. IEEE Spectrum, 1(11):60–66, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pk

[Ano64-290] Anonymous. People. IEEE Spectrum, 1(12):38–43, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Pl

[Ano64-291] Anonymous. People. IEEE Spectrum, 1(12):131–133, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PCFa

[Ano64-292] Anonymous. People: C. F. Horne selected to receive EIA Medal of Honor. IEEE Spectrum, 1 (2):41–56, February 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PCSb

[Ano64-293] Anonymous. PICA conference seeks papers for May 19–21 meeting. IEEE Spectrum, 1 (10):38, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PJP

[Ano64-294] Anonymous. Plasma jets proposed as satellite steering systems. IEEE Spectrum, 1(11):

166, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PBE

[Ano64-295] Anonymous. Post-baccalaureate education in industry: A report of the IEEE Committee on Education. IEEE Spectrum, 1 (6):120–122, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PVP

[Ano64-296] Anonymous. Preprint volume of papers for 1964 JACC is now available. IEEE Spectrum, 1 (11):47, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PAB

[Ano64-297] Anonymous. Proceedings available on biomedical engineering. IEEE Spectrum, 1(11): 52, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PCSa

[Ano64-298] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 1(1):227, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PDK

[Ano64-299] Anonymous. Professor Dow keynotes National Automotive Conference. IEEE Spectrum, 1 (11):46–47, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PBC

[Ano64-300] Anonymous. Programming the bureaucratic computer. IEEE Spectrum, 1(12):72–92, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PRQ

[Ano64-301] Anonymous. Progress report on quasioptical components. IEEE Spectrum, 1(7):5–6, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PNG

[Ano64-302] Anonymous. Propagation in nonuniform gyrotropic media. IEEE Spectrum, 1(10):5–6, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:PBO

[Ano64-303] Anonymous. Push-button operation takes over manual tasks as recently launched U.S. ship joins Merchant Marine. IEEE Spectrum, 1(10):144–146, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:RCM

[Ano64-304] Anonymous. Railroad conference to meet April 7–8 in Pittsburgh. IEEE Spectrum, 1 (11):52, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:RLE

[Ano64-305] Anonymous. The Raman laser: An electromechanical parametric oscillator. IEEE Spectrum, 1(8):5–6, August 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Rb

[Ano64-306] Anonymous. Reflection. IEEE Spectrum, 1(3):22–36, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ra

[Ano64-307] Anonymous. Reflections. IEEE Spectrum, 1(1):70–82, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Rc

[Ano64-308] Anonymous. Reflections. IEEE Spectrum, 1(6):8–17, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Rd

[Ano64-309] Anonymous. Reflections. IEEE Spectrum, 1(7):8–15, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Re

[Ano64-310] Anonymous. Reflections. IEEE Spectrum, 1(8):10–19, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Rf

[Ano64-311] Anonymous. Reflections. IEEE Spectrum, 1(9):12–21, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Rg

[Ano64-312] Anonymous. Reflections. IEEE Spectrum, 1(10):10–17, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Rh

[Ano64-313] Anonymous. Reflections. IEEE Spectrum, 1(11):14–23, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:Ri

[Ano64-314] Anonymous. Reflections. IEEE Spectrum, 1(12):8–18, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:RYAb

[Ano64-315] Anonymous. Reflections: 75 years age. IEEE Spectrum, 1 (4):12–19, April 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:RYAa

[Ano64-316] Anonymous. Reflections: 75 years ago. IEEE Spectrum, 1 (2):10–19, February 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:RYAc

[Ano64-317] Anonymous. Reflections: 75 years ago. IEEE Spectrum, 1(5):12–27, May 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:RMPa

[Ano64-318] Anonymous. Reviewed this month in proceedings of the IEEE. IEEE Spectrum, 1(6): 204–205, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:RMPb

[Ano64-319] Anonymous. Reviewed this month in proceedings of the IEEE. IEEE Spectrum, 1(8): 156–157, August 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SIa

[Ano64-320] Anonymous. Scanning the issues. IEEE Spectrum, 1(1): 141–142, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SSE

[Ano64-321] Anonymous. Seismic sounding of the Earth’s interior. IEEE Spectrum, 1(6):5–6, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPP

[Ano64-322] Anonymous. Soviet papers presented at IFAC meeting are available. IEEE Spectrum, 1 (11):51, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPa

[Ano64-323] Anonymous. Special publications. IEEE Spectrum, 1 (1):162, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPb

[Ano64-324] Anonymous. Special publications. IEEE Spectrum, 1(6): 176–179, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPc

[Ano64-325] Anonymous. Special publications. IEEE Spectrum, 1 (7):162–165, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPd

[Ano64-326] Anonymous. Special publications. IEEE Spectrum, 1(8): 128–131, August 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPe

[Ano64-327] Anonymous. Special publications. IEEE Spectrum, 1(9): 160–165, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPf

[Ano64-328] Anonymous. Special publications. IEEE Spectrum, 1(10): 140–143, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPg

[Ano64-329] Anonymous. Special publications. IEEE Spectrum, 1(11): 162–163, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPh

[Ano64-330] Anonymous. Special publications. IEEE Spectrum, 1(12): 118–120, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SLa

[Ano64-331] Anonymous. Spectral lines. IEEE Spectrum, 1(1):85, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SLb

[Ano64-332] Anonymous. Spectral lines. IEEE Spectrum, 1(3):181, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SMW

[Ano64-333] Anonymous. Study of Mars will start with the launch of Mariner-C. IEEE Spectrum, 1 (11):166, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SIb

[Ano64-334] Anonymous. Subject index.

IEEE Spectrum, 1(12):121–124, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SPY

[Ano64-335] Anonymous. Sun-pumped YAG laser requires no other power source. IEEE Spectrum, 1(10): 158–160, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SSW

[Ano64-336] Anonymous. SWE, SFPE, WSE, and HEC elected EJC members. IEEE Spectrum, 1(10):40, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SES

[Ano64-337] Anonymous. Symposium on exploration of solar system seeks papers. IEEE Spectrum, 1 (10):40, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:SRQ

[Ano64-338] Anonymous. Symposium on reliability and quality control scheduled. IEEE Spectrum, 1(10): 39–40, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TC

[Ano64-339] Anonymous. Table of contents.

IEEE Spectrum, 1(1):1, January 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TPIb

[Ano64-340] Anonymous. Technical papers invited for intermag conference. IEEE Spectrum, 1(11): 48–50, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TPIa

[Ano64-341] Anonymous. Technical program issued for symposium on switching circuit theory and logical design. IEEE Spectrum, 1(11): 36–38, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TYP

[Ano64-342] Anonymous. Ten years of progress in radio standards is celebrated by NBS laboratory at Boulder. IEEE Spectrum, 1(11): 165–166, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:IEE

[Ano64-343] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 1(1):2, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TFR

[Ano64-344] Anonymous. Thin-film rods give computer fast memory. IEEE Spectrum, 1(10):152, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TCI

[Ano64-345] Anonymous. Third congress of IFAC issues call for technical papers. IEEE Spectrum, 1 (10):38–39, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TRM

[Ano64-346] Anonymous. Time reversal may affect basic physics law. IEEE Spectrum, 1(10):160, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TJa

[Ano64-347] Anonymous. Translated journals. IEEE Spectrum, 1(1): 156–157, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TJb

[Ano64-348] Anonymous. Translated journals. IEEE Spectrum, 1(6): 166–175, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TJc

[Ano64-349] Anonymous. Translated journals. IEEE Spectrum, 1(7): 154–161, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TJd

[Ano64-350] Anonymous. Translated journals. IEEE Spectrum, 1(8):

122–127, August 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TJe

[Ano64-351] Anonymous. Translated journals. IEEE Spectrum, 1(9): 154–159, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TJf

[Ano64-352] Anonymous. Translated journals. IEEE Spectrum, 1(10): 134–139, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TJg

[Ano64-353] Anonymous. Translated journals. IEEE Spectrum, 1(11): 152–161, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:TJh

[Ano64-354] Anonymous. Translated journals. IEEE Spectrum, 1(12): 114–117, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:UET

[Ano64-355] Anonymous. Use of engineering talent is subject of White House release. IEEE Spectrum, 1 (10):152–158, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1964:ULB

[Ano64-356] Anonymous. Use of laser beam provides high accuracy in interferometer for industrial applications. IEEE Spectrum, 1(11): 164–165, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PPM

[Ano65a] Anonymous. 1 PPM: Potentiometric measuring system [advertisement]. IEEE Spectrum, 2(10):3, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BEV

[Ano65b] Anonymous. A 10-billion-electron-volt accelerator is being built. IEEE Spectrum, 2(8):23, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IICa

[Ano65c] Anonymous. 1964 IEEE International Convention Record. IEEE Spectrum, 2(1):10–11, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IICb

[Ano65d] Anonymous. 1965 IEEE International Convention Record. IEEE Spectrum, 2(7):124, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NCA

[Ano65e] Anonymous. 20th National Conference of ACM slated August 24–26. IEEE Spectrum, 2(8):

122–123, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:JHW

[Ano65f] Anonymous. A. S. Jerrems to head ’66 Winter Convention on AES. IEEE Spectrum, 2 (9):132, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAa

[Ano65g] Anonymous. Advance abstracts. IEEE Spectrum, 2(1): 132–141, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAb

[Ano65h] Anonymous. Advance abstracts. IEEE Spectrum, 2(2): 136–148, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAc

[Ano65i] Anonymous. Advance abstracts. IEEE Spectrum, 2(3):227–242, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAd

[Ano65j] Anonymous. Advance abstracts. IEEE Spectrum, 2(4):150–168, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAe

[Ano65k] Anonymous. Advance abstracts. IEEE Spectrum, 2(5):149–158, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAf

[Ano65l] Anonymous. Advance abstracts.

IEEE Spectrum, 2(6):122–137, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAg

[Ano65m] Anonymous. Advance abstracts.

IEEE Spectrum, 2(7):157–171, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAh

[Ano65n] Anonymous. Advance abstracts.

IEEE Spectrum, 2(8):140–147, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AAi

[Ano65o] Anonymous. Advance abstracts. IEEE Spectrum, 2(9): 146–163, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aa

[Ano65p] Anonymous. [advertisement].

IEEE Spectrum, 2(1):1, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ab

[Ano65q] Anonymous. [advertisement].

IEEE Spectrum, 2(1):3–4, January 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ac

[Ano65r] Anonymous. [advertisement]. IEEE Spectrum, 2(1):6–8, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ad

[Ano65s] Anonymous. [advertisement]. IEEE Spectrum, 2(1):12–13, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ae

[Ano65t] Anonymous. [advertisement]. IEEE Spectrum, 2(1):15–21, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Af

[Ano65u] Anonymous. [advertisement]. IEEE Spectrum, 2(1):25–48, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ah

[Ano65v] Anonymous. [advertisement]. IEEE Spectrum, 2(1):95–97, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ai

[Ano65w] Anonymous. [advertisement]. IEEE Spectrum, 2(1):103–117, January 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aj

[Ano65x] Anonymous. [advertisement].

IEEE Spectrum, 2(1):120–123, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ak

[Ano65y] Anonymous. [advertisement].

IEEE Spectrum, 2(1):124–125, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Al

[Ano65z] Anonymous. [advertisement].

IEEE Spectrum, 2(1):133–141, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Am

[Ano65-27] Anonymous. [advertisement].

IEEE Spectrum, 2(1):143–145, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:An

[Ano65-28] Anonymous. [advertisement].

IEEE Spectrum, 2(1):154–156, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ao

[Ano65-29] Anonymous. [advertisement].

IEEE Spectrum, 2(1):160, January 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ap

[Ano65-30] Anonymous. [advertisement]. IEEE Spectrum, 2(1):2, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aq

[Ano65-31] Anonymous. [advertisement]. IEEE Spectrum, 2(2):1, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ar

[Ano65-32] Anonymous. [advertisement]. IEEE Spectrum, 2(2):3, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:As

[Ano65-33] Anonymous. [advertisement]. IEEE Spectrum, 2(2):6–8, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:At

[Ano65-34] Anonymous. [advertisement]. IEEE Spectrum, 2(2):11–12, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Au

[Ano65-35] Anonymous. [advertisement]. IEEE Spectrum, 2(2):13, February 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Av

[Ano65-36] Anonymous. [advertisement].

IEEE Spectrum, 2(2):15–16, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aw

[Ano65-37] Anonymous. [advertisement].

IEEE Spectrum, 2(2):17, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ax

[Ano65-38] Anonymous. [advertisement].

IEEE Spectrum, 2(2):23, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ay

[Ano65-39] Anonymous. [advertisement].

IEEE Spectrum, 2(2):25–27, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Az

[Ano65-40] Anonymous. [advertisement].

IEEE Spectrum, 2(2):29–30, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aba

[Ano65-41] Anonymous. [advertisement].

IEEE Spectrum, 2(2):31, February 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abb

[Ano65-42] Anonymous. [advertisement]. IEEE Spectrum, 2(2):33, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abc

[Ano65-43] Anonymous. [advertisement]. IEEE Spectrum, 2(2):37, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abd

[Ano65-44] Anonymous. [advertisement]. IEEE Spectrum, 2(2):39–40, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abe

[Ano65-45] Anonymous. [advertisement]. IEEE Spectrum, 2(2):41–42, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abf

[Ano65-46] Anonymous. [advertisement]. IEEE Spectrum, 2(2):45–46, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abg

[Ano65-47] Anonymous. [advertisement]. IEEE Spectrum, 2(2):47–48, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abh

[Ano65-48] Anonymous. [advertisement].

IEEE Spectrum, 2(2):79, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abi

[Ano65-49] Anonymous. [advertisement].

IEEE Spectrum, 2(2):82, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abk

[Ano65-50] Anonymous. [advertisement].

IEEE Spectrum, 2(2):113, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abl

[Ano65-51] Anonymous. [advertisement].

IEEE Spectrum, 2(2):133, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abm

[Ano65-52] Anonymous. [advertisement].

IEEE Spectrum, 2(2):141, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abn

[Ano65-53] Anonymous. [advertisement].

IEEE Spectrum, 2(2):2, February 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abo

[Ano65-54] Anonymous. [advertisement]. IEEE Spectrum, 2(3):1, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abp

[Ano65-55] Anonymous. [advertisement]. IEEE Spectrum, 2(3):3, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abq

[Ano65-56] Anonymous. [advertisement]. IEEE Spectrum, 2(3):6–8, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abr

[Ano65-57] Anonymous. [advertisement]. IEEE Spectrum, 2(3):10–17, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abs

[Ano65-58] Anonymous. [advertisement]. IEEE Spectrum, 2(3):19–32, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abt

[Ano65-59] Anonymous. [advertisement]. IEEE Spectrum, 2(3):37–103, March 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abu

[Ano65-60] Anonymous. [advertisement].

IEEE Spectrum, 2(3):105–112, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abv

[Ano65-61] Anonymous. [advertisement].

IEEE Spectrum, 2(3):141–144, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abx

[Ano65-62] Anonymous. [advertisement].

IEEE Spectrum, 2(3):205–211, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aby

[Ano65-63] Anonymous. [advertisement].

IEEE Spectrum, 2(3):213–216, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abz

[Ano65-64] Anonymous. [advertisement].

IEEE Spectrum, 2(3):231–243, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aca

[Ano65-65] Anonymous. [advertisement].

IEEE Spectrum, 2(3):244–248, March 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acb

[Ano65-66] Anonymous. [advertisement]. IEEE Spectrum, 2(3):252–254, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acc

[Ano65-67] Anonymous. [advertisement]. IEEE Spectrum, 2(3):341, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acd

[Ano65-68] Anonymous. [advertisement]. IEEE Spectrum, 2(3):2, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ace

[Ano65-69] Anonymous. [advertisement]. IEEE Spectrum, 2(4):1, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acf

[Ano65-70] Anonymous. [advertisement]. IEEE Spectrum, 2(4):3, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acg

[Ano65-71] Anonymous. [advertisement]. IEEE Spectrum, 2(4):6–8, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ach

[Ano65-72] Anonymous. [advertisement].

IEEE Spectrum, 2(4):11–17, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aci

[Ano65-73] Anonymous. [advertisement].

IEEE Spectrum, 2(4):21–27, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acj

[Ano65-74] Anonymous. [advertisement].

IEEE Spectrum, 2(4):28–32, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ack

[Ano65-75] Anonymous. [advertisement].

IEEE Spectrum, 2(4):91–93, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aco

[Ano65-76] Anonymous. [advertisement].

IEEE Spectrum, 2(4):103–105, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acp

[Ano65-77] Anonymous. [advertisement].

IEEE Spectrum, 2(4):123–133, April 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acq

[Ano65-78] Anonymous. [advertisement]. IEEE Spectrum, 2(4):138–140, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acr

[Ano65-79] Anonymous. [advertisement]. IEEE Spectrum, 2(4):143, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acs

[Ano65-80] Anonymous. [advertisement]. IEEE Spectrum, 2(4):153–169, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Act

[Ano65-81] Anonymous. [advertisement]. IEEE Spectrum, 2(4):170–173, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acu

[Ano65-82] Anonymous. [advertisement]. IEEE Spectrum, 2(4):174–175, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acv

[Ano65-83] Anonymous. [advertisement]. IEEE Spectrum, 2(4):178–180, April 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acw

[Ano65-84] Anonymous. [advertisement].

IEEE Spectrum, 2(4):182–184, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acx

[Ano65-85] Anonymous. [advertisement].

IEEE Spectrum, 2(4):2, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acy

[Ano65-86] Anonymous. [advertisement].

IEEE Spectrum, 2(5):1, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acz

[Ano65-87] Anonymous. [advertisement].

IEEE Spectrum, 2(5):3, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ada

[Ano65-88] Anonymous. [advertisement].

IEEE Spectrum, 2(5):8, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adb

[Ano65-89] Anonymous. [advertisement].

IEEE Spectrum, 2(5):11–13, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adc

[Ano65-90] Anonymous. [advertisement]. IEEE Spectrum, 2(5):19–45, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Add

[Ano65-91] Anonymous. [advertisement]. IEEE Spectrum, 2(5):47–54, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ade

[Ano65-92] Anonymous. [advertisement]. IEEE Spectrum, 2(5):85–87, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adg

[Ano65-93] Anonymous. [advertisement]. IEEE Spectrum, 2(5):159, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adh

[Ano65-94] Anonymous. [advertisement]. IEEE Spectrum, 2(5):165, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adi

[Ano65-95] Anonymous. [advertisement]. IEEE Spectrum, 2(5):2, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adj

[Ano65-96] Anonymous. [advertisement].

IEEE Spectrum, 2(6):3, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adk

[Ano65-97] Anonymous. [advertisement].

IEEE Spectrum, 2(6):6–8, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adl

[Ano65-98] Anonymous. [advertisement].

IEEE Spectrum, 2(6):11–15, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adm

[Ano65-99] Anonymous. [advertisement].

IEEE Spectrum, 2(6):18–27, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adn

[Ano65-100] Anonymous. [advertisement].

IEEE Spectrum, 2(6):28–34, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ado

[Ano65-101] Anonymous. [advertisement].

IEEE Spectrum, 2(6):73–76, June 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adq

[Ano65-102] Anonymous. [advertisement]. IEEE Spectrum, 2(6):96–97, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adr

[Ano65-103] Anonymous. [advertisement]. IEEE Spectrum, 2(6):101, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ads

[Ano65-104] Anonymous. [advertisement]. IEEE Spectrum, 2(6):110–113, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adt

[Ano65-105] Anonymous. [advertisement]. IEEE Spectrum, 2(6):114–115, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adu

[Ano65-106] Anonymous. [advertisement]. IEEE Spectrum, 2(6):121, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adv

[Ano65-107] Anonymous. [advertisement]. IEEE Spectrum, 2(6):125–137, June 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adw

[Ano65-108] Anonymous. [advertisement].

IEEE Spectrum, 2(6):138–141, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adx

[Ano65-109] Anonymous. [advertisement].

IEEE Spectrum, 2(6):142, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ady

[Ano65-110] Anonymous. [advertisement].

IEEE Spectrum, 2(6):2, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adz

[Ano65-111] Anonymous. [advertisement].

IEEE Spectrum, 2(7):1, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aea

[Ano65-112] Anonymous. [advertisement].

IEEE Spectrum, 2(7):3, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aeb

[Ano65-113] Anonymous. [advertisement].

IEEE Spectrum, 2(7):6, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aec

[Ano65-114] Anonymous. [advertisement]. IEEE Spectrum, 2(7):9–15, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aed

[Ano65-115] Anonymous. [advertisement]. IEEE Spectrum, 2(7):17–25, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aee

[Ano65-116] Anonymous. [advertisement]. IEEE Spectrum, 2(7):26–31, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aef

[Ano65-117] Anonymous. [advertisement]. IEEE Spectrum, 2(7):91–94, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aeh

[Ano65-118] Anonymous. [advertisement]. IEEE Spectrum, 2(7):125, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aei

[Ano65-119] Anonymous. [advertisement]. IEEE Spectrum, 2(7):131–142, July 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aej

[Ano65-120] Anonymous. [advertisement].

IEEE Spectrum, 2(7):146–149, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aek

[Ano65-121] Anonymous. [advertisement].

IEEE Spectrum, 2(7):150–151, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ael

[Ano65-122] Anonymous. [advertisement].

IEEE Spectrum, 2(7):159–167, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aem

[Ano65-123] Anonymous. [advertisement].

IEEE Spectrum, 2(7):172–177, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aen

[Ano65-124] Anonymous. [advertisement].

IEEE Spectrum, 2(7):178–180, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aeo

[Ano65-125] Anonymous. [advertisement].

IEEE Spectrum, 2(7):2, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ACL

[Ano65-126] Anonymous. Advertisement — a closer look at the surfaces of solids [report from Bell Laboratories]. IEEE Spectrum, 2 (2):21, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AESa

[Ano65-127] Anonymous. Advertisement — Engineering Societies Personnel Service, Inc. (agency). IEEE Spectrum, 2(1):152–153, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AESb

[Ano65-128] Anonymous. Advertisement — Engineering Societies Personnel Service, Inc. (agency). IEEE Spectrum, 2(3):250–251, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AESc

[Ano65-129] Anonymous. Advertisement — Engineering Societies Personnel Service, Inc. (agency). IEEE Spectrum, 2(4):176–177, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AHD

[Ano65-130] Anonymous. Advertisement — how does a driver handle an automobile? [general motors research laboratories]. IEEE Spectrum, 2(4):94, April 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AOP

[Ano65-131] Anonymous. Advertisement — obviously photography is the neat way to capture information and record data. IEEE Spectrum, 2(2):35, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AOS

[Ano65-132] Anonymous. Advertisement — our Stirling engine: Just add heat [General Motors Research Laboratories]. IEEE Spectrum, 2(5):88, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:APCa

[Ano65-133] Anonymous. Advertisement — professional consulting services. IEEE Spectrum, 2(1):157, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:APCb

[Ano65-134] Anonymous. Advertisement — professional consulting services. IEEE Spectrum, 2(3):334, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:APCc

[Ano65-135] Anonymous. Advertisement — professional consulting services. IEEE Spectrum, 2(4):181, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:APCd

[Ano65-136] Anonymous. Advertisement — professional consulting services. IEEE Spectrum, 2(5):167, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:APCe

[Ano65-137] Anonymous. Advertisement — professional consulting services. IEEE Spectrum, 2(6):143–144, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:APCf

[Ano65-138] Anonymous. Advertisement — professional consulting services. IEEE Spectrum, 2(7):181, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ASV

[Ano65-139] Anonymous. Advertisement — Saturn V openings for engineers and scientists. IEEE Spectrum, 2(5):119, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ASL

[Ano65-140] Anonymous. Advertisement — side-looking radar systems analysts [hughes]. IEEE Spectrum, 2(3):217–221, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AUC

[Ano65-141] Anonymous. Advertisement — uninterruptible a-c power — without moving parts. IEEE Spectrum, 2(6):74, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aep

[Ano65-142] Anonymous. [advertisements].

IEEE Spectrum, 2(8):3–6, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aer

[Ano65-143] Anonymous. [advertisements].

IEEE Spectrum, 2(9):3–8, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aev

[Ano65-144] Anonymous. [advertisements].

IEEE Spectrum, 2(10):155– 156, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aew

[Ano65-145] Anonymous. [advertisements].

IEEE Spectrum, 2(11):3, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:AII

[Ano65-146] Anonymous. Annual index to IEEE Spectrum for volume 2, 1965. IEEE Spectrum, 2(12):

131–144, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ACM

[Ano65-147] Anonymous. Antenna-coupling measurements and transmission lines. IEEE Spectrum, 2(2): 5, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:APD

[Ano65-148] Anonymous. ASCE Power Division to sponsor first convention. IEEE Spectrum, 2(9): 24–26, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ag

[Ano65-149] Anonymous. Authors. IEEE Spectrum, 2(1):93–94, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abj

[Ano65-150] Anonymous. Authors. IEEE Spectrum, 2(2):111–112, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Abw

[Ano65-151] Anonymous. Authors. IEEE Spectrum, 2(3):196–197, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Acn

[Ano65-152] Anonymous. Authors. IEEE Spectrum, 2(4):102, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adf

[Ano65-153] Anonymous. Authors. IEEE Spectrum, 2(5):117–118, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Adp

[Ano65-154] Anonymous. Authors. IEEE Spectrum, 2(6):95, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aeg

[Ano65-155] Anonymous. Authors. IEEE Spectrum, 2(7):122–123, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aeq

[Ano65-156] Anonymous. Authors. IEEE Spectrum, 2(8):116–117, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aet

[Ano65-157] Anonymous. Authors. IEEE Spectrum, 2(9):120–123, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aeu

[Ano65-158] Anonymous. Authors. IEEE Spectrum, 2(10):98–99, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aex

[Ano65-159] Anonymous. Authors. IEEE Spectrum, 2(11):107–109, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Aey

[Ano65-160] Anonymous. Authors. IEEE Spectrum, 2(12):83, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BCa

[Ano65-161] Anonymous. [back cover]. IEEE Spectrum, 2(8):c4, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BCb

[Ano65-162] Anonymous. [back cover]. IEEE Spectrum, 2(9):c4, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BCc

[Ano65-163] Anonymous. [back cover]. IEEE Spectrum, 2(10):c4, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BCd

[Ano65-164] Anonymous. [back cover]. IEEE Spectrum, 2(11):c4, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BCe

[Ano65-165] Anonymous. [back cover]. IEEE Spectrum, 2(12):c4, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BEH

[Ano65-166] Anonymous. Biennial electric heating conference slated for September 28–29 in Cleveland. IEEE Spectrum, 2(9): 126–128, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRa

[Ano65-167] Anonymous. Book reviews.

IEEE Spectrum, 2(1):154–156, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRb

[Ano65-168] Anonymous. Book reviews.

IEEE Spectrum, 2(2):156–157, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRc

[Ano65-169] Anonymous. Book reviews.

IEEE Spectrum, 2(3):252–253, March 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRd

[Ano65-170] Anonymous. Book reviews. IEEE Spectrum, 2(4):178–181, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRe

[Ano65-171] Anonymous. Book reviews. IEEE Spectrum, 2(5):166, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRf

[Ano65-172] Anonymous. Book reviews. IEEE Spectrum, 2(7):178–180, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRg

[Ano65-173] Anonymous. Book reviews. IEEE Spectrum, 2(10):149– 151, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRh

[Ano65-174] Anonymous. Book reviews. IEEE Spectrum, 2(11):152–155, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BRi

[Ano65-175] Anonymous. Book reviews. IEEE Spectrum, 2(12):126–129, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IS

[Ano65-176] Anonymous. IEEE Spectrum.

IEEE Spectrum, 2(2):4, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BPa

[Ano65-177] Anonymous. [breaker page].

IEEE Spectrum, 2(1):4, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BPb

[Ano65-178] Anonymous. [breaker page].

IEEE Spectrum, 2(3):4, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BPc

[Ano65-179] Anonymous. [breaker page].

IEEE Spectrum, 2(4):4, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BPd

[Ano65-180] Anonymous. [breaker page].

IEEE Spectrum, 2(5):4, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BPe

[Ano65-181] Anonymous. [breaker page].

IEEE Spectrum, 2(7):4, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BPO

[Ano65-182] Anonymous. Brookyln polytech offers courses in new techniques. IEEE Spectrum, 2(9): 28–29, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:BPAa

[Ano65-183] Anonymous. The business principles of an aerospace industry. IEEE Spectrum, 2(6):31, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cb

[Ano65-184] Anonymous. Calendar. IEEE Spectrum, 2(1):120–123, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cd

[Ano65-185] Anonymous. Calendar. IEEE Spectrum, 2(2):127–129, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ce

[Ano65-186] Anonymous. Calendar. IEEE Spectrum, 2(3):213–216, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cg

[Ano65-187] Anonymous. Calendar. IEEE Spectrum, 2(4):138–139, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ci

[Ano65-188] Anonymous. Calendar. IEEE Spectrum, 2(5):134–137, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ck

[Ano65-189] Anonymous. Calendar. IEEE Spectrum, 2(6):110–113, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cm

[Ano65-190] Anonymous. Calendar. IEEE Spectrum, 2(7):146–149, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cp

[Ano65-191] Anonymous. Calendar. IEEE Spectrum, 2(8):128–131, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cr

[Ano65-192] Anonymous. Calendar. IEEE Spectrum, 2(9):135–139, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ct

[Ano65-193] Anonymous. Calendar. IEEE Spectrum, 2(10):116–119, October 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cw

[Ano65-194] Anonymous. Calendar. IEEE Spectrum, 2(11):126–129, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cz

[Ano65-195] Anonymous. Calendar. IEEE Spectrum, 2(12):92–95, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CSU

[Ano65-196] Anonymous. Camera system uses solid-state imaging device. IEEE Spectrum, 2(9):20, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CET

[Ano65-197] Anonymous. Canadian electronics throughout world to be studied. IEEE Spectrum, 2(9): 128–130, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CCP

[Ano65-198] Anonymous. Carbon-coated particles may reduce generating costs. IEEE Spectrum, 2(9): 16–18, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CAa

[Ano65-199] Anonymous. Classified advertising. IEEE Spectrum, 2(1): 128–129, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CAb

[Ano65-200] Anonymous. Classified advertising. IEEE Spectrum, 2(3): 198–199, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CAc

[Ano65-201] Anonymous. Classified advertising. IEEE Spectrum, 2(4):141, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CAd

[Ano65-202] Anonymous. Classified advertising. IEEE Spectrum, 2(6): 120–121, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CAe

[Ano65-203] Anonymous. Classified advertising. IEEE Spectrum, 2(7):154, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CAf

[Ano65-204] Anonymous. Classified advertising. IEEE Spectrum, 2(8): 118–119, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CAg

[Ano65-205] Anonymous. Classified advertising. IEEE Spectrum, 2(9): 142–143, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CFP

[Ano65-206] Anonymous. Conference to feature papers on EHV transmission. IEEE Spectrum, 2(8):127, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Co

[Ano65-207] Anonymous. Contents. IEEE Spectrum, 2(8):1–2, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cq

[Ano65-208] Anonymous. Contents. IEEE Spectrum, 2(9):1–2, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cs

[Ano65-209] Anonymous. Contents. IEEE Spectrum, 2(10):1–2, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cu

[Ano65-210] Anonymous. Contents. IEEE Spectrum, 2(11):1–2, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cx

[Ano65-211] Anonymous. Contents. IEEE Spectrum, 2(12):1–2, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:CN

[Ano65-212] Anonymous. Correction notice. IEEE Spectrum, 2(7):119, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cc

[Ano65-213] Anonymous. Correspondence.

IEEE Spectrum, 2(2):44–46, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cf

[Ano65-214] Anonymous. Correspondence.

IEEE Spectrum, 2(4):28–31, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ch

[Ano65-215] Anonymous. Correspondence.

IEEE Spectrum, 2(5):46–50, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cj

[Ano65-216] Anonymous. Correspondence.

IEEE Spectrum, 2(6):28–32, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cl

[Ano65-217] Anonymous. Correspondence. IEEE Spectrum, 2(7):26–31, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cv

[Ano65-218] Anonymous. Correspondence. IEEE Spectrum, 2(11):42–48, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Cy

[Ano65-219] Anonymous. Correspondence. IEEE Spectrum, 2(12):36–42, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:DFa

[Ano65-220] Anonymous. Designer’s file. IEEE Spectrum, 2(2):80–81, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:DFb

[Ano65-221] Anonymous. Designer’s file. IEEE Spectrum, 2(5):6–7, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:DHM

[Ano65-222] Anonymous. Developments in handling of materials will be featured. IEEE Spectrum, 2(8):127, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:DSA

[Ano65-223] Anonymous. E. D. Shanken appointed to EJC’s education program. IEEE Spectrum, 2(8):127, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ESSa

[Ano65-224] Anonymous. Economics of static standby a-c power systems. IEEE Spectrum, 2(5):22, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ESSb

[Ano65-225] Anonymous. Economics of static standby a-c power systems. IEEE Spectrum, 2(7):11, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:EMO

[Ano65-226] Anonymous. Electronic maps of ocean will help guide spacecraft. IEEE Spectrum, 2(9): 26–28, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:EPTa

[Ano65-227] Anonymous. An electronic piano-tuner based on digital techniques. Part I. IEEE Spectrum, 2(5):5, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:EPTb

[Ano65-228] Anonymous. An electronic piano-tuner based on digital techniques: Part II. IEEE Spectrum, 2(6):5, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:EFW

[Ano65-229] Anonymous. Engineering foundation will hold research conference. IEEE Spectrum, 2(8):123, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ESP

[Ano65-230] Anonymous. Engineering Societies Personnel Service, Inc. IEEE Spectrum, 2(2):154–155, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ESD

[Ano65-231] Anonymous. Exact shape determined for the saccharine molecule. IEEE Spectrum, 2 (9):26, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ECA

[Ano65-232] Anonymous. Executive committee approves IEEE chapters. IEEE Spectrum, 2(8):127, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:EA

[Ano65-233] Anonymous. Exhibitors [advertisement]. IEEE Spectrum, 2 (3):255–335, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FFS

[Ano65-234] Anonymous. Fighting fatigue with stress. IEEE Spectrum, 2(3):90, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FET

[Ano65-235] Anonymous. First EHV transmission conference scheduled for October 4–6 in Richmond, Va. IEEE Spectrum, 2(9): 125, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPd

[Ano65-236] Anonymous. Floor plan. IEEE Spectrum, 2(3):335–339, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FGL

[Ano65-237] Anonymous. Flowing gas laser has highest reported efficiency. IEEE Spectrum, 2(8):21–22, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPa

[Ano65-238] Anonymous. Focal points. IEEE Spectrum, 2(1):24–46, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPb

[Ano65-239] Anonymous. Focal points. IEEE Spectrum, 2(2):18–40, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPc

[Ano65-240] Anonymous. Focal points. IEEE Spectrum, 2(3):104–111, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPe

[Ano65-241] Anonymous. Focal points. IEEE Spectrum, 2(4):18–27, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPf

[Ano65-242] Anonymous. Focal points. IEEE Spectrum, 2(7):16–24, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPg

[Ano65-243] Anonymous. Focal points. IEEE Spectrum, 2(10):22–40, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPh

[Ano65-244] Anonymous. Focal points. IEEE Spectrum, 2(11):24–41, November 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPi

[Ano65-245] Anonymous. Focal points. IEEE Spectrum, 2(12):18–35, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPAa

[Ano65-246] Anonymous. Focal points:

American scientist offers unique theory and proof to solve Stonehenge ‘mystery of the megaliths’. IEEE Spectrum, 2(5):16–42, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPN

[Ano65-247] Anonymous. Focal points: New EHV line across St. Lawrence River will provide 735,000-volt power supply for Quebec. IEEE Spectrum, 2(9):16, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPTa

[Ano65-248] Anonymous. Focal points: Twophoton effect with laser light causes a specific chemical reaction for first time. IEEE Spectrum, 2(6):16–24, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FPTb

[Ano65-249] Anonymous. Focal points: Two satellites designed at M.I.T.: a radar calibration sphere and a communications satellite have been launched. IEEE Spectrum, 2(8):14, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCa

[Ano65-250] Anonymous. [front cover]. IEEE Spectrum, 2(1):c1, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCb

[Ano65-251] Anonymous. [front cover]. IEEE Spectrum, 2(1):c2, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCc

[Ano65-252] Anonymous. [front cover]. IEEE Spectrum, 2(2):c1, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCd

[Ano65-253] Anonymous. [front cover]. IEEE Spectrum, 2(2):c2, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCe

[Ano65-254] Anonymous. [front cover]. IEEE Spectrum, 2(3):c1, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCf

[Ano65-255] Anonymous. [front cover]. IEEE Spectrum, 2(3):c2, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCg

[Ano65-256] Anonymous. [front cover]. IEEE Spectrum, 2(4):c1, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCh

[Ano65-257] Anonymous. [front cover]. IEEE Spectrum, 2(4):c2, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCi

[Ano65-258] Anonymous. [front cover]. IEEE Spectrum, 2(5):c1, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCj

[Ano65-259] Anonymous. [front cover]. IEEE Spectrum, 2(5):c2, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCk

[Ano65-260] Anonymous. [front cover]. IEEE Spectrum, 2(6):c1, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCl

[Ano65-261] Anonymous. [front cover]. IEEE Spectrum, 2(6):c2, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCm

[Ano65-262] Anonymous. [front cover]. IEEE Spectrum, 2(7):c1, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCn

[Ano65-263] Anonymous. [front cover]. IEEE Spectrum, 2(7):c2, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCo

[Ano65-264] Anonymous. [front cover]. IEEE Spectrum, 2(8):c1, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCp

[Ano65-265] Anonymous. [front cover]. IEEE Spectrum, 2(9):c1, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCq

[Ano65-266] Anonymous. [front cover]. IEEE Spectrum, 2(10):c1, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCr

[Ano65-267] Anonymous. [front cover]. IEEE Spectrum, 2(11):c1, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:FCs

[Ano65-268] Anonymous. [front cover]. IEEE Spectrum, 2(12):c1, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:GCL

[Ano65-269] Anonymous. A good close look at corrosion mechanisms. IEEE Spectrum, 2(2):36, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:GMA

[Ano65-270] Anonymous. Graphical methods [advertisement]. IEEE Spectrum, 2(11):5–8, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HER

[Ano65-271] Anonymous. H. E. Edgerton receives 1965 Morris E. Leeds Award. IEEE Spectrum, 2(9): 131–132, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HURa

[Ano65-272] Anonymous. How to use reader service card! IEEE Spectrum, 2(1):34, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HURc

[Ano65-273] Anonymous. How to use reader service card! IEEE Spectrum, 2(2):156, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HURd

[Ano65-274] Anonymous. How to use reader service card! IEEE Spectrum, 2(3):344, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HURe

[Ano65-275] Anonymous. How to use reader service card! IEEE Spectrum, 2(4):142, April 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HURf

[Ano65-276] Anonymous. How to use reader service card! IEEE Spectrum, 2(5):170, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HURg

[Ano65-277] Anonymous. How to use reader service card! IEEE Spectrum, 2(7):181, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HURb

[Ano65-278] Anonymous. How to use reader service page! IEEE Spectrum, 2(1):142, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IWT

[Ano65-279] Anonymous. IEDM will take place October 20–22 in Washington. IEEE Spectrum, 2(8):124, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASb

[Ano65-280] Anonymous. IEEE and American Standards. IEEE Spectrum, 2(1):146–150, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IPa

[Ano65-281] Anonymous. IEEE publications. IEEE Spectrum, 2(10): 128–148, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IPb

[Ano65-282] Anonymous. IEEE publications. IEEE Spectrum, 2(11): 136–151, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IPc

[Ano65-283] Anonymous. IEEE publications. IEEE Spectrum, 2(12): 104–125, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IPSa

[Ano65-284] Anonymous. IEEE publications: Scanning the issues. IEEE Spectrum, 2(8):138–139, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IPSb

[Ano65-285] Anonymous. IEEE publications: Scanning the issues. IEEE Spectrum, 2(9):144–145, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IIS

[Ano65-286] Anonymous. IFAC–IFIP symposium slated for October in Munich. IEEE Spectrum, 2(9): 130–131, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAa

[Ano65-287] Anonymous. Index to advertisers. IEEE Spectrum, 2(1): 158–159, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAb

[Ano65-288] Anonymous. Index to advertisers. IEEE Spectrum, 2(2): 158–159, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAc

[Ano65-289] Anonymous. Index to advertisers. IEEE Spectrum, 2(3): 340–343, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAd

[Ano65-290] Anonymous. Index to advertisers. IEEE Spectrum, 2(4): 182–183, April 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAe

[Ano65-291] Anonymous. Index to advertisers. IEEE Spectrum, 2 (5):168–169, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAf

[Ano65-292] Anonymous. Index to advertisers. IEEE Spectrum, 2 (6):144–146, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAg

[Ano65-293] Anonymous. Index to advertisers. IEEE Spectrum, 2 (7):182–184, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAh

[Ano65-294] Anonymous. Index to advertisers. IEEE Spectrum, 2(8): 162–164, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAi

[Ano65-295] Anonymous. Index to advertisers. IEEE Spectrum, 2(9): 182–184, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAj

[Ano65-296] Anonymous. Index to advertisers. IEEE Spectrum, 2(10): 152–154, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAk

[Ano65-297] Anonymous. Index to advertisers. IEEE Spectrum, 2(11): 156–158, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IAl

[Ano65-298] Anonymous. Index to advertisers. IEEE Spectrum, 2(12): 145–146, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ISP

[Ano65-299] Anonymous. Industrial static power conversion conference to be held November 1–3 at Benjamin Franklin Hotel in Philadelphia. IEEE Spectrum, 2(9): 126, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASa

[Ano65-300] Anonymous. Information about the society. IEEE Spectrum, 2 (1):22–23, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASc

[Ano65-301] Anonymous. Information about the society. IEEE Spectrum, 2 (1):159–160, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASd

[Ano65-302] Anonymous. Information about the society. IEEE Spectrum, 2 (3):96–97, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASe

[Ano65-303] Anonymous. Information about the society. IEEE Spectrum, 2 (3):343–344, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASf

[Ano65-304] Anonymous. Information about the society. IEEE Spectrum, 2 (4):183–184, April 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASg

[Ano65-305] Anonymous. Information about the society. IEEE Spectrum, 2 (6):145–146, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASh

[Ano65-306] Anonymous. Information about the society. IEEE Spectrum, 2 (7):126–127, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IASi

[Ano65-307] Anonymous. Information about the society. IEEE Spectrum, 2 (7):183–184, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IIA

[Ano65-308] Anonymous. Information for IEEE authors. IEEE Spectrum, 2(8):111–115, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IBCa

[Ano65-309] Anonymous. [inside back cover]. IEEE Spectrum, 2(8):c3, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IBCb

[Ano65-310] Anonymous. [inside back cover]. IEEE Spectrum, 2(9):c3, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IBCc

[Ano65-311] Anonymous. [inside back cover]. IEEE Spectrum, 2(10):c3, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IBCd

[Ano65-312] Anonymous. [inside back cover]. IEEE Spectrum, 2 (11):c3, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IBCe

[Ano65-313] Anonymous. [inside back cover]. IEEE Spectrum, 2(12):c3, December 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IFCa

[Ano65-314] Anonymous. [inside front cover].

IEEE Spectrum, 2(8):c2, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IFCb

[Ano65-315] Anonymous. [inside front cover].

IEEE Spectrum, 2(9):c2, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IFCc

[Ano65-316] Anonymous. [inside front cover].

IEEE Spectrum, 2(10):c2, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IFCd

[Ano65-317] Anonymous. [inside front cover]. IEEE Spectrum, 2 (11):c2, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IFCe

[Ano65-318] Anonymous. [inside front cover].

IEEE Spectrum, 2(12):c2, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ISE

[Ano65-319] Anonymous. International space electronics symposium plans program of 12 technical sessions and panel discussion. IEEE Spectrum, 2(9):128, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ISI

[Ano65-320] Anonymous. International symposium on information theory planned. IEEE Spectrum, 2 (8):126, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:JES

[Ano65-321] Anonymous. Joint engineering societies plan management conference. IEEE Spectrum, 2 (8):123, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:KRG

[Ano65-322] Anonymous. Kodak research grants will total $250,000 for 1965. IEEE Spectrum, 2(8): 20–21, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:LEC

[Ano65-323] Anonymous. Laser experiment confirms 32-year-old prediction concerning electron scattering. IEEE Spectrum, 2(8):16–18, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:LSa

[Ano65-324] Anonymous. [list of staff]. IEEE Spectrum, 2(1):4, January 1965.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:LSb

[Ano65-325] Anonymous. [list of staff]. IEEE Spectrum, 2(3):4, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:LSc

[Ano65-326] Anonymous. List of staff. IEEE Spectrum, 2(10):4, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:LSd

[Ano65-327] Anonymous. List of staff.

IEEE Spectrum, 2(11):4, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:LSe

[Ano65-328] Anonymous. List of staff.

IEEE Spectrum, 2(12):4, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ma

[Ano65-329] Anonymous. [masthead]. IEEE Spectrum, 2(2):4, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Mb

[Ano65-330] Anonymous. [masthead]. IEEE Spectrum, 2(6):4, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:MGT

[Ano65-331] Anonymous. Maxwell’s greatest tribute to Faraday. IEEE Spectrum, 2(10):5–8, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:MDA

[Ano65-332] Anonymous. Membership directory to be available late this year. IEEE Spectrum, 2(9): 132, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:MBF

[Ano65-333] Anonymous. Microwave behavior of ferrimagnetics to be studied. IEEE Spectrum, 2 (9):134, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:MMI

[Ano65-334] Anonymous. Microwave miniaturization: An image-phasing front end. IEEE Spectrum, 2(1):5, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:MMF

[Ano65-335] Anonymous. Microwave miniaturization: Ferrite circulators. IEEE Spectrum, 2(4):5, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:MEC

[Ano65-336] Anonymous. Military electronics convention slated for September 22–24 in Washington, D.C. IEEE Spectrum, 2(9):125–126, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NSD

[Ano65-337] Anonymous. NBS seminars discuss microwave high-power tubes. IEEE Spectrum, 2(9): 24, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NPL

[Ano65-338] Anonymous. New PAR lock-in amplifier measures signals in the presence of noise by crosscorrelation [advertisement]. IEEE Spectrum, 2(12):3, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NTM

[Ano65-339] Anonymous. New techniques for making nearly perfect mirrors. IEEE Spectrum, 2(4):19, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIa

[Ano65-340] Anonymous. News of the IEEE. IEEE Spectrum, 2(1): 98–119, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIb

[Ano65-341] Anonymous. News of the IEEE. IEEE Spectrum, 2(2): 114–126, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIc

[Ano65-342] Anonymous. News of the IEEE. IEEE Spectrum, 2(3):200–212, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NId

[Ano65-343] Anonymous. News of the IEEE. IEEE Spectrum, 2(4):106–137, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIe

[Ano65-344] Anonymous. News of the IEEE. IEEE Spectrum, 2(5):120–133, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIf

[Ano65-345] Anonymous. News of the IEEE. IEEE Spectrum, 2(6):98–109, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIg

[Ano65-346] Anonymous. News of the IEEE. IEEE Spectrum, 2(7):128–145, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIh

[Ano65-347] Anonymous. News of the IEEE. IEEE Spectrum, 2(10): 100–115, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIi

[Ano65-348] Anonymous. News of the IEEE. IEEE Spectrum, 2(11): 110–125, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIj

[Ano65-349] Anonymous. News of the IEEE. IEEE Spectrum, 2(12): 84–91, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIN

[Ano65-350] Anonymous. News of the IEEE:

1965 National Power Conference, sponsored by IEEE ASME, to be held in Albany. IEEE Spectrum, 2(8):120, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NIS

[Ano65-351] Anonymous. News of the IEEE:

Summer power conference attended by 1200 registrants: 25 technical sessions held. IEEE Spectrum, 2(9):124, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:NSG

[Ano65-352] Anonymous. Nuclear science group elects new slate of officers. IEEE Spectrum, 2(9): 132, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Oa

[Ano65-353] Anonymous. Obituaries. IEEE Spectrum, 2(1):127, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ob

[Ano65-354] Anonymous. Obituaries. IEEE Spectrum, 2(2):131–132, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Oc

[Ano65-355] Anonymous. Obituaries. IEEE Spectrum, 2(3):222–224, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Od

[Ano65-356] Anonymous. Obituaries. IEEE Spectrum, 2(4):145–147, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Oe

[Ano65-357] Anonymous. Obituaries. IEEE Spectrum, 2(5):144–145, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Of

[Ano65-358] Anonymous. Obituaries. IEEE Spectrum, 2(6):117, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Og

[Ano65-359] Anonymous. Obituaries. IEEE Spectrum, 2(7):152–153, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:OPN

[Ano65-360] Anonymous. Obviously photography is the neat way to capture information and record data. IEEE Spectrum, 2(8):86–89, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:ORP

[Ano65-361] Anonymous. Oceanographic research program to use meteorburst communication system to transmit data from ocean floor. IEEE Spectrum, 2(8):14–16, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:OTO

[Ano65-362] Anonymous. Oscillator is tunable at optical frequencies. IEEE Spectrum, 2(9):23–24, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PSP

[Ano65-363] Anonymous. Papers are solicited for Pan American Congress. IEEE Spectrum, 2(8):124, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PAM

[Ano65-364] Anonymous. Papers on applied magnetics and related phenomena invited. IEEE Spectrum, 2(9):132–134, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PEM

[Ano65-365] Anonymous. Papers on engineering in medicine and biology sought. IEEE Spectrum, 2(8):127, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PDV

[Ano65-366] Anonymous. Patrolling for disabled vehicles. IEEE Spectrum, 2(3):5, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pa

[Ano65-367] Anonymous. People. IEEE Spectrum, 2(1):124–127, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pb

[Ano65-368] Anonymous. People. IEEE Spectrum, 2(2):130–131, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pc

[Ano65-369] Anonymous. People. IEEE Spectrum, 2(3):217–221, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pd

[Ano65-370] Anonymous. People. IEEE Spectrum, 2(4):142–145, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pe

[Ano65-371] Anonymous. People. IEEE Spectrum, 2(5):138–143, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pf

[Ano65-372] Anonymous. People. IEEE Spectrum, 2(6):114–117, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pg

[Ano65-373] Anonymous. People. IEEE Spectrum, 2(7):150–152, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ph

[Ano65-374] Anonymous. People. IEEE Spectrum, 2(8):132–137, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pi

[Ano65-375] Anonymous. People. IEEE Spectrum, 2(9):140–141, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pj

[Ano65-376] Anonymous. People. IEEE Spectrum, 2(10):120–127, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pk

[Ano65-377] Anonymous. People. IEEE Spectrum, 2(11):126–135, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Pl

[Ano65-378] Anonymous. People. IEEE Spectrum, 2(12):96–103, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PCT

[Ano65-379] Anonymous. Petroleum conference to take place September 13–15. IEEE Spectrum, 2 (9):132, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PDC

[Ano65-380] Anonymous. A philosophical discourse concerning speech, … [facsimile of 1668 cover page]. IEEE Spectrum, 2(4):44, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PP

[Ano65-381] Anonymous. Photon propulsion. IEEE Spectrum, 2(7):5, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PFS

[Ano65-382] Anonymous. Physics of failure to be symposium subject. IEEE Spectrum, 2(9):28, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PPW

[Ano65-383] Anonymous. Plans progress for winter power meeting. IEEE Spectrum, 2(8):127, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PSS

[Ano65-384] Anonymous. Pre-specification selling. IEEE Spectrum, 2 (2):32, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PCS

[Ano65-385] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 2(2):160, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PIP

[Ano65-386] Anonymous. The professor, the industrialist, and the painter. IEEE Spectrum, 2(11):96–100, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:PNP

[Ano65-387] Anonymous. Pulse network problems solved by experienced systems engineers [advertisement]. IEEE Spectrum, 2(5):3, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:RSU

[Ano65-388] Anonymous. Radio spectrum utilization [advertisement]. IEEE Spectrum, 2(12):130, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:RSPa

[Ano65-389] Anonymous. Reader service page. IEEE Spectrum, 2(2): 159–160, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:RSPb

[Ano65-390] Anonymous. Reader service page. IEEE Spectrum, 2(5): 169–170, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:RUI

[Ano65-391] Anonymous. Recorder uses 1/4inch sound tape for both picture and sound. IEEE Spectrum, 2(8):20, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ra

[Ano65-392] Anonymous. Reflections. IEEE Spectrum, 2(1):14–20, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rb

[Ano65-393] Anonymous. Reflections. IEEE Spectrum, 2(2):10–16, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rc

[Ano65-394] Anonymous. Reflections. IEEE Spectrum, 2(3):18–28, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rd

[Ano65-395] Anonymous. Reflections. IEEE Spectrum, 2(4):10–16, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Re

[Ano65-396] Anonymous. Reflections. IEEE Spectrum, 2(5):10–15, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rf

[Ano65-397] Anonymous. Reflections. IEEE Spectrum, 2(6):10–14, June 1965.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rg

[Ano65-398] Anonymous. Reflections. IEEE Spectrum, 2(7):8–12, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rh

[Ano65-399] Anonymous. Reflections. IEEE Spectrum, 2(8):8–13, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Ri

[Ano65-400] Anonymous. Reflections. IEEE Spectrum, 2(9):10–15, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rj

[Ano65-401] Anonymous. Reflections. IEEE Spectrum, 2(10):10–21, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rk

[Ano65-402] Anonymous. Reflections. IEEE Spectrum, 2(11):12–23, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:Rl

[Ano65-403] Anonymous. Reflections. IEEE Spectrum, 2(12):10–17, December 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:RSC

[Ano65-404] Anonymous. Region six conference is planned for April 26–28. IEEE Spectrum, 2(8):121–122, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:REE

[Ano65-405] Anonymous. Rise of the engineer executive. IEEE Spectrum, 2(9): 113–116, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:RES

[Ano65-406] Anonymous. Rural electricity studied at 9th annual IEEE conference. IEEE Spectrum, 2 (8):122, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:HRA

[Ano65-407] Anonymous. S. E. Harris is recipient of 1965 Alfred Noble Prize. IEEE Spectrum, 2(8):126, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SSH

[Ano65-408] Anonymous. Satisfaction and success in the highly organized society. IEEE Spectrum, 2(12): 64–69, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SIa

[Ano65-409] Anonymous. Scanning the issues. IEEE Spectrum, 2(1): 130–131, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SIb

[Ano65-410] Anonymous. Scanning the issues. IEEE Spectrum, 2(2): 134–135, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SIc

[Ano65-411] Anonymous. Scanning the issues. IEEE Spectrum, 2(3):225–226, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SId

[Ano65-412] Anonymous. Scanning the issues. IEEE Spectrum, 2(4):148–149, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SIe

[Ano65-413] Anonymous. Scanning the issues. IEEE Spectrum, 2(5):146–148, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SIf

[Ano65-414] Anonymous. Scanning the issues. IEEE Spectrum, 2(6):118–119, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SIg

[Ano65-415] Anonymous. Scanning the issues.

IEEE Spectrum, 2(7):155–156, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SCD

[Ano65-416] Anonymous. Scientists to collect dust particles in near space. IEEE Spectrum, 2(9): 22–23, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SAC

[Ano65-417] Anonymous. Second annual conference on industrial and commercial power systems slated for October 5–7. IEEE Spectrum, 2(8):120, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SBC

[Ano65-418] Anonymous. Second biophysics congress to be held in Vienna in 1966. IEEE Spectrum, 2(9): 20–22, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SFL

[Ano65-419] Anonymous. Single-frequency laser produced without loss of power. IEEE Spectrum, 2(9): 18–20, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SSCb

[Ano65-420] Anonymous. Solid-state circuits conference scheduled for February 9–11. IEEE Spectrum, 2 (8):123–124, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPa

[Ano65-421] Anonymous. Special publications. IEEE Spectrum, 2 (1):151, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPb

[Ano65-422] Anonymous. Special publications. IEEE Spectrum, 2 (2):153, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPc

[Ano65-423] Anonymous. Special publications. IEEE Spectrum, 2(3):249, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPd

[Ano65-424] Anonymous. Special publications. IEEE Spectrum, 2(4):174, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPe

[Ano65-425] Anonymous. Special publications. IEEE Spectrum, 2(5):164, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPf

[Ano65-426] Anonymous. Special publications. IEEE Spectrum, 2(6):142, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPg

[Ano65-427] Anonymous. Special publications. IEEE Spectrum, 2(7):177, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPh

[Ano65-428] Anonymous. Special publications. IEEE Spectrum, 2(8): 154–155, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPi

[Ano65-429] Anonymous. Special publications. IEEE Spectrum, 2(9): 174–175, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLa

[Ano65-430] Anonymous. Spectral lines.

IEEE Spectrum, 2(1):49, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLb

[Ano65-431] Anonymous. Spectral lines.

IEEE Spectrum, 2(2):49, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLc

[Ano65-432] Anonymous. Spectral lines. IEEE Spectrum, 2(3):113, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLd

[Ano65-433] Anonymous. Spectral lines. IEEE Spectrum, 2(4):33, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLe

[Ano65-434] Anonymous. Spectral lines. IEEE Spectrum, 2(5):55, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLf

[Ano65-435] Anonymous. Spectral lines. IEEE Spectrum, 2(6):35, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLg

[Ano65-436] Anonymous. Spectral lines. IEEE Spectrum, 2(7):61, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLR

[Ano65-437] Anonymous. Spectral lines: Role of the proceedings. IEEE Spectrum, 2(10):41, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLAa

[Ano65-438] Anonymous. Spectral lines: The ASEE goals report. IEEE Spectrum, 2(12):43, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SLN

[Ano65-439] Anonymous. Spectral lines: The new publications board. IEEE Spectrum, 2(11):49, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SHN

[Ano65-440] Anonymous. Stimulation of hearing nerve produces sound for deaf person. IEEE Spectrum, 2 (8):22–23, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPCa

[Ano65-441] Anonymous. Stored-program control — flexibility for telephone switching systems. IEEE Spectrum, 2(6):17, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SRF

[Ano65-442] Anonymous. Superconducting radio-frequency circuits [advertisement]. IEEE Spectrum, 2 (12):5–8, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SPAa

[Ano65-443] Anonymous. Symposium planned on antennas and propagation. IEEE Spectrum, 2(8):124–126, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:SSCa

[Ano65-444] Anonymous. Systems science conference is scheduled for October 18–19. IEEE Spectrum, 2(8):122, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TCa

[Ano65-445] Anonymous. Table of contents. IEEE Spectrum, 2(1):1–2, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TCb

[Ano65-446] Anonymous. Table of contents. IEEE Spectrum, 2(2):1–2, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TCc

[Ano65-447] Anonymous. Table of contents. IEEE Spectrum, 2(3):1–2, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TCd

[Ano65-448] Anonymous. Table of contents. IEEE Spectrum, 2(4):1–2, April 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TCe

[Ano65-449] Anonymous. Table of contents. IEEE Spectrum, 2(5):1– 2, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TCf

[Ano65-450] Anonymous. Table of contents. IEEE Spectrum, 2(6):1–2, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TCg

[Ano65-451] Anonymous. Table of contents. IEEE Spectrum, 2(7):1– 2, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TRN

[Ano65-452] Anonymous. Tales of restless nuclei. IEEE Spectrum, 2(6):1, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TP

[Ano65-453] Anonymous. Technical program. IEEE Spectrum, 2(3): 33–102, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IEEa

[Ano65-454] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 2(1):4, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IEEb

[Ano65-455] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 2(2):4, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IEEc

[Ano65-456] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 2(3):4, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IEEd

[Ano65-457] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 2(4): 4, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IEEe

[Ano65-458] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 2(5): 4, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IEEf

[Ano65-459] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 2(6): 4, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:IEEg

[Ano65-460] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 2(7): 4, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TCM

[Ano65-461] Anonymous. Thermal contour map of the Moon’s Sea of Tranquillity shows wide temperature variation during eclipse. IEEE Spectrum, 2(8):18–20, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TIG

[Ano65-462] Anonymous. Tokyo: 30th IEC general meeting: October 10– 23, 1965: Provisional timetable of meetings. IEEE Spectrum, 2 (9):50–60, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTa

[Ano65-463] Anonymous. Transients and trends. IEEE Spectrum, 2 (1):9, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTb

[Ano65-464] Anonymous. Transients and trends. IEEE Spectrum, 2 (2):9, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTc

[Ano65-465] Anonymous. Transients and trends. IEEE Spectrum, 2(3):9, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTd

[Ano65-466] Anonymous. Transients and trends. IEEE Spectrum, 2(4):9, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTe

[Ano65-467] Anonymous. Transients and trends. IEEE Spectrum, 2(5):9, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTf

[Ano65-468] Anonymous. Transients and trends. IEEE Spectrum, 2(6):9, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTg

[Ano65-469] Anonymous. Transients and trends. IEEE Spectrum, 2(7):7, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTh

[Ano65-470] Anonymous. Transients and trends. IEEE Spectrum, 2(8):7, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTi

[Ano65-471] Anonymous. Transients and trends. IEEE Spectrum, 2 (9):9, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTj

[Ano65-472] Anonymous. Transients and trends. IEEE Spectrum, 2 (10):9, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTk

[Ano65-473] Anonymous. Transients and trends. IEEE Spectrum, 2(11): 9–11, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TTl

[Ano65-474] Anonymous. Transients and trends. IEEE Spectrum, 2 (12):9, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJa

[Ano65-475] Anonymous. Translated journals. IEEE Spectrum, 2(1): 142–145, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJb

[Ano65-476] Anonymous. Translated journals. IEEE Spectrum, 2(2): 149–152, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJc

[Ano65-477] Anonymous. Translated journals. IEEE Spectrum, 2(3): 244–248, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJd

[Ano65-478] Anonymous. Translated journals. IEEE Spectrum, 2(4): 170–172, April 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJe

[Ano65-479] Anonymous. Translated journals. IEEE Spectrum, 2(5): 160–163, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJf

[Ano65-480] Anonymous. Translated journals. IEEE Spectrum, 2(6): 138–141, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJg

[Ano65-481] Anonymous. Translated journals. IEEE Spectrum, 2(7): 172–175, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJh

[Ano65-482] Anonymous. Translated journals. IEEE Spectrum, 2(8): 148–153, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TJi

[Ano65-483] Anonymous. Translated journals. IEEE Spectrum, 2(9): 164–173, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:TSE

[Ano65-484] Anonymous. Tutorial symposium on electron devices to be held. IEEE Spectrum, 2(8):126, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:UCL

[Ano65-485] Anonymous. University of California links medicine, engineering. IEEE Spectrum, 2(9): 20, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:WKL

[Ano65-486] Anonymous. W. K. Lewis designated John Fritz medalist for 1966. IEEE Spectrum, 2(8):127, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:WYP

[Ano65-487] Anonymous. Was your 1954 project this successful? [advertisement]. IEEE Spectrum, 2(2):43, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:WEL

[Ano65-488] Anonymous. WESCON/65 Exhibitors list. IEEE Spectrum, 2(7):32–60, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1965:XGM

[Ano65-489] Anonymous. XXXth General Meeting of the International Electrotechnical Commission. IEEE Spectrum, 2(9): 48–49, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IIC

[Ano66a] Anonymous. 1966 IEEE International Convention Record. IEEE Spectrum, 3(2):6, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IWP

[Ano66b] Anonymous. 1966 IEEE Winter Power Meeting. IEEE Spectrum, 3(5):92–95, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:DGH

[Ano66c] Anonymous. A 3-D glimpse of the hearing process. IEEE Spectrum, 3(11):8, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAa

[Ano66d] Anonymous. Advance abstracts. IEEE Spectrum, 3(1): 110–128, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAb

[Ano66e] Anonymous. Advance abstracts. IEEE Spectrum, 3(2): 126–143, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAc

[Ano66f] Anonymous. Advance abstracts.

IEEE Spectrum, 3(3):186–202, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAd

[Ano66g] Anonymous. Advance abstracts.

IEEE Spectrum, 3(4):122–137, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAe

[Ano66h] Anonymous. Advance abstracts.

IEEE Spectrum, 3(5):98–110, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAf

[Ano66i] Anonymous. Advance abstracts.

IEEE Spectrum, 3(6):112–132, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAg

[Ano66j] Anonymous. Advance abstracts.

IEEE Spectrum, 3(7):125–140, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAh

[Ano66k] Anonymous. Advance abstracts. IEEE Spectrum, 3(8):120–149, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAi

[Ano66l]  Anonymous.  Advance ab-

stracts. IEEE Spectrum, 3(9): 152–169, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAj

[Ano66m] Anonymous. Advance abstracts. IEEE Spectrum, 3(10): 124–143, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAk

[Ano66n] Anonymous. Advance abstracts. IEEE Spectrum, 3(11): 96–106, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AAl

[Ano66o] Anonymous. Advance abstracts. IEEE Spectrum, 3(12): 88–113, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aa

[Ano66p] Anonymous. [advertisement]. IEEE Spectrum, 3(1):1, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ab

[Ano66q] Anonymous. [advertisement].

IEEE Spectrum, 3(1):3, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ac

[Ano66r] Anonymous. [advertisement].

IEEE Spectrum, 3(1):6–8, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ad

[Ano66s] Anonymous. [advertisement].

IEEE Spectrum, 3(1):13–29, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ae

[Ano66t] Anonymous. [advertisement].

IEEE Spectrum, 3(1):35, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ag

[Ano66u] Anonymous. [advertisement].

IEEE Spectrum, 3(1):107, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ah

[Ano66v] Anonymous. [advertisement].

IEEE Spectrum, 3(1):111–129, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ai

[Ano66w] Anonymous. [advertisement]. IEEE Spectrum, 3(1):137–147, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aj

[Ano66x] Anonymous. [advertisement]. IEEE Spectrum, 3(1):149–158, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ak

[Ano66y] Anonymous. [advertisement]. IEEE Spectrum, 3(1):2, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Al

[Ano66z] Anonymous. [advertisement]. IEEE Spectrum, 3(2):1, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Am

[Ano66-27] Anonymous. [advertisement]. IEEE Spectrum, 3(2):3, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:An

[Ano66-28] Anonymous. [advertisement]. IEEE Spectrum, 3(2):10–11, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ao

[Ano66-29] Anonymous. [advertisement].

IEEE Spectrum, 3(2):13–38, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ap

[Ano66-30] Anonymous. [advertisement].

IEEE Spectrum, 3(2):39–50, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ar

[Ano66-31] Anonymous. [advertisement].

IEEE Spectrum, 3(2):81–83, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:As

[Ano66-32] Anonymous. [advertisement].

IEEE Spectrum, 3(2):131, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:At

[Ano66-33] Anonymous. [advertisement].

IEEE Spectrum, 3(2):144–149, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Au

[Ano66-34] Anonymous. [advertisement].

IEEE Spectrum, 3(2):163, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Av

[Ano66-35] Anonymous. [advertisement]. IEEE Spectrum, 3(2):2, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aw

[Ano66-36] Anonymous. [advertisement]. IEEE Spectrum, 3(3):1, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ax

[Ano66-37] Anonymous. [advertisement]. IEEE Spectrum, 3(3):3, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ay

[Ano66-38] Anonymous. [advertisement]. IEEE Spectrum, 3(3):6–7, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Az

[Ano66-39] Anonymous. [advertisement]. IEEE Spectrum, 3(3):11–19, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aba

[Ano66-40] Anonymous. [advertisement]. IEEE Spectrum, 3(3):21–23, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abb

[Ano66-41] Anonymous. [advertisement].

IEEE Spectrum, 3(3):32, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abc

[Ano66-42] Anonymous. [advertisement].

IEEE Spectrum, 3(3):37–112, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abe

[Ano66-43] Anonymous. [advertisement].

IEEE Spectrum, 3(3):143–146, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abf

[Ano66-44] Anonymous. [advertisement].

IEEE Spectrum, 3(3):181, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abg

[Ano66-45] Anonymous. [advertisement].

IEEE Spectrum, 3(3):189–203, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abh

[Ano66-46] Anonymous. [advertisement].

IEEE Spectrum, 3(3):205–209, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abi

[Ano66-47] Anonymous. [advertisement]. IEEE Spectrum, 3(3):211, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abj

[Ano66-48] Anonymous. [advertisement]. IEEE Spectrum, 3(3):213–219, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abk

[Ano66-49] Anonymous. [advertisement]. IEEE Spectrum, 3(3):221–229, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abl

[Ano66-50] Anonymous. [advertisement]. IEEE Spectrum, 3(3):230–303, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abm

[Ano66-51] Anonymous. [advertisement]. IEEE Spectrum, 3(3):2, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abn

[Ano66-52] Anonymous. [advertisement]. IEEE Spectrum, 3(4):1, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abo

[Ano66-53] Anonymous. [advertisement].

IEEE Spectrum, 3(4):3, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abp

[Ano66-54] Anonymous. [advertisement].

IEEE Spectrum, 3(4):6–7, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abq

[Ano66-55] Anonymous. [advertisement].

IEEE Spectrum, 3(4):10–11, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abr

[Ano66-56] Anonymous. [advertisement].

IEEE Spectrum, 3(4):13–31, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abs

[Ano66-57] Anonymous. [advertisement].

IEEE Spectrum, 3(4):34–39, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abt

[Ano66-58] Anonymous. [advertisement].

IEEE Spectrum, 3(4):41–45, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abu

[Ano66-59] Anonymous. [advertisement]. IEEE Spectrum, 3(4):50–52, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abw

[Ano66-60] Anonymous. [advertisement]. IEEE Spectrum, 3(4):119, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abx

[Ano66-61] Anonymous. [advertisement]. IEEE Spectrum, 3(4):125–136, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aby

[Ano66-62] Anonymous. [advertisement]. IEEE Spectrum, 3(4):139–141, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abz

[Ano66-63] Anonymous. [advertisement]. IEEE Spectrum, 3(4):149–152, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aca

[Ano66-64] Anonymous. [advertisement]. IEEE Spectrum, 3(4):155, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acb

[Ano66-65] Anonymous. [advertisement].

IEEE Spectrum, 3(4):157–161, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acc

[Ano66-66] Anonymous. [advertisement].

IEEE Spectrum, 3(4):163–165, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acd

[Ano66-67] Anonymous. [advertisement].

IEEE Spectrum, 3(4):2, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ace

[Ano66-68] Anonymous. [advertisement].

IEEE Spectrum, 3(5):1, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acf

[Ano66-69] Anonymous. [advertisement].

IEEE Spectrum, 3(5):3, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acg

[Ano66-70] Anonymous. [advertisement].

IEEE Spectrum, 3(5):11–25, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ach

[Ano66-71] Anonymous. [advertisement]. IEEE Spectrum, 3(5):27–31, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aci

[Ano66-72] Anonymous. [advertisement]. IEEE Spectrum, 3(5):33–40, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ack

[Ano66-73] Anonymous. [advertisement]. IEEE Spectrum, 3(5):101–112, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acl

[Ano66-74] Anonymous. [advertisement]. IEEE Spectrum, 3(5):114–121, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acm

[Ano66-75] Anonymous. [advertisement]. IEEE Spectrum, 3(5):123, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acn

[Ano66-76] Anonymous. [advertisement]. IEEE Spectrum, 3(5):125–136, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aco

[Ano66-77] Anonymous. [advertisement].

IEEE Spectrum, 3(5):138–145, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acp

[Ano66-78] Anonymous. [advertisement].

IEEE Spectrum, 3(5):2, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acq

[Ano66-79] Anonymous. [advertisement].

IEEE Spectrum, 3(6):1, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acr

[Ano66-80] Anonymous. [advertisement].

IEEE Spectrum, 3(6):3, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acs

[Ano66-81] Anonymous. [advertisement].

IEEE Spectrum, 3(6):6–7, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Act

[Ano66-82] Anonymous. [advertisement].

IEEE Spectrum, 3(6):11–23, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acu

[Ano66-83] Anonymous. [advertisement]. IEEE Spectrum, 3(6):25–29, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acv

[Ano66-84] Anonymous. [advertisement]. IEEE Spectrum, 3(6):31–40, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acx

[Ano66-85] Anonymous. [advertisement]. IEEE Spectrum, 3(6):79–82, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acy

[Ano66-86] Anonymous. [advertisement]. IEEE Spectrum, 3(6):115–133, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acz

[Ano66-87] Anonymous. [advertisement]. IEEE Spectrum, 3(6):135, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ada

[Ano66-88] Anonymous. [advertisement]. IEEE Spectrum, 3(6):143–145, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adb

[Ano66-89] Anonymous. [advertisement].

IEEE Spectrum, 3(6):147, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adc

[Ano66-90] Anonymous. [advertisement].

IEEE Spectrum, 3(6):151–157, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Add

[Ano66-91] Anonymous. [advertisement].

IEEE Spectrum, 3(6):2, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ade

[Ano66-92] Anonymous. [advertisement].

IEEE Spectrum, 3(7):1, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adf

[Ano66-93] Anonymous. [advertisement].

IEEE Spectrum, 3(7):3, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adg

[Ano66-94] Anonymous. [advertisement].

IEEE Spectrum, 3(7):9–15, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adh

[Ano66-95] Anonymous. [advertisement]. IEEE Spectrum, 3(7):17–38, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adi

[Ano66-96] Anonymous. [advertisement]. IEEE Spectrum, 3(7):41, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adj

[Ano66-97] Anonymous. [advertisement]. IEEE Spectrum, 3(7):43, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adk

[Ano66-98] Anonymous. [advertisement]. IEEE Spectrum, 3(7):47–54, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adm

[Ano66-99] Anonymous. [advertisement]. IEEE Spectrum, 3(7):89–92, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adn

[Ano66-100] Anonymous. [advertisement]. IEEE Spectrum, 3(7):124, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ado

[Ano66-101] Anonymous. [advertisement].

IEEE Spectrum, 3(7):127–141, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adp

[Ano66-102] Anonymous. [advertisement].

IEEE Spectrum, 3(7):143–145, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adq

[Ano66-103] Anonymous. [advertisement].

IEEE Spectrum, 3(7):149–153, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adr

[Ano66-104] Anonymous. [advertisement].

IEEE Spectrum, 3(7):155–158, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ads

[Ano66-105] Anonymous. [advertisement].

IEEE Spectrum, 3(7):161, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adt

[Ano66-106] Anonymous. [advertisement].

IEEE Spectrum, 3(7):164–177, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adu

[Ano66-107] Anonymous. [advertisement]. IEEE Spectrum, 3(7):2, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adv

[Ano66-108] Anonymous. [advertisement]. IEEE Spectrum, 3(8):1, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adw

[Ano66-109] Anonymous. [advertisement]. IEEE Spectrum, 3(8):3, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adx

[Ano66-110] Anonymous. [advertisement]. IEEE Spectrum, 3(8):6–7, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ady

[Ano66-111] Anonymous. [advertisement]. IEEE Spectrum, 3(8):11–19, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adz

[Ano66-112] Anonymous. [advertisement]. IEEE Spectrum, 3(8):21–36, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aea

[Ano66-113] Anonymous. [advertisement].

IEEE Spectrum, 3(8):39–40, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aeb

[Ano66-114] Anonymous. [advertisement].

IEEE Spectrum, 3(8):43–50, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aed

[Ano66-115] Anonymous. [advertisement].

IEEE Spectrum, 3(8):121–149, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aee

[Ano66-116] Anonymous. [advertisement].

IEEE Spectrum, 3(8):151–154, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aef

[Ano66-117] Anonymous. [advertisement].

IEEE Spectrum, 3(8):157–159, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aeg

[Ano66-118] Anonymous. [advertisement].

IEEE Spectrum, 3(8):163–165, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aeh

[Ano66-119] Anonymous. [advertisement]. IEEE Spectrum, 3(8):167–171, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aei

[Ano66-120] Anonymous. [advertisement]. IEEE Spectrum, 3(8):2, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aej

[Ano66-121] Anonymous. [advertisement]. IEEE Spectrum, 3(9):1, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aek

[Ano66-122] Anonymous. [advertisement]. IEEE Spectrum, 3(9):3, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ael

[Ano66-123] Anonymous. [advertisement]. IEEE Spectrum, 3(9):6–7, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aem

[Ano66-124] Anonymous. [advertisement]. IEEE Spectrum, 3(9):10–11, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aen

[Ano66-125] Anonymous. [advertisement].

IEEE Spectrum, 3(9):14–24, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aeo

[Ano66-126] Anonymous. [advertisement].

IEEE Spectrum, 3(9):29–40, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aep

[Ano66-127] Anonymous. [advertisement].

IEEE Spectrum, 3(9):43, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aeq

[Ano66-128] Anonymous. [advertisement].

IEEE Spectrum, 3(9):45–56, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aer

[Ano66-129] Anonymous. [advertisement].

IEEE Spectrum, 3(9):63–74, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aet

[Ano66-130] Anonymous. [advertisement].

IEEE Spectrum, 3(9):113–116, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aeu

[Ano66-131] Anonymous. [advertisement]. IEEE Spectrum, 3(9):159, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aev

[Ano66-132] Anonymous. [advertisement]. IEEE Spectrum, 3(9):175, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aew

[Ano66-133] Anonymous. [advertisement]. IEEE Spectrum, 3(9):179, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aex

[Ano66-134] Anonymous. [advertisement]. IEEE Spectrum, 3(9):187–192, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aey

[Ano66-135] Anonymous. [advertisement]. IEEE Spectrum, 3(9):195–197, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aez

[Ano66-136] Anonymous. [advertisement]. IEEE Spectrum, 3(9):217–219, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afa

[Ano66-137] Anonymous. [advertisement].

IEEE Spectrum, 3(9):221, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afb

[Ano66-138] Anonymous. [advertisement].

IEEE Spectrum, 3(9):2, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afc

[Ano66-139] Anonymous. [advertisement].

IEEE Spectrum, 3(10):1–3, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afd

[Ano66-140] Anonymous. [advertisement].

IEEE Spectrum, 3(10):3, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afe

[Ano66-141] Anonymous. [advertisement].

IEEE Spectrum, 3(10):6–7, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aff

[Ano66-142] Anonymous. [advertisement].

IEEE Spectrum, 3(10):11–15, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afg

[Ano66-143] Anonymous. [advertisement]. IEEE Spectrum, 3(10):17–38, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afh

[Ano66-144] Anonymous. [advertisement]. IEEE Spectrum, 3(10):41, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afi

[Ano66-145] Anonymous. [advertisement]. IEEE Spectrum, 3(10):43–54, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afk

[Ano66-146] Anonymous. [advertisement]. IEEE Spectrum, 3(10):85–88, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afl

[Ano66-147] Anonymous. [advertisement]. IEEE Spectrum, 3(10):121, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afm

[Ano66-148] Anonymous. [advertisement]. IEEE Spectrum, 3(10):127– 139, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afn

[Ano66-149] Anonymous. [advertisement].

IEEE Spectrum, 3(10):145, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afo

[Ano66-150] Anonymous. [advertisement].

IEEE Spectrum, 3(10):148, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afp

[Ano66-151] Anonymous. [advertisement].

IEEE Spectrum, 3(10):153, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afq

[Ano66-152] Anonymous. [advertisement].

IEEE Spectrum, 3(10):155– 157, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afr

[Ano66-153] Anonymous. [advertisement].

IEEE Spectrum, 3(10):159– 167, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afs

[Ano66-154] Anonymous. [advertisement].

IEEE Spectrum, 3(10):4, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aft

[Ano66-155] Anonymous. [advertisement]. IEEE Spectrum, 3(11):1, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afu

[Ano66-156] Anonymous. [advertisement]. IEEE Spectrum, 3(11):3, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afv

[Ano66-157] Anonymous. [advertisement]. IEEE Spectrum, 3(11):6–7, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afw

[Ano66-158] Anonymous. [advertisement]. IEEE Spectrum, 3(11):11, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afx

[Ano66-159] Anonymous. [advertisement]. IEEE Spectrum, 3(11):15–26, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afy

[Ano66-160] Anonymous. [advertisement]. IEEE Spectrum, 3(11):29, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afz

[Ano66-161] Anonymous. [advertisement].

IEEE Spectrum, 3(11):31–38, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agb

[Ano66-162] Anonymous. [advertisement].

IEEE Spectrum, 3(11):69–72, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agc

[Ano66-163] Anonymous. [advertisement].

IEEE Spectrum, 3(11):97–107, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agd

[Ano66-164] Anonymous. [advertisement].

IEEE Spectrum, 3(11):109–111, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Age

[Ano66-165] Anonymous. [advertisement].

IEEE Spectrum, 3(11):115–121, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agf

[Ano66-166] Anonymous. [advertisement].

IEEE Spectrum, 3(11):123, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agg

[Ano66-167] Anonymous. [advertisement]. IEEE Spectrum, 3(11):125–137, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agh

[Ano66-168] Anonymous. [advertisement]. IEEE Spectrum, 3(11):2, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agi

[Ano66-169] Anonymous. [advertisement]. IEEE Spectrum, 3(12):1, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agj

[Ano66-170] Anonymous. [advertisement]. IEEE Spectrum, 3(12):3, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agk

[Ano66-171] Anonymous. [advertisement]. IEEE Spectrum, 3(12):6–7, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agl

[Ano66-172] Anonymous. [advertisement]. IEEE Spectrum, 3(12):11, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agm

[Ano66-173] Anonymous. [advertisement].

IEEE Spectrum, 3(12):13–20, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agn

[Ano66-174] Anonymous. [advertisement].

IEEE Spectrum, 3(12):23–24, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ago

[Ano66-175] Anonymous. [advertisement].

IEEE Spectrum, 3(12):27–32, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agq

[Ano66-176] Anonymous. [advertisement].

IEEE Spectrum, 3(12):71–74, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agr

[Ano66-177] Anonymous. [advertisement].

IEEE Spectrum, 3(12):89–109, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ags

[Ano66-178] Anonymous. [advertisement].

IEEE Spectrum, 3(12):122–123, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agt

[Ano66-179] Anonymous. [advertisement]. IEEE Spectrum, 3(12):127, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agu

[Ano66-180] Anonymous. [advertisement]. IEEE Spectrum, 3(12):2, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:APP

[Ano66-181] Anonymous. Advertisement — pergamon press. IEEE Spectrum, 3(3):62, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:AIa

[Ano66-182] Anonymous. Annual index. IEEE Spectrum, 3(12):128–142, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Af

[Ano66-183] Anonymous. Authors. IEEE Spectrum, 3(1):105–106, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aq

[Ano66-184] Anonymous. Authors. IEEE Spectrum, 3(2):52–53, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abd

[Ano66-185] Anonymous. Authors. IEEE Spectrum, 3(3):114–115, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Abv

[Ano66-186] Anonymous. Authors. IEEE Spectrum, 3(4):54–55, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acj

[Ano66-187] Anonymous. Authors. IEEE Spectrum, 3(5):42, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Acw

[Ano66-188] Anonymous. Authors. IEEE Spectrum, 3(6):47–48, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Adl

[Ano66-189] Anonymous. Authors. IEEE Spectrum, 3(7):56–57, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aec

[Ano66-190] Anonymous. Authors. IEEE Spectrum, 3(8):52–53, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aes

[Ano66-191] Anonymous. Authors. IEEE Spectrum, 3(9):76–78, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Afj

[Ano66-192] Anonymous. Authors. IEEE Spectrum, 3(10):56–58, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Aga

[Ano66-193] Anonymous. Authors. IEEE Spectrum, 3(11):40–42, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Agp

[Ano66-194] Anonymous. Authors. IEEE Spectrum, 3(12):34, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BLT

[Ano66-195] Anonymous. “Beam lead” technique for fabricating solid-state devices. IEEE Spectrum, 3(7): 6, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRa

[Ano66-196] Anonymous. Book reviews. IEEE Spectrum, 3(1):148–155, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRb

[Ano66-197] Anonymous. Book reviews.

IEEE Spectrum, 3(2):156–161, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRc

[Ano66-198] Anonymous. Book reviews.

IEEE Spectrum, 3(3):220–229, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRd

[Ano66-199] Anonymous. Book reviews.

IEEE Spectrum, 3(4):162–164, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRe

[Ano66-200] Anonymous. Book reviews.

IEEE Spectrum, 3(5):137–144, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRf

[Ano66-201] Anonymous. Book reviews.

IEEE Spectrum, 3(6):150–157, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRg

[Ano66-202] Anonymous. Book reviews.

IEEE Spectrum, 3(7):159–163, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRh

[Ano66-203] Anonymous. Book reviews. IEEE Spectrum, 3(8):166–170, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRi

[Ano66-204] Anonymous. Book reviews. IEEE Spectrum, 3(9):186–190, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRj

[Ano66-205] Anonymous. Book reviews. IEEE Spectrum, 3(10):158– 167, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRk

[Ano66-206] Anonymous. Book reviews. IEEE Spectrum, 3(11):124–135, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BRl

[Ano66-207] Anonymous. Book reviews. IEEE Spectrum, 3(12):123–127, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPa

[Ano66-208] Anonymous. [breaker page]. IEEE Spectrum, 3(1):4, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPb

[Ano66-209] Anonymous. [breaker page].

IEEE Spectrum, 3(2):4, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPc

[Ano66-210] Anonymous. [breaker page].

IEEE Spectrum, 3(3):4, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPd

[Ano66-211] Anonymous. [breaker page].

IEEE Spectrum, 3(4):4, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPe

[Ano66-212] Anonymous. [breaker page].

IEEE Spectrum, 3(5):4, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPf

[Ano66-213] Anonymous. [breaker page].

IEEE Spectrum, 3(6):4, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPg

[Ano66-214] Anonymous. [breaker page].

IEEE Spectrum, 3(7):4, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPh

[Ano66-215] Anonymous. [breaker page]. IEEE Spectrum, 3(8):4, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPi

[Ano66-216] Anonymous. [breaker page]. IEEE Spectrum, 3(9):4, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPj

[Ano66-217] Anonymous. [breaker page]. IEEE Spectrum, 3(10):4, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPk

[Ano66-218] Anonymous. [breaker page]. IEEE Spectrum, 3(11):4, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:BPl

[Ano66-219] Anonymous. [breaker page]. IEEE Spectrum, 3(12):4, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ca

[Ano66-220] Anonymous. Calendar. IEEE Spectrum, 3(1):30–32, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cb

[Ano66-221] Anonymous. Calendar. IEEE Spectrum, 3(2):12–25, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cc

[Ano66-222] Anonymous. Calendar. IEEE Spectrum, 3(2):28–34, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cd

[Ano66-223] Anonymous. Calendar. IEEE Spectrum, 3(3):20–24, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ce

[Ano66-224] Anonymous. Calendar. IEEE Spectrum, 3(4):33–38, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cg

[Ano66-225] Anonymous. Calendar. IEEE Spectrum, 3(5):26–30, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ch

[Ano66-226] Anonymous. Calendar. IEEE Spectrum, 3(6):24–28, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cj

[Ano66-227] Anonymous. Calendar. IEEE Spectrum, 3(7):39–41, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ck

[Ano66-228] Anonymous. Calendar. IEEE Spectrum, 3(8):37–39, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cl

[Ano66-229] Anonymous. Calendar. IEEE Spectrum, 3(9):41–43, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cm

[Ano66-230] Anonymous. Calendar. IEEE Spectrum, 3(10):39–41, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cn

[Ano66-231] Anonymous. Calendar. IEEE Spectrum, 3(11):27–29, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Co

[Ano66-232] Anonymous. Calendar. IEEE Spectrum, 3(12):21–23, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:CAa

[Ano66-233] Anonymous. Classified advertising. IEEE Spectrum, 3(3): 184–185, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:CAb

[Ano66-234] Anonymous. Classified advertising. IEEE Spectrum, 3(9): 149, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:CNP

[Ano66-235] Anonymous. Comments on the northeast power failure. IEEE Spectrum, 3(5):84–90, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Cf

[Ano66-236] Anonymous. Correspondence.

IEEE Spectrum, 3(4):156–161, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ci

[Ano66-237] Anonymous. Correspondence.

IEEE Spectrum, 3(6):146–149, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:DYU

[Ano66-238] Anonymous. Do you use your engineering societies library as much as you might? IEEE Spectrum, 3(5):80–83, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:EPNa

[Ano66-239] Anonymous. Electronic product news. IEEE Spectrum, 3(3):25, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:EPNb

[Ano66-240] Anonymous. Electronic product news. IEEE Spectrum, 3(5):12, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:EPNc

[Ano66-241] Anonymous. Electronic product news. IEEE Spectrum, 3(7):45– 52, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:EPNd

[Ano66-242] Anonymous. Electronic product news. IEEE Spectrum, 3 (9):13, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPa

[Ano66-243] Anonymous. Focal points. IEEE Spectrum, 3(1):135–142, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPb

[Ano66-244] Anonymous. Focal points. IEEE Spectrum, 3(2):150–155, February 1966. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPc

[Ano66-245] Anonymous. Focal points. IEEE Spectrum, 3(3):212–218, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPd

[Ano66-246] Anonymous. Focal points. IEEE Spectrum, 3(4):144–152, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPe

[Ano66-247] Anonymous. Focal points. IEEE Spectrum, 3(5):124–134, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPf

[Ano66-248] Anonymous. Focal points. IEEE Spectrum, 3(6):140–145, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPg

[Ano66-249] Anonymous. Focal points. IEEE Spectrum, 3(7):147–152, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPh

[Ano66-250] Anonymous. Focal points. IEEE Spectrum, 3(8):156–161, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPi

[Ano66-251] Anonymous. Focal points. IEEE Spectrum, 3(9):176–181, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPj

[Ano66-252] Anonymous. Focal points. IEEE Spectrum, 3(10):149–152, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPk

[Ano66-253] Anonymous. Focal points. IEEE Spectrum, 3(11):114–120, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FPl

[Ano66-254] Anonymous. Focal points. IEEE Spectrum, 3(12):118–120, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FRR

[Ano66-255] Anonymous. Found: A rapid route to the shortest path. IEEE Spectrum, 3(10):8, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCS

[Ano66-256] Anonymous. Fourier coefficients and such. IEEE Spectrum, 3(8):5, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCa

[Ano66-257] Anonymous. [front cover]. IEEE Spectrum, 3(1):c1, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCb

[Ano66-258] Anonymous. [front cover]. IEEE Spectrum, 3(1):c2, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCc

[Ano66-259] Anonymous. [front cover].

IEEE Spectrum, 3(2):c1, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCd

[Ano66-260] Anonymous. [front cover].

IEEE Spectrum, 3(2):c2, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCe

[Ano66-261] Anonymous. [front cover]. IEEE Spectrum, 3(3):c1, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCf

[Ano66-262] Anonymous. [front cover]. IEEE Spectrum, 3(3):c2, March 1966.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCg

[Ano66-263] Anonymous. [front cover]. IEEE Spectrum, 3(4):c1, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCh

[Ano66-264] Anonymous. [front cover]. IEEE Spectrum, 3(4):c2, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCi

[Ano66-265] Anonymous. [front cover]. IEEE Spectrum, 3(5):c1, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCj

[Ano66-266] Anonymous. [front cover]. IEEE Spectrum, 3(5):c2, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCk

[Ano66-267] Anonymous. [front cover]. IEEE Spectrum, 3(6):c1, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCl

[Ano66-268] Anonymous. [front cover]. IEEE Spectrum, 3(6):c2, June 1966.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCm

[Ano66-269] Anonymous. [front cover]. IEEE Spectrum, 3(7):c1, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCn

[Ano66-270] Anonymous. [front cover]. IEEE Spectrum, 3(7):c2, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCo

[Ano66-271] Anonymous. [front cover]. IEEE Spectrum, 3(8):c1, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCp

[Ano66-272] Anonymous. [front cover]. IEEE Spectrum, 3(8):c2, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCq

[Ano66-273] Anonymous. [front cover]. IEEE Spectrum, 3(9):c1, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCr

[Ano66-274] Anonymous. [front cover]. IEEE Spectrum, 3(9):c2, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCt

[Ano66-275] Anonymous. [front cover]. IEEE Spectrum, 3(10):c1, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCu

[Ano66-276] Anonymous. [front cover]. IEEE Spectrum, 3(10):c2, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCv

[Ano66-277] Anonymous. [front cover]. IEEE Spectrum, 3(11):c1, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCw

[Ano66-278] Anonymous. [front cover]. IEEE Spectrum, 3(11):c2, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCx

[Ano66-279] Anonymous. [front cover]. IEEE Spectrum, 3(12):c1, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:FCy

[Ano66-280] Anonymous. [front cover]. IEEE Spectrum, 3(12):c2, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:HPA

[Ano66-281] Anonymous. High-performance amplifiers for coaxial transmission. IEEE Spectrum, 3(9): 8, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:HPT

[Ano66-282] Anonymous. High-purity thin films. IEEE Spectrum, 3(5):6, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:HWC

[Ano66-283] Anonymous. How we compressed a year into two minutes. IEEE Spectrum, 3(2):84, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IFAa

[Ano66-284] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 3(7):8–15, July 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IFAb

[Ano66-285] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 3 (8):10–19, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IFAc

[Ano66-286] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 3 (9):12–22, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IFAd

[Ano66-287] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 3 (10):10–14, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IFAe

[Ano66-288] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 3 (11):10, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IFAf

[Ano66-289] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 3 (12):10, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IJD

[Ano66-290] Anonymous. IEEE Japan directory 1966–67. IEEE Spectrum, 3 (9):57–199, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IP

[Ano66-291] Anonymous. IEEE publications. IEEE Spectrum, 3(6):110–111, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IRP

[Ano66-292] Anonymous. IEEE recommended practice for units in published scientific and technical work. IEEE Spectrum, 3 (3):169–173, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:ITP

[Ano66-293] Anonymous. IEEE Technical Program. IEEE Spectrum, 3 (3):33–110, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAa

[Ano66-294] Anonymous. Index to advertisers. IEEE Spectrum, 3(1): 156–158, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAb

[Ano66-295] Anonymous. Index to advertisers. IEEE Spectrum, 3(2): 162–164, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAc

[Ano66-296] Anonymous. Index to advertisers. IEEE Spectrum, 3(3): 310–312, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAd

[Ano66-297] Anonymous. Index to advertisers. IEEE Spectrum, 3(4): 166–168, April 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAe

[Ano66-298] Anonymous. Index to advertisers. IEEE Spectrum, 3 (5):146–148, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAf

[Ano66-299] Anonymous. Index to advertisers. IEEE Spectrum, 3 (6):158–160, June 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAg

[Ano66-300] Anonymous. Index to advertisers. IEEE Spectrum, 3 (7):178–180, July 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAh

[Ano66-301] Anonymous. Index to advertisers. IEEE Spectrum, 3(8): 172–174, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAi

[Ano66-302] Anonymous. Index to advertisers. IEEE Spectrum, 3(9): 222–224, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAj

[Ano66-303] Anonymous. Index to advertisers. IEEE Spectrum, 3(10):

168–170, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAk

[Ano66-304] Anonymous. Index to advertisers. IEEE Spectrum, 3(11): 138–140, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAl

[Ano66-305] Anonymous. Index to advertisers. IEEE Spectrum, 3(12): 143–144, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IAS

[Ano66-306] Anonymous. Information about the society. IEEE Spectrum, 3 (3):304–307, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IC

[Ano66-307] Anonymous. Instant careers.

IEEE Spectrum, 3(4):153–154, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:ICA

[Ano66-308] Anonymous. Instant careers [advertisement]. IEEE Spectrum, 3(6):109, June 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:ICM

[Ano66-309] Anonymous. Integrated circuits at microwave frequencies. IEEE Spectrum, 3(1):15, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEb

[Ano66-310] Anonymous. International exhibition of electronic components exhibitors list [advertisement]. IEEE Spectrum, 3(1):42–50, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:ICB

[Ano66-311] Anonymous. Interstellar communication is bunk. IEEE Spectrum, 3(1):5, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:JES

[Ano66-312] Anonymous. Japan Electronics Show 1966. IEEE Spectrum, 3(9):200–215, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:LQO

[Ano66-313] Anonymous. Lodge on “Quasi-Optics”. IEEE Spectrum, 3(4):5, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:LTM

[Ano66-314] Anonymous. Low-temperature millimeter and submillimeter wave antenna noise temperature definition. IEEE Spectrum, 3(5): 5, May 1966. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NTM

[Ano66-315] Anonymous. New type of microscopy reveals internal structures of crystals. IEEE Spectrum, 3(4):8, April 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NW

[Ano66-316] Anonymous. New worlds. IEEE Spectrum, 3(3):5, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIa

[Ano66-317] Anonymous. News of the IEEE.

IEEE Spectrum, 3(1):10–28, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIb

[Ano66-318] Anonymous. News of the IEEE.

IEEE Spectrum, 3(3):10–17, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIc

[Ano66-319] Anonymous. News of the IEEE.

IEEE Spectrum, 3(4):12–32, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NId

[Ano66-320] Anonymous. News of the IEEE.

IEEE Spectrum, 3(5):8–24, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIe

[Ano66-321] Anonymous. News of the IEEE. IEEE Spectrum, 3(6):10–22, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIf

[Ano66-322] Anonymous. News of the IEEE. IEEE Spectrum, 3(7):16–37, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIg

[Ano66-323] Anonymous. News of the IEEE. IEEE Spectrum, 3(8):20–34, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIh

[Ano66-324] Anonymous. News of the IEEE. IEEE Spectrum, 3(9): 25–38, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIi

[Ano66-325] Anonymous. News of the IEEE. IEEE Spectrum, 3(10): 16–36, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIj

[Ano66-326] Anonymous. News of the IEEE. IEEE Spectrum, 3(11):

12–25, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:NIk

[Ano66-327] Anonymous. News of the IEEE. IEEE Spectrum, 3(12): 12–18, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Oa

[Ano66-328] Anonymous. Obituaries. IEEE Spectrum, 3(1):39–41, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ob

[Ano66-329] Anonymous. Obituaries. IEEE Spectrum, 3(2):48–49, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Oc

[Ano66-330] Anonymous. Obituaries. IEEE Spectrum, 3(3):29–31, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Od

[Ano66-331] Anonymous. Obituaries. IEEE Spectrum, 3(4):47–49, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Oe

[Ano66-332] Anonymous. Obituaries. IEEE Spectrum, 3(6):38–39, June 1966.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Of

[Ano66-333] Anonymous. Obituaries. IEEE Spectrum, 3(7):52–53, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Og

[Ano66-334] Anonymous. Obituaries. IEEE Spectrum, 3(8):48–49, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Oh

[Ano66-335] Anonymous. Obituaries. IEEE Spectrum, 3(9):54–55, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Oi

[Ano66-336] Anonymous. Obituaries. IEEE Spectrum, 3(10):53, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Oj

[Ano66-337] Anonymous. Obituaries. IEEE Spectrum, 3(11):37, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ok

[Ano66-338] Anonymous. Obituaries. IEEE Spectrum, 3(12):31, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:JNS

[Ano66-339] Anonymous. On joining “nonstick” plastics. IEEE Spectrum, 3(12):8, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pa

[Ano66-340] Anonymous. People. IEEE Spectrum, 3(1):33–39, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pb

[Ano66-341] Anonymous. People. IEEE Spectrum, 3(2):38–48, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pc

[Ano66-342] Anonymous. People. IEEE Spectrum, 3(3):26–29, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pd

[Ano66-343] Anonymous. People. IEEE Spectrum, 3(4):40–47, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pe

[Ano66-344] Anonymous. People. IEEE Spectrum, 3(5):32–39, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pf

[Ano66-345] Anonymous. People. IEEE Spectrum, 3(6):30–38, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pg

[Ano66-346] Anonymous. People. IEEE Spectrum, 3(7):42–44, July 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Ph

[Ano66-347] Anonymous. People. IEEE Spectrum, 3(8):41–48, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pi

[Ano66-348] Anonymous. People. IEEE Spectrum, 3(9):44–54, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pj

[Ano66-349] Anonymous. People. IEEE Spectrum, 3(10):42–53, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pk

[Ano66-350] Anonymous. People. IEEE Spectrum, 3(11):30–37, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:Pl

[Ano66-351] Anonymous. People. IEEE Spectrum, 3(12):25–31, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PRP

[Ano66-352] Anonymous. The period of rotation of the planet mercury. IEEE Spectrum, 3(11):5, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PMA

[Ano66-353] Anonymous. The planet Mars its atmosphere and ionosphere. IEEE Spectrum, 3(12):5, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSa

[Ano66-354] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(2):164, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSb

[Ano66-355] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(3):308–309, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSc

[Ano66-356] Anonymous. Professional consulting services [advertisements].

IEEE Spectrum, 3(4):167–168, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSd

[Ano66-357] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(5):147–148, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSe

[Ano66-358] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(6):159–160, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSf

[Ano66-359] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(7):179, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSg

[Ano66-360] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(8):173–174, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSh

[Ano66-361] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(9):223, September 1966. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSi

[Ano66-362] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(10):169– 170, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSj

[Ano66-363] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(11):139–140, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:PCSk

[Ano66-364] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 3(12):144, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:REC

[Ano66-365] Anonymous. A relay for every communications need. IEEE Spectrum, 3(6):8, June 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:RPL

[Ano66-366] Anonymous. A research professor leaves his classroom and laboratory to become an astronaut. IEEE Spectrum, 3(2): 100–103, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIa

[Ano66-367] Anonymous. Scanning the issues. IEEE Spectrum, 3(1): 108–109, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIb

[Ano66-368] Anonymous. Scanning the issues. IEEE Spectrum, 3(2): 124–125, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIc

[Ano66-369] Anonymous. Scanning the issues. IEEE Spectrum, 3(3):182–183, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SId

[Ano66-370] Anonymous. Scanning the issues. IEEE Spectrum, 3(4):120–121, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIe

[Ano66-371] Anonymous. Scanning the issues. IEEE Spectrum, 3(5):96–97, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIf

[Ano66-372] Anonymous. Scanning the issues. IEEE Spectrum, 3(7):122–123, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIg

[Ano66-373] Anonymous. Scanning the issues. IEEE Spectrum, 3(8): 118–119, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIh

[Ano66-374] Anonymous. Scanning the issues. IEEE Spectrum, 3(9): 150–151, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIj

[Ano66-375] Anonymous. Scanning the issues. IEEE Spectrum, 3(10): 122–123, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIk

[Ano66-376] Anonymous. Scanning the issues. IEEE Spectrum, 3(11): 94–95, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SIl

[Ano66-377] Anonymous. Scanning the issues. IEEE Spectrum, 3(12): 86–87, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:ST

[Ano66-378] Anonymous. A shepherd’s tale.

IEEE Spectrum, 3(9):5, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPa

[Ano66-379] Anonymous. Special publications. IEEE Spectrum, 3 (1):134, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPb

[Ano66-380] Anonymous. Special publications. IEEE Spectrum, 3 (2):149, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPc

[Ano66-381] Anonymous. Special publications. IEEE Spectrum, 3(3):210, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPd

[Ano66-382] Anonymous. Special publications. IEEE Spectrum, 3(4):143, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPe

[Ano66-383] Anonymous. Special publications. IEEE Spectrum, 3(5):122, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPf

[Ano66-384] Anonymous. Special publications. IEEE Spectrum, 3(6): 138–139, June 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPg

[Ano66-385] Anonymous. Special publications. IEEE Spectrum, 3(7):146, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPh

[Ano66-386] Anonymous. Special publications. IEEE Spectrum, 3(8):155, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPi

[Ano66-387] Anonymous. Special publications. IEEE Spectrum, 3(9): 174, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPj

[Ano66-388] Anonymous. Special publications. IEEE Spectrum, 3 (10):147, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPk

[Ano66-389] Anonymous. Special publications. IEEE Spectrum, 3(11): 113, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SPl

[Ano66-390] Anonymous. Special publications. IEEE Spectrum, 3(12): 117, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLa

[Ano66-391] Anonymous. Spectral lines. IEEE Spectrum, 3(1):51, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLb

[Ano66-392] Anonymous. Spectral lines. IEEE Spectrum, 3(2):51, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLc

[Ano66-393] Anonymous. Spectral lines. IEEE Spectrum, 3(3):113, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLd

[Ano66-394] Anonymous. Spectral lines. IEEE Spectrum, 3(4):53, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLe

[Ano66-395] Anonymous. Spectral lines. IEEE Spectrum, 3(5):41, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLf

[Ano66-396] Anonymous. Spectral lines. IEEE Spectrum, 3(6):41, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLg

[Ano66-397] Anonymous. Spectral lines.

IEEE Spectrum, 3(7):55, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLh

[Ano66-398] Anonymous. Spectral lines.

IEEE Spectrum, 3(8):51, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLi

[Ano66-399] Anonymous. Spectral lines.

IEEE Spectrum, 3(9):75, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLj

[Ano66-400] Anonymous. Spectral lines.

IEEE Spectrum, 3(10):55, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLk

[Ano66-401] Anonymous. Spectral lines.

IEEE Spectrum, 3(11):39, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SLl

[Ano66-402] Anonymous. Spectral lines.

IEEE Spectrum, 3(12):33, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SSE

[Ano66-403] Anonymous. Starting salaries of engineers are deceptively high. IEEE Spectrum, 3(2):7, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SSN

[Ano66-404] Anonymous. “Subjective Space” — a new method of uncovering meaning. IEEE Spectrum, 3(2):8, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SSC

[Ano66-405] Anonymous. Superfluidity and superconductivity: a common origin. IEEE Spectrum, 3(8):8, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:SII

[Ano66-406] Anonymous. A supplement to “Information for IEEE authors”. IEEE Spectrum, 3(5):91, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCa

[Ano66-407] Anonymous. Table of contents. IEEE Spectrum, 3(1):1–2, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCb

[Ano66-408] Anonymous. Table of contents. IEEE Spectrum, 3(2):1–2, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCc

[Ano66-409] Anonymous. Table of contents. IEEE Spectrum, 3(3):1–2, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCd

[Ano66-410] Anonymous. Table of contents. IEEE Spectrum, 3(4):1–2, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCe

[Ano66-411] Anonymous. Table of contents. IEEE Spectrum, 3(5):1– 2, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCf

[Ano66-412] Anonymous. Table of contents. IEEE Spectrum, 3(6):1–2, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCg

[Ano66-413] Anonymous. Table of contents. IEEE Spectrum, 3(7):1– 2, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCi

[Ano66-414] Anonymous. Table of contents.

IEEE Spectrum, 3(8):1–2, August 1966. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCk

[Ano66-415] Anonymous. Table of contents. IEEE Spectrum, 3(9): 1–2, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCm

[Ano66-416] Anonymous. Table of contents. IEEE Spectrum, 3(10):1–2, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCo

[Ano66-417] Anonymous. Table of contents. IEEE Spectrum, 3(11): 1–2, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCq

[Ano66-418] Anonymous. Table of contents. IEEE Spectrum, 3(12):1–2, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCh

[Ano66-419] Anonymous. Technical correspondence. IEEE Spectrum, 3 (7):154–157, July 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCj

[Ano66-420] Anonymous. Technical correspondence. IEEE Spectrum, 3 (8):162–165, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCl

[Ano66-421] Anonymous. Technical correspondence. IEEE Spectrum, 3(9): 182–185, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCn

[Ano66-422] Anonymous. Technical correspondence. IEEE Spectrum, 3 (10):154–156, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCp

[Ano66-423] Anonymous. Technical correspondence. IEEE Spectrum, 3 (11):122, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TCr

[Ano66-424] Anonymous. Technical correspondence. IEEE Spectrum, 3 (12):121, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TO

[Ano66-425] Anonymous. Technical obsolescence. IEEE Spectrum, 3 (2):5, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TRT

[Ano66-426] Anonymous. A technique for the reduction of triangulation data. IEEE Spectrum, 3(7):5, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEa

[Ano66-427] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(1):4, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEc

[Ano66-428] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(2):4, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEd

[Ano66-429] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(3):4, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEe

[Ano66-430] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(4): 4, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEf

[Ano66-431] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(5): 4, May 1966. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEg

[Ano66-432] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(6): 4, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEh

[Ano66-433] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(7): 4, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEi

[Ano66-434] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(8):4, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEj

[Ano66-435] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3 (9):4, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEk

[Ano66-436] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3(10):4, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEl

[Ano66-437] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3 (11):4, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:IEEm

[Ano66-438] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 3 (12):4, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTa

[Ano66-439] Anonymous. Transients and trends. IEEE Spectrum, 3 (1):9, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTb

[Ano66-440] Anonymous. Transients and trends. IEEE Spectrum, 3 (2):9, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTc

[Ano66-441] Anonymous. Transients and trends. IEEE Spectrum, 3(3):9, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTd

[Ano66-442] Anonymous. Transients and trends. IEEE Spectrum, 3(4):9, April 1966. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTe

[Ano66-443] Anonymous. Transients and trends. IEEE Spectrum, 3(5):7, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTf

[Ano66-444] Anonymous. Transients and trends. IEEE Spectrum, 3(6):9, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTg

[Ano66-445] Anonymous. Transients and trends. IEEE Spectrum, 3(7):7, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTh

[Ano66-446] Anonymous. Transients and trends. IEEE Spectrum, 3(8):9, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTi

[Ano66-447] Anonymous. Transients and trends. IEEE Spectrum, 3 (9):9, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTj

[Ano66-448] Anonymous. Transients and trends. IEEE Spectrum, 3 (10):9, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTk

[Ano66-449] Anonymous. Transients and trends. IEEE Spectrum, 3 (11):9, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TTl

[Ano66-450] Anonymous. Transients and trends. IEEE Spectrum, 3 (12):9, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJa

[Ano66-451] Anonymous. Translated journals. IEEE Spectrum, 3(1): 130–133, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJb

[Ano66-452] Anonymous. Translated journals. IEEE Spectrum, 3(2): 144–147, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJc

[Ano66-453] Anonymous. Translated journals. IEEE Spectrum, 3(3): 204–209, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJd

[Ano66-454] Anonymous. Translated journals. IEEE Spectrum, 3(4):

138–143, April 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJe

[Ano66-455] Anonymous. Translated journals. IEEE Spectrum, 3(5): 113–120, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJf

[Ano66-456] Anonymous. Translated journals. IEEE Spectrum, 3(6): 134–137, June 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJg

[Ano66-457] Anonymous. Translated journals. IEEE Spectrum, 3(7): 142–145, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJh

[Ano66-458] Anonymous. Translated journals. IEEE Spectrum, 3(8): 150–152, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJi

[Ano66-459] Anonymous. Translated journals. IEEE Spectrum, 3(9): 170–173, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJj

[Ano66-460] Anonymous. Translated journals. IEEE Spectrum, 3(10):

144–146, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJk

[Ano66-461] Anonymous. Translated journals. IEEE Spectrum, 3(11): 108–112, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TJl

[Ano66-462] Anonymous. Translated journals. IEEE Spectrum, 3(12): 114–116, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TSL

[Ano66-463] Anonymous. A tunable source of “laser” light. IEEE Spectrum, 3(3):8, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:TSI

[Ano66-464] Anonymous. Turning a science into an industry. IEEE Spectrum, 3(1):99–102, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1966:WFM

[Ano66-465] Anonymous. Which filter has the minimum insertion loss? IEEE Spectrum, 3(6):5, June 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAa

[Ano67a] Anonymous. Advance abstracts. IEEE Spectrum, 4(1):

99–114, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAb

[Ano67b] Anonymous. Advance abstracts. IEEE Spectrum, 4(2): 98–106, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAc

[Ano67c] Anonymous. Advance abstracts.

IEEE Spectrum, 4(3):152–179, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAd

[Ano67d] Anonymous. Advance abstracts.

IEEE Spectrum, 4(4):118–139, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAe

[Ano67e] Anonymous. Advance abstracts.

IEEE Spectrum, 4(5):96–103, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAf

[Ano67f] Anonymous. Advance abstracts.

IEEE Spectrum, 4(6):105–119, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAg

[Ano67g] Anonymous. Advance abstracts.

IEEE Spectrum, 4(7):112–147, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAh

[Ano67h] Anonymous. Advance abstracts. IEEE Spectrum, 4(8):126–147, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAi

[Ano67i]  Anonymous.  Advance ab-

stracts. IEEE Spectrum, 4(9): 112–129, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAj

[Ano67j]  Anonymous.  Advance ab-

stracts. IEEE Spectrum, 4(10): 106–119, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAk

[Ano67k] Anonymous. Advance abstracts. IEEE Spectrum, 4(11): 108–121, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAl

[Ano67l]  Anonymous.  Advance ab-

stracts. IEEE Spectrum, 4(12): 106–129, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aa

[Ano67m] Anonymous. [advertisement]. IEEE Spectrum, 4(1):1, January 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ab

[Ano67n] Anonymous. [advertisement].

IEEE Spectrum, 4(1):3, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ac

[Ano67o] Anonymous. [advertisement].

IEEE Spectrum, 4(1):6–8, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ad

[Ano67p] Anonymous. [advertisement].

IEEE Spectrum, 4(1):11–14, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ae

[Ano67q] Anonymous. [advertisement].

IEEE Spectrum, 4(1):17, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Af

[Ano67r] Anonymous. [advertisement].

IEEE Spectrum, 4(1):21–34, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ag

[Ano67s] Anonymous. [advertisement].

IEEE Spectrum, 4(1):37, January 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ah

[Ano67t] Anonymous. [advertisement]. IEEE Spectrum, 4(1):39–44, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aj

[Ano67u] Anonymous. [advertisement]. IEEE Spectrum, 4(1):71–74, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ak

[Ano67v] Anonymous. [advertisement]. IEEE Spectrum, 4(1):114, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Al

[Ano67w] Anonymous. [advertisement]. IEEE Spectrum, 4(1):117, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Am

[Ano67x] Anonymous. [advertisement]. IEEE Spectrum, 4(1):119, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:An

[Ano67y] Anonymous. [advertisement]. IEEE Spectrum, 4(1):121–123, January 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ao

[Ano67z] Anonymous. [advertisement].

IEEE Spectrum, 4(1):129, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ap

[Ano67-27] Anonymous. [advertisement].

IEEE Spectrum, 4(1):131–141, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aq

[Ano67-28] Anonymous. [advertisement].

IEEE Spectrum, 4(1):2, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ar

[Ano67-29] Anonymous. [advertisement].

IEEE Spectrum, 4(2):nil1, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:As

[Ano67-30] Anonymous. [advertisement].

IEEE Spectrum, 4(2):3, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:At

[Ano67-31] Anonymous. [advertisement].

IEEE Spectrum, 4(2):6–8, February 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Au

[Ano67-32] Anonymous. [advertisement]. IEEE Spectrum, 4(2):10–11, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Av

[Ano67-33] Anonymous. [advertisement]. IEEE Spectrum, 4(2):15–30, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aw

[Ano67-34] Anonymous. [advertisement]. IEEE Spectrum, 4(2):33, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ax

[Ano67-35] Anonymous. [advertisement]. IEEE Spectrum, 4(2):35–44, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Az

[Ano67-36] Anonymous. [advertisement]. IEEE Spectrum, 4(2):69–72, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aba

[Ano67-37] Anonymous. [advertisement]. IEEE Spectrum, 4(2):95, February 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abb

[Ano67-38] Anonymous. [advertisement].

IEEE Spectrum, 4(2):101–107, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abc

[Ano67-39] Anonymous. [advertisement].

IEEE Spectrum, 4(2):109–111, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abd

[Ano67-40] Anonymous. [advertisement].

IEEE Spectrum, 4(2):113, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abe

[Ano67-41] Anonymous. [advertisement].

IEEE Spectrum, 4(2):115–119, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abf

[Ano67-42] Anonymous. [advertisement].

IEEE Spectrum, 4(2):121–123, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abg

[Ano67-43] Anonymous. [advertisement].

IEEE Spectrum, 4(2):125–137, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abh

[Ano67-44] Anonymous. [advertisement]. IEEE Spectrum, 4(2):nil2, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abi

[Ano67-45] Anonymous. [advertisement]. IEEE Spectrum, 4(3):1, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abj

[Ano67-46] Anonymous. [advertisement]. IEEE Spectrum, 4(3):3, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abk

[Ano67-47] Anonymous. [advertisement]. IEEE Spectrum, 4(3):6, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abl

[Ano67-48] Anonymous. [advertisement]. IEEE Spectrum, 4(3):7–8, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abm

[Ano67-49] Anonymous. [advertisement]. IEEE Spectrum, 4(3):11–13, March 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abn

[Ano67-50] Anonymous. [advertisement].

IEEE Spectrum, 4(3):20, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abo

[Ano67-51] Anonymous. [advertisement].

IEEE Spectrum, 4(3):23, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abp

[Ano67-52] Anonymous. [advertisement].

IEEE Spectrum, 4(3):25–32, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abq

[Ano67-53] Anonymous. [advertisement].

IEEE Spectrum, 4(3):39–100, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abs

[Ano67-54] Anonymous. [advertisement].

IEEE Spectrum, 4(3):129–132, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abt

[Ano67-55] Anonymous. [advertisement].

IEEE Spectrum, 4(3):155–175, March 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abu

[Ano67-56] Anonymous. [advertisement]. IEEE Spectrum, 4(3):181–183, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abv

[Ano67-57] Anonymous. [advertisement]. IEEE Spectrum, 4(3):185, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abw

[Ano67-58] Anonymous. [advertisement]. IEEE Spectrum, 4(3):188–191, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abx

[Ano67-59] Anonymous. [advertisement]. IEEE Spectrum, 4(3):193–197, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aby

[Ano67-60] Anonymous. [advertisement]. IEEE Spectrum, 4(3):199–206, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abz

[Ano67-61] Anonymous. [advertisement]. IEEE Spectrum, 4(3):207–293, March 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aca

[Ano67-62] Anonymous. [advertisement].

IEEE Spectrum, 4(3):2, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acb

[Ano67-63] Anonymous. [advertisement].

IEEE Spectrum, 4(4):1, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acc

[Ano67-64] Anonymous. [advertisement].

IEEE Spectrum, 4(4):3, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acd

[Ano67-65] Anonymous. [advertisement].

IEEE Spectrum, 4(4):6, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ace

[Ano67-66] Anonymous. [advertisement].

IEEE Spectrum, 4(4):7–10, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acf

[Ano67-67] Anonymous. [advertisement].

IEEE Spectrum, 4(4):12–13, April 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acg

[Ano67-68] Anonymous. [advertisement]. IEEE Spectrum, 4(4):15–19, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ach

[Ano67-69] Anonymous. [advertisement]. IEEE Spectrum, 4(4):21–32, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aci

[Ano67-70] Anonymous. [advertisement]. IEEE Spectrum, 4(4):35, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acj

[Ano67-71] Anonymous. [advertisement]. IEEE Spectrum, 4(4):37–44, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acl

[Ano67-72] Anonymous. [advertisement]. IEEE Spectrum, 4(4):83–86, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acm

[Ano67-73] Anonymous. [advertisement]. IEEE Spectrum, 4(4):115, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acn

[Ano67-74] Anonymous. [advertisement].

IEEE Spectrum, 4(4):122–133, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aco

[Ano67-75] Anonymous. [advertisement].

IEEE Spectrum, 4(4):141–145, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acp

[Ano67-76] Anonymous. [advertisement].

IEEE Spectrum, 4(4):147–148, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acq

[Ano67-77] Anonymous. [advertisement].

IEEE Spectrum, 4(4):153, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acr

[Ano67-78] Anonymous. [advertisement].

IEEE Spectrum, 4(4):155, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acs

[Ano67-79] Anonymous. [advertisement].

IEEE Spectrum, 4(4):159–165, April 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Act

[Ano67-80] Anonymous. [advertisement]. IEEE Spectrum, 4(4):2, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acu

[Ano67-81] Anonymous. [advertisement]. IEEE Spectrum, 4(5):1, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acv

[Ano67-82] Anonymous. [advertisement]. IEEE Spectrum, 4(5):3, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acw

[Ano67-83] Anonymous. [advertisement]. IEEE Spectrum, 4(5):6–8, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acx

[Ano67-84] Anonymous. [advertisement]. IEEE Spectrum, 4(5):11–13, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acy

[Ano67-85] Anonymous. [advertisement]. IEEE Spectrum, 4(5):17, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Acz

[Ano67-86] Anonymous. [advertisement].

IEEE Spectrum, 4(5):20–28, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ada

[Ano67-87] Anonymous. [advertisement].

IEEE Spectrum, 4(5):31, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adb

[Ano67-88] Anonymous. [advertisement].

IEEE Spectrum, 4(5):33–38, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Add

[Ano67-89] Anonymous. [advertisement].

IEEE Spectrum, 4(5):67–70, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ade

[Ano67-90] Anonymous. [advertisement].

IEEE Spectrum, 4(5):93, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adf

[Ano67-91] Anonymous. [advertisement].

IEEE Spectrum, 4(5):99–103, May 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adg

[Ano67-92] Anonymous. [advertisement]. IEEE Spectrum, 4(5):105–109, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adh

[Ano67-93] Anonymous. [advertisement]. IEEE Spectrum, 4(5):111, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adi

[Ano67-94] Anonymous. [advertisement]. IEEE Spectrum, 4(5):113–124, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adj

[Ano67-95] Anonymous. [advertisement]. IEEE Spectrum, 4(5):127–129, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adk

[Ano67-96] Anonymous. [advertisement]. IEEE Spectrum, 4(5):131–133, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adl

[Ano67-97] Anonymous. [advertisement]. IEEE Spectrum, 4(5):2, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adm

[Ano67-98] Anonymous. [advertisement].

IEEE Spectrum, 4(6):1, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adn

[Ano67-99] Anonymous. [advertisement].

IEEE Spectrum, 4(6):3, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ado

[Ano67-100] Anonymous. [advertisement].

IEEE Spectrum, 4(6):6, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adp

[Ano67-101] Anonymous. [advertisement].

IEEE Spectrum, 4(6):7–8, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adq

[Ano67-102] Anonymous. [advertisement].

IEEE Spectrum, 4(6):10, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adr

[Ano67-103] Anonymous. [advertisement].

IEEE Spectrum, 4(6):12–17, June 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ads

[Ano67-104] Anonymous. [advertisement]. IEEE Spectrum, 4(6):21–27, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adt

[Ano67-105] Anonymous. [advertisement]. IEEE Spectrum, 4(6):31–33, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adu

[Ano67-106] Anonymous. [advertisement]. IEEE Spectrum, 4(6):35–46, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adw

[Ano67-107] Anonymous. [advertisement]. IEEE Spectrum, 4(6):71–74, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adx

[Ano67-108] Anonymous. [advertisement]. IEEE Spectrum, 4(6):104, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ady

[Ano67-109] Anonymous. [advertisement]. IEEE Spectrum, 4(6):109–117, June 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adz

[Ano67-110] Anonymous. [advertisement].

IEEE Spectrum, 4(6):121–125, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aea

[Ano67-111] Anonymous. [advertisement].

IEEE Spectrum, 4(6):131, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aeb

[Ano67-112] Anonymous. [advertisement].

IEEE Spectrum, 4(6):133–135, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aec

[Ano67-113] Anonymous. [advertisement].

IEEE Spectrum, 4(6):137–139, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aed

[Ano67-114] Anonymous. [advertisement].

IEEE Spectrum, 4(6):2, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aee

[Ano67-115] Anonymous. [advertisement].

IEEE Spectrum, 4(7):1, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aef

[Ano67-116] Anonymous. [advertisement]. IEEE Spectrum, 4(7):3, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aeg

[Ano67-117] Anonymous. [advertisement]. IEEE Spectrum, 4(7):6, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aeh

[Ano67-118] Anonymous. [advertisement]. IEEE Spectrum, 4(7):7–8, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aei

[Ano67-119] Anonymous. [advertisement]. IEEE Spectrum, 4(7):11–13, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aej

[Ano67-120] Anonymous. [advertisement]. IEEE Spectrum, 4(7):17–28, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aek

[Ano67-121] Anonymous. [advertisement]. IEEE Spectrum, 4(7):31, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ael

[Ano67-122] Anonymous. [advertisement].

IEEE Spectrum, 4(7):33–35, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aen

[Ano67-123] Anonymous. [advertisement].

IEEE Spectrum, 4(7):87–90, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aeo

[Ano67-124] Anonymous. [advertisement].

IEEE Spectrum, 4(7):115–137, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aep

[Ano67-125] Anonymous. [advertisement].

IEEE Spectrum, 4(7):149–151, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aeq

[Ano67-126] Anonymous. [advertisement].

IEEE Spectrum, 4(7):153, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aer

[Ano67-127] Anonymous. [advertisement].

IEEE Spectrum, 4(7):155–159, July 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aes

[Ano67-128] Anonymous. [advertisement]. IEEE Spectrum, 4(7):163, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aet

[Ano67-129] Anonymous. [advertisement]. IEEE Spectrum, 4(7):167–171, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aeu

[Ano67-130] Anonymous. [advertisement]. IEEE Spectrum, 4(7):2, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aev

[Ano67-131] Anonymous. [advertisement]. IEEE Spectrum, 4(8):1, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aew

[Ano67-132] Anonymous. [advertisement]. IEEE Spectrum, 4(8):3, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aex

[Ano67-133] Anonymous. [advertisement]. IEEE Spectrum, 4(8):6, August 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aey

[Ano67-134] Anonymous. [advertisement].

IEEE Spectrum, 4(8):7–8, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aez

[Ano67-135] Anonymous. [advertisement].

IEEE Spectrum, 4(8):13–32, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afa

[Ano67-136] Anonymous. [advertisement].

IEEE Spectrum, 4(8):35, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afb

[Ano67-137] Anonymous. [advertisement].

IEEE Spectrum, 4(8):37–40, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afc

[Ano67-138] Anonymous. [advertisement].

IEEE Spectrum, 4(8):42–68, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afe

[Ano67-139] Anonymous. [advertisement].

IEEE Spectrum, 4(8):95–98, August 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aff

[Ano67-140] Anonymous. [advertisement]. IEEE Spectrum, 4(8):131, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afg

[Ano67-141] Anonymous. [advertisement]. IEEE Spectrum, 4(8):151, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afh

[Ano67-142] Anonymous. [advertisement]. IEEE Spectrum, 4(8):153, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afi

[Ano67-143] Anonymous. [advertisement]. IEEE Spectrum, 4(8):155–159, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afj

[Ano67-144] Anonymous. [advertisement]. IEEE Spectrum, 4(8):161–163, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afk

[Ano67-145] Anonymous. [advertisement]. IEEE Spectrum, 4(8):167–171, August 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afl

[Ano67-146] Anonymous. [advertisement].

IEEE Spectrum, 4(8):173–181, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afm

[Ano67-147] Anonymous. [advertisement].

IEEE Spectrum, 4(8):186–187, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afn

[Ano67-148] Anonymous. [advertisement].

IEEE Spectrum, 4(8):2, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afo

[Ano67-149] Anonymous. [advertisement].

IEEE Spectrum, 4(9):1, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afp

[Ano67-150] Anonymous. [advertisement].

IEEE Spectrum, 4(9):3, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afq

[Ano67-151] Anonymous. [advertisement].

IEEE Spectrum, 4(9):6, September 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afr

[Ano67-152] Anonymous. [advertisement]. IEEE Spectrum, 4(9):7–8, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afs

[Ano67-153] Anonymous. [advertisement]. IEEE Spectrum, 4(9):12–32, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aft

[Ano67-154] Anonymous. [advertisement]. IEEE Spectrum, 4(9):35, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afu

[Ano67-155] Anonymous. [advertisement]. IEEE Spectrum, 4(9):37–42, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afw

[Ano67-156] Anonymous. [advertisement]. IEEE Spectrum, 4(9):77–80, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afx

[Ano67-157] Anonymous. [advertisement]. IEEE Spectrum, 4(9):123–129, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afy

[Ano67-158] Anonymous. [advertisement].

IEEE Spectrum, 4(9):131–133, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afz

[Ano67-159] Anonymous. [advertisement].

IEEE Spectrum, 4(9):135, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aga

[Ano67-160] Anonymous. [advertisement].

IEEE Spectrum, 4(9):137–141, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agb

[Ano67-161] Anonymous. [advertisement].

IEEE Spectrum, 4(9):143, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agc

[Ano67-162] Anonymous. [advertisement].

IEEE Spectrum, 4(9):145–153, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agd

[Ano67-163] Anonymous. [advertisement].

IEEE Spectrum, 4(9):2, September 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Age

[Ano67-164] Anonymous. [advertisement]. IEEE Spectrum, 4(10):1, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agf

[Ano67-165] Anonymous. [advertisement]. IEEE Spectrum, 4(10):3, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agg

[Ano67-166] Anonymous. [advertisement]. IEEE Spectrum, 4(10):6, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agh

[Ano67-167] Anonymous. [advertisement]. IEEE Spectrum, 4(10):7–8, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agj

[Ano67-168] Anonymous. [advertisement]. IEEE Spectrum, 4(10):13–32, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agk

[Ano67-169] Anonymous. [advertisement]. IEEE Spectrum, 4(10):35, October 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agl

[Ano67-170] Anonymous. [advertisement].

IEEE Spectrum, 4(10):37–40, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agn

[Ano67-171] Anonymous. [advertisement].

IEEE Spectrum, 4(10):69–72, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ago

[Ano67-172] Anonymous. [advertisement].

IEEE Spectrum, 4(10):103, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agp

[Ano67-173] Anonymous. [advertisement].

IEEE Spectrum, 4(10):111, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agq

[Ano67-174] Anonymous. [advertisement].

IEEE Spectrum, 4(10):127, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agr

[Ano67-175] Anonymous. [advertisement].

IEEE Spectrum, 4(10):129– 131, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ags

[Ano67-176] Anonymous. [advertisement]. IEEE Spectrum, 4(10):135, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agt

[Ano67-177] Anonymous. [advertisement]. IEEE Spectrum, 4(10):2, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agu

[Ano67-178] Anonymous. [advertisement]. IEEE Spectrum, 4(11):1, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agv

[Ano67-179] Anonymous. [advertisement]. IEEE Spectrum, 4(11):3, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agw

[Ano67-180] Anonymous. [advertisement]. IEEE Spectrum, 4(11):4–7, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agx

[Ano67-181] Anonymous. [advertisement]. IEEE Spectrum, 4(11):11–15, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agy

[Ano67-182] Anonymous. [advertisement].

IEEE Spectrum, 4(11):18–30, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agz

[Ano67-183] Anonymous. [advertisement].

IEEE Spectrum, 4(11):33, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aha

[Ano67-184] Anonymous. [advertisement].

IEEE Spectrum, 4(11):35–38, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahc

[Ano67-185] Anonymous. [advertisement].

IEEE Spectrum, 4(11):71–74, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahd

[Ano67-186] Anonymous. [advertisement].

IEEE Spectrum, 4(11):105, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahe

[Ano67-187] Anonymous. [advertisement].

IEEE Spectrum, 4(11):111–119, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahf

[Ano67-188] Anonymous. [advertisement]. IEEE Spectrum, 4(11):125, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahg

[Ano67-189] Anonymous. [advertisement]. IEEE Spectrum, 4(11):131, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahh

[Ano67-190] Anonymous. [advertisement]. IEEE Spectrum, 4(11):133–135, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahi

[Ano67-191] Anonymous. [advertisement]. IEEE Spectrum, 4(11):139, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahj

[Ano67-192] Anonymous. [advertisement]. IEEE Spectrum, 4(11):2, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahk

[Ano67-193] Anonymous. [advertisement]. IEEE Spectrum, 4(12):1, December 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahl

[Ano67-194] Anonymous. [advertisement].

IEEE Spectrum, 4(12):3, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahm

[Ano67-195] Anonymous. [advertisement].

IEEE Spectrum, 4(12):6, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahn

[Ano67-196] Anonymous. [advertisement].

IEEE Spectrum, 4(12):7–8, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aho

[Ano67-197] Anonymous. [advertisement].

IEEE Spectrum, 4(12):12–26, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahp

[Ano67-198] Anonymous. [advertisement].

IEEE Spectrum, 4(12):29, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahq

[Ano67-199] Anonymous. [advertisement].

IEEE Spectrum, 4(12):31–38, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahs

[Ano67-200] Anonymous. [advertisement]. IEEE Spectrum, 4(12):83–86, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aht

[Ano67-201] Anonymous. [advertisement]. IEEE Spectrum, 4(12):103, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahu

[Ano67-202] Anonymous. [advertisement]. IEEE Spectrum, 4(12):114–125, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahv

[Ano67-203] Anonymous. [advertisement]. IEEE Spectrum, 4(12):131–135, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahw

[Ano67-204] Anonymous. [advertisement]. IEEE Spectrum, 4(12):2, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ALS

[Ano67-205] Anonymous. AIL Logic Symbols Standard. IEEE Spectrum, 4(2):5, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AIa

[Ano67-206] Anonymous. Annual index.

IEEE Spectrum, 4(12):146–166, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AAAa

[Ano67-207] Anonymous. Art at AIL [advertisement]. IEEE Spectrum, 4(5): 5, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ai

[Ano67-208] Anonymous. Authors. IEEE Spectrum, 4(1):46–48, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ay

[Ano67-209] Anonymous. Authors. IEEE Spectrum, 4(2):46–47, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Abr

[Ano67-210] Anonymous. Authors. IEEE Spectrum, 4(3):102–103, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ack

[Ano67-211] Anonymous. Authors. IEEE Spectrum, 4(4):46–48, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adc

[Ano67-212] Anonymous. Authors. IEEE Spectrum, 4(5):40–41, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Adv

[Ano67-213] Anonymous. Authors. IEEE Spectrum, 4(6):48, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Aem

[Ano67-214] Anonymous. Authors. IEEE Spectrum, 4(7):60–62, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afd

[Ano67-215] Anonymous. Authors. IEEE Spectrum, 4(8):70–72, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Afv

[Ano67-216] Anonymous. Authors. IEEE Spectrum, 4(9):44–46, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Agm

[Ano67-217] Anonymous. Authors. IEEE Spectrum, 4(10):42–43, October 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahb

[Ano67-218] Anonymous. Authors. IEEE Spectrum, 4(11):40–41, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ahr

[Ano67-219] Anonymous. Authors. IEEE Spectrum, 4(12):40–41, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:AGI

[Ano67-220] Anonymous. An authors’ guide to IEEE publications. IEEE Spectrum, 4(1):15–16, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRa

[Ano67-221] Anonymous. Book reviews.

IEEE Spectrum, 4(1):130–140, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRb

[Ano67-222] Anonymous. Book reviews.

IEEE Spectrum, 4(2):124–136, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRc

[Ano67-223] Anonymous. Book reviews.

IEEE Spectrum, 4(3):198–204, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRd

[Ano67-224] Anonymous. Book reviews. IEEE Spectrum, 4(4):158–165, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRe

[Ano67-225] Anonymous. Book reviews. IEEE Spectrum, 4(5):112–122, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRf

[Ano67-226] Anonymous. Book reviews. IEEE Spectrum, 4(6):136–141, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRg

[Ano67-227] Anonymous. Book reviews. IEEE Spectrum, 4(7):165–171, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRh

[Ano67-228] Anonymous. Book reviews. IEEE Spectrum, 4(8):166–172, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRi

[Ano67-229] Anonymous. Book reviews. IEEE Spectrum, 4(9):144–153, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRj

[Ano67-230] Anonymous. Book reviews.

IEEE Spectrum, 4(10):132– 138, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRk

[Ano67-231] Anonymous. Book reviews.

IEEE Spectrum, 4(11):136–141, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BRl

[Ano67-232] Anonymous. Book reviews.

IEEE Spectrum, 4(12):142–145, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISa

[Ano67-233] Anonymous. IEEE Spectrum.

IEEE Spectrum, 4(3):6, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISb

[Ano67-234] Anonymous. IEEE Spectrum.

IEEE Spectrum, 4(4):6, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISc

[Ano67-235] Anonymous. IEEE Spectrum.

IEEE Spectrum, 4(5):6, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISd

[Ano67-236] Anonymous. IEEE Spectrum. IEEE Spectrum, 4(6):6, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISe

[Ano67-237] Anonymous. IEEE Spectrum. IEEE Spectrum, 4(7):6, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISf

[Ano67-238] Anonymous. IEEE Spectrum. IEEE Spectrum, 4(8):6, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISg

[Ano67-239] Anonymous. IEEE Spectrum. IEEE Spectrum, 4(9):6, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISh

[Ano67-240] Anonymous. IEEE Spectrum. IEEE Spectrum, 4(10):6, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISi

[Ano67-241] Anonymous. IEEE Spectrum. IEEE Spectrum, 4(11):6, November 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISj

[Ano67-242] Anonymous. IEEE Spectrum.

IEEE Spectrum, 4(12):6, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPa

[Ano67-243] Anonymous. [breaker page].

IEEE Spectrum, 4(1):4, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPb

[Ano67-244] Anonymous. [breaker page].

IEEE Spectrum, 4(2):4, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPc

[Ano67-245] Anonymous. [breaker page].

IEEE Spectrum, 4(3):4, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPd

[Ano67-246] Anonymous. [breaker page].

IEEE Spectrum, 4(4):4, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPe

[Ano67-247] Anonymous. [breaker page].

IEEE Spectrum, 4(5):4, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPf

[Ano67-248] Anonymous. [breaker page]. IEEE Spectrum, 4(6):4, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPg

[Ano67-249] Anonymous. [breaker page]. IEEE Spectrum, 4(7):4, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPh

[Ano67-250] Anonymous. [breaker page]. IEEE Spectrum, 4(8):4, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPi

[Ano67-251] Anonymous. [breaker page]. IEEE Spectrum, 4(9):4, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPj

[Ano67-252] Anonymous. [breaker page]. IEEE Spectrum, 4(10):4, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPk

[Ano67-253] Anonymous. [breaker page]. IEEE Spectrum, 4(11):4, November 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BPl

[Ano67-254] Anonymous. [breaker page].

IEEE Spectrum, 4(12):4, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:BHA

[Ano67-255] Anonymous. Brewster and his angle. IEEE Spectrum, 4(7):5, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ca

[Ano67-256] Anonymous. Calendar. IEEE Spectrum, 4(1):35–37, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Cb

[Ano67-257] Anonymous. Calendar. IEEE Spectrum, 4(2):31–33, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Cc

[Ano67-258] Anonymous. Calendar. IEEE Spectrum, 4(3):21–23, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Cd

[Ano67-259] Anonymous. Calendar. IEEE Spectrum, 4(4):33–35, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ce

[Ano67-260] Anonymous. Calendar. IEEE Spectrum, 4(5):29–30, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Cf

[Ano67-261] Anonymous. Calendar. IEEE Spectrum, 4(6):29–30, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Cg

[Ano67-262] Anonymous. Calendar. IEEE Spectrum, 4(7):29–30, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ch

[Ano67-263] Anonymous. Calendar. IEEE Spectrum, 4(8):33–35, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ci

[Ano67-264] Anonymous. Calendar. IEEE Spectrum, 4(9):33–35, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Cj

[Ano67-265] Anonymous. Calendar. IEEE Spectrum, 4(10):33–34, October 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ck

[Ano67-266] Anonymous. Calendar. IEEE Spectrum, 4(11):31–32, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Cl

[Ano67-267] Anonymous. Calendar. IEEE Spectrum, 4(12):27–29, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:CTA

[Ano67-268] Anonymous. A case for “talking” automobiles. IEEE Spectrum, 4(4):5, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:CAa

[Ano67-269] Anonymous. Classified advertising. IEEE Spectrum, 4(6):127, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:CAb

[Ano67-270] Anonymous. Classified advertising. IEEE Spectrum, 4(11): 127, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:CIC

[Ano67-271] Anonymous. Compact infrared CO_{2} lasers. IEEE Spectrum, 4(1):5, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:CHT

[Ano67-272] Anonymous. The course of highway traffic theory. IEEE Spectrum, 4(3):5, March 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ECP

[Ano67-273] Anonymous. Experiments in the control of physiological function by radio-frequency transmission across the intact skin boundary. IEEE Spectrum, 4(6):5, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FAD

[Ano67-274] Anonymous. A fast-acting detector of infrared light. IEEE Spectrum, 4(7):15, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FF

[Ano67-275] Anonymous. Feedback from the field. IEEE Spectrum, 4(12): 77–82, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPa

[Ano67-276] Anonymous. Focal points. IEEE Spectrum, 4(1):120–125, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPb

[Ano67-277] Anonymous. Focal points. IEEE Spectrum, 4(2):114–119, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPc

[Ano67-278] Anonymous. Focal points. IEEE Spectrum, 4(3):186–191, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPd

[Ano67-279] Anonymous. Focal points. IEEE Spectrum, 4(4):149–153, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPe

[Ano67-280] Anonymous. Focal points. IEEE Spectrum, 4(5):125–129, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPf

[Ano67-281] Anonymous. Focal points. IEEE Spectrum, 4(6):128–131, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPg

[Ano67-282] Anonymous. Focal points. IEEE Spectrum, 4(7):154–158, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPh

[Ano67-283] Anonymous. Focal points. IEEE Spectrum, 4(8):154–159, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPi

[Ano67-284] Anonymous. Focal points. IEEE Spectrum, 4(9):136–141, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPj

[Ano67-285] Anonymous. Focal points. IEEE Spectrum, 4(10):124–127, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPk

[Ano67-286] Anonymous. Focal points. IEEE Spectrum, 4(11):128–131, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FPl

[Ano67-287] Anonymous. Focal points. IEEE Spectrum, 4(12):136–138, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:RLG

[Ano67-288] Anonymous. From Rayleigh to the LaGuardia Glideslope. IEEE Spectrum, 4(8):5, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCa

[Ano67-289] Anonymous. [front cover]. IEEE Spectrum, 4(1):c1, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCb

[Ano67-290] Anonymous. [front cover]. IEEE Spectrum, 4(1):c2, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCc

[Ano67-291] Anonymous. [front cover].

IEEE Spectrum, 4(2):c1, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCd

[Ano67-292] Anonymous. [front cover].

IEEE Spectrum, 4(2):c2, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCe

[Ano67-293] Anonymous. [front cover]. IEEE Spectrum, 4(3):c1, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCf

[Ano67-294] Anonymous. [front cover]. IEEE Spectrum, 4(3):c2, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCg

[Ano67-295] Anonymous. [front cover]. IEEE Spectrum, 4(4):c1, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCh

[Ano67-296] Anonymous. [front cover]. IEEE Spectrum, 4(4):c2, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCi

[Ano67-297] Anonymous. [front cover]. IEEE Spectrum, 4(5):c1, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCj

[Ano67-298] Anonymous. [front cover]. IEEE Spectrum, 4(5):c2, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCk

[Ano67-299] Anonymous. [front cover]. IEEE Spectrum, 4(6):c1, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCl

[Ano67-300] Anonymous. [front cover]. IEEE Spectrum, 4(6):c2, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCm

[Ano67-301] Anonymous. [front cover]. IEEE Spectrum, 4(7):c1, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCn

[Ano67-302] Anonymous. [front cover]. IEEE Spectrum, 4(7):c2, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCo

[Ano67-303] Anonymous. [front cover]. IEEE Spectrum, 4(8):c1, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCp

[Ano67-304] Anonymous. [front cover]. IEEE Spectrum, 4(8):c2, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCq

[Ano67-305] Anonymous. [front cover]. IEEE Spectrum, 4(9):c1, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCr

[Ano67-306] Anonymous. [front cover]. IEEE Spectrum, 4(9):c2, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCs

[Ano67-307] Anonymous. [front cover]. IEEE Spectrum, 4(10):c1, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCt

[Ano67-308] Anonymous. [front cover]. IEEE Spectrum, 4(10):c2, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCu

[Ano67-309] Anonymous. [front cover]. IEEE Spectrum, 4(11):c1, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCv

[Ano67-310] Anonymous. [front cover]. IEEE Spectrum, 4(11):c2, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCw

[Ano67-311] Anonymous. [front cover]. IEEE Spectrum, 4(12):c1, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FCx

[Ano67-312] Anonymous. [front cover]. IEEE Spectrum, 4(12):c2, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:FTI

[Ano67-313] Anonymous. “Functional” tantalum integrated circuits. IEEE Spectrum, 4(10):11, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEA

[Ano67-314] Anonymous. IEEE educational activities: an ad hoc committee report. IEEE Spectrum, 4(11): 101–104, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IFAa

[Ano67-315] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 4 (1):10–12, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IFAb

[Ano67-316] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 4(3):10–13, March 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IFAc

[Ano67-317] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 4(4):14–19, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IFAd

[Ano67-318] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 4(5):10–12, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IFAe

[Ano67-319] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 4(6):11–16, June 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IFAf

[Ano67-320] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 4(7):10–12, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IFAg

[Ano67-321] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 4 (11):10–14, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IJD

[Ano67-322] Anonymous. IEEE Japan directory 1967–1968. IEEE Spectrum, 4(8):41–64, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:INR

[Ano67-323] Anonymous. IEEE’s new Region Ten. IEEE Spectrum, 4(8):185, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAa

[Ano67-324] Anonymous. Index to advertisers. IEEE Spectrum, 4(1):

142–144, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAb

[Ano67-325] Anonymous. Index to advertisers. IEEE Spectrum, 4(2): 138–140, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAc

[Ano67-326] Anonymous. Index to advertisers. IEEE Spectrum, 4(3): 294–298, March 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAd

[Ano67-327] Anonymous. Index to advertisers. IEEE Spectrum, 4(4): 166–168, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAe

[Ano67-328] Anonymous. Index to advertisers. IEEE Spectrum, 4 (5):134–136, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAf

[Ano67-329] Anonymous. Index to advertisers. IEEE Spectrum, 4 (6):142–144, June 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAg

[Ano67-330] Anonymous. Index to advertisers. IEEE Spectrum, 4 (7):172–174, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAh

[Ano67-331] Anonymous. Index to advertisers. IEEE Spectrum, 4(8): 189–191, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAi

[Ano67-332] Anonymous. Index to advertisers. IEEE Spectrum, 4(9): 154–156, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAj

[Ano67-333] Anonymous. Index to advertisers. IEEE Spectrum, 4(10): 139–140, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAk

[Ano67-334] Anonymous. Index to advertisers. IEEE Spectrum, 4(11): 142–144, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IAl

[Ano67-335] Anonymous. Index to advertisers. IEEE Spectrum, 4(12): 167–168, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISMa

[Ano67-336] Anonymous. Inside solidifying metals. IEEE Spectrum, 4(3):17, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ISMb

[Ano67-337] Anonymous. Inside solidifying metals. IEEE Spectrum, 4 (9):11, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEb

[Ano67-338] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(2):4, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:ITC

[Ano67-339] Anonymous. International Technical Conferences Japan 1967. IEEE Spectrum, 4(8):184, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:JPE

[Ano67-340] Anonymous. Japan Pavilion at Expo 67 in Montreal. IEEE Spectrum, 4(8):188, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:MVD

[Ano67-341] Anonymous. Making voices from the depths sound deeper. IEEE Spectrum, 4(5):19, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:MWR

[Ano67-342] Anonymous. Millimeter wave research in the 1890’s. IEEE Spectrum, 4(12):5, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:MGL

[Ano67-343] Anonymous. Molecular-gas lasers. IEEE Spectrum, 4(1): 19, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NPI

[Ano67-344] Anonymous. New procedure for IEEE membership application. IEEE Spectrum, 4(11):9, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIa

[Ano67-345] Anonymous. News of the IEEE. IEEE Spectrum, 4(1):18–33, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIb

[Ano67-346] Anonymous. News of the IEEE. IEEE Spectrum, 4(2): 12–28, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIc

[Ano67-347] Anonymous. News of the IEEE. IEEE Spectrum, 4(3):14–20, March 1967. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NId

[Ano67-348] Anonymous. News of the IEEE.

IEEE Spectrum, 4(4):20–30, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIe

[Ano67-349] Anonymous. News of the IEEE.

IEEE Spectrum, 4(5):14–27, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIf

[Ano67-350] Anonymous. News of the IEEE.

IEEE Spectrum, 4(6):18–28, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIg

[Ano67-351] Anonymous. News of the IEEE.

IEEE Spectrum, 4(7):14–27, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIh

[Ano67-352] Anonymous. News of the IEEE.

IEEE Spectrum, 4(8):10–29, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIi

[Ano67-353] Anonymous. News of the IEEE. IEEE Spectrum, 4(9):

10–28, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIj

[Ano67-354] Anonymous. News of the IEEE. IEEE Spectrum, 4(10): 10–30, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIk

[Ano67-355] Anonymous. News of the IEEE. IEEE Spectrum, 4(11): 16–30, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:NIl

[Ano67-356] Anonymous. News of the IEEE. IEEE Spectrum, 4(12): 10–24, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Oa

[Ano67-357] Anonymous. Obituaries. IEEE Spectrum, 4(1):43, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ob

[Ano67-358] Anonymous. Obituaries. IEEE Spectrum, 4(2):42–43, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Oc

[Ano67-359] Anonymous. Obituaries. IEEE Spectrum, 4(3):30–31, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Od

[Ano67-360] Anonymous. Obituaries. IEEE Spectrum, 4(4):42–43, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Oe

[Ano67-361] Anonymous. Obituaries. IEEE Spectrum, 4(5):37, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Of

[Ano67-362] Anonymous. Obituaries. IEEE Spectrum, 4(6):44–45, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Og

[Ano67-363] Anonymous. Obituaries. IEEE Spectrum, 4(7):37–39, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Oh

[Ano67-364] Anonymous. Obituaries. IEEE Spectrum, 4(8):39, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Oi

[Ano67-365] Anonymous. Obituaries. IEEE Spectrum, 4(9):39, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Oj

[Ano67-366] Anonymous. Obituaries. IEEE Spectrum, 4(10):39, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ok

[Ano67-367] Anonymous. Obituaries. IEEE Spectrum, 4(11):36–37, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ol

[Ano67-368] Anonymous. Obituaries. IEEE Spectrum, 4(12):37, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:OHY

[Ano67-369] Anonymous. One hundred years of modern japan. IEEE Spectrum, 4(8):182–183, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:POB

[Ano67-370] Anonymous. Passive octavebandwidth HF helicon-wave isolator. IEEE Spectrum, 4(9): 5, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pa

[Ano67-371] Anonymous. People. IEEE Spectrum, 4(1):38–43, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pb

[Ano67-372] Anonymous. People. IEEE Spectrum, 4(2):34–42, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pc

[Ano67-373] Anonymous. People. IEEE Spectrum, 4(3):24–30, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pd

[Ano67-374] Anonymous. People. IEEE Spectrum, 4(4):36–42, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pe

[Ano67-375] Anonymous. People. IEEE Spectrum, 4(5):32–37, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pf

[Ano67-376] Anonymous. People. IEEE Spectrum, 4(6):34–44, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Ph

[Ano67-377] Anonymous. People. IEEE Spectrum, 4(7):32–37, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pi

[Ano67-378] Anonymous. People. IEEE Spectrum, 4(8):36–39, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pj

[Ano67-379] Anonymous. People. IEEE Spectrum, 4(9):36–39, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pk

[Ano67-380] Anonymous. People. IEEE Spectrum, 4(10):36–38, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pl

[Ano67-381] Anonymous. People. IEEE Spectrum, 4(11):34–36, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:Pm

[Ano67-382] Anonymous. People. IEEE Spectrum, 4(12):30–37, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PG

[Ano67-383] Anonymous. Photodiodes with gain. IEEE Spectrum, 4(2): 13, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PR

[Ano67-384] Anonymous. Photos of a rainstorm. IEEE Spectrum, 4 (11):21, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSa

[Ano67-385] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(1):143–144, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSb

[Ano67-386] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(2):139–140, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSc

[Ano67-387] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(3):295–297, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSd

[Ano67-388] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(4):167–168, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSe

[Ano67-389] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(5):135–136, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSf

[Ano67-390] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(6):143–144, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSg

[Ano67-391] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(7):173–174, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSh

[Ano67-392] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(8):191–192, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSi

[Ano67-393] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(9):155–156, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSj

[Ano67-394] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(10):140, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSk

[Ano67-395] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(11):143–144, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCSl

[Ano67-396] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 4(12):168, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:PCP

[Ano67-397] Anonymous. Programming complex problems simply. IEEE Spectrum, 4(6):19, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:RSO

[Ano67-398] Anonymous. Remotely synchronizing oscillators over time varying paths. IEEE Spectrum, 4 (10):5, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIa

[Ano67-399] Anonymous. Scanning the issues. IEEE Spectrum, 4(1):97–98, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIb

[Ano67-400] Anonymous. Scanning the issues. IEEE Spectrum, 4(2): 96–97, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIc

[Ano67-401] Anonymous. Scanning the issues. IEEE Spectrum, 4(3):150–151, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SId

[Ano67-402] Anonymous. Scanning the issues. IEEE Spectrum, 4(4):116–117, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIe

[Ano67-403] Anonymous. Scanning the issues. IEEE Spectrum, 4(5):94–95, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIf

[Ano67-404] Anonymous. Scanning the issues. IEEE Spectrum, 4(6):102–103, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIg

[Ano67-405] Anonymous. Scanning the issues.

IEEE Spectrum, 4(7):110–111, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIh

[Ano67-406] Anonymous. Scanning the issues. IEEE Spectrum, 4(8): 124–125, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIi

[Ano67-407] Anonymous. Scanning the issues. IEEE Spectrum, 4(9): 110–111, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIj

[Ano67-408] Anonymous. Scanning the issues. IEEE Spectrum, 4(10): 104–105, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIk

[Ano67-409] Anonymous. Scanning the issues. IEEE Spectrum, 4(11): 106–107, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SIl

[Ano67-410] Anonymous. Scanning the issues. IEEE Spectrum, 4(12): 104–105, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SCT

[Ano67-411] Anonymous. See a computer talk. IEEE Spectrum, 4(8):11, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPa

[Ano67-412] Anonymous. Special publications. IEEE Spectrum, 4 (1):118, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPb

[Ano67-413] Anonymous. Special publications. IEEE Spectrum, 4 (2):112, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPc

[Ano67-414] Anonymous. Special publications. IEEE Spectrum, 4(3):184, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPd

[Ano67-415] Anonymous. Special publications. IEEE Spectrum, 4(4):146, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPe

[Ano67-416] Anonymous. Special publications. IEEE Spectrum, 4(5):110, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPf

[Ano67-417] Anonymous. Special publications. IEEE Spectrum, 4(6):126, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPg

[Ano67-418] Anonymous. Special publications. IEEE Spectrum, 4(7):152, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPh

[Ano67-419] Anonymous. Special publications. IEEE Spectrum, 4(8):152, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPi

[Ano67-420] Anonymous. Special publications. IEEE Spectrum, 4(9): 134, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPj

[Ano67-421] Anonymous. Special publications. IEEE Spectrum, 4 (10):123, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPk

[Ano67-422] Anonymous. Special publications. IEEE Spectrum, 4(11): 126, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SPl

[Ano67-423] Anonymous. Special publications. IEEE Spectrum, 4(12): 135, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SLa

[Ano67-424] Anonymous. Spectral lines. IEEE Spectrum, 4(1):45, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SLb

[Ano67-425] Anonymous. Spectral lines. IEEE Spectrum, 4(2):45, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SLc

[Ano67-426] Anonymous. Spectral lines. IEEE Spectrum, 4(3):101, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SLd

[Ano67-427] Anonymous. Spectral lines. IEEE Spectrum, 4(7):59, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SLe

[Ano67-428] Anonymous. Spectral lines. IEEE Spectrum, 4(9):43, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SLf

[Ano67-429] Anonymous. Spectral lines.

IEEE Spectrum, 4(10):41, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SLg

[Ano67-430] Anonymous. Spectral lines.

IEEE Spectrum, 4(11):39, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SLh

[Ano67-431] Anonymous. Spectral lines.

IEEE Spectrum, 4(12):39, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:SMN

[Ano67-432] Anonymous. Speech: Man’s natural communication. IEEE Spectrum, 4(6):75–86, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCa

[Ano67-433] Anonymous. Table of contents.

IEEE Spectrum, 4(1):1–2, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCc

[Ano67-434] Anonymous. Table of contents.

IEEE Spectrum, 4(2):1–2, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCe

[Ano67-435] Anonymous. Table of contents. IEEE Spectrum, 4(3):1–2, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCg

[Ano67-436] Anonymous. Table of contents. IEEE Spectrum, 4(4):1–2, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCi

[Ano67-437] Anonymous. Table of contents. IEEE Spectrum, 4(5):1– 2, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCk

[Ano67-438] Anonymous. Table of contents. IEEE Spectrum, 4(6):1–2, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCm

[Ano67-439] Anonymous. Table of contents. IEEE Spectrum, 4(7):1– 2, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCo

[Ano67-440] Anonymous. Table of contents. IEEE Spectrum, 4(8):1–2, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCq

[Ano67-441] Anonymous. Table of contents. IEEE Spectrum, 4(9): 1–2, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCs

[Ano67-442] Anonymous. Table of contents.

IEEE Spectrum, 4(10):1–2, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCu

[Ano67-443] Anonymous. Table of contents. IEEE Spectrum, 4(11): 1–2, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCw

[Ano67-444] Anonymous. Table of contents.

IEEE Spectrum, 4(12):1–2, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCb

[Ano67-445] Anonymous. Technical correspondence. IEEE Spectrum, 4 (1):126–129, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCd

[Ano67-446] Anonymous. Technical correspondence. IEEE Spectrum, 4 (2):120–122, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCf

[Ano67-447] Anonymous. Technical correspondence. IEEE Spectrum, 4 (3):192–197, March 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCh

[Ano67-448] Anonymous. Technical correspondence. IEEE Spectrum, 4 (4):154–157, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCj

[Ano67-449] Anonymous. Technical correspondence. IEEE Spectrum, 4 (5):130–133, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCl

[Ano67-450] Anonymous. Technical correspondence. IEEE Spectrum, 4 (6):132–135, June 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCn

[Ano67-451] Anonymous. Technical correspondence. IEEE Spectrum, 4 (7):160–164, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCp

[Ano67-452] Anonymous. Technical correspondence. IEEE Spectrum, 4 (8):160–165, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCr

[Ano67-453] Anonymous. Technical correspondence. IEEE Spectrum, 4(9): 142–143, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCt

[Ano67-454] Anonymous. Technical correspondence. IEEE Spectrum, 4 (10):128–130, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCv

[Ano67-455] Anonymous. Technical correspondence. IEEE Spectrum, 4 (11):132–135, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TCx

[Ano67-456] Anonymous. Technical correspondence. IEEE Spectrum, 4(12):139–141, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TP

[Ano67-457] Anonymous. Technical program. IEEE Spectrum, 4(3): 33–98, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TYV

[Ano67-458] Anonymous. Thanks for your vote of confidence. IEEE Spectrum, 4(11):5, November 1967.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEa

[Ano67-459] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(1):4, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEc

[Ano67-460] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(3):4, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEd

[Ano67-461] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(4): 4, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEe

[Ano67-462] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(5): 4, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEf

[Ano67-463] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(6): 4, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEg

[Ano67-464] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(7): 4, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEh

[Ano67-465] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(8):4, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEi

[Ano67-466] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4 (9):4, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEj

[Ano67-467] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4(10):4, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEk

[Ano67-468] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4 (11):4, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:IEEl

[Ano67-469] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 4 (12):4, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TFF

[Ano67-470] Anonymous. Thin films: faster by computer. IEEE Spectrum, 4(12):11, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTa

[Ano67-471] Anonymous. Transients and trends. IEEE Spectrum, 4 (1):9, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTb

[Ano67-472] Anonymous. Transients and trends. IEEE Spectrum, 4 (2):9, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTc

[Ano67-473] Anonymous. Transients and trends. IEEE Spectrum, 4(3):9, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTd

[Ano67-474] Anonymous. Transients and trends. IEEE Spectrum, 4(4):11, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTe

[Ano67-475] Anonymous. Transients and trends. IEEE Spectrum, 4(5):9, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTf

[Ano67-476] Anonymous. Transients and trends. IEEE Spectrum, 4(6):9, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTg

[Ano67-477] Anonymous. Transients and trends. IEEE Spectrum, 4(7):9, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTh

[Ano67-478] Anonymous. Transients and trends. IEEE Spectrum, 4(8):9, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTi

[Ano67-479] Anonymous. Transients and trends. IEEE Spectrum, 4 (9):9, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTj

[Ano67-480] Anonymous. Transients and trends. IEEE Spectrum, 4 (10):9, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTk

[Ano67-481] Anonymous. Transients and trends. IEEE Spectrum, 4 (11):8, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TTl

[Ano67-482] Anonymous. Transients and trends. IEEE Spectrum, 4 (12):9, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJa

[Ano67-483] Anonymous. Translated journals. IEEE Spectrum, 4(1): 115–117, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJb

[Ano67-484] Anonymous. Translated journals. IEEE Spectrum, 4(2): 108–111, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJc

[Ano67-485] Anonymous. Translated journals. IEEE Spectrum, 4(3): 180–183, March 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJd

[Ano67-486] Anonymous. Translated journals. IEEE Spectrum, 4(4): 140–144, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJe

[Ano67-487] Anonymous. Translated journals. IEEE Spectrum, 4(5): 104–109, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJf

[Ano67-488] Anonymous. Translated journals. IEEE Spectrum, 4(6): 120–125, June 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJg

[Ano67-489] Anonymous. Translated journals. IEEE Spectrum, 4(7): 148–151, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJh

[Ano67-490] Anonymous. Translated journals. IEEE Spectrum, 4(8): 148–151, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJi

[Ano67-491] Anonymous. Translated journals. IEEE Spectrum, 4(9): 130–132, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJj

[Ano67-492] Anonymous. Translated journals. IEEE Spectrum, 4(10): 120–122, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJk

[Ano67-493] Anonymous. Translated journals. IEEE Spectrum, 4(11): 122–125, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:TJl

[Ano67-494] Anonymous. Translated journals. IEEE Spectrum, 4(12): 130–133, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1967:WEL

[Ano67-495] Anonymous. WESCON 68 Exhibitors list. IEEE Spectrum, 4(7):40–58, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAa

[Ano68a] Anonymous. Advance abstracts. IEEE Spectrum, 5(1): 98–107, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAb

[Ano68b] Anonymous. Advance abstracts. IEEE Spectrum, 5(2): 112–126, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAc

[Ano68c] Anonymous. Advance abstracts. IEEE Spectrum, 5(3):106–119, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAd

[Ano68d] Anonymous. Advance abstracts.

IEEE Spectrum, 5(4):122–149, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAe

[Ano68e] Anonymous. Advance abstracts.

IEEE Spectrum, 5(5):96–111, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAf

[Ano68f] Anonymous. Advance abstracts.

IEEE Spectrum, 5(6):98–117, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAg

[Ano68g] Anonymous. Advance abstracts.

IEEE Spectrum, 5(7):108–123, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAh

[Ano68h] Anonymous. Advance abstracts.

IEEE Spectrum, 5(8):101–113, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAi

[Ano68i]  Anonymous.  Advance ab-

stracts. IEEE Spectrum, 5(9): 108–123, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAj

[Ano68j]  Anonymous.  Advance ab-

stracts. IEEE Spectrum, 5(10): 104–123, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAk

[Ano68k] Anonymous. Advance abstracts. IEEE Spectrum, 5(11): 137–154, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AAl

[Ano68l]  Anonymous.  Advance ab-

stracts. IEEE Spectrum, 5(12): 86–106, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aa

[Ano68m] Anonymous. [advertisement]. IEEE Spectrum, 5(1):nil1, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ab

[Ano68n] Anonymous. [advertisement]. IEEE Spectrum, 5(1):3, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ac

[Ano68o] Anonymous. [advertisement]. IEEE Spectrum, 5(1):6, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ad

[Ano68p] Anonymous. [advertisement].

IEEE Spectrum, 5(1):7–10, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ae

[Ano68q] Anonymous. [advertisement].

IEEE Spectrum, 5(1):12–13, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Af

[Ano68r] Anonymous. [advertisement].

IEEE Spectrum, 5(1):15, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ag

[Ano68s] Anonymous. [advertisement].

IEEE Spectrum, 5(1):19–32, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ah

[Ano68t] Anonymous. [advertisement].

IEEE Spectrum, 5(1):35, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ai

[Ano68u] Anonymous. [advertisement].

IEEE Spectrum, 5(1):37–42, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ak

[Ano68v] Anonymous. [advertisement]. IEEE Spectrum, 5(1):71–74, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Al

[Ano68w] Anonymous. [advertisement]. IEEE Spectrum, 5(1):99–107, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Am

[Ano68x] Anonymous. [advertisement]. IEEE Spectrum, 5(1):109–113, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:An

[Ano68y] Anonymous. [advertisement]. IEEE Spectrum, 5(1):115–116, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ao

[Ano68z] Anonymous. [advertisement]. IEEE Spectrum, 5(1):119–123, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ap

[Ano68-27] Anonymous. [advertisement]. IEEE Spectrum, 5(1):125–131, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aq

[Ano68-28] Anonymous. [advertisement].

IEEE Spectrum, 5(1):133–137, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ar

[Ano68-29] Anonymous. [advertisement].

IEEE Spectrum, 5(1):nil2, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:As

[Ano68-30] Anonymous. [advertisement].

IEEE Spectrum, 5(2):nil1, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:At

[Ano68-31] Anonymous. [advertisement].

IEEE Spectrum, 5(2):3, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Au

[Ano68-32] Anonymous. [advertisement].

IEEE Spectrum, 5(2):6, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Av

[Ano68-33] Anonymous. [advertisement].

IEEE Spectrum, 5(2):7–10, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aw

[Ano68-34] Anonymous. [advertisement]. IEEE Spectrum, 5(2):13, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ax

[Ano68-35] Anonymous. [advertisement]. IEEE Spectrum, 5(2):17–36, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ay

[Ano68-36] Anonymous. [advertisement]. IEEE Spectrum, 5(2):39, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Az

[Ano68-37] Anonymous. [advertisement]. IEEE Spectrum, 5(2):41, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abb

[Ano68-38] Anonymous. [advertisement]. IEEE Spectrum, 5(2):75–78, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abc

[Ano68-39] Anonymous. [advertisement]. IEEE Spectrum, 5(2):109, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abd

[Ano68-40] Anonymous. [advertisement].

IEEE Spectrum, 5(2):117–127, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abe

[Ano68-41] Anonymous. [advertisement].

IEEE Spectrum, 5(2):130–131, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abf

[Ano68-42] Anonymous. [advertisement].

IEEE Spectrum, 5(2):133, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abg

[Ano68-43] Anonymous. [advertisement].

IEEE Spectrum, 5(2):135, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abh

[Ano68-44] Anonymous. [advertisement].

IEEE Spectrum, 5(2):139–141, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abi

[Ano68-45] Anonymous. [advertisement].

IEEE Spectrum, 5(2):143–148, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abj

[Ano68-46] Anonymous. [advertisement]. IEEE Spectrum, 5(2):nil2, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abk

[Ano68-47] Anonymous. [advertisement]. IEEE Spectrum, 5(3):nil1, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abl

[Ano68-48] Anonymous. [advertisement]. IEEE Spectrum, 5(3):3, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abm

[Ano68-49] Anonymous. [advertisement]. IEEE Spectrum, 5(3):6, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abn

[Ano68-50] Anonymous. [advertisement]. IEEE Spectrum, 5(3):7–9, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abo

[Ano68-51] Anonymous. [advertisement]. IEEE Spectrum, 5(3):11–12, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abp

[Ano68-52] Anonymous. [advertisement].

IEEE Spectrum, 5(3):17–37, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abq

[Ano68-53] Anonymous. [advertisement].

IEEE Spectrum, 5(3):39–50, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abr

[Ano68-54] Anonymous. [advertisement].

IEEE Spectrum, 5(3):53, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abs

[Ano68-55] Anonymous. [advertisement].

IEEE Spectrum, 5(3):55–58, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abu

[Ano68-56] Anonymous. [advertisement].

IEEE Spectrum, 5(3):89–92, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abv

[Ano68-57] Anonymous. [advertisement].

IEEE Spectrum, 5(3):121–123, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abw

[Ano68-58] Anonymous. [advertisement]. IEEE Spectrum, 5(3):127–129, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abx

[Ano68-59] Anonymous. [advertisement]. IEEE Spectrum, 5(3):137–148, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aby

[Ano68-60] Anonymous. [advertisement]. IEEE Spectrum, 5(3):nil2, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abz

[Ano68-61] Anonymous. [advertisement]. IEEE Spectrum, 5(4):nil1, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aca

[Ano68-62] Anonymous. [advertisement]. IEEE Spectrum, 5(4):3, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acb

[Ano68-63] Anonymous. [advertisement]. IEEE Spectrum, 5(4):6–7, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acc

[Ano68-64] Anonymous. [advertisement].

IEEE Spectrum, 5(4):9–11, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acd

[Ano68-65] Anonymous. [advertisement].

IEEE Spectrum, 5(4):15–34, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ace

[Ano68-66] Anonymous. [advertisement].

IEEE Spectrum, 5(4):37, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acf

[Ano68-67] Anonymous. [advertisement].

IEEE Spectrum, 5(4):41–44, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ach

[Ano68-68] Anonymous. [advertisement].

IEEE Spectrum, 5(4):85–88, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aci

[Ano68-69] Anonymous. [advertisement].

IEEE Spectrum, 5(4):121, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acj

[Ano68-70] Anonymous. [advertisement]. IEEE Spectrum, 5(4):131–139, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ack

[Ano68-71] Anonymous. [advertisement]. IEEE Spectrum, 5(4):151, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acl

[Ano68-72] Anonymous. [advertisement]. IEEE Spectrum, 5(4):161, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acm

[Ano68-73] Anonymous. [advertisement]. IEEE Spectrum, 5(4):nil2, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acn

[Ano68-74] Anonymous. [advertisement]. IEEE Spectrum, 5(5):nil1, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aco

[Ano68-75] Anonymous. [advertisement]. IEEE Spectrum, 5(5):3, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acq

[Ano68-76] Anonymous. [advertisement].

IEEE Spectrum, 5(5):6, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acr

[Ano68-77] Anonymous. [advertisement].

IEEE Spectrum, 5(5):7–9, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acs

[Ano68-78] Anonymous. [advertisement].

IEEE Spectrum, 5(5):17–32, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Act

[Ano68-79] Anonymous. [advertisement].

IEEE Spectrum, 5(5):35, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acu

[Ano68-80] Anonymous. [advertisement].

IEEE Spectrum, 5(5):37–40, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acw

[Ano68-81] Anonymous. [advertisement].

IEEE Spectrum, 5(5):67–70, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acx

[Ano68-82] Anonymous. [advertisement]. IEEE Spectrum, 5(5):93, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acy

[Ano68-83] Anonymous. [advertisement]. IEEE Spectrum, 5(5):101–109, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acz

[Ano68-84] Anonymous. [advertisement]. IEEE Spectrum, 5(5):112–115, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ada

[Ano68-85] Anonymous. [advertisement]. IEEE Spectrum, 5(5):117, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adb

[Ano68-86] Anonymous. [advertisement]. IEEE Spectrum, 5(5):123–127, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adc

[Ano68-87] Anonymous. [advertisement]. IEEE Spectrum, 5(5):129–131, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Add

[Ano68-88] Anonymous. [advertisement].

IEEE Spectrum, 5(5):nil2, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ade

[Ano68-89] Anonymous. [advertisement].

IEEE Spectrum, 5(6):nil1, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adf

[Ano68-90] Anonymous. [advertisement].

IEEE Spectrum, 5(6):3, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adg

[Ano68-91] Anonymous. [advertisement].

IEEE Spectrum, 5(6):6, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adh

[Ano68-92] Anonymous. [advertisement].

IEEE Spectrum, 5(6):7–8, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adi

[Ano68-93] Anonymous. [advertisement].

IEEE Spectrum, 5(6):11–24, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adj

[Ano68-94] Anonymous. [advertisement]. IEEE Spectrum, 5(6):27, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adk

[Ano68-95] Anonymous. [advertisement]. IEEE Spectrum, 5(6):30–31, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adl

[Ano68-96] Anonymous. [advertisement]. IEEE Spectrum, 5(6):34–36, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adn

[Ano68-97] Anonymous. [advertisement]. IEEE Spectrum, 5(6):67–70, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ado

[Ano68-98] Anonymous. [advertisement]. IEEE Spectrum, 5(6):111, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adp

[Ano68-99] Anonymous. [advertisement]. IEEE Spectrum, 5(6):125, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adq

[Ano68-100] Anonymous. [advertisement].

IEEE Spectrum, 5(6):nil2, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adr

[Ano68-101] Anonymous. [advertisement].

IEEE Spectrum, 5(7):nil1, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ads

[Ano68-102] Anonymous. [advertisement].

IEEE Spectrum, 5(7):3, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adt

[Ano68-103] Anonymous. [advertisement].

IEEE Spectrum, 5(7):6, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adu

[Ano68-104] Anonymous. [advertisement].

IEEE Spectrum, 5(7):7, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adv

[Ano68-105] Anonymous. [advertisement].

IEEE Spectrum, 5(7):11–30, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adw

[Ano68-106] Anonymous. [advertisement]. IEEE Spectrum, 5(7):33, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adx

[Ano68-107] Anonymous. [advertisement]. IEEE Spectrum, 5(7):35–40, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adz

[Ano68-108] Anonymous. [advertisement]. IEEE Spectrum, 5(7):75–78, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aea

[Ano68-109] Anonymous. [advertisement]. IEEE Spectrum, 5(7):105, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aeb

[Ano68-110] Anonymous. [advertisement]. IEEE Spectrum, 5(7):114–115, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aec

[Ano68-111] Anonymous. [advertisement]. IEEE Spectrum, 5(7):127, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aed

[Ano68-112] Anonymous. [advertisement].

IEEE Spectrum, 5(7):133–135, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aee

[Ano68-113] Anonymous. [advertisement].

IEEE Spectrum, 5(7):139, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aef

[Ano68-114] Anonymous. [advertisement].

IEEE Spectrum, 5(7):nil2, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aeg

[Ano68-115] Anonymous. [advertisement].

IEEE Spectrum, 5(8):nil1, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aeh

[Ano68-116] Anonymous. [advertisement].

IEEE Spectrum, 5(8):3, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aei

[Ano68-117] Anonymous. [advertisement].

IEEE Spectrum, 5(8):6, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aej

[Ano68-118] Anonymous. [advertisement]. IEEE Spectrum, 5(8):7–8, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aek

[Ano68-119] Anonymous. [advertisement]. IEEE Spectrum, 5(8):13–19, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ael

[Ano68-120] Anonymous. [advertisement]. IEEE Spectrum, 5(8):23, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aem

[Ano68-121] Anonymous. [advertisement]. IEEE Spectrum, 5(8):25–30, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aeo

[Ano68-122] Anonymous. [advertisement]. IEEE Spectrum, 5(8):67–70, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aep

[Ano68-123] Anonymous. [advertisement]. IEEE Spectrum, 5(8):127, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aeq

[Ano68-124] Anonymous. [advertisement].

IEEE Spectrum, 5(8):nil2, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aer

[Ano68-125] Anonymous. [advertisement].

IEEE Spectrum, 5(9):nil1, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aes

[Ano68-126] Anonymous. [advertisement].

IEEE Spectrum, 5(9):3, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aet

[Ano68-127] Anonymous. [advertisement].

IEEE Spectrum, 5(9):6, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aeu

[Ano68-128] Anonymous. [advertisement].

IEEE Spectrum, 5(9):7–8, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aev

[Ano68-129] Anonymous. [advertisement].

IEEE Spectrum, 5(9):11–26, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aew

[Ano68-130] Anonymous. [advertisement]. IEEE Spectrum, 5(9):29, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aex

[Ano68-131] Anonymous. [advertisement]. IEEE Spectrum, 5(9):31–36, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aez

[Ano68-132] Anonymous. [advertisement]. IEEE Spectrum, 5(9):73–76, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afa

[Ano68-133] Anonymous. [advertisement]. IEEE Spectrum, 5(9):114–119, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afb

[Ano68-134] Anonymous. [advertisement]. IEEE Spectrum, 5(9):125, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afc

[Ano68-135] Anonymous. [advertisement]. IEEE Spectrum, 5(9):133–135, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afd

[Ano68-136] Anonymous. [advertisement].

IEEE Spectrum, 5(9):139–143, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afe

[Ano68-137] Anonymous. [advertisement].

IEEE Spectrum, 5(9):nil2, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aff

[Ano68-138] Anonymous. [advertisement].

IEEE Spectrum, 5(10):nil1, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afg

[Ano68-139] Anonymous. [advertisement].

IEEE Spectrum, 5(10):3, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afh

[Ano68-140] Anonymous. [advertisement].

IEEE Spectrum, 5(10):6, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afi

[Ano68-141] Anonymous. [advertisement].

IEEE Spectrum, 5(10):7–8, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afj

[Ano68-142] Anonymous. [advertisement]. IEEE Spectrum, 5(10):13–34, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afk

[Ano68-143] Anonymous. [advertisement]. IEEE Spectrum, 5(10):37–38, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afm

[Ano68-144] Anonymous. [advertisement]. IEEE Spectrum, 5(10):73–76, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afn

[Ano68-145] Anonymous. [advertisement]. IEEE Spectrum, 5(10):108, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afo

[Ano68-146] Anonymous. [advertisement]. IEEE Spectrum, 5(10):125, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afp

[Ano68-147] Anonymous. [advertisement]. IEEE Spectrum, 5(10):131, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afq

[Ano68-148] Anonymous. [advertisement].

IEEE Spectrum, 5(10):137– 139, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afr

[Ano68-149] Anonymous. [advertisement].

IEEE Spectrum, 5(10):nil2, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afs

[Ano68-150] Anonymous. [advertisement].

IEEE Spectrum, 5(11):nil1, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aft

[Ano68-151] Anonymous. [advertisement].

IEEE Spectrum, 5(11):3, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afu

[Ano68-152] Anonymous. [advertisement].

IEEE Spectrum, 5(11):6, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afv

[Ano68-153] Anonymous. [advertisement].

IEEE Spectrum, 5(11):7–8, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afw

[Ano68-154] Anonymous. [advertisement]. IEEE Spectrum, 5(11):13–40, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afx

[Ano68-155] Anonymous. [advertisement]. IEEE Spectrum, 5(11):43, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afy

[Ano68-156] Anonymous. [advertisement]. IEEE Spectrum, 5(11):45–49, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aga

[Ano68-157] Anonymous. [advertisement]. IEEE Spectrum, 5(11):95–98, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agb

[Ano68-158] Anonymous. [advertisement]. IEEE Spectrum, 5(11):140–147, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agc

[Ano68-159] Anonymous. [advertisement]. IEEE Spectrum, 5(11):157, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agd

[Ano68-160] Anonymous. [advertisement].

IEEE Spectrum, 5(11):159, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Age

[Ano68-161] Anonymous. [advertisement].

IEEE Spectrum, 5(11):162, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agf

[Ano68-162] Anonymous. [advertisement].

IEEE Spectrum, 5(11):166–167, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agg

[Ano68-163] Anonymous. [advertisement].

IEEE Spectrum, 5(11):nil2, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agh

[Ano68-164] Anonymous. [advertisement].

IEEE Spectrum, 5(12):nil1, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agi

[Ano68-165] Anonymous. [advertisement].

IEEE Spectrum, 5(12):3, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agj

[Ano68-166] Anonymous. [advertisement]. IEEE Spectrum, 5(12):6, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agk

[Ano68-167] Anonymous. [advertisement]. IEEE Spectrum, 5(12):7–8, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agl

[Ano68-168] Anonymous. [advertisement]. IEEE Spectrum, 5(12):10–11, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agm

[Ano68-169] Anonymous. [advertisement]. IEEE Spectrum, 5(12):13–22, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agn

[Ano68-170] Anonymous. [advertisement]. IEEE Spectrum, 5(12):25–26, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agp

[Ano68-171] Anonymous. [advertisement]. IEEE Spectrum, 5(12):73–76, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agq

[Ano68-172] Anonymous. [advertisement].

IEEE Spectrum, 5(12):89–107, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agr

[Ano68-173] Anonymous. [advertisement].

IEEE Spectrum, 5(12):109, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ags

[Ano68-174] Anonymous. [advertisement].

IEEE Spectrum, 5(12):114–115, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Agt

[Ano68-175] Anonymous. [advertisement].

IEEE Spectrum, 5(12):nil2, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AVC

[Ano68-176] Anonymous. The anatomy of vibrating crystals. IEEE Spectrum, 5(4):13, April 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:AIa

[Ano68-177] Anonymous. Annual index.

IEEE Spectrum, 5(12):125–144, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aj

[Ano68-178] Anonymous. Authors. IEEE Spectrum, 5(1):44–45, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aba

[Ano68-179] Anonymous. Authors. IEEE Spectrum, 5(2):44, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Abt

[Ano68-180] Anonymous. Authors. IEEE Spectrum, 5(3):60–61, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acg

[Ano68-181] Anonymous. Authors. IEEE Spectrum, 5(4):46, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Acv

[Ano68-182] Anonymous. Authors. IEEE Spectrum, 5(5):42–43, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Adm

[Ano68-183] Anonymous. Authors. IEEE Spectrum, 5(6):38–39, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ady

[Ano68-184] Anonymous. Authors. IEEE Spectrum, 5(7):42–43, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aen

[Ano68-185] Anonymous. Authors. IEEE Spectrum, 5(8):32–33, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Aey

[Ano68-186] Anonymous. Authors. IEEE Spectrum, 5(9):38–39, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afl

[Ano68-187] Anonymous. Authors. IEEE Spectrum, 5(10):43–44, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Afz

[Ano68-188] Anonymous. Authors. IEEE Spectrum, 5(11):62, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ago

[Ano68-189] Anonymous. Authors. IEEE Spectrum, 5(12):32, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ACP

[Ano68-190] Anonymous. Automation comes to the printing and publishing industry newspaper manufacture and distribution; communications. IEEE Spectrum, 5(4):49–62, April 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRa

[Ano68-191] Anonymous. Book reviews. IEEE Spectrum, 5(1):132–141, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRb

[Ano68-192] Anonymous. Book reviews. IEEE Spectrum, 5(2):142–149, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRc

[Ano68-193] Anonymous. Book reviews. IEEE Spectrum, 5(3):136–147, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRd

[Ano68-194] Anonymous. Book reviews. IEEE Spectrum, 5(4):162–169, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRe

[Ano68-195] Anonymous. Book reviews. IEEE Spectrum, 5(5):128–133, May 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRf

[Ano68-196] Anonymous. Book reviews.

IEEE Spectrum, 5(6):130–134, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRg

[Ano68-197] Anonymous. Book reviews.

IEEE Spectrum, 5(7):136–141, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRh

[Ano68-198] Anonymous. Book reviews.

IEEE Spectrum, 5(8):125–131, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRi

[Ano68-199] Anonymous. Book reviews.

IEEE Spectrum, 5(9):138–143, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRj

[Ano68-200] Anonymous. Book reviews.

IEEE Spectrum, 5(10):136– 143, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRk

[Ano68-201] Anonymous. Book reviews.

IEEE Spectrum, 5(11):168–171, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BRl

[Ano68-202] Anonymous. Book reviews. IEEE Spectrum, 5(12):122–124, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISa

[Ano68-203] Anonymous. IEEE Spectrum. IEEE Spectrum, 5(1):6, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISb

[Ano68-204] Anonymous. IEEE Spectrum. IEEE Spectrum, 5(2):6, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISc

[Ano68-205] Anonymous. IEEE Spectrum. IEEE Spectrum, 5(3):6, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISd

[Ano68-206] Anonymous. IEEE Spectrum. IEEE Spectrum, 5(4):6, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISe

[Ano68-207] Anonymous. IEEE Spectrum. IEEE Spectrum, 5(5):6, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISf

[Ano68-208] Anonymous. IEEE Spectrum.

IEEE Spectrum, 5(6):6, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISg

[Ano68-209] Anonymous. IEEE Spectrum.

IEEE Spectrum, 5(7):6, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISh

[Ano68-210] Anonymous. IEEE Spectrum.

IEEE Spectrum, 5(8):6, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISi

[Ano68-211] Anonymous. IEEE Spectrum.

IEEE Spectrum, 5(9):6, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISj

[Ano68-212] Anonymous. IEEE Spectrum.

IEEE Spectrum, 5(10):6, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISk

[Ano68-213] Anonymous. IEEE Spectrum.

IEEE Spectrum, 5(11):6, November 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ISl

[Ano68-214] Anonymous. IEEE Spectrum. IEEE Spectrum, 5(12):6, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BMW

[Ano68-215] Anonymous. Breadboarding the modern way. IEEE Spectrum, 5(5):11, May 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPa

[Ano68-216] Anonymous. [breaker page]. IEEE Spectrum, 5(1):4, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPb

[Ano68-217] Anonymous. [breaker page]. IEEE Spectrum, 5(2):4, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPc

[Ano68-218] Anonymous. [breaker page]. IEEE Spectrum, 5(3):4, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPd

[Ano68-219] Anonymous. [breaker page]. IEEE Spectrum, 5(4):4, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPe

[Ano68-220] Anonymous. [breaker page].

IEEE Spectrum, 5(5):4, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPf

[Ano68-221] Anonymous. [breaker page].

IEEE Spectrum, 5(6):4, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPg

[Ano68-222] Anonymous. [breaker page].

IEEE Spectrum, 5(7):4, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPh

[Ano68-223] Anonymous. [breaker page].

IEEE Spectrum, 5(8):4, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPi

[Ano68-224] Anonymous. [breaker page].

IEEE Spectrum, 5(9):4, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPj

[Ano68-225] Anonymous. [breaker page].

IEEE Spectrum, 5(10):4, October 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPk

[Ano68-226] Anonymous. [breaker page]. IEEE Spectrum, 5(11):4, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BPl

[Ano68-227] Anonymous. [breaker page]. IEEE Spectrum, 5(12):4, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ca

[Ano68-228] Anonymous. Calendar. IEEE Spectrum, 5(1):33–35, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Cb

[Ano68-229] Anonymous. Calendar. IEEE Spectrum, 5(2):37–39, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Cc

[Ano68-230] Anonymous. Calendar. IEEE Spectrum, 5(3):51–53, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Cd

[Ano68-231] Anonymous. Calendar. IEEE Spectrum, 5(4):35–37, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ce

[Ano68-232] Anonymous. Calendar. IEEE Spectrum, 5(5):33–34, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Cf

[Ano68-233] Anonymous. Calendar. IEEE Spectrum, 5(6):25–26, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Cg

[Ano68-234] Anonymous. Calendar. IEEE Spectrum, 5(7):31–32, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ch

[Ano68-235] Anonymous. Calendar. IEEE Spectrum, 5(8):21–22, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ci

[Ano68-236] Anonymous. Calendar. IEEE Spectrum, 5(9):27–28, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Cj

[Ano68-237] Anonymous. Calendar. IEEE Spectrum, 5(10):35–36, October 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ck

[Ano68-238] Anonymous. Calendar. IEEE Spectrum, 5(11):41–42, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Cm

[Ano68-239] Anonymous. Calendar. IEEE Spectrum, 5(12):23–24, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CL

[Ano68-240] Anonymous. Ceramics from liquids. IEEE Spectrum, 5(6):28, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CRE

[Ano68-241] Anonymous. Challenge and a receptive ear [Gulf General Atomic, Inc. — job offerings]. IEEE Spectrum, 5(4):118, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAa

[Ano68-242] Anonymous. Classified advertising. IEEE Spectrum, 5(3): 131–135, March 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAb

[Ano68-243] Anonymous. Classified advertising. IEEE Spectrum, 5(4):

157–159, April 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAc

[Ano68-244] Anonymous. Classified advertising. IEEE Spectrum, 5(5):121, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAd

[Ano68-245] Anonymous. Classified advertising. IEEE Spectrum, 5(6):129, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAe

[Ano68-246] Anonymous. Classified advertising. IEEE Spectrum, 5(7):131, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAf

[Ano68-247] Anonymous. Classified advertising. IEEE Spectrum, 5(8):117, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAg

[Ano68-248] Anonymous. Classified advertising. IEEE Spectrum, 5(9): 137, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAh

[Ano68-249] Anonymous. Classified advertising. IEEE Spectrum, 5 (10):135, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAi

[Ano68-250] Anonymous. Classified advertising. IEEE Spectrum, 5(11): 169–173, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:CAj

[Ano68-251] Anonymous. Classified advertising. IEEE Spectrum, 5(12): 121, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Cn

[Ano68-252] Anonymous. Correction. IEEE Spectrum, 5(12):60, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:DLB

[Ano68-253] Anonymous. Delay-line bandwidth reducer. IEEE Spectrum, 5(11):5, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:DSR

[Ano68-254] Anonymous. Directions for speech research. IEEE Spectrum, 5(3):83–88, March 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:EP

[Ano68-255] Anonymous. Equations for plowing. IEEE Spectrum, 5 (1):17, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:EA

[Ano68-256] Anonymous. Exhibitors [advertisement]. IEEE Spectrum, 5 (3):149–249, March 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPa

[Ano68-257] Anonymous. Focal points. IEEE Spectrum, 5(1):117–122, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPb

[Ano68-258] Anonymous. Focal points. IEEE Spectrum, 5(2):134–137, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPc

[Ano68-259] Anonymous. Focal points. IEEE Spectrum, 5(3):125–128, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPd

[Ano68-260] Anonymous. Focal points. IEEE Spectrum, 5(4):154–157, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPe

[Ano68-261] Anonymous. Focal points. IEEE Spectrum, 5(5):118–120, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPf

[Ano68-262] Anonymous. Focal points. IEEE Spectrum, 5(6):122–126, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPg

[Ano68-263] Anonymous. Focal points. IEEE Spectrum, 5(7):128–130, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPh

[Ano68-264] Anonymous. Focal points. IEEE Spectrum, 5(8):118–120, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPi

[Ano68-265] Anonymous. Focal points. IEEE Spectrum, 5(9):128–131, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPj

[Ano68-266] Anonymous. Focal points. IEEE Spectrum, 5(10):128–130, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPk

[Ano68-267] Anonymous. Focal points. IEEE Spectrum, 5(11):160–164, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FPl

[Ano68-268] Anonymous. Focal points. IEEE Spectrum, 5(12):112–117, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCa

[Ano68-269] Anonymous. [front cover]. IEEE Spectrum, 5(1):c1, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCb

[Ano68-270] Anonymous. [front cover]. IEEE Spectrum, 5(1):c2, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCc

[Ano68-271] Anonymous. [front cover].

IEEE Spectrum, 5(2):c1, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCd

[Ano68-272] Anonymous. [front cover].

IEEE Spectrum, 5(2):c2, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCe

[Ano68-273] Anonymous. [front cover]. IEEE Spectrum, 5(3):c1, March 1968.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCf

[Ano68-274] Anonymous. [front cover]. IEEE Spectrum, 5(3):c2, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCg

[Ano68-275] Anonymous. [front cover]. IEEE Spectrum, 5(4):c1, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCh

[Ano68-276] Anonymous. [front cover]. IEEE Spectrum, 5(4):c2, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCi

[Ano68-277] Anonymous. [front cover]. IEEE Spectrum, 5(5):c1, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCj

[Ano68-278] Anonymous. [front cover]. IEEE Spectrum, 5(5):c2, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCk

[Ano68-279] Anonymous. [front cover]. IEEE Spectrum, 5(6):c1, June 1968.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCl

[Ano68-280] Anonymous. [front cover]. IEEE Spectrum, 5(6):c2, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCm

[Ano68-281] Anonymous. [front cover]. IEEE Spectrum, 5(7):c1, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCn

[Ano68-282] Anonymous. [front cover]. IEEE Spectrum, 5(7):c2, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCo

[Ano68-283] Anonymous. [front cover]. IEEE Spectrum, 5(8):c1, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCp

[Ano68-284] Anonymous. [front cover]. IEEE Spectrum, 5(8):c2, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCq

[Ano68-285] Anonymous. [front cover]. IEEE Spectrum, 5(9):c1, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCr

[Ano68-286] Anonymous. [front cover]. IEEE Spectrum, 5(9):c2, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCs

[Ano68-287] Anonymous. [front cover]. IEEE Spectrum, 5(10):c1, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCt

[Ano68-288] Anonymous. [front cover]. IEEE Spectrum, 5(10):c2, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCu

[Ano68-289] Anonymous. [front cover]. IEEE Spectrum, 5(11):c1, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCv

[Ano68-290] Anonymous. [front cover]. IEEE Spectrum, 5(11):c2, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCw

[Ano68-291] Anonymous. [front cover]. IEEE Spectrum, 5(12):c1, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FCx

[Ano68-292] Anonymous. [front cover]. IEEE Spectrum, 5(12):c2, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:FWE

[Ano68-293] Anonymous. Future Wescon exhibitors [advertisement]. IEEE Spectrum, 5(8):132–133, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:GCM

[Ano68-294] Anonymous. The generation of correlated multivariate samples for Monte Carlo simulation. IEEE Spectrum, 5(2):5, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:GEP

[Ano68-295] Anonymous. Graphical evaluation of polynomial functions of an imaginary variable. IEEE Spectrum, 5(3):5, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:HSIa

[Ano68-296] Anonymous. High-sensitivity infrared (10.6 micron) heterodyne receivers with gigahertz IF capability. Part I: Sensitivity and bandwidth. IEEE Spectrum, 5(6):5, June 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:HSIb

[Ano68-297] Anonymous. High-sensitivity infrared 10.6-micron heterodyne receivers with large IF bandwidth. Part II: Receiver development. IEEE Spectrum, 5(7): 5, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IFAa

[Ano68-298] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 5(1):14, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IFAb

[Ano68-299] Anonymous. IEEE forum [advertisement]. IEEE Spectrum, 5(2):12, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IJD

[Ano68-300] Anonymous. IEEE Japan Directory 1969 [advertisement]. IEEE Spectrum, 5(11):50–185, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IPS

[Ano68-301] Anonymous. IEEE position statement on the ASEE Goals of Engineering Education Report. IEEE Spectrum, 5(12): 81–83, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAa

[Ano68-302] Anonymous. Index to advertisers. IEEE Spectrum, 5(1): 142–144, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAb

[Ano68-303] Anonymous. Index to advertisers. IEEE Spectrum, 5(2): 150–152, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAc

[Ano68-304] Anonymous. Index to advertisers. IEEE Spectrum, 5(3): 254–260, March 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAd

[Ano68-305] Anonymous. Index to advertisers. IEEE Spectrum, 5(4): 170–172, April 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAe

[Ano68-306] Anonymous. Index to advertisers. IEEE Spectrum, 5 (5):134–136, May 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAf

[Ano68-307] Anonymous. Index to advertisers. IEEE Spectrum, 5 (6):135–136, June 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAg

[Ano68-308] Anonymous. Index to advertisers. IEEE Spectrum, 5 (7):150–152, July 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAh

[Ano68-309] Anonymous. Index to advertisers. IEEE Spectrum, 5(8): 134–136, August 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAi

[Ano68-310] Anonymous. Index to advertisers. IEEE Spectrum, 5(9): 144–146, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAj

[Ano68-311] Anonymous. Index to advertisers. IEEE Spectrum, 5(10): 144–146, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAk

[Ano68-312] Anonymous. Index to advertisers. IEEE Spectrum, 5(11): 186–188, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAl

[Ano68-313] Anonymous. Index to advertisers. IEEE Spectrum, 5(12):

145–146, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IAS

[Ano68-314] Anonymous. Information about the society. IEEE Spectrum, 5 (3):250–253, March 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:BFS

[Ano68-315] Anonymous. L-band frequency synthesizer. IEEE Spectrum, 5 (9):5, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:LC

[Ano68-316] Anonymous. List of contributors. IEEE Spectrum, 5(5):36–38, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:ME

[Ano68-317] Anonymous. Microbridges for electrons. IEEE Spectrum, 5(7): 8, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:MF

[Ano68-318] Anonymous. A monolithic filter. IEEE Spectrum, 5(8):11, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:MC

[Ano68-319] Anonymous. Movies via computer. IEEE Spectrum, 5(3):189, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NUS

[Ano68-320] Anonymous. A new use for the sun. IEEE Spectrum, 5 (11):11, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIa

[Ano68-321] Anonymous. News of the IEEE. IEEE Spectrum, 5(1):16–30, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIb

[Ano68-322] Anonymous. News of the IEEE. IEEE Spectrum, 5(2): 14–34, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIc

[Ano68-323] Anonymous. News of the IEEE. IEEE Spectrum, 5(3):38–48, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NId

[Ano68-324] Anonymous. News of the IEEE. IEEE Spectrum, 5(4):12–32, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIe

[Ano68-325] Anonymous. News of the IEEE. IEEE Spectrum, 5(5):12–31, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIf

[Ano68-326] Anonymous. News of the IEEE.

IEEE Spectrum, 5(6):10–23, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIg

[Ano68-327] Anonymous. News of the IEEE.

IEEE Spectrum, 5(7):10–28, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIh

[Ano68-328] Anonymous. News of the IEEE.

IEEE Spectrum, 5(8):10–20, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIi

[Ano68-329] Anonymous. News of the IEEE. IEEE Spectrum, 5(9): 10–25, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIj

[Ano68-330] Anonymous. News of the IEEE. IEEE Spectrum, 5(10): 10–33, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIk

[Ano68-331] Anonymous. News of the IEEE. IEEE Spectrum, 5(11):

10–38, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:NIl

[Ano68-332] Anonymous. News of the IEEE. IEEE Spectrum, 5(12): 12–21, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Oa

[Ano68-333] Anonymous. Obituaries. IEEE Spectrum, 5(1):40–41, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ob

[Ano68-334] Anonymous. Obituaries. IEEE Spectrum, 5(2):42, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Oc

[Ano68-335] Anonymous. Obituaries. IEEE Spectrum, 5(3):57, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Od

[Ano68-336] Anonymous. Obituaries. IEEE Spectrum, 5(4):42–43, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Oe

[Ano68-337] Anonymous. Obituaries. IEEE Spectrum, 5(5):38–39, May 1968.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Of

[Ano68-338] Anonymous. Obituaries. IEEE Spectrum, 5(6):33, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Og

[Ano68-339] Anonymous. Obituaries. IEEE Spectrum, 5(7):39, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Oh

[Ano68-340] Anonymous. Obituaries. IEEE Spectrum, 5(8):28–29, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Oi

[Ano68-341] Anonymous. Obituaries. IEEE Spectrum, 5(9):35, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Oj

[Ano68-342] Anonymous. Obituaries. IEEE Spectrum, 5(10):41, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ok

[Ano68-343] Anonymous. Obituaries. IEEE Spectrum, 5(11):47, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ol

[Ano68-344] Anonymous. Obituaries. IEEE Spectrum, 5(12):30, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pa

[Ano68-345] Anonymous. People. IEEE Spectrum, 5(1):36–40, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pb

[Ano68-346] Anonymous. People. IEEE Spectrum, 5(2):40–42, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pc

[Ano68-347] Anonymous. People. IEEE Spectrum, 5(3):54–57, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pd

[Ano68-348] Anonymous. People. IEEE Spectrum, 5(4):38–42, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pe

[Ano68-349] Anonymous. People. IEEE Spectrum, 5(6):29–33, June 1968.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pf

[Ano68-350] Anonymous. People. IEEE Spectrum, 5(7):34–39, July 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pg

[Ano68-351] Anonymous. People. IEEE Spectrum, 5(8):24–28, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Ph

[Ano68-352] Anonymous. People. IEEE Spectrum, 5(9):30–35, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pi

[Ano68-353] Anonymous. People. IEEE Spectrum, 5(10):39–41, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pj

[Ano68-354] Anonymous. People. IEEE Spectrum, 5(11):44–47, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:Pk

[Ano68-355] Anonymous. People. IEEE Spectrum, 5(12):27–30, December 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PDM

[Ano68-356] Anonymous. A practical detector for the millimeter-to-infrared “gap”. IEEE Spectrum, 5 (10):5, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PIM

[Ano68-357] Anonymous. Procedure for IEEE membership application. IEEE Spectrum, 5(11):29–32, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSa

[Ano68-358] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(1):143–144, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSb

[Ano68-359] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(2):151–152, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSc

[Ano68-360] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(3):255–260, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSd

[Ano68-361] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(4):172, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSe

[Ano68-362] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(5):135–136, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSf

[Ano68-363] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(6):135–136, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSg

[Ano68-364] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(7):151–152, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSh

[Ano68-365] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(8):135–136, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSi

[Ano68-366] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(9):145–146, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSj

[Ano68-367] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(10):145– 146, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSk

[Ano68-368] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(11):187, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PCSl

[Ano68-369] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 5(12):146, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PHS

[Ano68-370] Anonymous. Progress in HF superconducting filters. IEEE Spectrum, 5(8):5, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:PNA

[Ano68-371] Anonymous. Proper-name antennas. IEEE Spectrum, 5(5):5, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SET

[Ano68-372] Anonymous. Satellite educational television for underdeveloped countries. IEEE Spectrum, 5(12):5, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIa

[Ano68-373] Anonymous. Scanning the issues.

IEEE Spectrum, 5(1):96–97, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIb

[Ano68-374] Anonymous. Scanning the issues. IEEE Spectrum, 5(2): 110–111, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIc

[Ano68-375] Anonymous. Scanning the issues.

IEEE Spectrum, 5(3):104–105, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SId

[Ano68-376] Anonymous. Scanning the issues.

IEEE Spectrum, 5(4):119–120, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIe

[Ano68-377] Anonymous. Scanning the issues. IEEE Spectrum, 5(5):94–95, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIf

[Ano68-378] Anonymous. Scanning the issues. IEEE Spectrum, 5(6): 96–97, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIg

[Ano68-379] Anonymous. Scanning the issues. IEEE Spectrum, 5(7):106–107, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIh

[Ano68-380] Anonymous. Scanning the issues. IEEE Spectrum, 5(8): 99–100, August 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIi

[Ano68-381] Anonymous. Scanning the issues. IEEE Spectrum, 5(9): 106–107, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIj

[Ano68-382] Anonymous. Scanning the issues. IEEE Spectrum, 5(10): 102–103, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIk

[Ano68-383] Anonymous. Scanning the issues. IEEE Spectrum, 5(11): 135–136, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SIl

[Ano68-384] Anonymous. Scanning the issues. IEEE Spectrum, 5(12): 84–85, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SCT

[Ano68-385] Anonymous. See a computer talk. IEEE Spectrum, 5(2): 15, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPL

[Ano68-386] Anonymous. “Self-portrait” of a laser signal. IEEE Spectrum, 5(10):11, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SCE

[Ano68-387] Anonymous. Short courses in electrical and electronics engineering fall 1968. IEEE Spectrum, 5(7):91–94, July 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPa

[Ano68-388] Anonymous. Special publications. IEEE Spectrum, 5 (1):114, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPb

[Ano68-389] Anonymous. Special publications. IEEE Spectrum, 5 (2):132, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPc

[Ano68-390] Anonymous. Special publications. IEEE Spectrum, 5(3):124, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPd

[Ano68-391] Anonymous. Special publications. IEEE Spectrum, 5(4):153, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPe

[Ano68-392] Anonymous. Special publications. IEEE Spectrum, 5(5):116, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPf

[Ano68-393] Anonymous. Special publications. IEEE Spectrum, 5(6):121, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPg

[Ano68-394] Anonymous. Special publications. IEEE Spectrum, 5(7):126, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPh

[Ano68-395] Anonymous. Special publications. IEEE Spectrum, 5(8):116, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPi

[Ano68-396] Anonymous. Special publications. IEEE Spectrum, 5(9): 127, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPj

[Ano68-397] Anonymous. Special publications. IEEE Spectrum, 5 (10):127, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPk

[Ano68-398] Anonymous. Special publications. IEEE Spectrum, 5(11): 158, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SPm

[Ano68-399] Anonymous. Special publications. IEEE Spectrum, 5(12): 111, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLa

[Ano68-400] Anonymous. Spectral lines.

IEEE Spectrum, 5(1):43, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLb

[Ano68-401] Anonymous. Spectral lines. IEEE Spectrum, 5(2):43, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLc

[Ano68-402] Anonymous. Spectral lines. IEEE Spectrum, 5(3):59, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLd

[Ano68-403] Anonymous. Spectral lines. IEEE Spectrum, 5(4):45, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLe

[Ano68-404] Anonymous. Spectral lines. IEEE Spectrum, 5(5):41, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLf

[Ano68-405] Anonymous. Spectral lines. IEEE Spectrum, 5(6):37, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLg

[Ano68-406] Anonymous. Spectral lines. IEEE Spectrum, 5(7):41, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLh

[Ano68-407] Anonymous. Spectral lines.

IEEE Spectrum, 5(8):31, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLi

[Ano68-408] Anonymous. Spectral lines.

IEEE Spectrum, 5(9):37, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLj

[Ano68-409] Anonymous. Spectral lines.

IEEE Spectrum, 5(10):42, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLk

[Ano68-410] Anonymous. Spectral lines.

IEEE Spectrum, 5(11):61, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SLl

[Ano68-411] Anonymous. Spectral lines.

IEEE Spectrum, 5(12):31, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SRSa

[Ano68-412] Anonymous. Spectrum reader service card. IEEE Spectrum, 5 (11):1a–1b, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SRSb

[Ano68-413] Anonymous. Spectrum reader service card. IEEE Spectrum, 5(11):189–190, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:SMN

[Ano68-414] Anonymous. The spiral map: A new form of Karnaugh map. IEEE Spectrum, 5(1):5, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCa

[Ano68-415] Anonymous. Table of contents. IEEE Spectrum, 5(1):1–2, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCc

[Ano68-416] Anonymous. Table of contents. IEEE Spectrum, 5(2):1–2, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCe

[Ano68-417] Anonymous. Table of contents. IEEE Spectrum, 5(3):1–2, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCg

[Ano68-418] Anonymous. Table of contents. IEEE Spectrum, 5(4):1–2, April 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCi

[Ano68-419] Anonymous. Table of contents. IEEE Spectrum, 5(5):1– 2, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCk

[Ano68-420] Anonymous. Table of contents. IEEE Spectrum, 5(6):1–2, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCm

[Ano68-421] Anonymous. Table of contents. IEEE Spectrum, 5(7):1– 2, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCo

[Ano68-422] Anonymous. Table of contents.

IEEE Spectrum, 5(8):1–2, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCq

[Ano68-423] Anonymous. Table of contents. IEEE Spectrum, 5(9): 1–2, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCs

[Ano68-424] Anonymous. Table of contents.

IEEE Spectrum, 5(10):1–2, October 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCu

[Ano68-425] Anonymous. Table of contents. IEEE Spectrum, 5(11): 1c–2, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCw

[Ano68-426] Anonymous. Table of contents. IEEE Spectrum, 5(12):1–2, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCb

[Ano68-427] Anonymous. Technical correspondence. IEEE Spectrum, 5 (1):124–130, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCd

[Ano68-428] Anonymous. Technical correspondence. IEEE Spectrum, 5 (2):138–140, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCf

[Ano68-429] Anonymous. Technical correspondence. IEEE Spectrum, 5 (3):130–134, March 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCh

[Ano68-430] Anonymous. Technical correspondence. IEEE Spectrum, 5 (4):159–161, April 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCj

[Ano68-431] Anonymous. Technical correspondence. IEEE Spectrum, 5 (5):122–126, May 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCl

[Ano68-432] Anonymous. Technical correspondence. IEEE Spectrum, 5 (6):127–129, June 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCn

[Ano68-433] Anonymous. Technical correspondence. IEEE Spectrum, 5 (7):132–135, July 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCp

[Ano68-434] Anonymous. Technical correspondence. IEEE Spectrum, 5 (8):121–124, August 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCr

[Ano68-435] Anonymous. Technical correspondence. IEEE Spectrum, 5(9): 132–136, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCt

[Ano68-436] Anonymous. Technical correspondence. IEEE Spectrum, 5 (10):132–134, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCv

[Ano68-437] Anonymous. Technical correspondence. IEEE Spectrum, 5 (11):165–167, November 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TCx

[Ano68-438] Anonymous. Technical correspondence. IEEE Spectrum, 5(12):118–121, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TP

[Ano68-439] Anonymous. Technical program. IEEE Spectrum, 5(3): 13–36, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEa

[Ano68-440] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(1):4, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEb

[Ano68-441] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(2):4, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEc

[Ano68-442] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(3):4, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEd

[Ano68-443] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(4): 4, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEe

[Ano68-444] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(5): 4, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEf

[Ano68-445] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(6): 4, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEg

[Ano68-446] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(7): 4, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEh

[Ano68-447] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(8):4, August 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEi

[Ano68-448] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5 (9):4, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEj

[Ano68-449] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5(10):4, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEk

[Ano68-450] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5 (11):4, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:IEEl

[Ano68-451] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 5 (12):4, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTa

[Ano68-452] Anonymous. Transient and trends. IEEE Spectrum, 5 (1):11, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTb

[Ano68-453] Anonymous. Transients and trends. IEEE Spectrum, 5 (2):11, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTc

[Ano68-454] Anonymous. Transients and trends. IEEE Spectrum, 5(3):10, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTd

[Ano68-455] Anonymous. Transients and trends. IEEE Spectrum, 5(4):8, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTe

[Ano68-456] Anonymous. Transients and trends. IEEE Spectrum, 5(5):10, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTf

[Ano68-457] Anonymous. Transients and trends. IEEE Spectrum, 5(6):9, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTg

[Ano68-458] Anonymous. Transients and trends. IEEE Spectrum, 5(7):9, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTh

[Ano68-459] Anonymous. Transients and trends. IEEE Spectrum, 5(8):9, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTi

[Ano68-460] Anonymous. Transients and trends. IEEE Spectrum, 5 (9):9, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTj

[Ano68-461] Anonymous. Transients and trends. IEEE Spectrum, 5 (10):9, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTk

[Ano68-462] Anonymous. Transients and trends. IEEE Spectrum, 5 (11):9, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TTl

[Ano68-463] Anonymous. Transients and trends. IEEE Spectrum, 5 (12):9, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJa

[Ano68-464] Anonymous. Translated journals. IEEE Spectrum, 5(1): 108–112, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJb

[Ano68-465] Anonymous. Translated journals. IEEE Spectrum, 5(2): 128–130, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJc

[Ano68-466] Anonymous. Translated journals. IEEE Spectrum, 5(3): 120–122, March 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJd

[Ano68-467] Anonymous. Translated journals. IEEE Spectrum, 5(4): 150–152, April 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJe

[Ano68-468] Anonymous. Translated journals. IEEE Spectrum, 5(5): 112–114, May 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJf

[Ano68-469] Anonymous. Translated journals. IEEE Spectrum, 5(6): 118–120, June 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJg

[Ano68-470] Anonymous. Translated journals. IEEE Spectrum, 5(7): 124–125, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJh

[Ano68-471] Anonymous. Translated journals. IEEE Spectrum, 5(8): 114–115, August 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJi

[Ano68-472] Anonymous. Translated journals. IEEE Spectrum, 5(9): 124–126, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJj

[Ano68-473] Anonymous. Translated journals. IEEE Spectrum, 5(10): 124–126, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJk

[Ano68-474] Anonymous. Translated journals. IEEE Spectrum, 5(11): 155–157, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:TJl

[Ano68-475] Anonymous. Translated journals. IEEE Spectrum, 5(12): 108–110, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:VA

[Ano68-476] Anonymous. The variety of antennas. IEEE Spectrum, 5(4):5, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1968:WEL

[Ano68-477] Anonymous. WESCON 68 Exhibitors list. IEEE Spectrum, 5(7):142–149, July 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1969:ATCa

[Ano69a] Anonymous. Advance tables of contents. IEEE Spectrum, 6(1): 100–105, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCb

[Ano69b] Anonymous. Advance tables of contents. IEEE Spectrum, 6(2): 81–84, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCc

[Ano69c] Anonymous. Advance tables of contents. IEEE Spectrum, 6(3): 122–127, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCd

[Ano69d] Anonymous. Advance tables of contents. IEEE Spectrum, 6(4): 98–104, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCe

[Ano69e] Anonymous. Advance tables of contents. IEEE Spectrum, 6(5): 108–111, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCf

[Ano69f] Anonymous. Advance tables of contents. IEEE Spectrum, 6(6): 102–107, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCg

[Ano69g] Anonymous. Advance tables of contents. IEEE Spectrum, 6(7): 97–101, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCh

[Ano69h] Anonymous. Advance tables of contents. IEEE Spectrum, 6(8): 84–88, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCi

[Ano69i] Anonymous. Advance tables of contents. IEEE Spectrum, 6(9): 95–98, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCj

[Ano69j] Anonymous. Advance tables of contents. IEEE Spectrum, 6(10): 98–102, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCk

[Ano69k] Anonymous. Advance tables of contents. IEEE Spectrum, 6(11): 100–106, November 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ATCl

[Ano69l] Anonymous. Advance tables of contents. IEEE Spectrum, 6(12): 99–103, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aa

[Ano69m] Anonymous. [advertisement].

IEEE Spectrum, 6(1):nil1, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ab

[Ano69n] Anonymous. [advertisement].

IEEE Spectrum, 6(1):1, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ac

[Ano69o] Anonymous. [advertisement].

IEEE Spectrum, 6(1):6–7, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ad

[Ano69p] Anonymous. [advertisement].

IEEE Spectrum, 6(1):9–10, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ae

[Ano69q] Anonymous. [advertisement].

IEEE Spectrum, 6(1):11–12, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Af

[Ano69r] Anonymous. [advertisement].

IEEE Spectrum, 6(1):16–24, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ag

[Ano69s] Anonymous. [advertisement].

IEEE Spectrum, 6(1):27–28, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ah

[Ano69t] Anonymous. [advertisement]. IEEE Spectrum, 6(1):75–77, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ai

[Ano69u] Anonymous. [advertisement]. IEEE Spectrum, 6(1):99, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aj

[Ano69v] Anonymous. [advertisement]. IEEE Spectrum, 6(1):111–116, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ak

[Ano69w] Anonymous. [advertisement]. IEEE Spectrum, 6(1):125, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Al

[Ano69x] Anonymous. [advertisement]. IEEE Spectrum, 6(1):127–142, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Am

[Ano69y] Anonymous. [advertisement]. IEEE Spectrum, 6(1):144–145, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:An

[Ano69z] Anonymous. [advertisement]. IEEE Spectrum, 6(1):nil2, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ao

[Ano69-27] Anonymous. [advertisement].

IEEE Spectrum, 6(2):nil1–1, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ap

[Ano69-28] Anonymous. [advertisement].

IEEE Spectrum, 6(2):15–16, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aq

[Ano69-29] Anonymous. [advertisement].

IEEE Spectrum, 6(2):61–64, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ar

[Ano69-30] Anonymous. [advertisement].

IEEE Spectrum, 6(2):83–85, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:As

[Ano69-31] Anonymous. [advertisement].

IEEE Spectrum, 6(2):87, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:At

[Ano69-32] Anonymous. [advertisement].

IEEE Spectrum, 6(2):91–99, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Au

[Ano69-33] Anonymous. [advertisement].

IEEE Spectrum, 6(2):102–113, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Av

[Ano69-34] Anonymous. [advertisement]. IEEE Spectrum, 6(2):115–117, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aw

[Ano69-35] Anonymous. [advertisement]. IEEE Spectrum, 6(2):nil2, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ax

[Ano69-36] Anonymous. [advertisement]. IEEE Spectrum, 6(3):nil1–1, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ay

[Ano69-37] Anonymous. [advertisement]. IEEE Spectrum, 6(3):6–7, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Az

[Ano69-38] Anonymous. [advertisement]. IEEE Spectrum, 6(3):9, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aba

[Ano69-39] Anonymous. [advertisement]. IEEE Spectrum, 6(3):11–14, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abb

[Ano69-40] Anonymous. [advertisement]. IEEE Spectrum, 6(3):21–50, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abc

[Ano69-41] Anonymous. [advertisement].

IEEE Spectrum, 6(3):91–94, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abd

[Ano69-42] Anonymous. [advertisement].

IEEE Spectrum, 6(3):123–127, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abe

[Ano69-43] Anonymous. [advertisement].

IEEE Spectrum, 6(3):129–131, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abf

[Ano69-44] Anonymous. [advertisement].

IEEE Spectrum, 6(3):133, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abg

[Ano69-45] Anonymous. [advertisement].

IEEE Spectrum, 6(3):135–143, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abh

[Ano69-46] Anonymous. [advertisement].

IEEE Spectrum, 6(3):157–162, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abi

[Ano69-47] Anonymous. [advertisement].

IEEE Spectrum, 6(3):242, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abj

[Ano69-48] Anonymous. [advertisement]. IEEE Spectrum, 6(3):nil2, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abk

[Ano69-49] Anonymous. [advertisement]. IEEE Spectrum, 6(4):nil1, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abl

[Ano69-50] Anonymous. [advertisement]. IEEE Spectrum, 6(4):1, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abm

[Ano69-51] Anonymous. [advertisement]. IEEE Spectrum, 6(4):11–20, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abn

[Ano69-52] Anonymous. [advertisement]. IEEE Spectrum, 6(4):23–24, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abo

[Ano69-53] Anonymous. [advertisement]. IEEE Spectrum, 6(4):75–78, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abp

[Ano69-54] Anonymous. [advertisement]. IEEE Spectrum, 6(4):101–105, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abq

[Ano69-55] Anonymous. [advertisement].

IEEE Spectrum, 6(4):107–109, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abr

[Ano69-56] Anonymous. [advertisement].

IEEE Spectrum, 6(4):111, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abs

[Ano69-57] Anonymous. [advertisement].

IEEE Spectrum, 6(4):115–119, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abt

[Ano69-58] Anonymous. [advertisement].

IEEE Spectrum, 6(4):121–144, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abu

[Ano69-59] Anonymous. [advertisement].

IEEE Spectrum, 6(4):146–147, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abv

[Ano69-60] Anonymous. [advertisement].

IEEE Spectrum, 6(4):nil2, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abw

[Ano69-61] Anonymous. [advertisement].

IEEE Spectrum, 6(5):nil1, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abx

[Ano69-62] Anonymous. [advertisement]. IEEE Spectrum, 6(5):1, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aby

[Ano69-63] Anonymous. [advertisement]. IEEE Spectrum, 6(5):7–9, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Abz

[Ano69-64] Anonymous. [advertisement]. IEEE Spectrum, 6(5):11–30, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aca

[Ano69-65] Anonymous. [advertisement]. IEEE Spectrum, 6(5):34, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acb

[Ano69-66] Anonymous. [advertisement]. IEEE Spectrum, 6(5):75–78, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acc

[Ano69-67] Anonymous. [advertisement]. IEEE Spectrum, 6(5):121, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acd

[Ano69-68] Anonymous. [advertisement]. IEEE Spectrum, 6(5):nil2, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ace

[Ano69-69] Anonymous. [advertisement].

IEEE Spectrum, 6(6):nil1, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acf

[Ano69-70] Anonymous. [advertisement].

IEEE Spectrum, 6(6):1b–1, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acg

[Ano69-71] Anonymous. [advertisement].

IEEE Spectrum, 6(6):7, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ach

[Ano69-72] Anonymous. [advertisement].

IEEE Spectrum, 6(6):9–17, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aci

[Ano69-73] Anonymous. [advertisement].

IEEE Spectrum, 6(6):21–24, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acj

[Ano69-74] Anonymous. [advertisement].

IEEE Spectrum, 6(6):71–74, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ack

[Ano69-75] Anonymous. [advertisement].

IEEE Spectrum, 6(6):109, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acl

[Ano69-76] Anonymous. [advertisement]. IEEE Spectrum, 6(6):112–113, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acm

[Ano69-77] Anonymous. [advertisement]. IEEE Spectrum, 6(6):123–129, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acn

[Ano69-78] Anonymous. [advertisement]. IEEE Spectrum, 6(6):133–135, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aco

[Ano69-79] Anonymous. [advertisement]. IEEE Spectrum, 6(6):141–142, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acp

[Ano69-80] Anonymous. [advertisement]. IEEE Spectrum, 6(6):nil2, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acq

[Ano69-81] Anonymous. [advertisement]. IEEE Spectrum, 6(7):nil1, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acr

[Ano69-82] Anonymous. [advertisement]. IEEE Spectrum, 6(7):1, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acs

[Ano69-83] Anonymous. [advertisement].

IEEE Spectrum, 6(7):5–7, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Act

[Ano69-84] Anonymous. [advertisement].

IEEE Spectrum, 6(7):9, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acu

[Ano69-85] Anonymous. [advertisement].

IEEE Spectrum, 6(7):11–22, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acv

[Ano69-86] Anonymous. [advertisement].

IEEE Spectrum, 6(7):25–28, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acw

[Ano69-87] Anonymous. [advertisement].

IEEE Spectrum, 6(7):67–70, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acx

[Ano69-88] Anonymous. [advertisement].

IEEE Spectrum, 6(7):103, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acy

[Ano69-89] Anonymous. [advertisement].

IEEE Spectrum, 6(7):107–113, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Acz

[Ano69-90] Anonymous. [advertisement]. IEEE Spectrum, 6(7):nil2, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ada

[Ano69-91] Anonymous. [advertisement]. IEEE Spectrum, 6(8):nil1, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adb

[Ano69-92] Anonymous. [advertisement]. IEEE Spectrum, 6(8):1, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adc

[Ano69-93] Anonymous. [advertisement]. IEEE Spectrum, 6(8):7–9, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Add

[Ano69-94] Anonymous. [advertisement]. IEEE Spectrum, 6(8):11–18, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ade

[Ano69-95] Anonymous. [advertisement]. IEEE Spectrum, 6(8):22, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adf

[Ano69-96] Anonymous. [advertisement]. IEEE Spectrum, 6(8):59–62, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adg

[Ano69-97] Anonymous. [advertisement].

IEEE Spectrum, 6(8):94–95, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adh

[Ano69-98] Anonymous. [advertisement].

IEEE Spectrum, 6(8):97–103, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adi

[Ano69-99] Anonymous. [advertisement].

IEEE Spectrum, 6(8):105–113, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adj

[Ano69-100] Anonymous. [advertisement].

IEEE Spectrum, 6(8):115, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adk

[Ano69-101] Anonymous. [advertisement].

IEEE Spectrum, 6(8):nil2, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adl

[Ano69-102] Anonymous. [advertisement].

IEEE Spectrum, 6(9):nil1, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adm

[Ano69-103] Anonymous. [advertisement].

IEEE Spectrum, 6(9):1c–1c, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adn

[Ano69-104] Anonymous. [advertisement]. IEEE Spectrum, 6(9):7–9, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ado

[Ano69-105] Anonymous. [advertisement]. IEEE Spectrum, 6(9):11–16, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adp

[Ano69-106] Anonymous. [advertisement]. IEEE Spectrum, 6(9):21–22, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adq

[Ano69-107] Anonymous. [advertisement]. IEEE Spectrum, 6(9):63–66, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adr

[Ano69-108] Anonymous. [advertisement]. IEEE Spectrum, 6(9):99, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ads

[Ano69-109] Anonymous. [advertisement]. IEEE Spectrum, 6(9):101–103, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adt

[Ano69-110] Anonymous. [advertisement]. IEEE Spectrum, 6(9):104–105, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adu

[Ano69-111] Anonymous. [advertisement].

IEEE Spectrum, 6(9):120, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adv

[Ano69-112] Anonymous. [advertisement].

IEEE Spectrum, 6(9):nil2, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adw

[Ano69-113] Anonymous. [advertisement].

IEEE Spectrum, 6(10):nil1, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adx

[Ano69-114] Anonymous. [advertisement].

IEEE Spectrum, 6(10):1a–1, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ady

[Ano69-115] Anonymous. [advertisement].

IEEE Spectrum, 6(10):6–7, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Adz

[Ano69-116] Anonymous. [advertisement].

IEEE Spectrum, 6(10):9–15, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aea

[Ano69-117] Anonymous. [advertisement].

IEEE Spectrum, 6(10):21, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aeb

[Ano69-118] Anonymous. [advertisement]. IEEE Spectrum, 6(10):25–26, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aec

[Ano69-119] Anonymous. [advertisement]. IEEE Spectrum, 6(10):69–72, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aed

[Ano69-120] Anonymous. [advertisement]. IEEE Spectrum, 6(10):97, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aee

[Ano69-121] Anonymous. [advertisement]. IEEE Spectrum, 6(10):101– 103, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aef

[Ano69-122] Anonymous. [advertisement]. IEEE Spectrum, 6(10):105– 107, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aeg

[Ano69-123] Anonymous. [advertisement]. IEEE Spectrum, 6(10):108– 109, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aeh

[Ano69-124] Anonymous. [advertisement]. IEEE Spectrum, 6(10):111– 115, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aei

[Ano69-125] Anonymous. [advertisement].

IEEE Spectrum, 6(10):133, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aej

[Ano69-126] Anonymous. [advertisement].

IEEE Spectrum, 6(10):139, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aek

[Ano69-127] Anonymous. [advertisement].

IEEE Spectrum, 6(10):nil2, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ael

[Ano69-128] Anonymous. [advertisement].

IEEE Spectrum, 6(11):nil1, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aem

[Ano69-129] Anonymous. [advertisement].

IEEE Spectrum, 6(11):1a–1a, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aen

[Ano69-130] Anonymous. [advertisement].

IEEE Spectrum, 6(11):1, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aeo

[Ano69-131] Anonymous. [advertisement].

IEEE Spectrum, 6(11):6–7, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aep

[Ano69-132] Anonymous. [advertisement]. IEEE Spectrum, 6(11):11–18, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aeq

[Ano69-133] Anonymous. [advertisement]. IEEE Spectrum, 6(11):23, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aer

[Ano69-134] Anonymous. [advertisement]. IEEE Spectrum, 6(11):25–34, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aes

[Ano69-135] Anonymous. [advertisement]. IEEE Spectrum, 6(11):38, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aet

[Ano69-136] Anonymous. [advertisement]. IEEE Spectrum, 6(11):73–76, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aeu

[Ano69-137] Anonymous. [advertisement]. IEEE Spectrum, 6(11):107, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aev

[Ano69-138] Anonymous. [advertisement]. IEEE Spectrum, 6(11):112–113, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aew

[Ano69-139] Anonymous. [advertisement].

IEEE Spectrum, 6(11):117–120, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aex

[Ano69-140] Anonymous. [advertisement].

IEEE Spectrum, 6(11):123–127, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aey

[Ano69-141] Anonymous. [advertisement].

IEEE Spectrum, 6(11):137, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aez

[Ano69-142] Anonymous. [advertisement].

IEEE Spectrum, 6(11):143–144, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afa

[Ano69-143] Anonymous. [advertisement].

IEEE Spectrum, 6(11):nil2, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afb

[Ano69-144] Anonymous. [advertisement].

IEEE Spectrum, 6(12):nil1, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afc

[Ano69-145] Anonymous. [advertisement].

IEEE Spectrum, 6(12):1a–1a, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afd

[Ano69-146] Anonymous. [advertisement]. IEEE Spectrum, 6(12):1, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afe

[Ano69-147] Anonymous. [advertisement]. IEEE Spectrum, 6(12):6–7, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Aff

[Ano69-148] Anonymous. [advertisement]. IEEE Spectrum, 6(12):17–20, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afg

[Ano69-149] Anonymous. [advertisement]. IEEE Spectrum, 6(12):23, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afh

[Ano69-150] Anonymous. [advertisement]. IEEE Spectrum, 6(12):25–29, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afi

[Ano69-151] Anonymous. [advertisement]. IEEE Spectrum, 6(12):33, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afj

[Ano69-152] Anonymous. [advertisement]. IEEE Spectrum, 6(12):34–46, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afk

[Ano69-153] Anonymous. [advertisement].

IEEE Spectrum, 6(12):81–84, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afl

[Ano69-154] Anonymous. [advertisement].

IEEE Spectrum, 6(12):107, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afm

[Ano69-155] Anonymous. [advertisement].

IEEE Spectrum, 6(12):125–135, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afn

[Ano69-156] Anonymous. [advertisement].

IEEE Spectrum, 6(12):137, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afo

[Ano69-157] Anonymous. [advertisement].

IEEE Spectrum, 6(12):159, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Afp

[Ano69-158] Anonymous. [advertisement].

IEEE Spectrum, 6(12):nil2, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:AIa

[Ano69-159] Anonymous. Annual index.

IEEE Spectrum, 6(12):140–158, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:AGI

[Ano69-160] Anonymous. An author’s guide to IEEE publications. IEEE Spectrum, 6(1):10a–10b, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRa

[Ano69-161] Anonymous. Book reviews. IEEE Spectrum, 6(1):110–115, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRb

[Ano69-162] Anonymous. Book reviews. IEEE Spectrum, 6(2):90–99, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRc

[Ano69-163] Anonymous. Book reviews. IEEE Spectrum, 6(3):134–143, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRd

[Ano69-164] Anonymous. Book reviews. IEEE Spectrum, 6(4):112–119, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRe

[Ano69-165] Anonymous. Book reviews. IEEE Spectrum, 6(5):117–125, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRf

[Ano69-166] Anonymous. Book reviews. IEEE Spectrum, 6(6):114–121, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRg

[Ano69-167] Anonymous. Book reviews.

IEEE Spectrum, 6(7):106–115, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRh

[Ano69-168] Anonymous. Book reviews.

IEEE Spectrum, 6(8):96–102, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRi

[Ano69-169] Anonymous. Book reviews.

IEEE Spectrum, 6(9):106–112, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRj

[Ano69-170] Anonymous. Book reviews.

IEEE Spectrum, 6(10):110– 114, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRk

[Ano69-171] Anonymous. Book reviews.

IEEE Spectrum, 6(11):114–120, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BRl

[Ano69-172] Anonymous. Book reviews.

IEEE Spectrum, 6(12):108–118, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRa

[Ano69-173] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(5):1a–1b, May 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRb

[Ano69-174] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(5):153–154, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRc

[Ano69-175] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(6):1a–1b, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRd

[Ano69-176] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(6):141–142, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRe

[Ano69-177] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(7):1a–1b, July 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRf

[Ano69-178] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(7):137–138, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRg

[Ano69-179] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(8):1a–1b, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRh

[Ano69-180] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(8):121–122, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRi

[Ano69-181] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(9):1a–1b, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRj

[Ano69-182] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(9):129–130, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRk

[Ano69-183] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(10):1a–1b, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRl

[Ano69-184] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(10):139–140, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRm

[Ano69-185] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(11):1a–1b, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRn

[Ano69-186] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(11):143–144, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRo

[Ano69-187] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(12):1a–1b, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:ISRp

[Ano69-188] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 6(12):163–164, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPa

[Ano69-189] Anonymous. [breaker page]. IEEE Spectrum, 6(1):4, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPb

[Ano69-190] Anonymous. [breaker page]. IEEE Spectrum, 6(2):4, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPc

[Ano69-191] Anonymous. [breaker page]. IEEE Spectrum, 6(3):4, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPd

[Ano69-192] Anonymous. [breaker page]. IEEE Spectrum, 6(4):4, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPe

[Ano69-193] Anonymous. [breaker page]. IEEE Spectrum, 6(5):4, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPf

[Ano69-194] Anonymous. [breaker page].

IEEE Spectrum, 6(6):4, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPg

[Ano69-195] Anonymous. [breaker page].

IEEE Spectrum, 6(7):4, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPh

[Ano69-196] Anonymous. [breaker page].

IEEE Spectrum, 6(8):4, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPi

[Ano69-197] Anonymous. [breaker page].

IEEE Spectrum, 6(9):4, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPj

[Ano69-198] Anonymous. [breaker page].

IEEE Spectrum, 6(10):4, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPk

[Ano69-199] Anonymous. [breaker page].

IEEE Spectrum, 6(11):4, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:BPl

[Ano69-200] Anonymous. [breaker page].

IEEE Spectrum, 6(12):4, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CSM

[Ano69-201] Anonymous. C-Scan — a milestone for carrier landing. IEEE Spectrum, 6(9):5, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ca

[Ano69-202] Anonymous. Calendar. IEEE Spectrum, 6(1):25–26, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cb

[Ano69-203] Anonymous. Calendar. IEEE Spectrum, 6(2):11–12, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cc

[Ano69-204] Anonymous. Calendar. IEEE Spectrum, 6(3):15–16, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cd

[Ano69-205] Anonymous. Calendar. IEEE Spectrum, 6(4):21–22, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ce

[Ano69-206] Anonymous. Calendar. IEEE Spectrum, 6(5):31–33, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cf

[Ano69-207] Anonymous. Calendar. IEEE Spectrum, 6(6):18–20, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cg

[Ano69-208] Anonymous. Calendar. IEEE Spectrum, 6(7):23–26, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ch

[Ano69-209] Anonymous. Calendar. IEEE Spectrum, 6(8):19–21, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ci

[Ano69-210] Anonymous. Calendar. IEEE Spectrum, 6(9):18–20, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ck

[Ano69-211] Anonymous. Calendar. IEEE Spectrum, 6(10):22–24, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cm

[Ano69-212] Anonymous. Calendar. IEEE Spectrum, 6(11):35–37, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cn

[Ano69-213] Anonymous. Calendar. IEEE Spectrum, 6(12):30–32, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CAa

[Ano69-214] Anonymous. Classified advertising. IEEE Spectrum, 6(3): 145–155, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CAb

[Ano69-215] Anonymous. Classified advertising. IEEE Spectrum, 6(5): 115–116, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CAc

[Ano69-216] Anonymous. Classified advertising. IEEE Spectrum, 6(6): 115–121, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CAd

[Ano69-217] Anonymous. Classified advertising. IEEE Spectrum, 6(7): 118–123, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CAe

[Ano69-218] Anonymous. Classified advertising. IEEE Spectrum, 6(9):107– 113, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CAf

[Ano69-219] Anonymous. Classified advertising. IEEE Spectrum, 6(10): 119–131, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CAg

[Ano69-220] Anonymous. Classified advertising. IEEE Spectrum, 6(11): 111, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:CAh

[Ano69-221] Anonymous. Classified advertising. IEEE Spectrum, 6(12):111– 121, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cj

[Ano69-222] Anonymous. Correction. IEEE Spectrum, 6(10):15, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Cl

[Ano69-223] Anonymous. Corrections. IEEE Spectrum, 6(11):17, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:DFT

[Ano69-224] Anonymous. Design for the future in terminal air traffic control. IEEE Spectrum, 6(2): 5, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:E

[Ano69-225] Anonymous. Expo ’70. IEEE Spectrum, 6(12):34–45, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FMA

[Ano69-226] Anonymous. Figures of merit for antenna patterns. IEEE Spectrum, 6(10):5, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPa

[Ano69-227] Anonymous. Focal points. IEEE Spectrum, 6(1):13–23, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPb

[Ano69-228] Anonymous. Focal points. IEEE Spectrum, 6(2):7–10, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPc

[Ano69-229] Anonymous. Focal points. IEEE Spectrum, 6(3):10–13, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPd

[Ano69-230] Anonymous. Focal points. IEEE Spectrum, 6(4):10–18, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPe

[Ano69-231] Anonymous. Focal points. IEEE Spectrum, 6(5):10–29, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPf

[Ano69-232] Anonymous. Focal points. IEEE Spectrum, 6(6):8–16, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPg

[Ano69-233] Anonymous. Focal points. IEEE Spectrum, 6(7):10–20, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPh

[Ano69-234] Anonymous. Focal points. IEEE Spectrum, 6(8):10–17, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPi

[Ano69-235] Anonymous. Focal points. IEEE Spectrum, 6(9):10–17, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPj

[Ano69-236] Anonymous. Focal points. IEEE Spectrum, 6(10):16–20, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPk

[Ano69-237] Anonymous. Focal points. IEEE Spectrum, 6(11):24–30, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FPl

[Ano69-238] Anonymous. Focal points. IEEE Spectrum, 6(12):24–28, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fa

[Ano69-239] Anonymous. Forum. IEEE Spectrum, 6(1):8, January 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fb

[Ano69-240] Anonymous. Forum. IEEE Spectrum, 6(2):6, February 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fd

[Ano69-241] Anonymous. Forum. IEEE Spectrum, 6(3):8–9, March 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fe

[Ano69-242] Anonymous. Forum. IEEE Spectrum, 6(4):8–9, April 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ff

[Ano69-243] Anonymous. Forum. IEEE Spectrum, 6(5):6–8, May 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fg

[Ano69-244] Anonymous. Forum. IEEE Spectrum, 6(6):6, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fh

[Ano69-245] Anonymous. Forum. IEEE Spectrum, 6(7):8, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fi

[Ano69-246] Anonymous. Forum. IEEE Spectrum, 6(8):6–8, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fj

[Ano69-247] Anonymous. Forum. IEEE Spectrum, 6(9):6–9, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fk

[Ano69-248] Anonymous. Forum. IEEE Spectrum, 6(10):8–15, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fl

[Ano69-249] Anonymous. Forum. IEEE Spectrum, 6(11):8–17, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Fm

[Ano69-250] Anonymous. Forum. IEEE Spectrum, 6(12):8–19, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCa

[Ano69-251] Anonymous. [front cover]. IEEE Spectrum, 6(1):c1, January 1969.

CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCb

[Ano69-252] Anonymous. [front cover]. IEEE Spectrum, 6(1):c2, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCc

[Ano69-253] Anonymous. [front cover]. IEEE Spectrum, 6(2):c1, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCd

[Ano69-254] Anonymous. [front cover]. IEEE Spectrum, 6(2):c2, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCe

[Ano69-255] Anonymous. [front cover]. IEEE Spectrum, 6(3):c1, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCf

[Ano69-256] Anonymous. [front cover]. IEEE Spectrum, 6(3):c2, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCg

[Ano69-257] Anonymous. [front cover]. IEEE Spectrum, 6(4):c1, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCh

[Ano69-258] Anonymous. [front cover]. IEEE Spectrum, 6(4):c2, April 1969. CODEN IEESAM. ISSN 1939-

Anonymous:1969:FCi

[Ano69-259] Anonymous. [front cover]. IEEE Spectrum, 6(5):c1, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCj

[Ano69-260] Anonymous. [front cover]. IEEE Spectrum, 6(5):c2, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCk

[Ano69-261] Anonymous. [front cover]. IEEE Spectrum, 6(6):c1, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCl

[Ano69-262] Anonymous. [front cover]. IEEE Spectrum, 6(6):c2, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCm

[Ano69-263] Anonymous. [front cover]. IEEE Spectrum, 6(7):c1, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCn

[Ano69-264] Anonymous. [front cover]. IEEE Spectrum, 6(7):c2, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCo

[Ano69-265] Anonymous. [front cover]. IEEE Spectrum, 6(8):c1, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCp

[Ano69-266] Anonymous. [front cover]. IEEE Spectrum, 6(8):c2, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCq

[Ano69-267] Anonymous. [front cover]. IEEE Spectrum, 6(9):c1, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCr

[Ano69-268] Anonymous. [front cover]. IEEE Spectrum, 6(9):c2, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCs

[Ano69-269] Anonymous. [front cover]. IEEE Spectrum, 6(10):c1, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCt

[Ano69-270] Anonymous. [front cover]. IEEE Spectrum, 6(10):c2, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCu

[Ano69-271] Anonymous. [front cover]. IEEE Spectrum, 6(11):c1, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCv

[Ano69-272] Anonymous. [front cover]. IEEE Spectrum, 6(11):c2, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCw

[Ano69-273] Anonymous. [front cover]. IEEE Spectrum, 6(12):c1, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:FCx

[Ano69-274] Anonymous. [front cover]. IEEE Spectrum, 6(12):c2, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:GDN

[Ano69-275] Anonymous. Gas discharge noise generators: The calibration problem. IEEE Spectrum, 6(6): 5, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IR

[Ano69-276] Anonymous. IEEE reports for 1968. IEEE Spectrum, 6(6): 85–98, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IIC

[Ano69-277] Anonymous. IEEE’69 Internatonal Conventon and Exhibition, March 24–27 — Exhibitors. IEEE Spectrum, 6(3):163–237, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAa

[Ano69-278] Anonymous. Index to advertisers. IEEE Spectrum, 6(1):148– 152, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAb

[Ano69-279] Anonymous. Index to advertisers. IEEE Spectrum, 6(2):120– 122, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAc

[Ano69-280] Anonymous. Index to advertisers. IEEE Spectrum, 6(3): 243–250, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAd

[Ano69-281] Anonymous. Index to advertisers. IEEE Spectrum, 6(4): 150–152, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAe

[Ano69-282] Anonymous. Index to advertisers. IEEE Spectrum, 6(5): 150–152, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAf

[Ano69-283] Anonymous. Index to advertisers. IEEE Spectrum, 6(6): 138–140, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAg

[Ano69-284] Anonymous. Index to advertisers. IEEE Spectrum, 6(7): 134–136, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAh

[Ano69-285] Anonymous. Index to advertisers. IEEE Spectrum, 6(8): 118–120, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAi

[Ano69-286] Anonymous. Index to advertisers. IEEE Spectrum, 6(9):126– 128, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAj

[Ano69-287] Anonymous. Index to advertisers. IEEE Spectrum, 6(10): 136–138, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAk

[Ano69-288] Anonymous. Index to advertisers. IEEE Spectrum, 6(11):140– 142, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IAl

[Ano69-289] Anonymous. Index to advertisers. IEEE Spectrum, 6(12):160– 162, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IASa

[Ano69-290] Anonymous. Information about the society. IEEE Spectrum, 6(2): 13–14, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IASb

[Ano69-291] Anonymous. Information about the society. IEEE Spectrum, 6 (3):33–34, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IASc

[Ano69-292] Anonymous. Information about the society. IEEE Spectrum, 6 (3):225, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IASd

[Ano69-293] Anonymous. Information about the society. IEEE Spectrum, 6(3): 238–242, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IASe

[Ano69-294] Anonymous. Information about the society. IEEE Spectrum, 6 (11):21–22, November 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IASf

[Ano69-295] Anonymous. Information about the society. IEEE Spectrum, 6 (11):40, November 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IPU

[Ano69-296] Anonymous. Infrared parametric upconversion. IEEE Spectrum, 6(3):5, March 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:MCO

[Ano69-297] Anonymous. Maximum channel occupancy. IEEE Spectrum, 6 (11):5, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:MBI

[Ano69-298] Anonymous. A microminiature S-band integrated receiver. IEEE Spectrum, 6(1):5, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NBM

[Ano69-299] Anonymous. Narrow bandwidth microwave elliptic-function filter. IEEE Spectrum, 6(5):5, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NPAa

[Ano69-300] Anonymous. New product applications. IEEE Spectrum, 6(9): 87–92, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NPAb

[Ano69-301] Anonymous. New product applications. IEEE Spectrum, 6(10): 80–88, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NPAc

[Ano69-302] Anonymous. New product applications. IEEE Spectrum, 6(11): 92–97, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NPAd

[Ano69-303] Anonymous. New product applications. IEEE Spectrum, 6(12): 85–90, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NUS

[Ano69-304] Anonymous. A new use for the sun. IEEE Spectrum, 6(3):41, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NWa

[Ano69-305] Anonymous. News from washington. IEEE Spectrum, 6(11): 19–20, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NWb

[Ano69-306] Anonymous. News from washington. IEEE Spectrum, 6(12): 21–22, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIa

[Ano69-307] Anonymous. News of the IEEE. IEEE Spectrum, 6(1): 126–140, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIb

[Ano69-308] Anonymous. News of the IEEE. IEEE Spectrum, 6(2):100– 113, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIc

[Ano69-309] Anonymous. News of the IEEE.

IEEE Spectrum, 6(3):144–155, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NId

[Ano69-310] Anonymous. News of the IEEE.

IEEE Spectrum, 6(4):120–143, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIe

[Ano69-311] Anonymous. News of the IEEE.

IEEE Spectrum, 6(5):136–145, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIf

[Ano69-312] Anonymous. News of the IEEE.

IEEE Spectrum, 6(6):122–128, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIg

[Ano69-313] Anonymous. News of the IEEE.

IEEE Spectrum, 6(7):116–123, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIh

[Ano69-314] Anonymous. News of the IEEE.

IEEE Spectrum, 6(8):104–113, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIi

[Ano69-315] Anonymous. News of the IEEE. IEEE Spectrum, 6(9):114– 119, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIj

[Ano69-316] Anonymous. News of the IEEE. IEEE Spectrum, 6(10): 116–128, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIk

[Ano69-317] Anonymous. News of the IEEE. IEEE Spectrum, 6(11): 121–135, November 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:NIl

[Ano69-318] Anonymous. News of the IEEE. IEEE Spectrum, 6(12): 122–135, December 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Oa

[Ano69-319] Anonymous. Obituaries. IEEE Spectrum, 6(1):147, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ob

[Ano69-320] Anonymous. Obituaries. IEEE Spectrum, 6(2):116–119, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Oc

[Ano69-321] Anonymous. Obituaries. IEEE Spectrum, 6(3):160–161, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Od

[Ano69-322] Anonymous. Obituaries. IEEE Spectrum, 6(4):149, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Oe

[Ano69-323] Anonymous. Obituaries. IEEE Spectrum, 6(5):148–149, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Of

[Ano69-324] Anonymous. Obituaries. IEEE Spectrum, 6(6):136–137, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Og

[Ano69-325] Anonymous. Obituaries. IEEE Spectrum, 6(7):127, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Oh

[Ano69-326] Anonymous. Obituaries. IEEE Spectrum, 6(8):117, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Oi

[Ano69-327] Anonymous. Obituaries. IEEE Spectrum, 6(9):124–125, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Oj

[Ano69-328] Anonymous. Obituaries. IEEE Spectrum, 6(10):134–135, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ok

[Ano69-329] Anonymous. Obituaries. IEEE Spectrum, 6(11):138–139, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ol

[Ano69-330] Anonymous. Obituaries. IEEE Spectrum, 6(12):139, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:OGS

[Ano69-331] Anonymous. Operation GIT a success. IEEE Spectrum, 6(4):86, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PAM

[Ano69-332] Anonymous. Parametric amplifiers and masers 1969. IEEE Spectrum, 6(8):5, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PSF

[Ano69-333] Anonymous. Parametric sum– frequency upconverter. IEEE Spectrum, 6(12):5, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pa

[Ano69-334] Anonymous. People. IEEE Spectrum, 6(1):143–147, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pb

[Ano69-335] Anonymous. People. IEEE Spectrum, 6(2):114–116, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pc

[Ano69-336] Anonymous. People. IEEE Spectrum, 6(3):156–160, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pd

[Ano69-337] Anonymous. People. IEEE Spectrum, 6(4):145–149, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pe

[Ano69-338] Anonymous. People. IEEE Spectrum, 6(5):146–148, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pf

[Ano69-339] Anonymous. People. IEEE Spectrum, 6(6):130–136, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pg

[Ano69-340] Anonymous. People. IEEE Spectrum, 6(7):124–127, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Ph

[Ano69-341] Anonymous. People. IEEE Spectrum, 6(8):114–117, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pi

[Ano69-342] Anonymous. People. IEEE Spectrum, 6(9):121–124, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pj

[Ano69-343] Anonymous. People. IEEE Spectrum, 6(10):132–134, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pk

[Ano69-344] Anonymous. People. IEEE Spectrum, 6(11):136–138, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:Pl

[Ano69-345] Anonymous. People. IEEE Spectrum, 6(12):136–139, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSa

[Ano69-346] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(1):149–152, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSb

[Ano69-347] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(2):121–122, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSc

[Ano69-348] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(3):244–250, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSd

[Ano69-349] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(4):151–152, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSe

[Ano69-350] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(5):151–152, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSf

[Ano69-351] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(6):139–140, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSg

[Ano69-352] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(7):135, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSh

[Ano69-353] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(8):119, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSi

[Ano69-354] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(9):127, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSj

[Ano69-355] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(10):137, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSk

[Ano69-356] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(11):141, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:PCSl

[Ano69-357] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 6(12):161, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIa

[Ano69-358] Anonymous. Scanning the issues. IEEE Spectrum, 6(1):96–98, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIb

[Ano69-359] Anonymous. Scanning the issues. IEEE Spectrum, 6(2):79– 80, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIc

[Ano69-360] Anonymous. Scanning the issues. IEEE Spectrum, 6(3): 119–121, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SId

[Ano69-361] Anonymous. Scanning the issues. IEEE Spectrum, 6(4): 94–97, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIe

[Ano69-362] Anonymous. Scanning the issues. IEEE Spectrum, 6(5): 106–107, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIf

[Ano69-363] Anonymous. Scanning the issues. IEEE Spectrum, 6(6): 99–101, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIg

[Ano69-364] Anonymous. Scanning the issues.

IEEE Spectrum, 6(7):95–96, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIh

[Ano69-365] Anonymous. Scanning the issues.

IEEE Spectrum, 6(8):82–83, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIi

[Ano69-366] Anonymous. Scanning the issues. IEEE Spectrum, 6(9):93– 94, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIj

[Ano69-367] Anonymous. Scanning the issues. IEEE Spectrum, 6(10): 95–96, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIk

[Ano69-368] Anonymous. Scanning the issues. IEEE Spectrum, 6(11):98– 99, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SIl

[Ano69-369] Anonymous. Scanning the issues. IEEE Spectrum, 6(12):97– 98, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SCEc

[Ano69-370] Anonymous. Short courses in electrical and electronics engineering: Fall-winter 1969-1970. IEEE Spectrum, 6(10): 89–94, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SCEb

[Ano69-371] Anonymous. Short courses in electrical and electronics engineering: Summer-fall 1969. IEEE Spectrum, 6(5):126–135, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SCEd

[Ano69-372] Anonymous. Short courses in electrical and electronics engineering winter-spring 1970. IEEE Spectrum, 6(12):91–96, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SCEa

[Ano69-373] Anonymous. Short courses in electrical and electronics engineering winter, spring, summer 1969. IEEE Spectrum, 6(1):117– 124, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SBM

[Ano69-374] Anonymous. A simple, better microphone. IEEE Spectrum, 6 (1):78, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SYW

[Ano69-375] Anonymous. Sixty years of wireless and radio autobiographical reminiscence by haraden pratt. IEEE Spectrum, 6(11):

41–47, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPa

[Ano69-376] Anonymous. Special publications. IEEE Spectrum, 6(1): 109, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPb

[Ano69-377] Anonymous. Special publications. IEEE Spectrum, 6(2): 89, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPc

[Ano69-378] Anonymous. Special publications. IEEE Spectrum, 6(3):132, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPd

[Ano69-379] Anonymous. Special publications. IEEE Spectrum, 6(4):110, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPe

[Ano69-380] Anonymous. Special publications. IEEE Spectrum, 6(5):114, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPf

[Ano69-381] Anonymous. Special publications. IEEE Spectrum, 6(6): 112–113, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPg

[Ano69-382] Anonymous. Special publications. IEEE Spectrum, 6(7):105, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPh

[Ano69-383] Anonymous. Special publications. IEEE Spectrum, 6(8): 94–95, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPi

[Ano69-384] Anonymous. Special publications. IEEE Spectrum, 6(9):104– 105, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPj

[Ano69-385] Anonymous. Special publications. IEEE Spectrum, 6(10): 108–109, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPk

[Ano69-386] Anonymous. Special publications. IEEE Spectrum, 6(11): 112–113, November 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SPl

[Ano69-387] Anonymous. Special publications. IEEE Spectrum, 6(12): 106, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLa

[Ano69-388] Anonymous. Spectral lines. IEEE Spectrum, 6(1):29, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLb

[Ano69-389] Anonymous. Spectral lines. IEEE Spectrum, 6(2):17, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLc

[Ano69-390] Anonymous. Spectral lines.

IEEE Spectrum, 6(3):51, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLd

[Ano69-391] Anonymous. Spectral lines.

IEEE Spectrum, 6(4):25, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLe

[Ano69-392] Anonymous. Spectral lines.

IEEE Spectrum, 6(5):35, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLf

[Ano69-393] Anonymous. Spectral lines.

IEEE Spectrum, 6(6):25, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLg

[Ano69-394] Anonymous. Spectral lines.

IEEE Spectrum, 6(7):29, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLh

[Ano69-395] Anonymous. Spectral lines.

IEEE Spectrum, 6(8):23, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLi

[Ano69-396] Anonymous. Spectral lines.

IEEE Spectrum, 6(9):23, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLj

[Ano69-397] Anonymous. Spectral lines. IEEE Spectrum, 6(10):27, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLk

[Ano69-398] Anonymous. Spectral lines. IEEE Spectrum, 6(11):39, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SLl

[Ano69-399] Anonymous. Spectral lines. IEEE Spectrum, 6(12):47, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SRSa

[Ano69-400] Anonymous. Spectrum reader service card. IEEE Spectrum, 6 (1):1a–1b, January 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SRSb

[Ano69-401] Anonymous. Spectrum reader service card. IEEE Spectrum, 6 (1):153–154, January 1969. CO-DEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SRSc

[Ano69-402] Anonymous. Spectrum reader service card. IEEE Spectrum, 6 (4):1a–1b, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:SRSd

[Ano69-403] Anonymous. Spectrum reader service card. IEEE Spectrum, 6 (4):153–154, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:STI

[Ano69-404] Anonymous. Strategy and tactics for integrated electronics. IEEE Spectrum, 6(6):26–33, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCa

[Ano69-405] Anonymous. Table of contents.

IEEE Spectrum, 6(1):2–3, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCb

[Ano69-406] Anonymous. Table of contents.

IEEE Spectrum, 6(2):2–3, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCc

[Ano69-407] Anonymous. Table of contents.

IEEE Spectrum, 6(3):2–3, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCd

[Ano69-408] Anonymous. Table of contents.

IEEE Spectrum, 6(4):2–3, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCe

[Ano69-409] Anonymous. Table of contents.

IEEE Spectrum, 6(5):2–3, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCf

[Ano69-410] Anonymous. Table of contents.

IEEE Spectrum, 6(6):2–3, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCg

[Ano69-411] Anonymous. Table of contents. IEEE Spectrum, 6(7):2–3, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCh

[Ano69-412] Anonymous. Table of contents. IEEE Spectrum, 6(8):2–3, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCi

[Ano69-413] Anonymous. Table of contents. IEEE Spectrum, 6(9):2– 3, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCj

[Ano69-414] Anonymous. Table of contents. IEEE Spectrum, 6(10):2–3, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCk

[Ano69-415] Anonymous. Table of contents. IEEE Spectrum, 6(11): 2–3, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TCl

[Ano69-416] Anonymous. Table of contents. IEEE Spectrum, 6(12):2–3, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TP

[Ano69-417] Anonymous. Technical program. IEEE Spectrum, 6(3): 17–48, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TAR

[Ano69-418] Anonymous. Technological advances reflected in new wileyinterscience books [advertisement]. IEEE Spectrum, 6(4):6–7, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEa

[Ano69-419] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6 (1):4, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEb

[Ano69-420] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6 (2):4, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEc

[Ano69-421] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6 (3):4, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEd

[Ano69-422] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6(4): 4, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEe

[Ano69-423] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6(5): 4, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEf

[Ano69-424] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6(6): 4, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEg

[Ano69-425] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6(7): 4, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEh

[Ano69-426] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6 (8):4, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEi

[Ano69-427] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6 (9):4, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEj

[Ano69-428] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6 (10):4, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEk

[Ano69-429] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6 (11):4, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:IEEl

[Ano69-430] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 6 (12):4, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTa

[Ano69-431] Anonymous. Transients and trends. IEEE Spectrum, 6(1): 14, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTb

[Ano69-432] Anonymous. Transients and trends. IEEE Spectrum, 6(2): 9, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTc

[Ano69-433] Anonymous. Transients and trends. IEEE Spectrum, 6(3):12, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTd

[Ano69-434] Anonymous. Transients and trends. IEEE Spectrum, 6(4):12, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTe

[Ano69-435] Anonymous. Transients and trends. IEEE Spectrum, 6(5):12, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTf

[Ano69-436] Anonymous. Transients and trends. IEEE Spectrum, 6(6):10, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTg

[Ano69-437] Anonymous. Transients and trends. IEEE Spectrum, 6(7):12, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTh

[Ano69-438] Anonymous. Transients and trends. IEEE Spectrum, 6(8):12, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTi

[Ano69-439] Anonymous. Transients and trends. IEEE Spectrum, 6(9): 12, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TTj

[Ano69-440] Anonymous. Transients and trends. IEEE Spectrum, 6(10): 18, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJa

[Ano69-441] Anonymous. Translated journals. IEEE Spectrum, 6(1):106– 108, January 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJb

[Ano69-442] Anonymous. Translated journals. IEEE Spectrum, 6(2):86– 88, February 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJc

[Ano69-443] Anonymous. Translated journals. IEEE Spectrum, 6(3): 128–131, March 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJd

[Ano69-444] Anonymous. Translated journals. IEEE Spectrum, 6(4): 106–108, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJe

[Ano69-445] Anonymous. Translated journals. IEEE Spectrum, 6(5): 112–113, May 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJf

[Ano69-446] Anonymous. Translated journals. IEEE Spectrum, 6(6): 108–111, June 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJg

[Ano69-447] Anonymous. Translated journals. IEEE Spectrum, 6(7): 102–104, July 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJh

[Ano69-448] Anonymous. Translated journals. IEEE Spectrum, 6(8):89–93, August 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJi

[Ano69-449] Anonymous. Translated journals. IEEE Spectrum, 6(9):100– 103, September 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJj

[Ano69-450] Anonymous. Translated journals. IEEE Spectrum, 6(10):104– 107, October 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJk

[Ano69-451] Anonymous. Translated journals. IEEE Spectrum, 6(11):108– 110, November 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:TJl

[Ano69-452] Anonymous. Translated journals. IEEE Spectrum, 6(12):104– 105, December 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:VPD

[Ano69-453] Anonymous. Video probability distributions. IEEE Spectrum, 6(4):5, April 1969. CODEN IEESAM. ISSN 1939-9340.

Anonymous:1969:WEL

[Ano69-454] Anonymous. WESCON/69 Exhibitors list. IEEE Spectrum, 6 (7):128–133, July 1969. CODEN IEESAM. ISSN 1939-9340.

Aiken:1964:PAC

[AOB64] H. H. Aiken, A. G. Oettinger, and T. C. Bartee. Proposed automatic calculating machine. IEEE Spectrum, 1(8):62–69, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Asija:1968:ILC

[Asi68] S. P. Asija. Instant logic conversion. IEEE Spectrum, 5(12): 77–80, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Atherton:1964:SDG

[Ath64] D. L. Atherton. Superconducting dc generators and motors. IEEE Spectrum, 1(12):67–71, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Atkins:1965:WCC

[Atk65] J. B. Atkins. Worst-case circuit design. IEEE Spectrum, 2 (3):152–161, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Aboudara:1967:TMR

[AWSB67] D. N. Aboudara, C. W. Woods, R. S. Silver, and J. C. Beckett. Tomorrow’s mass rapid transitavailable today! IEEE Spectrum, 4(1):60–70, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baker:1968:LDT

[Bak68] C. E. Baker. Laser display technology. IEEE Spectrum, 5(12): 39–50, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baron:1966:ED

[Bar66a] S. Baron. The economics of desalination. IEEE Spectrum, 3 (12):63–70, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barrow:1966:ITS

[Bar66b] B. B. Barrow. IEEE takes a stand on units. IEEE Spectrum, 3(3):164–168, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Basore:1964:BR

[BDS^{+}64] B. L. Basore, J. K. Dillard, J. E. Sherman, N. M. Blachman, H. A. Samulon, J. R. Singer, and L. K. Lee. Book reviews. IEEE Spectrum, 1(11): 180–185, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bennett:1968:PPL

[Ben68] R. R. Bennett. Planning for power a look at tomorrow’s station sizes. IEEE Spectrum, 5 (9):67–72, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berkner:1964:GEB

[Ber64] L. V. Berkner. Graduate education: A basic national resource. IEEE Spectrum, 1(2): 66–73, February 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bergland:1969:GTF

[Ber69] G. D. Bergland. A guided tour of the Fast Fourier Transform. IEEE Spectrum, 6(7):41–52, July 1969. CODEN IEESAM. ISSN 1939-9340.

Bruckner:1968:EOE

[BFS68] A. Bruckner, W. J. Fabrycky, and J. E. Shamblin. Economic optimization of energy conversion with storage. IEEE Spectrum, 5(4):101–107, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1964:RMP

[BG64] W. C. Brown and R. H.

George. Rectification of microwave power. IEEE Spectrum, 1(10):92–97, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1964:EAD

[BH64] J. J. W. Brown and E. M.

Hunter. EHV AC and DC transmission. IEEE Spectrum, 1(1): 123–128, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bickmore:1964:AAAb

[Bic64a] R. W. Bickmore. Acknowledgment to “Adaptive antenna arrays”. IEEE Spectrum, 1 (12):66, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bickmore:1964:AAAa

[Bic64b] R. W. Bickmore. Adaptive antenna arrays. IEEE Spectrum, 1(8):78–88, August 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Biedenbach:1968:CEE

[Bie68] J. M. Biedenbach. Continuing education for the engineering manager. IEEE Spectrum, 5(11): 99–101, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bishop:1966:RPC

[Bis66] A. S. Bishop. Roadblocks in the path of controlled fusion.

IEEE Spectrum, 3(7):106–111, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brookner:1967:DSO

[BKW67] E. Brookner, M. Kolker, and R. M. Wilmotte. Deep-space optical communications. IEEE Spectrum, 4(1):75–82, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Booker:1965:AREa

[BL65a] H. G. Booker and C. G. Little. Atmospheric research and electromagnetic telecommunication: Part I. IEEE Spectrum, 2(8):44–52, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Booker:1965:AREb

[BL65b] H. G. Booker and C. G. Little. Atmospheric research and electromagnetic telecommunication. Part II: Telecommunicationoriented research programs in fiscal year 1964. IEEE Spectrum, 2 (9):98–103, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brodey:1967:HET

[BL67] W. M. Brodey and N. Lindgren. Human enhancement through evolutionary technology. IEEE Spectrum, 4(9):87–97, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brodey:1968:HEB

[BL68] W. M. Brodey and N. Lindgren.

Human enhancement: Beyond the machine age. IEEE Spectrum, 5(2):79–93, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Black:1964:CME

[Bla64] G. Black. The changing market for electrical engineers. IEEE Spectrum, 1(8):70–77, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Blanchard:1965:SCR

[Bla65] J. H. Blanchard. Special conference report: Power industry computer application conference. IEEE Spectrum, 2(8):109–110, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bloembergen:1967:NHQ

[Blo67] N. Bloembergen. New horizons in quantum electronics. IEEE Spectrum, 4(7):82–86, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1964:MPG

[BM64] W. C. Brown and T. Moreno. Microwave power generation. IEEE Spectrum, 1(10):77–81, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brigham:1967:FFT

[BM67] E. O. Brigham and R. E. Morrow. The Fast Fourier Transform. IEEE Spectrum, 4(12): 63–70, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baptist:1969:EFG

[BN69] J. V. Baptist and R. Y. Nitta. Electrical features of the grand coulee third power plant. IEEE Spectrum, 6(1):89–95, January 1969. CODEN IEESAM. ISSN 1939-9340.

Boelter:1964:RU

[BO64] L. M. K. Boelter and R. R. O’Neill. The role of the university. IEEE Spectrum, 1 (6):114–115, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bonn:1964:BR

[Bon64] T. H. Bonn. Book reviews. IEEE Spectrum, 1(2):168–171, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Booker:1964:BRC

[Boo64] H. G. Booker. Book review: Clerk Maxwell and modern science, C. Domb, ed. IEEE Spectrum, 1(6):204, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brockman:1969:PRD

[BP69a] M. H. Brockman and E. C. Posner. Power requirements for deep-space telecommunication links. IEEE Spectrum, 6 (3):95–99, March 1969. CODEN IEESAM. ISSN 1939-9340.

Brown:1969:ISA

[BP69b] W. M. Brown and L. J. Porcello.

An introduction to syntheticaperture radar. IEEE Spectrum, 6(9):52–62, September 1969. CO-DEN IEESAM. ISSN 1939-9340.

Bradner:1964:GMO

[Bra64] H. Bradner. Geophysical measurements at ocean bottom. IEEE Spectrum, 1(11):108–114, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brewer:1965:PEP

[Bre65] G. R. Brewer. Physical electronic phenomena in ion propulsion engines. IEEE Spectrum, 2(8):65–79, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brennan:1969:A

[Bre69] D. Brennan. For the ABM.

IEEE Spectrum, 6(8):46–50, August 1969. CODEN IEESAM. ISSN 1939-9340.

Brogan:1964:MPG

[Bro64a] T. R. Brogan. MHD power generation. IEEE Spectrum, 1 (2):58–65, February 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bronwell:1964:PL

[Bro64b] A. B. Bronwell. Is philosophy off limits? IEEE Spectrum, 1(5):110–114, May 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1964:ICC

[Bro64c] J. E. Brown. Integrated circuits in consumer products. IEEE Spectrum, 1(6):75, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1964:FST

[Bro64d] W. C. Brown. Free-space transmission. IEEE Spectrum, 1 (10):86–91, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brooks:1967:DEE

[Bro67a] H. Brooks. Dilemmas of engineering education. IEEE Spectrum, 4(2):89–91, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1967:RFS

[Bro67b] C. J. Brown. Resolution of flyingspot scanner systems. IEEE Spectrum, 4(8):73–78, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1967:TRT

[Bro67c] G. H. Brown. Television’s role in tomorrow’s world. IEEE Spectrum, 4(10):56–58, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bromfield:1969:ETE

[Bro69] E. A. Bromfield. The emergence of the technician engineer in electrical and electronic engineering. IEEE Spectrum, 6 (6):46–51, June 1969. CODEN IEESAM. ISSN 1939-9340.

Bary:1964:HSE

[BRP64] C. W. Bary, R. E. Rice, and J. F.

Paquette. Heat storage for electric house heating. IEEE Spectrum, 1(4):109–112, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Burns:1967:HPL

[Bur67] F. P. Burns. High-power lasers their performance, limitations, and future. IEEE Spectrum, 4 (3):115–120, March 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Balet:1966:RPC

[BW66] W. J. Balet and R. L. Webb. Reactive power control in a metro system. IEEE Spectrum, 3(9): 141–144, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Camras:1965:ISD

[Cam65] M. Camras. Information storage density. IEEE Spectrum, 2(7):98–105, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carter:1969:TST

[Car69] C. M. Carter. Trade secrets and the technical man. IEEE Spectrum, 6(2):51–55, February 1969. CODEN IEESAM. ISSN 1939-9340.

Caswell:1964:TFC

[Cas64] H. L. Caswell. Thin-film circuit technology: Part II: Thin-film superconducting devices. IEEE Spectrum, 1(5):84–99, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cornog:1966:SMU

[CB66] J. R. Cornog and H. L. Bryan. Search methods used with transistor patent applications. IEEE Spectrum, 3(2):116–121, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Concordia:1966:PMR

[CdKS66] C. Concordia, F. P. deMello, L. K. Kirchmayer, and R. P. Schulz. Prime-mover response and system dynamic performance. IEEE Spectrum, 3(10): 106–111, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chambers:1967:RNG

[Cha67]  R. P. Chambers.  Random-

number generation on digital computers. IEEE Spectrum, 4 (2):48–56, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cheston:1968:PAR

[Che68] T. C. Cheston. Phased arrays for radars. IEEE Spectrum, 5(11):

102–111, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chipp:1966:CAT

[Chi66] R. D. Chipp. Community antenna television systems. IEEE Spectrum, 3(7):58–66, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chodorow:1964:BRM

[Cho64] M. Chodorow. Book review: Microwave tubes and semiconductor devices, G. D. Sims and I. M. Stephenson. IEEE Spectrum, 1(7):184–187, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clark:1964:RT

[Cla64] W. R. Clark. Report of the treasurer. IEEE Spectrum, 1 (7):136–139, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clark:1965:JAC

[Cla65] R. N. Clark. 1965 joint automatic control conference. IEEE Spectrum, 2(12):81–82, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carlin:1969:CME

[CLM^{+}69] J. F. Carlin, R. H. Lyons, J. R.

Mitiguy, E. F. Murphy, C. D. Galloway, M. A. Sager, and A. J. Wood. Corporate model of an electric utility. IEEE Spectrum, 6(6):75–84, June 1969. CODEN IEESAM. ISSN 1939-9340.

Chittenden:1967:ANP

[CN67] W. A. Chittenden and A. Nathan. Approaching nuclear power. IEEE Spectrum, 4(7):105–109, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cohn:1965:ACE

[Coh65] N. Cohn. The automatic control of electric power in the United States. IEEE Spectrum, 2(11): 67–77, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Collier:1966:SCV

[Col66a] R. J. Collier. Some current views on holography. IEEE Spectrum, 3(7):67–74, July 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Collis:1966:CAT

[Col66b] R. T. H. Collis. Clear air turbulence detection. IEEE Spectrum, 3(4):56–61, April 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

CEE:1968:CSE

[Com68] Commission on Engineering Education. Computer science in electrical engineering. IEEE Spectrum, 5(3):96–103, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Concordia:1965:PIS

[Con65] C. Concordia. Performance of interconnected systems following disturbances. IEEE Spectrum, 2(6):68–80, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Concordia:1968:CPR

[Con68] C. Concordia. Considerations in planning for reliable electric service. IEEE Spectrum, 5(8):71–76, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cooper:1964:C

[Coo64] P. W. Cooper. Correspondence.

IEEE Spectrum, 1(7):188–189, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cooper:1966:OCE

[Coo66] B. Cooper. Optical communications in the Earth’s atmosphere. IEEE Spectrum, 3(7):83– 88, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Copeland:1967:BNR

[Cop67] J. A. Copeland. Bulk negativeresistance semiconductor devices. IEEE Spectrum, 4(5): 71–77, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cotellessa:1965:RCP

[Cot65] R. F. Cotellessa. Report on the 1964 conference on precision electromagnetic measurements. IEEE Spectrum, 2(2): 103–110, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cottony:1966:USE

[Cot66] H. V. Cottony. 1965 URSI Symposium on Electromagnetic Wave Theory, Delft, Netherlands. IEEE Spectrum, 3(4): 115–118, April 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cox:1965:LMR

[Cox65] K. A. Cox. The land mobile radio services. IEEE Spectrum, 2 (10):42–43, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chance:1967:WGE

[CPH67] B. Chance, K. Pye, and J. Higgins. Waveform generation by enzymatic oscillators. IEEE Spectrum, 4(8):79–86, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christ:1969:ATM

[CR69] J. G. Christ and J. N. Ramsey. Analysis tools for microminiaturized circuits. IEEE Spectrum, 6 (3):108–118, March 1969. CO-DEN IEESAM. ISSN 1939-9340.

Crewe:1965:WWT

[Cre65] A. V. Crewe. What will they think of next? [future of electronics and electronic devices].

IEEE Spectrum, 2(1):84–86, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Critchlow:1964:MM

[Cri64]  A. J. Critchlow.  Multipro-

gramming and multiprocessing. IEEE Spectrum, 1(3):193–198, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cutler:1964:BRI

[Cut64a] C. C. Cutler. Book review: Introduction to space communication systems, George N. Krassner and Jackson V. Michaels. IEEE Spectrum, 1(6):204, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cutrona:1964:OCT

[Cut64b] L. J. Cutrona. Optical computing techniques. IEEE Spectrum, 1(10):101–108, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cutler:1967:SLa

[Cut67a] C. C. Cutler. Spectral lines.

IEEE Spectrum, 4(4):45, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cutler:1967:SLb

[Cut67b] C. C. Cutler. Spectral lines.

IEEE Spectrum, 4(5):39, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cutler:1967:SLc

[Cut67c] C. C. Cutler. Spectral lines. IEEE Spectrum, 4(6):47, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cutler:1967:SLd

[Cut67d] C. C. Cutler. Spectral lines. IEEE Spectrum, 4(8):69, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Daggett:1968:PSP

[Dag68] E. H. Daggett. A power supply for pulsed power welding. IEEE Spectrum, 5(1):67–70, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Damon:1967:SSM

[Dam67] R. W. Damon. Solid-state microwave delay lines. IEEE Spectrum, 4(6):87–92, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dillard:1969:FRB

[DBW69] J. K. Dillard, C. J. Baldwin, and N. H. Woodley. The future role of breeder reactors in utility planning. IEEE Spectrum, 6(3): 100–107, March 1969. CODEN IEESAM. ISSN 1939-9340.

DeSimone:1968:EI

[De 68] D. V. De Simone. Education for innovation. IEEE Spectrum, 5 (1):83–89, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeLange:1968:OHD

[DeL68] O. E. DeLange. Optical heterodyne detection. IEEE Spectrum, 5(10):77–85, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dent:1967:DE

[Den67] J. Dent. Diagnostic engineering.

IEEE Spectrum, 4(7):99–104, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dimeff:1967:SDD

[DGH67] J. Dimeff, W. D. Gunter, and R. J. Hruby. Spectral dependence of deep-space communications capability. IEEE Spectrum, 4(9): 98–104, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Diamond:1969:DPS

[Dia69] J. M. Diamond. A double phasesensitive detector for bridge balancing. IEEE Spectrum, 6(6): 62–70, June 1969. CODEN IEESAM. ISSN 1939-9340.

Dibner:1964:JCM

[Dib64]  B. Dibner.  James Clerk

Maxwell. IEEE Spectrum, 1(12): 50–56, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dibner:1967:MFC

[Dib67] B. Dibner. Michael Faraday — a centennial. IEEE Spectrum, 4 (8):114–119, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dix:1968:GTA

[Dix68] J. P. Dix. Game-theoretic applications. IEEE Spectrum, 5 (4):108–117, April 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dalziel:1969:LEC

[DL69] C. F. Dalziel and W. R. Lee. Lethal electric currents. IEEE Spectrum, 6(2):44–50, February 1969. CODEN IEESAM. ISSN 1939-9340.

Donald:1964:C

[DMMS64] D. K. Donald, H. F. Matare, B. F. Miessner, and J. A. Spady. Correspondence. IEEE Spectrum, 1(12):156–157, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dobrin:1968:OPE

[Dob68] M. B. Dobrin. Optical processing in the earth sciences. IEEE Spectrum, 5(9):59–66, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dobkin:1969:LC

[Dob69] R. C. Dobkin. Logarithmic converters. IEEE Spectrum, 6(11): 69–72, November 1969. CODEN IEESAM. ISSN 1939-9340.

Dominy:1969:EAP

[Dom69] F. E. Dominy. Economic aspects of the Pacific Northwest– Southwest Intertie. IEEE Spectrum, 6(2):65–71, February 1969. CODEN IEESAM. ISSN 1939-9340.

Dougal:1966:IDC

[Dou66] A. A. Dougal. International developments in controlled thermonuclear fusion. IEEE Spectrum, 3(2):87–93, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Deloraine:1965:APC

[DR65] E. M. Deloraine and A. H.

Reeves. The 25th anniversary of pulse code modulation. IEEE Spectrum, 2(5):56–63, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dunning:1965:DSE

[Dun65]  J. R. Dunning.  Do scien-

tists and engineers need each other? IEEE Spectrum, 2 (10):70–72, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Durkee:1965:BRI

[Dur65] A. L. Durkee. Book review: Ionospheric Radio Propagation, Kenneth Davies, Ed. IEEE Spectrum, 2(9):178–181, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ewald:1964:RIE

[EA64] E. Ewald and D. W. Angland. Regional integration of electric power systems. IEEE Spectrum, 1(4):96–101, April 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Erdelyi:1967:CUU

[EB67] E. A. Erdelyi and F. S. Barnes. Cooperation of universities and utilities for the education of power-system engineers. IEEE Spectrum, 4(12):71–76, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ebeling:1964:MEP

[Ebe64] D. G. Ebeling. Maintaining engineering proficiency. IEEE Spectrum, 1(6):118–119, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Everitt:1964:SCE

[EBO^{+}64] W. L. Everitt, L. M. K. Boelter, R. R. O’Neill, G. E. Moore, E. C. Gentle, D. G. Ebeling, and H. K. Work. A symposium on continuing education. IEEE Spectrum, 1(6):112, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Edwards:1964:BVC

[EC64] P. G. Edwards and J. Clapper. Better vocoders are coming. IEEE Spectrum, 1(9):119– 129, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ehrenreich:1965:OPM

[Ehr65] H. Ehrenreich. The optical properties of metals. IEEE Spectrum, 2(3):162–170, March 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Einspruch:1966:USS

[Ein66] N. G. Einspruch. Ultrasonic studies of solids. IEEE Spectrum, 3(3):116–124, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eckert:1969:PSR

[EK69] R. P. Eckert and P. M. Kelly.

Public safety, the radio spectrum, and the President’s Task Force on Communications Policy. IEEE Spectrum, 6(1):37–46, January 1969. CODEN IEESAM. ISSN 1939-9340.

Eleccion:1968:XRC

[Ele68]  M. Eleccion.  X radiation

from color television receivers. IEEE Spectrum, 5(7):95–104, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Elfant:1965:RIC

[Elf65] R. F. Elfant. Report on the 1965 INTERMAG Conference. IEEE Spectrum, 2(7):118–119, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Elliott:1966:RE

[Ell66] R. S. Elliott. Relativity and electricity. IEEE Spectrum, 3 (3):140–152, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Edgerton:1969:IEN

[EMK69] H. E. Edgerton, V. E. Mac-Roberts, and M. Khanna. Improvements in electronics for nature photography. IEEE Spectrum, 6(7):89–94, July 1969. CO-DEN IEESAM. ISSN 1939-9340.

Ende:1966:RCI

[End66] A. H. Ende. The role of comsat in international communications. IEEE Spectrum, 3(10): 77–79, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Esaki:1966:NHS

[Esa66] L. Esaki. New horizons in semimetal alloys. IEEE Spectrum, 3(2):74–86, February 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Evanzia:1969:LAE

[Eva69] W. J. Evanzia. A look at Apollo electronics. IEEE Spectrum, 6 (9):81–86, September 1969. CO-DEN IEESAM. ISSN 1939-9340.

Everitt:1964:I

[Eve64] W. L. Everitt. Introduction. IEEE Spectrum, 1(6):112–114, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frosch:1964:ISF

[FAB64] R. A. Frosch, V. C. Anderson, and H. Bradner. Instrumenting the sea floor. IEEE Spectrum, 1(11):101–104, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fano:1965:MSC

[Fan65] R. M. Fano. The MAC system:

the computer utility approach. IEEE Spectrum, 2(1):56–64, January 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fein:1964:AI

[Fei64a] L. Fein. The artificial intelligentsia. IEEE Spectrum, 1 (2):74–89, February 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Feist:1964:RTE

[Fei64b] W. Feist. Research in tunnel emission. IEEE Spectrum, 1(12): 57–66, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Felix:1964:GEC

[Fel64] F. Felix. Growth of energy consumption and national income throughout the world. IEEE Spectrum, 1(7):81–102, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frame:1964:BR

[FFC64] J. S. Frame, L. Fein, and M. Chodorow. Book reviews. IEEE Spectrum, 1(12): 154–155, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Finn:1966:CGE

[Fin66] B. S. Finn. Controversy and the growth of the electrical art. IEEE Spectrum, 3(1):52–56, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fink:1967:IHP

[Fin67] D. G. Fink. “IEEE Headquarters” — people, facilities, and functions. IEEE Spectrum, 4 (12):92–99, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fink:1969:MTE

[Fin69] D. G. Fink. Market trends in the electronics industry. IEEE Spectrum, 6(12):57–60, December 1969. CODEN IEESAM. ISSN 1939-9340.

Frame:1964:MFAb

[FK64] J. S. Frame and H. E. Koenig. Matrix functions and applications: Part III. Applications of matrices to systems analysis. IEEE Spectrum, 1(5):100–109, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frye:1966:ACA

[FK66] E. O. Frye and D. E. Killham. Aircraft collision avoidance systems. IEEE Spectrum, 3(1):72–80, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fine:1964:LIB

[FKS64] S. Fine, E. Klein, and R. E.

Scott. Laser irradiation of biological systems. IEEE Spectrum, 1(4):81–95, April 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Franklin:1966:TDC

[FL66] R. H. Franklin and H. B. Law.

Trends in digital communication by wire. IEEE Spectrum, 3(11): 52–58, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Flock:1968:MBM

[Flo68] W. L. Flock. Monitoring bird movements by radar. IEEE Spectrum, 5(6):62–66, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fenton:1969:EH

[FO69] R. E. Fenton and K. W. Olson.

The electronic highway. IEEE Spectrum, 6(7):60–66, July 1969. CODEN IEESAM. ISSN 1939-9340.

Forrester:1964:CFU

[For64]  J. W. Forrester.  Common

foundations underlying engineering and management. IEEE Spectrum, 1(9):66–77, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Forshay:1968:TNP

[For68] R. L. Forshay. TVA nondomestic purchases by a government utility. IEEE Spectrum, 5(9): 100–105, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fowler:1964:TFC

[Fow64] A. B. Fowler. Thin-film circuit technology part III. Active thinfilm devices. IEEE Spectrum, 1(6):102–111, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frame:1964:MFAa

[Fra64a] J. S. Frame. Matrix functions and applications: Part II. Functions of a matrix. IEEE Spectrum, 1(4):102–108, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frame:1964:MFAc

[Fra64b] J. S. Frame. Matrix functions and applications: Part IV. Matrix functions and constituent matrices. IEEE Spectrum, 1 (6):123–131, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frame:1964:MFAd

[Fra64c] J. S. Frame. Matrix functions and applications: Part V. Similarity reductions by rational or orthogonal matrices. IEEE Spectrum, 1(7):103–109, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frame:1964:PMO

[Fra64d] J. S. Frame. Part I: Matrix operations and generalized inverses. IEEE Spectrum, 1(3): 209–220, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:1964:BR

[Fra64e] B. H. Frank. Book reviews.

IEEE Spectrum, 1(5):168–171, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Franklin:1966:CWC

[Fra66] R. H. Franklin. Compatibility in world communications. IEEE Spectrum, 3(10):73–77, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1964:EHH

[Fri64a] G. D. Friedlander. Electric heating a hot line of systems. IEEE Spectrum, 1(11):84–100, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1964:PSA

[Fri64b] G. D. Friedlander. Pumped storage an answer to peaking power. IEEE Spectrum, 1(10): 58–75, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1964:QQT

[Fri64c] G. D. Friedlander. The ’Quoddy question time and tide. IEEE Spectrum, 1(9):96–118, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1965:AAI

[Fri65a] G. D. Friedlander. Airborne asphyxia an international problem. IEEE Spectrum, 2(10): 56–69, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1965:CCPa

[Fri65b] G. D. Friedlander. Computercontrolled power systems. Part I — boiler-turbine unit controls. IEEE Spectrum, 2(4): 60–81, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1965:CCPb

[Fri65c] G. D. Friedlander. Computercontrolled power systems. Part II — area controls and load dispatch. IEEE Spectrum, 2 (5):72–91, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1965:EEE

[Fri65d] G. D. Friedlander. Esthetics and electric energy. IEEE Spectrum, 2(6):46–55, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1965:GGG

[Fri65e] G. D. Friedlander. Giant generators for growing power demand. IEEE Spectrum, 2(2):

70–86, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1965:SSS

[Fri65f] G. D. Friedlander. Science and the salty sea. IEEE Spectrum, 2(8):53–64, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1966:NTS

[Fri66a] G. D. Friedlander. National transportation symposium 9th joint railroad conference. IEEE Spectrum, 3(7):116–121, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1966:NPF

[Fri66b] G. D. Friedlander. The northeast power failure a blanket of darkness. IEEE Spectrum, 3 (2):54–73, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1966:OEF

[Fri66c] G. D. Friedlander. Ocean engineering: food fish for a hungry world. IEEE Spectrum, 3 (11):59–68, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1966:OEP

[Fri66d] G. D. Friedlander. Ocean engineering: probing the depths of a wild frontier. IEEE Spectrum, 3 (10):94–105, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1966:SSI

[Fri66e] G. D. Friedlander. Seismic and sonic instrumentation for ‘ocean engineering’. IEEE Spectrum, 3(9):101–112, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1966:WWI

[Fri66f]  G. D. Friedlander.  World

War II radar: The yellow-green eye. IEEE Spectrum, 3(5):62–71, May 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1967:BNC

[Fri67a] G. D. Friedlander. Birth of the ‘new city’: an exciting creation. IEEE Spectrum, 4 (4):70–82, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1967:RVH

[Fri67b] G. D. Friedlander. Railway vs. highway: the zoom of things to come. IEEE Spectrum, 4(9): 62–76, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1967:WWIb

[Fri67c] G. D. Friedlander. World War II: Electronics and the U.S. Navy magnetic mines, acoustical and homing torpedoes, and proximity fuzes. IEEE Spectrum, 4 (12):46–56, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1967:WWIa

[Fri67d] G. D. Friedlander. World War II: Electronics and the U.S. Navy radar, sonar, loran, and infrared techniques. IEEE Spectrum, 4 (11):56–70, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1968:ACP

[Fri68a] G. D. Friedlander. Automation comes to the printing and publishing industry production and distribution of magazines, periodicals, and books. IEEE Spectrum, 5(5):53–62, May 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1968:PPF

[Fri68b] G. D. Friedlander. Prevention of power failures: The FPC report of 1967. IEEE Spectrum, 5 (2):53–61, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1968:REPb

[Fri68c] G. D. Friedlander. Railroad electrification: past, present, and future competition of the diesel; modern signaling, communications, and safety devices; prospects for the future. IEEE Spectrum, 5(8):56–66, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1968:REPc

[Fri68d] G. D. Friedlander. Railroad electrification: past, present, and future development of the great European systems. IEEE Spectrum, 5(9):77–90, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1968:REPa

[Fri68e] G. D. Friedlander. Railroad electrification: Past, present, and future history of systems in the United States. IEEE Spectrum, 5(7):50–65, July 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1968:SBPa

[Fri68f] G. D. Friedlander. The story of Bonneville Power: 1937–1968– 1987 … dams of the Columbia River Basin. IEEE Spectrum, 5 (11):78–94, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1968:SBPb

[Fri68g] G. D. Friedlander. The story of Bonneville Power: 1937–1968– 1987 … thermal generation, EHV transmission, telecommunications, automation, and service reliability. IEEE Spectrum, 5 (12):61–72, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1969:ATM

[Fri69a] G. D. Friedlander. Art and technology: a merger of disciplines. IEEE Spectrum, 6(10): 60–68, October 1969. CODEN IEESAM. ISSN 1939-9340.

Friedlander:1969:CCV

[Fri69b] G. D. Friedlander. Computercontrolled vehicular traffic. IEEE Spectrum, 6(2):30–43, February 1969. CODEN IEESAM. ISSN 1939-9340.

Fang:1968:EPS

[FRMH68] I. E. Fang, I. Roshwalb, J. Moshman, and D. G. Herzberg. Election projection seminar. IEEE Spectrum, 5(9):40–51, September 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frosch:1969:NLS

[Fro69] R. A. Frosch. A new look at systems engineering. IEEE Spectrum, 6(9):24–28, September 1969. CODEN IEESAM. ISSN 1939-9340.

Frysinger:1969:EFC

[Fry69] G. R. Frysinger. The economical fuel cell. IEEE Spectrum, 6 (3):83–90, March 1969. CODEN IEESAM. ISSN 1939-9340.

Ferrell:1967:SCR

[FS67a] W. R. Ferrell and T. B. Sheridan.

Supervisory control of remote manipulation. IEEE Spectrum, 4 (10):81–88, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fubini:1967:LSS

[FS67b] E. G. Fubini and M. G. Smith.

Limitations in solid-state technology. IEEE Spectrum, 4 (5):55–59, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gabriel:1967:WBG

[Gab67] R. P. Gabriel. Wired broadcasting in great britain. IEEE Spectrum, 4(4):97–105, April 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geher:1966:RPS

[Geh66] K. Geher. Report on Prague’s Summer School on Circuit Theory. IEEE Spectrum, 3(2): 122–123, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gentle:1964:CI

[Gen64] E. C. Gentle. A challenge to industry. IEEE Spectrum, 1 (6):117–118, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Giacoletto:1965:ESC

[Gia65] L. J. Giacoletto. Energy storage and conversion. IEEE Spectrum, 2(2):95–102, February 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Giaever:1966:DT

[Gia66]  I. Giaever.  A dc trans-

former. IEEE Spectrum, 3(9): 117–122, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gibson:1965:CEC

[Gib65] J. E. Gibson. From control engineering to control science. IEEE Spectrum, 2(5):69–71, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Glang:1969:PTT

[GJPM69] R. Glang, K. H. Jaeckel, M. H.

Perkins, and L. I. Maissel. Pulse trimming of thin-film resistors. IEEE Spectrum, 6(8):71–81, August 1969. CODEN IEESAM. ISSN 1939-9340.

Gerber:1969:EOU

[GN69] A. Gerber and B. C. Netschert.

The energy outlook for the United States. IEEE Spectrum, 6(6):38–45, June 1969. CODEN IEESAM. ISSN 1939-9340.

Goldsmith:1964:ISR

[Gol64]  A. N. Goldsmith.  IEEE

Spectrum — retrospect and prospect. IEEE Spectrum, 1(1): 110–113, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldsmith:1965:BE

[Gol65] A. N. Goldsmith. Biomedical engineering. IEEE Spectrum, 2 (12):46–56, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Golden:1966:RSC

[Gol66] A. Golden. Radar separation of closely spaced targets. IEEE Spectrum, 3(2):94–99, February 1966. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Gordon:1968:TGC

[Gor68] C. K. Gordon. The third great crisis in mathematics. IEEE Spectrum, 5(5):47–52, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gossick:1968:SCW

[Gos68] B. R. Gossick. Sir Charles Wheatstone: forgotten genius. IEEE Spectrum, 5(3):80–82, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goubau:1964:WG

[Gou64] G. Goubau. Waveguides general. IEEE Spectrum, 1(10): 81–84, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grems:1964:BR

[Gre64] M. Grems. Book reviews. IEEE Spectrum, 1(7):184, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greber:1966:PE

[Gre66] H. Greber. The philosophy of engineering. IEEE Spectrum, 3 (10):112–115, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greber:1967:SPL

[Gre67a] H. Greber. The stormproof power line. IEEE Spectrum, 4(4):61–69, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greenwood:1967:RSB

[Gre67b] T. L. Greenwood. The radio spectrum below 550 kHz. IEEE Spectrum, 4(3):121–123, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grebene:1969:SVC

[Gre69] A. B. Grebene. A sinusoidal voltage-controlled oscillator for integrated circuits. IEEE Spectrum, 6(3):79–82, March 1969. CODEN IEESAM. ISSN 1939-9340.

Grimes:1966:EQT

[Gri66] D. M. Grimes. Electromagnetism and quantum theory. IEEE Spectrum, 3(5):55–61, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grossman:1965:NIA

[Gro65] M. Grossman. On the nature of information an application of entropy. IEEE Spectrum, 2(12): 70–80, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grove:1966:ISI

[Gro66] A. C. Grove. International standardization interface with the future. IEEE Spectrum, 3(8): 91–101, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gunther:1966:TSC

[Gun66] F. A. Gunther. Tropospheric scatter communications past, present, and future. IEEE Spectrum, 3(9):79–100, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Guthart:1965:BRA

[Gut65] H. Guthart. Book review: Astrophysics and space science: an integration of sciences, Allen J. Mcmahon. IEEE Spectrum, 2 (9):176, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gutzwiller:1967:TRD

[Gut67] F. W. Gutzwiller. Thyristors and rectifier diodes-the semiconductor workhorses. IEEE Spectrum, 4(8):102–111, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Griffin:1964:C

[GW64] P. H. Griffin and E. G. Warren. Correspondence. IEEE Spectrum, 1(2):172–175, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hackborn:1965:DCF

[Hac65] R. Hackborn. Discounted-cash-flow analysis of the return on engineering investment. IEEE Spectrum, 2(10):86–95, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Haggerty:1964:EII

[Hag64a] P. E. Haggerty. The economic impact of integrated circuitry. IEEE Spectrum, 1 (6):80–82, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Haggerty:1964:I

[Hag64b] P. E. Haggerty. Introduction. IEEE Spectrum, 1(6):63, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Haggerty:1968:ET

[Hag68] P. E. Haggerty. Electronics in transition. IEEE Spectrum, 5(5):44–46, May 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Halpern:1967:FCN

[Hal67] M. Halpern. Foundations of the case for natural-language programming. IEEE Spectrum, 4 (3):140–149, March 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hamer:1969:CCH

[Ham69] D. W. Hamer. Ceramic capacitors for hybrid integrated circuits. IEEE Spectrum, 6(1): 79–84, January 1969. CODEN IEESAM. ISSN 1939-9340.

Harnden:1965:PDS

[Har65] J. D. Harnden. Present-day solid-state power switches. IEEE Spectrum, 2(9):107–112, September 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Harris:1966:NLN

[Har66] F. K. Harris. A nonmagnetic laboratory for the National Bureau of Standards. IEEE Spectrum, 3 (11):85–87, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harmuth:1969:AWF

[Har69] H. F. Harmuth. Applications of Walsh functions in communications. IEEE Spectrum, 6(11): 82–91, November 1969. CODEN IEESAM. ISSN 1939-9340.

Haun:1968:LA

[Hau68]  R. D. Haun.  Laser appli-

cations. IEEE Spectrum, 5 (5):82–92, May 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hays:1964:PCS

[Hay64] J. B. Hays. Protecting communication systems from EMP effects of nuclear explosions. IEEE Spectrum, 1(5):115–122, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hammond:1969:CRD

[HB69] D. L. Hammond and A. Benjaminson. The crystal resonatora digital transducer. IEEE Spectrum, 6(4):53–58, April 1969. CODEN IEESAM. ISSN 1939-9340.

Hays:1964:C

[HBK64] J. B. Hays, L. E. Bernier, and J. E. Karlin. Correspondence. IEEE Spectrum, 1(6):206–209, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herold:1968:THI

[HE68] E. W. Herold and M. Eleccion.

Technical highlights of the 1968 IEEE Convention. IEEE Spectrum, 5(3):62–69, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heartz:1964:AST

[Hea64] R. A. Heartz. Analog simulation techniques. IEEE Spectrum, 1 (3):221–230, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Healy:1969:CR

[Hea69] T. J. Healy. Convolution revisited. IEEE Spectrum, 6(4):87–93, April 1969. CODEN IEESAM. ISSN 1939-9340.

Heenan:1964:BW

[Hee64] N. I. Heenan. The beam waveguide. IEEE Spectrum, 1(10): 84–86, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hempstead:1966:MPU

[Hem66] C. F. Hempstead. Motion perception using oscilloscope display. IEEE Spectrum, 3(9):

128–135, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Henney:1965:BRA

[Hen65]  K. Henney.  Book review:

Alessandro Volta and the electric battery [by Bern Dibner, Franklin Watts, Inc., New York, NY, USA, 1964]. IEEE Spectrum, 2 (9):176, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herold:1965:FET

[Her65] E. W. Herold. The future of the electron tube. IEEE Spectrum, 2(1):50–55, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herwald:1968:IMA

[Her68a] S. W. Herwald. The IEEE membership attitude survey: what it means to you. IEEE Spectrum, 5 (12):33–38, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herwald:1968:IR

[Her68b] S. W. Herwald. IEEE reports for 1967. IEEE Spectrum, 5 (6):83–95, June 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herwald:1968:MP

[Her68c] S. W. Herwald. Message from the president. IEEE Spectrum, 5(1):46, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hubert:1965:HWP

[HG65] F. J. Hubert and M. R. Gent.

Half-wavelength power transmission lines. IEEE Spectrum, 2 (1):87–92, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Humphrey:1966:ISC

[HHA^{+}66] L. E. Humphrey, R. C. Hess, G. I.

Addis, A. E. Ruprecht, P. H. Ware, E. J. Steeve, J. A. Schneider, I. F. Matthysse, and E. M. Scoran. Insulated sodium conductors a future trend. IEEE Spectrum, 3(11):73–81, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hull:1964:C

[HHE^{+}64] R. W. Hull, R. P. Haviland, W. W. Everett, O. N. Minot, and H. O. Stasis. Correspondence. IEEE Spectrum, 1(4):164–173, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Haggerty:1964:IC

[HHN^{+}64] P. E. Haggerty, C. L. Hogan, R. N. Noyce, L. C. Maier, J. E. Brown, and C. H. Knowles. Integrated circuits. IEEE Spectrum, 1(6):62, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hicks:1966:FES

[Hic66] B. C. Hicks. The future of energy supply. IEEE Spectrum, 3 (10):82–84, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hicks:1967:DCC

[Hic67] K. L. Hicks. Disaster control coordination for large interconnected systems. IEEE Spectrum, 4(11):52–55, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Higinbotham:1966:PCI

[Hig66] W. A. Higinbotham. Purdue Conference on Instrumentation for High-Energy Physics. IEEE Spectrum, 3(1):103–104, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hintzman:1969:CTH

[Hin69] F. H. Hintzman. Computer tuning of hybrid audio oscillators. IEEE Spectrum, 6(2):56– 60, February 1969. CODEN IEESAM. ISSN 1939-9340.

Hirsch:1967:SAB

[Hir67] C. J. Hirsch. Some aspects of binaural sound. IEEE Spectrum, 4 (2):80–85, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hirsch:1968:CTS

[Hir68a] C. J. Hirsch. Color television standards for region 2. IEEE Spectrum, 5(2):62–67, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hirschfeld:1968:LIC

[Hir68b] R. A. Hirschfeld. Linear integrated circuits in communication systems. IEEE Spectrum, 5 (11):71–77, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hallows:1968:EH

[HK68] R. L. Hallows and R. J. Klensch.

Electronic halftones. IEEE Spectrum, 5(10):64–72, October 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hochman:1969:ABL

[Hoc69] H. T. Hochman. The art of building LSIs. IEEE Spectrum, 6(9): 29–36, September 1969. CODEN IEESAM. ISSN 1939-9340.

Hoeppner:1966:IRT

[Hoe66] C. H. Hoeppner. Industrial radio telemetry. IEEE Spectrum, 3 (10):89–93, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hoffman:1968:VSC

[Hof68] H. L. Hoffman. VITA — a sociotechnical challenge for the IEEE. IEEE Spectrum, 5(8): 82–89, August 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hogan:1964:TIC

[Hog64] C. L. Hogan. Types of integrated circuits. IEEE Spectrum, 1(6):63–71, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hollomon:1964:NST

[Hol64] J. H. Hollomon. A national science and technology policy. IEEE Spectrum, 1(8):89–97, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hollomon:1967:NPI

[Hol67] J. H. Hollomon. New patterns of industry-government partnership. IEEE Spectrum, 4(5):78–82, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hormann:1964:HCS

[Hor64] A. Hormann. How a computer system can learn. IEEE Spectrum, 1(7):110–119, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hornzig:1966:CC

[Hor66] D. F. Hornzig. Communications and civilization. IEEE Spectrum, 3(5):43–45, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horan:1968:SIIa

[Hor68a] J. J. Horan. Spacecraft infrared imaging principles and applications. IEEE Spectrum, 5 (6):71–75, June 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horan:1968:SIIb

[Hor68b] J. J. Horan. Spacecraft infrared imaging systems engineering aspects. IEEE Spectrum, 5(7):66– 74, July 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Howell:1969:PS

[How69] R. P. Howell. On professional salaries. IEEE Spectrum, 6(2): 22–29, February 1969. CODEN IEESAM. ISSN 1939-9340.

Herbstreit:1967:ISC

[HP67] J. W. Herbstreit and H. Pouliquen.

International standards for color television. IEEE Spectrum, 4 (3):104–111, March 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hillman:1968:HBL

[HS68] D. K. Hillman and G. E. Smith.

The high-brightness LED. IEEE Spectrum, 5(1):62–66, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Huang:1965:PPT

[Hua65] T. S. Huang. PCM picture transmission. IEEE Spectrum, 2(12): 57–63, December 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Huelsman:1969:DLA

[Hue69] L. P. Huelsman. The distributedlumped-active network its application to filtering problems. IEEE Spectrum, 6(8):51–58, August 1969. CODEN IEESAM. ISSN 1939-9340.

Huntley:1964:NCL

[Hun64]  W. H. Huntley.  New co-

herent light diffraction techniques. IEEE Spectrum, 1(1): 114–122, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ikegami:1969:SSM

[IN69] F. Ikegami and Y. Ninomiya. Solid-state microwave relay systems in japan. IEEE Spectrum, 6 (12):48–56, December 1969. CO-DEN IEESAM. ISSN 1939-9340.

Jacobson:1969:HPH

[Jac69] L. Jacobson. How to prototype hybrid-circuit patterns and screens at budget prices. IEEE Spectrum, 6(7):82–88, July 1969. CODEN IEESAM. ISSN 1939-9340.

James:1965:DTA

[Jam65] R. T. James. Data transmission — the art of moving information. IEEE Spectrum, 2 (1):65–83, January 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

James:1966:HSI

[Jam66] R. T. James. High-speed information channels. IEEE Spectrum, 3(4):79–95, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jansky:1964:CSS

[Jan64] C. M. Jansky. The current state of the science of marine navigation. IEEE Spectrum, 1(10):

109–118, October 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jordan:1964:DBA

[JDDM64] E. C. Jordan, G. A. Deschamps, J. D. Dyson, and P. E. Mayes. Developments in broadband antennas. IEEE Spectrum, 1 (4):58–71, April 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1967:ENT

[Joh67] H. W. Johnson. Educating for the new technology. IEEE Spectrum, 4(8):99–101, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaiser:1966:SSH

[Kai66] F. D. Kaiser. Solid-state HV dc transmission technology. IEEE Spectrum, 3(11):48–51, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaminow:1965:PPO

[Kam65] I. P. Kaminow. Parametric principles in optics. IEEE Spectrum, 2(4):35–43, April 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kawakami:1966:EFN

[Kaw66] T. Kawakami. Electrical features of the new tokaido line. IEEE Spectrum, 3(1):57–63, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaufman:1964:CI

[KDP64] A. B. Kaufman, L. Dwon, and A. F. Pacela. Correspondence [3 items]. IEEE Spectrum, 1(9): 196–201, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Klein:1968:SMT

[KE68] M. Klein and H. Edgerton. Sonar a modern technique for ocean exploitation. IEEE Spectrum, 5(6):40–46, June 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Keenleyside:1965:CAC

[Kee65]  H. L. Keenleyside.  Cana-

dian aspects of the Columbia Treaty. IEEE Spectrum, 2(9): 95–97, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kelly:1967:CPC

[Kel67] P. M. Kelly. Communications, the police, and the crime commission. IEEE Spectrum, 4 (5):83–92, May 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Keyes:1969:PPL

[Key69] R. W. Keyes. Physical problems and limits in computer logic. IEEE Spectrum, 6(5):36–45, May 1969. CODEN IEESAM. ISSN 1939-9340.

Kadet:1964:PE

[KF64] J. Kadet and B. H. Frank. PERT for the engineer. IEEE Spectrum, 1(11):131–137, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Killpatrick:1967:LG

[Kil67]  J. Killpatrick.  The laser

gyro. IEEE Spectrum, 4(10): 44–55, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

King:1965:LWS

[Kin65] D. D. King. Long waves or short perspectives on the millimeter region. IEEE Spectrum, 2(5):64–68, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Keister:1965:NES

[KKV65] W. Keister, R. W. Ketchledge, and H. E. Vaughan. A new electronic switching system. IEEE Spectrum, 2(2):87–94, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kennedy:1964:SDD

[KL64] R. S. Kennedy and I. L. Lebow.

Signal design for dispersive channels. IEEE Spectrum, 1(3): 231–237, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Knapp:1969:PSH

[Kna69] S. R. Knapp. Pumped storage the handmaiden of nuclear power. IEEE Spectrum, 6(4): 46–52, April 1969. CODEN IEESAM. ISSN 1939-9340.

Knowles:1964:RDI

[Kno64] C. H. Knowles. Research and development in integrated circuits. IEEE Spectrum, 1(6): 76–79, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kobayashi:1969:TEJ

[Kob69] K. Kobayashi. Telecommunications and electronics in japan. IEEE Spectrum, 6(1):85–88, January 1969. CODEN IEESAM. ISSN 1939-9340.

Koehler:1965:SSH

[Koe65]  D. Koehler.  Semiconduc-

tor switching at high pulse rates. IEEE Spectrum, 2(11): 50–66, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kohl:1969:CM

[Koh69] W. H. Kohl. The compatibility of materials. IEEE Spectrum, 6(1): 67–74, January 1969. CODEN IEESAM. ISSN 1939-9340.

Kompfner:1967:EDS

[Kom67] R. Kompfner. Electron devices in science and technology. IEEE Spectrum, 4(9):47–52, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kopplin:1969:SCE

[Kop69] J. O. Kopplin. Stimulating change in engineering education. IEEE Spectrum, 6(1):64–66, January 1969. CODEN IEESAM. ISSN 1939-9340.

Kortman:1967:DCR

[Kor67] C. M. Kortman. Data compression by redundancy reduction. IEEE Spectrum, 4(3):133–139, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Korpel:1968:AIH

[Kor68] A. Korpel. Acoustic imaging and holography. IEEE Spectrum, 5 (10):45–52, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kraus:1964:RAR

[Kra64] J. D. Kraus. Recent advances in radio astronomy. IEEE Spectrum, 1(9):78–95, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kroemer:1968:NCS

[Kro68] H. Kroemer. Negative conductance in semiconductors. IEEE Spectrum, 5(1):47–56, January 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Klensch:1969:CPS

[KS69] R. J. Klensch and E. D.

Simshauser. The CRT in phototypesetting systems. IEEE Spectrum, 6(9):75–80, September 1969. CODEN IEESAM. ISSN 1939-9340.

Kusko:1968:PPS

[Kus68] A. Kusko. A prediction of power system development, 1968 to 2030. IEEE Spectrum, 5 (4):75–80, April 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lambiotte:1966:CAW

[Lam66a] M. Lambiotte. The CCITT and automatic worldwide telephony. IEEE Spectrum, 3(10): 70–73, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lamm:1966:PHV

[Lam66b] A. U. Lamm. The peculiarities of high-voltage dc power transmission. IEEE Spectrum, 3(8):76–84, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Landauer:1967:WPP

[Lan67]  R. Landauer.  Wanted: a

physically possible theory of physics. IEEE Spectrum, 4(9): 105–109, September 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lasser:1966:DCO

[Las66] M. E. Lasser. Detection of coherent optical radiation. IEEE Spectrum, 3(4):73–78, April 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lesk:1964:HPS

[LAT^{+}64] I. A. Lesk, C. B. Ackerman, W. E. Taylor, J. C. Haenichen, W. C. Davis, and G. B. Finn. High-power solid-state devices. IEEE Spectrum, 1(1):

93–109, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Laube:1967:EIC

[Lau67] H. L. Laube. Economics of indoor climate control. IEEE Spectrum, 4(8):120–123, August 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Laverick:1968:HFS

[Lav68] C. Laverick. High-field superconductor technology. IEEE Spectrum, 5(4):63–74, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lawrence:1964:RUD

[Law64]  R. F. Lawrence.  Residen-

tial underground distribution. IEEE Spectrum, 1(3):199–207, March 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lax:1965:PSL

[Lax65] B. Lax. Progress in semiconductor lasers. IEEE Spectrum, 2(7):62–75, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levine:1965:BR

[LBCM65] S. Levine, J. L. Bernstein, B. R.

Cooper, and O. Myers. Book reviews. IEEE Spectrum, 2(8): 156–161, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lefkovitz:1964:ALC

[Lef64] D. Lefkovitz. Automation and the library of congress. IEEE Spectrum, 1(8):154, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lender:1966:CLC

[Len66] A. Lender. Correlative level coding for binary-data transmission. IEEE Spectrum, 3(2): 104–115, February 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levine:1966:PS

[Lev66] M. A. Levine. Plasmas in space. IEEE Spectrum, 3(11): 43–47, November 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lewis:1966:CCR

[Lew66] W. D. Lewis. Communications — concept and reality. IEEE Spectrum, 3(11):82–84, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lickson:1968:PCA

[Lic68] C. P. Lickson. Privacy and the computer age. IEEE Spectrum, 5(10):58–63, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Liebhafsky:1966:FCF

[Lie66] H. A. Liebhafsky. Fuel cells and fuel batteries an engineering view. IEEE Spectrum, 3 (12):48–56, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ligomenides:1967:WMU

[Lig67]  P. A. Ligomenides.  Wave-

mechanical uncertainty and speed limitations. IEEE Spectrum, 4(2):65–68, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Linder:1964:IR

[Lin64] C. H. Linder. IEEE reports for 1963. IEEE Spectrum, 1 (7):131–132, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1965:AHE

[Lin65a] N. Lindgren. The artificial heart exemplar of medical engineering enterprise. IEEE Spectrum, 2 (9):67–83, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1965:MRHa

[Lin65b] N. Lindgren. Machine recognition of human language: Part I — automatic speech recognition. IEEE Spectrum, 2(3):114–136, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1965:MRHb

[Lin65c] N. Lindgren. Machine recognition of human language. Part II — theoretical models of speech perception and language. IEEE Spectrum, 2(4):45–59, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1965:MRHc

[Lin65d] N. Lindgren. Machine recognition of human language: Part III — cursive script recognition. IEEE Spectrum, 2(5):104–116, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1966:EES

[Lin66a] N. Lindgren. Earthquake or explosion? The science of nuclear test detection. IEEE Spectrum, 3(8):66–75, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1966:HFEa

[Lin66b]  N. Lindgren.  Human fac-

tors in engineering part i man in the man-made environment. IEEE Spectrum, 3(3):132–139, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1966:HFEb

[Lin66c]  N. Lindgren.  Human fac-

tors in engineering. Part II. Advanced man-machine systems and concepts. IEEE Spectrum, 3(4):62–72, April 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1966:TYD

[Lin66d] N. Lindgren. Ten years of DEW-Line. IEEE Spectrum, 3(12):

57–62, December 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Linvill:1966:EES

[Lin66e] W. K. Linvill. Engineeringeconomic systems: a new profession. IEEE Spectrum, 3 (4):96–102, April 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1967:ECH

[Lin67a] N. Lindgren. Electric cars hope springs eternal. IEEE Spectrum, 4(4):49–60, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1967:FGE

[Lin67b]  N. Lindgren.  Future goals

of engineering in biology and medicine. IEEE Spectrum, 4(11): 93–100, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1968:PSE

[Lin68a] N. Lindgren. Purposive systems: The edge of knowledge. IEEE Spectrum, 5(4):89–100, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1968:UB

[Lin68b] N. Lindgren. To understand brains. IEEE Spectrum, 5(9): 52–58, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1969:ATS

[Lin69a] N. Lindgren. Art and technology I. Steps toward a new synergism. IEEE Spectrum, 6 (4):59–68, April 1969. CODEN IEESAM. ISSN 1939-9340.

Lindgren:1969:ATI

[Lin69b] N. Lindgren. Art and technology II. A call for collaboration. IEEE Spectrum, 6(5):46–56, May 1969. CODEN IEESAM. ISSN 1939-9340.

Lindgren:1969:UM

[Lin69c] N. Lindgren. Ultrasonics in medicine. IEEE Spectrum, 6(11): 48–57, November 1969. CODEN IEESAM. ISSN 1939-9340.

Luce:1965:PPP

[LL65] C. F. Luce and W. A. Lewis. Power progress in the Pacific Northwest. IEEE Spectrum, 2 (9):84–89, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lessor:1964:TFC

[LMT64] A. E. Lessor, L. I. Maissel, and R. E. Thun. Thin-film circuit technology: Part I: Thin-film R-C networks. IEEE Spectrum, 1(4):72–80, April 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lotspeich:1968:ELB

[Lot68] J. F. Lotspeich. Electrooptic light-beam deflection. IEEE Spectrum, 5(2):45–52, February 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Lorenz:1967:SDL

[LP67] M. R. Lorenz and M. H.

Pilkuhn. Semiconductor-diode light sources. IEEE Spectrum, 4(4):87–96, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Luce:1965:WEH

[LPK65] C. F. Luce, T. A. Phillips, and H. L. Keenleyside. Widening energy horizons in the Pacific Northwest. IEEE Spectrum, 2 (9):84, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lebedev:1967:ETU

[LR67] B. P. Lebedev and S. S. Rokotian. EHV transmission in the U.S.S.R. power grid. IEEE Spectrum, 4(2):73–79, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lemmond:1964:EBW

[LS64] C. Q. Lemmond and L. H. Stauffer. Energy beams as working tools. IEEE Spectrum, 1(7):66– 80, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

MacAdam:1964:SL

[Mac64] W. K. MacAdam. Spectral lines.

IEEE Spectrum, 1(10):57, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

MacAdam:1967:ILA

[Mac67a] W. K. MacAdam. IEEE in Latin America. IEEE Spectrum, 4(1):49, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

MacAdam:1967:IR

[Mac67b] W. K. MacAdam. IEEE reports for 1966. IEEE Spectrum, 4(6):93–101, June 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

MacAdam:1967:MYP

[Mac67c] W. K. MacAdam. A message from your president. IEEE Spectrum, 4(6):49, June 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maier:1964:ICI

[Mai64] L. C. Maier. Integrated circuits in industrial equipment. IEEE Spectrum, 1(6):72–75, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maloney:1969:AAR

[Mal69] W. T. Maloney. Acoustooptical approaches to radar signal processing. IEEE Spectrum, 6(10): 40–48, October 1969. CODEN IEESAM. ISSN 1939-9340.

Mark:1966:NAN

[Mar66] H. Mark. New applications for nuclear energy. IEEE Spectrum, 3(1):64–71, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Marko:1967:ITC

[Mar67] H. Marko. Information theory and cybernetics. IEEE Spectrum, 4(11):75–83, November 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mathison:1965:MMT

[Mat65] R. P. Mathison. Mariner Mars 1964 telemetry and command system. IEEE Spectrum, 2(7):76– 84, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

May:1965:UTW

[May65] J. E. May. Ultrasonic travelingwave devices for communications. IEEE Spectrum, 2(10): 73–85, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Murtaugh:1966:FPN

[MC66] S. A. Murtaugh and H. E.

Criel. Fundamentals of proportional navigation. IEEE Spectrum, 3(12):75–85, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McCalmont:1967:MAD

[McC67] A. M. McCalmont. Multipleaccess discrete-address communication systems. IEEE Spectrum, 4(8):87–94, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McLean:1966:WCT

[McL66] F. C. McLean. Worldwide color television standards. IEEE Spectrum, 3(6):59, June 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McLean:1968:EV

[McL68] F. McLean. The European viewpoint. IEEE Spectrum, 5(2): 68–74, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McPherson:1967:DCR

[McP67] J. C. McPherson. Data communication requirements of computer systems. IEEE Spectrum, 4 (12):42–45, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mermelstein:1965:CHR

[ME65] P. Mermelstein and M. Eden. Correspondence: Handwriting recognition. IEEE Spectrum, 2 (9):30, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Meadows:1965:SST

[Mea65] H. E. Meadows. Symposium on signal transmission and processing. IEEE Spectrum, 2 (10):96–97, October 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Melcher:1968:CW

[Mel68]  J. R. Melcher.  Complex

waves. IEEE Spectrum, 5(10):

86–101, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Muhleman:1965:RRAa

[MGC65a] D. O. Muhleman, R. Goldstein, and R. Carpenter. A review of radar astronomy. Part I. IEEE Spectrum, 2(10):44–55, October 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Muhleman:1965:RRAb

[MGC65b] D. O. Muhleman, R. Goldstein, and R. Carpenter. A review of radar astronomy. Part II. IEEE Spectrum, 2(11):78–89, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Michaels:1967:TNB

[Mic67] E. L. Michaels. Today’s need for balanced urban transit systems. IEEE Spectrum, 4(12): 87–91, December 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1965:PTS

[Mil65] D. B. Miller. Plasma thrustors for space propulsion. IEEE Spectrum, 2(6):36–45, June 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Millar:1968:DCS

[Mil68a] R. A. Millar. Digital control of shaft speed and position. IEEE Spectrum, 5(1):90–95, January 1968. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Milsum:1968:TBS

[Mil68b] J. H. Milsum. The technosphere, the biosphere, the sociosphere their systems modeling and optimization. IEEE Spectrum, 5 (6):76–82, June 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miramontes:1965:SCR

[Mir65] F. C. Miramontes. Special conference report: 1965 rural electrification conference. IEEE Spectrum, 2(9):117–119, September 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mitra:1969:SAF

[Mit69] S. K. Mitra. Synthesizing active filters. IEEE Spectrum, 6(1): 47–63, January 1969. CODEN IEESAM. ISSN 1939-9340.

Martin:1965:RIN

[ML65] J. A. Martin and R. S. Livingston. Report on the IEEE– NBS Particle Accelerator Conference. IEEE Spectrum, 2 (7):120–121, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:1967:CSC

[MM67] E. J. Martin and W. S. McKee. Commercial satellite communications experience. IEEE Spectrum, 4(7):63–69, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:1965:C

[MMGN65] R. K. Moore, A. J. Mare, R. W.

Gillette, and J. D. Nixon. Correspondence. IEEE Spectrum, 2 (8):24–28, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Montgomery:1964:SM

[Mon64] D. B. Montgomery. Superconducting magnets. IEEE Spectrum, 1(2):103–113, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Montgomery:1966:HSC

[Mon66]  D. B. Montgomery.  High-

strength conductors for supermagnets. IEEE Spectrum, 3 (8):111–114, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:1964:ES

[Moo64] G. E. Moore. Education for survival. IEEE Spectrum, 1 (6):115–117, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:1967:RCS

[Moo67] R. K. Moore. Radio communication in the sea. IEEE Spectrum, 4 (11):42–51, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:1968:MEB

[Moo68] A. D. Moore. The modern engineering bandwagon. IEEE Spectrum, 5(7):79–82, July 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morton:1965:PF

[Mor65] J. A. Morton. From physics to function. IEEE Spectrum, 2(9): 62–66, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mottram:1965:STC

[Mot65] E. T. Mottram. Submarine telephone cables. IEEE Spectrum, 2(5):96–103, May 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mundie:1966:RTD

[MOV66] J. R. Mundie, H. L. Oestreicher, and H. E. Von Gierke. Real-time digital analysis system for biological data. IEEE Spectrum, 3 (10):116–120, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mullins:1964:FP

[Mul64] F. G. Mullins. Focal points. IEEE Spectrum, 1(8):132–153, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Montgomery:1966:SLL

[MW66] D. R. Montgomery and F. J. Wolf. The surveyor lunar landing television system. IEEE Spectrum, 3(8):54–61, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nagy:1967:PRI

[Nag67] G. Nagy. Pattern Recognition 1966 IEEE Workshop. IEEE Spectrum, 4(2):92–94, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nasser:1968:SBA

[Nas68]  E. Nasser.  Spark break-

down in air at a positive point. IEEE Spectrum, 5(11): 127–134, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Netschert:1969:APE

[Net69] B. C. Netschert. Air pollution and electric power. IEEE Spectrum, 6(7):71–76, July 1969. CO-DEN IEESAM. ISSN 1939-9340.

Nievergelt:1968:CCP

[Nie68] J. Nievergelt. Computers and computing — past present future. IEEE Spectrum, 5(1): 57–61, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nodelman:1965:BRI

[Nod65] H. M. Nodelman. Book review: Introduction to Boolean algebra and logic design: a program for self-introduction, Gerhard E. Hoernes and Melvin F. Heilweil. IEEE Spectrum, 2(9): 176–178, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Noll:1967:DCC

[Nol67] A. M. Noll. The digital computer as a creative medium. IEEE Spectrum, 4(10):89–95, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nottingham:1964:BR

[Not64] W. B. Nottingham. Book reviews. IEEE Spectrum, 1(3): 300–303, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Noyce:1964:ICM

[Noy64] R. N. Noyce. Integrated circuits in military equipment. IEEE Spectrum, 1(6):71–72, June 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Noyce:1968:MIE

[Noy68] R. N. Noyce. Making integrated electronics technology work. IEEE Spectrum, 5(5): 63–66, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

North:1969:TFP

[NP69] H. Q. North and D. L. Pyke. Technological forecasting in planning for company growth. IEEE Spectrum, 6(1):30–36, January 1969. CODEN IEESAM. ISSN 1939-9340.

Okress:1964:MPE

[OBM^{+}64] E. C. Okress, W. C. Brown, T. Moreno, G. Goubau, N. I.

Heenan, and R. H. George. Microwave power engineering. IEEE Spectrum, 1(10):76, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

OConnell:1969:EEC

[OFM^{+}69] J. D. O’Connell, E. G. Fubini, K. G. McKay, J. Hillier, and J. H. Hollomon. Electronically expanding the citizen’s world. IEEE Spectrum, 6(7):30–40, July 1969. CODEN IEESAM. ISSN 1939-9340.

Okress:1964:I

[Okr64a] E. C. Okress. Introduction.

IEEE Spectrum, 1(10):76, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Okress:1964:S

[Okr64b] E. C. Okress. Summary. IEEE Spectrum, 1(10):98–100, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliver:1964:VSB

[Oli64] B. M. Oliver. The view from the starship bridge and other observations. IEEE Spectrum, 1 (1):86–92, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliver:1965:ILA

[Oli65a] B. M. Oliver. The IEEE looks ahead. IEEE Spectrum, 2(4):34, April 1965. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliver:1965:IR

[Oli65b] B. M. Oliver. IEEE reports for 1964. IEEE Spectrum, 2 (6):86–94, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliver:1967:MRU

[Oli67] B. M. Oliver. Major recommendations of the U.S. President’s Patent Commission. IEEE Spectrum, 4(2):57–64, February 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Olson:1964:SPS

[Ols64] H. F. Olson. Speech processing systems. IEEE Spectrum, 1 (2):90–102, February 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Omi:1966:PCE

[Omi66] H. Omi. Progress of the communications and electronics industry in japan. IEEE Spectrum, 3(9):220–221, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oudin:1966:SDC

[OT66] J. M. Oudin and R. A. Tellier. Submarine dc cables. IEEE Spectrum, 3(7):75–82, July 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Paananen:1966:PIA

[Paa66] R. A. Paananen. Progress in ionized-argon lasers. IEEE Spectrum, 3(6):88–99, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Page:1969:HEA

[Pag69] C. V. Page. Human experience in artificial intelligence. IEEE Spectrum, 6(9):67–74, September 1969. CODEN IEESAM. ISSN 1939-9340.

Paris:1969:FUT

[Par69] L. Paris. The future of UHV transmission lines. IEEE Spectrum, 6(9):44–51, September 1969. CODEN IEESAM. ISSN 1939-9340.

Peattie:1966:HAF

[Pea66] C. G. Peattie. Hydrocarbon air fuel cell systems. IEEE Spectrum, 3(6):69–76, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pease:1967:SEM

[Pea67] R. F. W. Pease. The scanning electron microscope. IEEE Spectrum, 4(10):96–102, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peden:1968:NFE

[Ped68] I. C. Peden. New faces of eve:

Women in electrical engineering. IEEE Spectrum, 5(4):81–84, April 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Petrick:1965:LMM

[Pet65] M. Petrick. Liquid-metal magnetohydrodynamics. IEEE Spectrum, 2(3):137–151, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Phillips:1964:MIC

[Phi64] A. B. Phillips. Monolithic integrated circuits. IEEE Spectrum, 1(6):83–101, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Phillips:1965:EGW

[Phi65] T. A. Phillips. Energy for the growing west. IEEE Spectrum, 2 (9):89–95, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pierce:1966:NMC

[Pie66] J. R. Pierce. New means of communication. IEEE Spectrum, 3 (8):62–65, August 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pierce:1968:MML

[Pie68a] J. R. Pierce. Men, machines, and languages. IEEE Spectrum, 5(7):44–49, July 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pierce:1968:SPA

[Pie68b] J. R. Pierce. Some practical aspects of digital transmission. IEEE Spectrum, 5(11): 63–70, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Popkov:1969:ETS

[Pop69] V. I. Popkov. EHV transmission in the Soviet Union. IEEE Spectrum, 6(2):18–21, February 1969. CODEN IEESAM. ISSN 1939-9340.

Porter:1965:SS

[Por65] R. W. Porter. Science in and for space. IEEE Spectrum, 2(9): 104–106, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pratt:1964:RS

[Pra64] H. Pratt. Report of the Secretary 1963. IEEE Spectrum, 1(7):132–136, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pratt:1968:BRR

[Pra68] H. Pratt. Book review: Radio ancestors: an autobiography by Robert H. Marriott. IEEE Spectrum, 5(6):52–61, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pritchard:1966:STF

[Pri66] J. P. Pritchard. Superconducting thin-film technology and applications. IEEE Spectrum, 3 (5):46–54, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Purington:1967:M

[Pur67] E. S. Purington. Mathematics 10. IEEE Spectrum, 4(8): 112–113, August 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Quittner:1964:C

[QBBB64] G. F. Quittner, A. B. Bronwell, E. W. Baer, and B. W. Bryan. Correspondence. IEEE Spectrum, 1(8):158–163, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Quigley:1964:C

[QDH64] F. V. Quigley, L. M. Duryee, and K. Holum. Correspondence. IEEE Spectrum, 1(11): 186–187, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rajchman:1965:MPF

[Raj65] J. A. Rajchman. Memories in present and future generations of computers. IEEE Spectrum, 2 (11):90–95, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rathjens:1969:AA

[Rat69] G. Rathjens. Against the ABM. IEEE Spectrum, 6(8):40–45, August 1969. CODEN IEESAM. ISSN 1939-9340.

Rawson:1969:VVM

[Raw69] E. G. Rawson. Vibrating varifocal mirrors for 3-D imaging. IEEE Spectrum, 6(9):37– 43, September 1969. CODEN IEESAM. ISSN 1939-9340.

Richardson:1964:UBE

[RB64] J. M. Richardson and J. F.

Brockman. The U.S. basis of electromagnetic measurements. IEEE Spectrum, 1(1): 129–138, January 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reiff:1966:IST

[Rei66] G. A. Reiff. Interplanetary spacecraft telecommunication systems. IEEE Spectrum, 3(4): 103–111, April 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Richards:1967:CRP

[Ric67] P. I. Richards. Computing reliable power spectra. IEEE Spectrum, 4(1):83–90, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rincliffe:1967:POL

[Rin67] R. G. Rincliffe. Planning and operation of a large power pool. IEEE Spectrum, 4(1):91–96, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ritchings:1968:RES

[Rit68] F. A. Ritchings. Raw energy sources for electric generation. IEEE Spectrum, 5(8):34–45, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rathbun:1968:NTI

[RJ68] D. K. Rathbun and H. J. Jensen. Nuclear test instrumentation with miniature superconductive cables. IEEE Spectrum, 5(9): 91–99, September 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reimann:1965:POC

[RK65] O. A. Reimann and W. F. Kosonocky. Progress in optical computer research. IEEE Spectrum, 2(3):181–195, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robinson:1965:RGE

[Rob65] T. A. Robinson. The role of grounding in eliminating electronic interference. IEEE Spectrum, 2(7):85–89, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roberts:1968:SDT

[Rob68a] D. H. Roberts. Silicon device technology. IEEE Spectrum, 5 (2):101–108, February 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Robinson:1968:HEC

[Rob68b] J. A. Robinson. How engineers can communicate more effectively with managers. IEEE Spectrum, 5(6):47–51, June 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roddis:1965:NPW

[Rod65a] L. H. Roddis. Nuclear power in the world today. IEEE Spectrum, 2(7):90–97, July 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rodrigue:1965:MWN

[Rod65b] G. P. Rodrigue. Magnetoelastic waves — new mechanisms for energy transport. IEEE Spectrum, 2(6):81–85, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roizen:1967:UCT

[Roi67] J. Roizen. Universal color television: an electronic fantasia. IEEE Spectrum, 4(3):112–114, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Romeira:1965:C

[Rom65] M. P. Romeira. Correspondence. IEEE Spectrum, 2(9): 30–32, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenberg:1965:ASS

[Ros65] M. Rosenberg. Another Steinmetz solution. IEEE Spectrum, 2 (9):32–47, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosberger:1967:NKK

[Ros67] P. Rosberger. A new keyboard without keys. IEEE Spectrum, 4 (2):86–88, February 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rudin:1967:AEU

[Rud67] H. Rudin. Automatic equalization using transversal filters. IEEE Spectrum, 4(1):53–59, January 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:GEP

[Ryd64a] J. D. Ryder. The genesis of an editorial policy. IEEE Spectrum, 1(6):147–150, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:SLa

[Ryd64b] J. D. Ryder. Spectral lines. IEEE Spectrum, 1(2):57, February 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:SLb

[Ryd64c] J. D. Ryder. Spectral lines. IEEE Spectrum, 1(4):57, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:SLc

[Ryd64d] J. D. Ryder. Spectral lines. IEEE Spectrum, 1(5):73–74, May 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:SLd

[Ryd64e] J. D. Ryder. Spectral lines. IEEE Spectrum, 1(6):61, June 1964.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:SLe

[Ryd64f] J. D. Ryder. Spectral lines. IEEE Spectrum, 1(7):65, July 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:SLf

[Ryd64g] J. D. Ryder. Spectral lines. IEEE Spectrum, 1(8):61, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:SLg

[Ryd64h] J. D. Ryder. Spectral lines. IEEE Spectrum, 1(11):67, November 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ryder:1964:SLh

[Ryd64i] J. D. Ryder. Spectral lines.

IEEE Spectrum, 1(12):49, December 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Salpeter:1965:NEPa

[Sal65a] J. L. Salpeter. On the nature of the electron — Part I. IEEE Spectrum, 2(3):171–180, March 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Salpeter:1965:NEPb

[Sal65b] J. L. Salpeter. On the nature of the electron — Part II. IEEE Spectrum, 2(4):96–101, April 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Salton:1965:PAI

[Sal65c] G. Salton. Progress in automatic information retrieval. IEEE Spectrum, 2(8):90–103, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Savas:1969:FCU

[Sav69] E. S. Savas. Feedback controls on urban air pollution. IEEE Spectrum, 6(7):77–81, July 1969. CO-DEN IEESAM. ISSN 1939-9340.

Sauerland:1968:CIF

[SB68] F. Sauerland and W. Blum. Ceramic IF filters for consumer products. IEEE Spectrum, 5(11): 112–126, November 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schultz:1965:ACH

[Sch65] T. J. Schultz. Acoustics of the concert hall. IEEE Spectrum, 2(6):56–67, June 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schnapf:1969:TD

[Sch69a] A. Schnapf. The tiros decade. IEEE Spectrum, 6(7):53–59, July 1969. CODEN IEESAM. ISSN 1939-9340.

Schroeder:1969:IC

[Sch69b] M. R. Schroeder. Images from computers. IEEE Spectrum, 6 (3):66–78, March 1969. CODEN IEESAM. ISSN 1939-9340.

Scott:1966:ESO

[Sco66] R. M. Scott. Electrooptics in space operation and research. IEEE Spectrum, 3(1):85–88, January 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Scott:1968:USE

[Sco68] R. L. Scott. Utility system effects of space plasmas. IEEE Spectrum, 5(3):93–95, March 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Seidel:1966:FRP

[Sei66]  H. Seidel.  Frequency rela-

tions of parametric interactions. IEEE Spectrum, 3(7):93–105, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Seifert:1968:T

[Sei68] W. W. Seifert. Transportation-1993. IEEE Spectrum, 5(10): 53–57, October 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selby:1966:ICM

[Sel66] M. C. Selby. International comparison of measurements at high frequencies. IEEE Spectrum, 3 (1):89–98, January 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shaw:1967:NPN

[Sha67] M. Shaw. Nuclear power the next decade of development. IEEE Spectrum, 4(10):73–80, October 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapiro:1968:PRA

[Sha68a] I. I. Shapiro. Planetary radar astronomy. IEEE Spectrum, 5 (3):70–79, March 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapiro:1968:TSL

[Sha68b] L. D. Shapiro. Time synchronization from Loran-C. IEEE Spectrum, 5(8):46–55, August 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shepherd:1966:IDI

[She66a] W. G. Shepherd. IEEE dues to be increased in 1967. IEEE Spectrum, 3(6):42–46, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shepherd:1966:ILP

[She66b] W. G. Shepherd. IEEE launches program for new and expanded services. IEEE Spectrum, 3 (10):59–63, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shepherd:1966:IR

[She66c] W. G. Shepherd. IEEE reports for 1965. IEEE Spectrum, 3 (6):100–108, June 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shepherd:1966:UPS

[She66d] W. G. Shepherd. The uses of a professional society. IEEE Spectrum, 3(12):35–38, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shepherd:1968:EEM

[She68] W. G. Shepherd. Education for the engineering mission. IEEE Spectrum, 5(2):94–100, February 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sherman:1969:SEH

[She69] H. Sherman. The Soviet engineer: his life and living. IEEE Spectrum, 6(12):75–80, December 1969. CODEN IEESAM. ISSN 1939-9340.

Shockley:1966:AST

[Sho66] W. Shockley. Articulated science teaching and balanced emphasis. IEEE Spectrum, 3(6):49–58, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Siders:1967:CAD

[Sid67] R. A. Siders. Computer-aided design. IEEE Spectrum, 4(11): 84–92, November 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Siegel:1968:EUC

[Sie68] S. N. Siegel. Electronics and the urban crisis. IEEE Spectrum, 5(5):78–81, May 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Siforov:1965:TDM

[Sif65] V. I. Siforov. Trends in the development of modern radioelectronics. IEEE Spectrum, 2(5):92–95, May 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Skifter:1964:MHR

[Ski64] H. R. Skifter. In memoriam: Hector R. Skifter. IEEE Spectrum, 1(9):5, September 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1966:CSP

[Smi66a] G. P. Smith. Chameleon in the sun: photochromic glass. IEEE Spectrum, 3(12):39–47, December 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1966:CP

[Smi66b] John Smith. The challenge of Pi. IEEE Spectrum, 3(10):5, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smith:1967:ETE

[Smi67] W. S. Smith. Electronics and the traffic engineer. IEEE Spectrum, 4(1):50–52, January 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugata:1969:ICT

[SN69] E. Sugata and T. Namekawa.

Integrated circuits for television receivers. IEEE Spectrum, 6 (5):64–74, May 1969. CODEN IEESAM. ISSN 1939-9340.

Sorensen:1965:ATC

[Sor65] A. A. Sorensen. Aerospace technical conference. IEEE Spectrum, 2(11):104–106, November 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spence:1966:FRS

[Spe66] R. E. Spence. Future role of satellites in air traffic control. IEEE Spectrum, 3(7):112–115, July 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sproul:1965:IAC

[Spr65]  P. T. Sproul.  IEEE An-

nual Communications Convention. IEEE Spectrum, 2(11): 101–103, November 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapiro:1964:PFFa

[SRLF64a] G. Shapiro, G. J. Rogers, O. B.

Laug, and P. M. Fulcomer. Project fist: Fault isolation by semiautomatic techniques: Part I. Basic concept and techniques. IEEE Spectrum, 1(8):98–111, August 1964. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Shapiro:1964:PFFb

[SRLF64b] G. Shapiro, G. J. Rogers, O. B. Laug, and P. M. Fulcomer. Project fist: Fault isolation by semiautomatic techniques: Part II: Detailed instrumentation. IEEE Spectrum, 1(9): 130–144, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sass:1967:CRA

[SSC67] A. R. Sass, W. C. Stewart, and L. S. Cosentino. Cryogenic random-access memories. IEEE Spectrum, 4(7):91–98, July 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Senders:1965:SCR

[SSF65] J. W. Senders, T. B. Sheridan, and W. R. Ferrell. Special conference report: Men and computers sixth annual symposium on human factors in electronics. IEEE Spectrum, 2(8): 104–108, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Simek:1967:HID

[ST67] J. G. Simek and C. J. Tunis. Handprinting input device for computer systems. IEEE Spectrum, 4(7):72–81, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Storm:1969:SSP

[Sto69] H. F. Storm. Solid-state power electronics in the U.S.A. IEEE Spectrum, 6(10):49–59, October 1969. CODEN IEESAM. ISSN 1939-9340.

Strock:1964:E

[Str64] L. W. Strock. Electroluminescence. IEEE Spectrum, 1(11): 68–83, November 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stuehler:1968:HST

[Stu68] J. E. Stuehler. Hardware software tradeoffs in testing. IEEE Spectrum, 5(12):51–56, December 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sundberg:1969:MTF

[Sun69] Y. Sundberg. Magnetic traveling fields for metallurgical processes. IEEE Spectrum, 6(5):79–88, May 1969. CODEN IEESAM. ISSN 1939-9340.

Susskind:1966:OTE

[Sus66] C. Susskind. The origin of the term ‘electronics’. IEEE Spectrum, 3(5):72–79, May 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1967:CED

[Sus67] C. Susskind. Conference on electron device research: A pattern to be copied? IEEE Spectrum, 4(12):100–102, December 1967.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1968:EHEa

[Sus68a] C. Susskind. The early history of electronics I. Electromagnetics before Hertz. IEEE Spectrum, 5(8):90–98, August 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1968:EHEb

[Sus68b] C. Susskind. The early history of electronics II. The experiments of Hertz. IEEE Spectrum, 5 (12):57–60, December 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1969:EHEa

[Sus69a] C. Susskind. The early history of electronics. III. Prehistory of radiotelegraphy. IEEE Spectrum, 6(4):69–74, April 1969. CODEN IEESAM. ISSN 1939-9340.

Susskind:1969:EHEb

[Sus69b] C. Susskind. The early history of electronics. IV. First radiotelegraphy experiments. IEEE Spectrum, 6(8):66–70, August 1969. CODEN IEESAM. ISSN 1939-9340.

Schwoerer:1964:NPT

[SW64] F. Schwoerer and W. F. Witzig. Nuclear power today and tomorrow. IEEE Spectrum, 1 (7):120–130, July 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schlicke:1967:CEF

[SW67] H. M. Schlicke and H. Weidmann. Compatible EMI filters. IEEE Spectrum, 4(10): 59–68, October 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Swinkels:1968:EVP

[Swi68] D. A. J. Swinkels. Electrochemical vehicle power plants. IEEE Spectrum, 5(5):71–77, May 1968. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tanner:1967:NSE

[TA67] R. L. Tanner and M. G. Andreasen. Numerical solution of electromagnetic problems. IEEE Spectrum, 4(9):53–61, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Taylor:1969:FCT

[Tay69] A. S. Taylor. The future of cable TV. IEEE Spectrum, 6(11): 77–81, November 1969. CODEN IEESAM. ISSN 1939-9340.

Teller:1967:PEE

[Tel67]  A. J. Teller.  The philoso-

phy of an engineering educator. IEEE Spectrum, 4(3):124–128, March 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Theile:1966:CCT

[The66] R. Theile. Compatible color transmission and European modulation methods. IEEE Spectrum, 3(6):60–68, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thun:1969:TFT

[Thu69] R. E. Thun. Thick films or thin? IEEE Spectrum, 6(10): 73–79, October 1969. CODEN IEESAM. ISSN 1939-9340.

Tilson:1969:AMR

[Til69a] S. Tilson. The antiballistic missile: Report on the ABM. IEEE Spectrum, 6(8):24–39, August 1969. CODEN IEESAM. ISSN 1939-9340.

Tilson:1969:EWMa

[Til69b]  S. Tilson.  Electricity and

weather modification. I. A survey of scientific relationships. IEEE Spectrum, 6(4):26–45, April 1969. CODEN IEESAM. ISSN 1939-9340.

Tilson:1969:EWMb

[Til69c]  S. Tilson.  Electricity and

weather modification. II. Plans and programs, prospects and purpose. IEEE Spectrum, 6(5): 89–101, May 1969. CODEN IEESAM. ISSN 1939-9340.

Tilson:1969:EWMc

[Til69d]  S. Tilson.  Electricity and

weather modification. III. Pollution problems and social impact. IEEE Spectrum, 6(6): 52–61, June 1969. CODEN IEESAM. ISSN 1939-9340.

Tomiyasu:1964:SCR

[Tom64] K. Tomiyasu. Special conference report: Report from Tokyo on the ICMCI. IEEE Spectrum, 1 (12):96–99, December 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tompkins:1969:PSM

[Tom69] H. E. Tompkins. The 1969 Popov Society Meeting: Some impressions from U.S. delegates. IEEE Spectrum, 6(12):69–74, December 1969. CODEN IEESAM. ISSN 1939-9340.

Townes:1965:NLP

[Tow65] Charles H. Townes. 1964 Nobel lecture: Production of coherent radiation by atoms and molecules. IEEE Spectrum, 2 (8):30–43, August 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tow:1969:SSA

[Tow69] J. Tow. A step-by-step activefilter design. IEEE Spectrum, 6 (12):64–68, December 1969. CO-DEN IEESAM. ISSN 1939-9340.

Trauboth:1969:SEW

[Tra69] H. Trauboth. Space exploration wisdom or folly? IEEE Spectrum, 6(10):36–39, October 1969. CODEN IEESAM. ISSN 1939-9340.

Truitt:1964:HCW

[Tru64] T. D. Truitt. Hybrid computation: What is it? Who needs it? IEEE Spectrum, 1 (6):132–146, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tsu:1967:MPG

[Tsu67] T. C. Tsu. MHD power generators in central stations. IEEE Spectrum, 4(6):59–65, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Turner:1967:HSE

[Tur67] E. M. Turner. The human side of engineering. IEEE Spectrum, 4(7):70–71, July 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tribus:1967:PEE

[TZW^{+}67] M. Tribus, L. A. Zadeh, E. Weber, J. H. Hollomon, T. F. Jones, and R. H. Lyon. Practical experience and engineering education. IEEE Spectrum, 4(5):60–66, May 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Uman:1966:QLS

[Uma66] M. A. Uman. Quantitative lightning spectroscopy. IEEE Spectrum, 3(8):102–110, August 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Umezu:1967:EPE

[UN67] T. Umezu and H. Nakamura. Electric power engineering in japan. IEEE Spectrum, 4(4): 106–114, April 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vadasz:1969:SGT

[VGRM69] L. L. Vadasz, A. S. Grove, T. A.

Rowe, and G. E. Moore. Silicongate technology. IEEE Spectrum, 6(10):28–35, October 1969. CO-DEN IEESAM. ISSN 1939-9340.

Vogely:1967:PUE

[VM67] W. A. Vogely and W. E. Morrison. Patterns of U.S. energy consumption to 1980. IEEE Spectrum, 4(9):81–86, September 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vogely:1969:AUE

[Vog69] W. A. Vogely. Analytical uses of energy balances. IEEE Spectrum, 6(5):57–63, May 1969. CODEN IEESAM. ISSN 1939-9340.

Vollmer:1967:AL

[Vol67] J. Vollmer. Applied lasers. IEEE Spectrum, 4(6):66–70, June 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

vonBraun:1965:TSE

[von65]  W. von Braun.  The time

for space exploration. IEEE Spectrum, 2(12):44–45, December 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vallbona:1966:ELM

[VSG^{+}66] C. Vallbona, W. A. Spencer, L. A. Geddes, W. F. Blose, and J. Canzoneri. Experience with on-line monitoring in critical illness. IEEE Spectrum, 3(9): 136–140, September 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wagner:1969:FNC

[Wag69] R. E. Wagner. Fluidics a new control tool. IEEE Spectrum, 6 (11):58–68, November 1969. CO-DEN IEESAM. ISSN 1939-9340.

Wallmark:1964:FET

[Wal64] J. T. Wallmark. The field-effect transistor an old device with new promise. IEEE Spectrum, 1 (3):182–192, March 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Walsh:1967:EEC

[Wal67] E. M. Walsh. Electrogasdynamic energy conversion. IEEE Spectrum, 4(12):57–62, December 1967. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wald:1969:GSF

[Wal69] G. Wald. A generation in search of a future. IEEE Spectrum, 6 (6):34–37, June 1969. CODEN IEESAM. ISSN 1939-9340.

Warner:1967:CMB

[War67]  R. M. Warner.  Comparing

MOS and bipolar integrated circuits. IEEE Spectrum, 4 (6):50–58, June 1967. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wagoner:1965:CPS

[WAS65] C. D. Wagoner, P. L. Alger, and C. P. Steinmetz. Charles Proteus Steinmetz. IEEE Spectrum, 2(4):82–95, April 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weber:1964:BR

[Web64a] E. Weber. Book reviews. IEEE Spectrum, 1(4):162–163, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weber:1964:TCE

[Web64b] E. Weber. Technological challenges to educating engineers. IEEE Spectrum, 1(10):119– 120, October 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weed:1966:TEE

[Wee66] H. R. Weed. Trends in European engineering education. IEEE Spectrum, 3(3):125–128, March 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:1965:HMR

[Wei65] H. G. Weiss. The haystack microwave research facility. IEEE Spectrum, 2(2):50–69, February 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wei:1966:NTN

[Wei66a] L. Y. Wei. A new theory of nerve conduction. IEEE Spectrum, 3(9):123–127, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiner:1966:UMW

[Wei66b] J. H. Weiner. U.S. military worldwide communications. IEEE Spectrum, 3(10): 80–81, October 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiss:1968:GD

[Wei68] H. Weiss. Galvanomagnetic devices. IEEE Spectrum, 5(1): 75–82, January 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Westfall:1966:CCU

[Wes66] T. B. Westfall. Compatibility — a challenge to universal communications global cables and satellite communication. IEEE Spectrum, 3(10):64–67, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Westfall:1968:ICP

[Wes68] T. B. Westfall. International communications: the problems of progress. IEEE Spectrum, 5 (8):77–81, August 1968. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wooster:1966:CEI

[WGC^{+}66] H. Wooster, P. L. Garvin, L. D. Callimahos, J. C. Lilly, W. O. Davis, and F. J. Heyden. Communication with extraterrestrial intelligence. IEEE Spectrum, 3 (3):153–163, March 1966. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

White:1969:OPE

[Whi69] A. R. White. Organization planning in engineering to meet change. IEEE Spectrum, 6(5): 102–105, May 1969. CODEN IEESAM. ISSN 1939-9340.

Willenrbock:1964:SL

[Wil64a] F. K. Willenrbock. Spectral lines. IEEE Spectrum, 1(9): 65, September 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wilson:1964:TPG

[Wil64b] V. C. Wilson. Thermionic power generation. IEEE Spectrum, 1(5):75–83, May 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Willenbrock:1965:RPS

[Wil65a] F. K. Willenbrock. Report on the A. S. Popov Society meeting. IEEE Spectrum, 2(7):116–117, July 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Willenbrock:1965:SLa

[Wil65b] F. K. Willenbrock. Spectral lines. IEEE Spectrum, 2(8):29, August 1965. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Willenbrock:1965:SLb

[Wil65c] F. K. Willenbrock. Spectral lines. IEEE Spectrum, 2(9): 61, September 1965. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wingert:1966:PWT

[Win66] L. F. Wingert. Progress in worldwide telephone dialing. IEEE Spectrum, 3(10):67–69, October 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weber:1964:RAH

[WL64] E. Weber and C. H. Linder. Report of the ad hoc committee on professional relations policy. IEEE Spectrum, 1(4):115–116, April 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Work:1964:AIC

[Wor64] H. K. Work. An American Institute for continuing education of engineers. IEEE Spectrum, 1(6):119–120, June 1964. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Williams:1969:IMP

[WP69] W. H. Williams and J. H. Parker. An IC medium power voltage regulator. IEEE Spectrum, 6(2): 72–78, February 1969. CODEN IEESAM. ISSN 1939-9340.

Weimer:1969:MSS

[WPS^{+}69] P. K. Weimer, W. S. Pike, G. Sadasiv, F. V. Shallcross, and L. Meray-Horvath. Multielement self-scanned mosaic sensors. IEEE Spectrum, 6(3):52–65, March 1969. CODEN IEESAM. ISSN 1939-9340.

Webb:1966:C

[WR66] G. N. Webb and R. E. Rogers.

The contourograph. IEEE Spectrum, 3(6):77–87, June 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yamada:1966:PPF

[Yam66] T. Yamada. Problems of power frequencies in japan. IEEE Spectrum, 3(9):216, September 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Young:1966:MNW

[YW66] L. R. Young and R. Winblade.

M.I.T. NASA Working Conference on Manual Control. IEEE Spectrum, 3(11):88–93, November 1966. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zimmerman:1969:HES

[Zim69] H. A. Zimmerman. How to extend sampling-oscilloscope versatility. IEEE Spectrum, 6(4): 79–85, April 1969. CODEN IEESAM. ISSN 1939-9340.

Zverev:1966:GAE

[Zve66] A. I. Zverev. The golden anniversary of electric wave filters. IEEE Spectrum, 3(3):129–131, March 1966. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).